Liquid War 6

A unique multiplayer wargame

Documentation for version 0.6.3902 codename “Goliath” 6 May 2015

by Christian Mauduit <ufoot@ufoot.org>

Liquid War 6, a unique multiplayer wargame. Copyright (c) 2005, 2006, 2007, 2008, 2009, 2010, 2011, 2012, 2013, 2014, 2015 Christian Mauduit <ufoot@ufoot.org>

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 In a nutshell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Project status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.2.1 What works, and what does not (yet). . . . . . . . . . . . . . . . . . . . . . 1 1.2.2 What has changed since Liquid War 5.x? . . . . . . . . . . . . . . . . . . 2 1.2.3 Revision history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2.4 Road map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3 How you can help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.1 Help GNU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.2 Todo list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 User’s manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 Mailing lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.1 General discussion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.2 Announcements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.3 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.4 IRC channel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.2 Getting the game . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.1 Download source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.2 Download binaries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.3 GIT repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.4 Daily snapshots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2.5 Check integrity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.1 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.2 Optional libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3.3 Optional tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.3.4 Installing requirements using RPM/DEB packages . . . . . . . . 11 2.3.5 Compiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.4 Extra maps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4.1 The extra maps package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4.2 Install extra maps on GNU/Linux and POSIX systems. … 13 2.4.3 Raw install of extra maps (all-platforms) . . . . . . . . . . . . . . . . . 13

2.5 Troubleshooting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.1 Compilation problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.2 Check installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.3 Problems running the game . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.6 Quick start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.6.1 Quick start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.7 Strategy tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.8 User interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.8.1 A reduced set of keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.8.2 Combining mouse, keyboard and joysticks . . . . . . . . . . . . . . . . 16

2.8.3 Quit with F10. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.9 Solo game. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.9.1 Current state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.9.2 Team profiles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.9.3 Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

2.10 Network games . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.10.1 Choose your “public url” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.10.2 Starting a node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.10.3 Connecting to a node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.10.4 Communities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.10.5 Firewall settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.10.6 Is the game secure? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

2.11 Graphics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.11.1 Standard, high and low resolution . . . . . . . . . . . . . . . . . . . . . . . 21 2.11.2 Display rate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

2.12 Sound & music . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.12.1 Current status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.12.2 The future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

2.13 Config file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.14 Logs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.15 Report bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

3 Hacker’s guide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.1 Designing levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.1 Why is level design so important? . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.2 Format overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.3 Resolution (map size) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.4 Metadata . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.1.5 map.png . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.1.6 layer2.png ... layer7.png . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.1.7 texture.png, texture.jpeg and texture-alpha.jpeg . . . . . . . . . . 25 3.1.8 glue.png and boost.png. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.1.9 danger.png and medicine.png . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.1.10 one-way-<direction>.png. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.1.11 cursor.png and cursor-color.png . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.1.12 rules.xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.1.13 hints.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.1.14 style.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.1.15 teams.xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.1.16 Resampling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.17 Music . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.18 Experience (“exp”). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

3.2 Translating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.2.1 Using gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.2.2 Formatted strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.2.3 Partial translation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

3.3 Architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.3.1 C + Guile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.3.2 Threading and SMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.3.3 Internal libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.4 Memory structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.5 100% predictable algorithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.6 Graphics backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

3.6.1 Modularity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.6.2 List of backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.6.3 How to write a new backend. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.7 Core algorithm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.7.2 Level, game struct, game state and pilot . . . . . . . . . . . . . . . . . . 39 3.7.3 Getting informations about where fighters are . . . . . . . . . . . . 40

3.8 Compilation tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 3.8.1 Advanced ./configure options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 3.8.2 Debian packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.8.3 Red Hat packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.8.4 Microsoft Windows msys/mingw32 port . . . . . . . . . . . . . . . . . . 42 3.8.5 Mac OS X port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.8.6 GP2X port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

3.9 Coding guidelines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.9.1 Project goals reminder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.9.2 Common sense . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.9.3 Unitary tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.9.4 Memory allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.9.5 Private and public interfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.9.6 Commit policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.9.7 Audit the code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

3.10 Using the console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 3.11 Advanced tweaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

3.11.1 Hacking ressources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.11.2 Optimize for speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

3.12 Writing modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.13 Use as a library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.14 Network protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

3.14.1 No server, no client, only nodes. . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.14.2 Out of band messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.14.3 Regular messages overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 3.14.4 Regular control messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 3.14.5 Regular MISS messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.14.6 Regular META messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.14.7 Regular DATA messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.14.8 Other raw technical stuff (WIP) . . . . . . . . . . . . . . . . . . . . . . . . . 59

3.15 Technical HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 3.15.1 Release check-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 3.15.2 Add a new option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 3.15.3 Add a new internal library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 3.15.4 Add a new module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

3.16 Using GNU Arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 3.16.1 About GNU Arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 3.16.2 Getting the latest version from the repository. . . . . . . . . . . . 64 3.16.3 Setting up your own arch repository . . . . . . . . . . . . . . . . . . . . . 64 3.16.4 Synchronizing your repository with upstream releases …. 65 3.16.5 Submitting patches. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.16.6 Recover from broken lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

3.17 Using GIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.17.1 About GIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.17.2 Getting the latest source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 3.17.3 Developper access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 3.17.4 Submitting patches. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

3.18 Jenkins builds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

4 Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

4.1 Basic options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.1 about . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.2 audit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.3 copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.4 credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.5 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 4.1.6 defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.7 help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.8 host. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.9 list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.10 modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.11 pedigree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.12 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.1.13 version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

4.2 Doc options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.1 example-hints-xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.2 example-rules-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.3 example-style-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.4 example-teams-xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.5 list-advanced . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.6 list-aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.7 list-doc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.8 list-funcs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.9 list-graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.2.10 list-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.11 list-input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.12 list-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.13 list-map-hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.14 list-map-rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.15 list-map-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.16 list-map-teams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.17 list-network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.18 list-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 4.2.19 list-players . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.20 list-quick. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.21 list-show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.22 list-sound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.23 list-team-colors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.24 list-weapons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

4.3 Show options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.3.1 show-build-abs-srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.3.2 show-build-bin-id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.3.3 show-build-bugs-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.3.4 show-build-cflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.5 show-build-codename. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.6 show-build-configure-args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.7 show-build-copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.8 show-build-datadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.9 show-build-date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.10 show-build-docdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.11 show-build-enable-allinone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.3.12 show-build-enable-console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.13 show-build-enable-fullstatic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.14 show-build-enable-gcov . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.15 show-build-enable-gprof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.16 show-build-enable-gtk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.17 show-build-enable-instrument. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.18 show-build-enable-mod-caca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.19 show-build-enable-mod-csound . . . . . . . . . . . . . . . . . . . . . . . . . . 73 4.3.20 show-build-enable-mod-gl1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.21 show-build-enable-mod-gles2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.22 show-build-enable-mod-http . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.23 show-build-enable-mod-ogg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.24 show-build-enable-mod-soft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.25 show-build-enable-openmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.26 show-build-enable-optimize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 4.3.27 show-build-enable-paranoid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.28 show-build-enable-profiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.29 show-build-enable-valgrind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.30 show-build-endianness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.31 show-build-gcc-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.32 show-build-gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.33 show-build-gp2x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.34 show-build-home-url. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.35 show-build-host-cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.3.36 show-build-host-os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.37 show-build-hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.38 show-build-includedir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.39 show-build-ldflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.40 show-build-libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.41 show-build-license . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.42 show-build-localedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.43 show-build-mac-os-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.44 show-build-md5sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 4.3.45 show-build-ms-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.46 show-build-package-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.47 show-build-package-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.48 show-build-package-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.49 show-build-package-tarname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.50 show-build-pointer-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.51 show-build-prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.52 show-build-stamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.53 show-build-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.54 show-build-top-srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.55 show-build-unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.56 show-build-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.57 show-build-version-base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.58 show-build-version-major . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.59 show-build-version-minor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.60 show-build-x86 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.61 show-config-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.62 show-cwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.63 show-data-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.64 show-default-config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.65 show-default-data-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.66 show-default-log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.67 show-default-map-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.68 show-default-map-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.69 show-default-mod-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.3.70 show-default-music-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.71 show-default-music-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.72 show-default-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.73 show-default-script-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.74 show-default-user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.75 show-log-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.76 show-map-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.77 show-map-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.3.78 show-mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.79 show-music-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.80 show-music-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.81 show-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.82 show-run-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.83 show-script-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.84 show-user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

4.4 Path options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.4.1 config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.2 data-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.3 log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.4 map-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.5 map-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.4.6 mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.4.7 music-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.4.8 music-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 4.4.9 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.4.10 script-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.4.11 user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

4.5 Players options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.5.1 player1-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 4.5.2 player1-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.5.3 player1-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.5.4 player2-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.5.5 player2-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.5.6 player2-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.5.7 player3-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.5.8 player3-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.5.9 player3-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.5.10 player4-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.5.11 player4-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.5.12 player4-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

4.6 Input options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.6.1 auto-release-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.6.2 click-to-focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.6.3 cursor-sensitivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.6.4 custom-alt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.6.5 custom-ctrl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.6.6 custom-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.6.7 custom-enter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.6.8 custom-esc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.6.9 custom-left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.6.10 custom-pgdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.6.11 custom-pgup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.6.12 custom-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.6.13 custom-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.6.14 double-click-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.6.15 max-cursor-speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.6.16 mouse-sensitivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.6.17 repeat-delay. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 4.6.18 repeat-interval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 4.6.19 use-double-click . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 4.6.20 use-esc-button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 4.6.21 zoom-step. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 4.6.22 zoom-stick-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

4.7 Graphics options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.7.1 capture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.7.2 fullscreen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.7.3 gfx-backend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.7.4 gfx-quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 4.7.5 height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.7.6 width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.7.7 windowed-mode-limit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

4.8 Sound options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.8.1 ambiance-exclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.8.2 ambiance-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.8.3 ambiance-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.8.4 fx-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.8.5 music-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.8.6 snd-backend. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 4.8.7 water-volume. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

4.9 Network options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 4.9.1 bind-ip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 4.9.2 bind-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 4.9.3 broadcast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 4.9.4 cli-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 4.9.5 known-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 4.9.6 node-description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 4.9.7 node-title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 4.9.8 password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 4.9.9 public-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 4.9.10 skip-network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 4.9.11 srv-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

4.10 Map parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.10.1 chosen-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.10.2 force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.10.3 use-cursor-texture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.10.4 use-hints-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 4.10.5 use-music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 4.10.6 use-rules-xml. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 4.10.7 use-style-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 4.10.8 use-teams-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 4.10.9 use-texture. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

4.11 Map rules.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 4.11.1 boost-power. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 4.11.2 color-conflict-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 4.11.3 cursor-pot-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 4.11.4 danger-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.11.5 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.11.6 fighter-attack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.11.7 fighter-defense . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.11.8 fighter-new-health. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 4.11.9 fighter-regenerate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 4.11.10 frags-fade-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 4.11.11 frags-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 4.11.12 frags-to-distribute. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 4.11.13 glue-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 4.11.14 highest-team-color-allowed . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 4.11.15 highest-weapon-allowed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 4.11.16 max-cursor-pot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 4.11.17 max-cursor-pot-offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 4.11.18 max-nb-cursors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 4.11.19 max-nb-nodes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 4.11.20 max-nb-teams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 4.11.21 max-round-delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 4.11.22 max-zone-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 4.11.23 medicine-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 4.11.24 moves-per-round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 4.11.25 nb-attack-tries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 4.11.26 nb-defense-tries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 4.11.27 nb-move-tries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 4.11.28 respawn-delay. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 4.11.29 respawn-position-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 4.11.30 respawn-team . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 4.11.31 round-delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 4.11.32 rounds-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 4.11.33 side-attack-factor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 4.11.34 side-defense-factor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 4.11.35 single-army-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 4.11.36 spread-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 4.11.37 spread-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 4.11.38 spreads-per-round. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 4.11.39 start-blue-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 4.11.40 start-blue-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 4.11.41 start-cyan-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 4.11.42 start-cyan-y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 4.11.43 start-green-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 4.11.44 start-green-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 4.11.45 start-lightblue-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 4.11.46 start-lightblue-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 4.11.47 start-magenta-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 4.11.48 start-magenta-y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 4.11.49 start-orange-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 4.11.50 start-orange-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 4.11.51 start-pink-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 4.11.52 start-pink-y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 4.11.53 start-position-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 4.11.54 start-purple-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 4.11.55 start-purple-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 4.11.56 start-red-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.11.57 start-red-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.11.58 start-yellow-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.11.59 start-yellow-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.11.60 team-profile-blue-aggressive . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.11.61 team-profile-blue-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 4.11.62 team-profile-blue-handicap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 4.11.63 team-profile-blue-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 4.11.64 team-profile-blue-vulnerable . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 4.11.65 team-profile-blue-weapon-alternate-id . . . . . . . . . . . . . . . . . 116 4.11.66 team-profile-blue-weapon-id . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 4.11.67 team-profile-blue-weapon-mode. . . . . . . . . . . . . . . . . . . . . . . . 116 4.11.68 team-profile-cyan-aggressive . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 4.11.69 team-profile-cyan-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 4.11.70 team-profile-cyan-handicap. . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 4.11.71 team-profile-cyan-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 4.11.72 team-profile-cyan-vulnerable . . . . . . . . . . . . . . . . . . . . . . . . . . 117 4.11.73 team-profile-cyan-weapon-alternate-id . . . . . . . . . . . . . . . . . 117 4.11.74 team-profile-cyan-weapon-id. . . . . . . . . . . . . . . . . . . . . . . . . . . 118 4.11.75 team-profile-cyan-weapon-mode . . . . . . . . . . . . . . . . . . . . . . . 118 4.11.76 team-profile-green-aggressive . . . . . . . . . . . . . . . . . . . . . . . . . . 118 4.11.77 team-profile-green-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 4.11.78 team-profile-green-handicap . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 4.11.79 team-profile-green-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 4.11.80 team-profile-green-vulnerable. . . . . . . . . . . . . . . . . . . . . . . . . . 119 4.11.81 team-profile-green-weapon-alternate-id . . . . . . . . . . . . . . . . 119 4.11.82 team-profile-green-weapon-id . . . . . . . . . . . . . . . . . . . . . . . . . . 119 4.11.83 team-profile-green-weapon-mode . . . . . . . . . . . . . . . . . . . . . . 120 4.11.84 team-profile-lightblue-aggressive . . . . . . . . . . . . . . . . . . . . . . . 120 4.11.85 team-profile-lightblue-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 4.11.86 team-profile-lightblue-handicap. . . . . . . . . . . . . . . . . . . . . . . . 120 4.11.87 team-profile-lightblue-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . 120 4.11.88 team-profile-lightblue-vulnerable . . . . . . . . . . . . . . . . . . . . . . 121 4.11.89 team-profile-lightblue-weapon-alternate-id . . . . . . . . . . . . . 121 4.11.90 team-profile-lightblue-weapon-id. . . . . . . . . . . . . . . . . . . . . . . 121 4.11.91 team-profile-lightblue-weapon-mode . . . . . . . . . . . . . . . . . . . 121 4.11.92 team-profile-magenta-aggressive . . . . . . . . . . . . . . . . . . . . . . . 122 4.11.93 team-profile-magenta-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 4.11.94 team-profile-magenta-handicap . . . . . . . . . . . . . . . . . . . . . . . . 122 4.11.95 team-profile-magenta-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . 122 4.11.96 team-profile-magenta-vulnerable. . . . . . . . . . . . . . . . . . . . . . . 122 4.11.97 team-profile-magenta-weapon-alternate-id . . . . . . . . . . . . . 123 4.11.98 team-profile-magenta-weapon-id . . . . . . . . . . . . . . . . . . . . . . . 123 4.11.99 team-profile-magenta-weapon-mode . . . . . . . . . . . . . . . . . . . 123 4.11.100 team-profile-orange-aggressive. . . . . . . . . . . . . . . . . . . . . . . . 123 4.11.101 team-profile-orange-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 4.11.102 team-profile-orange-handicap. . . . . . . . . . . . . . . . . . . . . . . . . 124 4.11.103 team-profile-orange-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 4.11.104 team-profile-orange-vulnerable . . . . . . . . . . . . . . . . . . . . . . . 124 4.11.105 team-profile-orange-weapon-alternate-id . . . . . . . . . . . . . . 124 4.11.106 team-profile-orange-weapon-id . . . . . . . . . . . . . . . . . . . . . . . 125 4.11.107 team-profile-orange-weapon-mode . . . . . . . . . . . . . . . . . . . . 125 4.11.108 team-profile-pink-aggressive . . . . . . . . . . . . . . . . . . . . . . . . . . 125 4.11.109 team-profile-pink-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 4.11.110 team-profile-pink-handicap . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 4.11.111 team-profile-pink-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 4.11.112 team-profile-pink-vulnerable . . . . . . . . . . . . . . . . . . . . . . . . . 126 4.11.113 team-profile-pink-weapon-alternate-id . . . . . . . . . . . . . . . . 126 4.11.114 team-profile-pink-weapon-id. . . . . . . . . . . . . . . . . . . . . . . . . . 126 4.11.115 team-profile-pink-weapon-mode . . . . . . . . . . . . . . . . . . . . . . 127 4.11.116 team-profile-purple-aggressive . . . . . . . . . . . . . . . . . . . . . . . . 127 4.11.117 team-profile-purple-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 4.11.118 team-profile-purple-handicap . . . . . . . . . . . . . . . . . . . . . . . . . 127 4.11.119 team-profile-purple-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 4.11.120 team-profile-purple-vulnerable. . . . . . . . . . . . . . . . . . . . . . . . 128 4.11.121 team-profile-purple-weapon-alternate-id . . . . . . . . . . . . . . 128 4.11.122 team-profile-purple-weapon-id. . . . . . . . . . . . . . . . . . . . . . . . 128 4.11.123 team-profile-purple-weapon-mode . . . . . . . . . . . . . . . . . . . . 128 4.11.124 team-profile-red-aggressive . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 4.11.125 team-profile-red-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 4.11.126 team-profile-red-handicap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 4.11.127 team-profile-red-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 4.11.128 team-profile-red-vulnerable. . . . . . . . . . . . . . . . . . . . . . . . . . . 130 4.11.129 team-profile-red-weapon-alternate-id . . . . . . . . . . . . . . . . . 130 4.11.130 team-profile-red-weapon-id . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 4.11.131 team-profile-red-weapon-mode . . . . . . . . . . . . . . . . . . . . . . . 130 4.11.132 team-profile-yellow-aggressive . . . . . . . . . . . . . . . . . . . . . . . . 130 4.11.133 team-profile-yellow-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 4.11.134 team-profile-yellow-handicap . . . . . . . . . . . . . . . . . . . . . . . . . 131 4.11.135 team-profile-yellow-mobile . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 4.11.136 team-profile-yellow-vulnerable. . . . . . . . . . . . . . . . . . . . . . . . 131 4.11.137 team-profile-yellow-weapon-alternate-id . . . . . . . . . . . . . . 132 4.11.138 team-profile-yellow-weapon-id . . . . . . . . . . . . . . . . . . . . . . . . 132 4.11.139 team-profile-yellow-weapon-mode . . . . . . . . . . . . . . . . . . . . 132 4.11.140 total-armies-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 4.11.141 total-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 4.11.142 use-team-profiles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 4.11.143 vertical-move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 4.11.144 weapon-charge-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 4.11.145 weapon-charge-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 4.11.146 weapon-duration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 4.11.147 weapon-tune-berzerk-power . . . . . . . . . . . . . . . . . . . . . . . . . . 134 4.11.148 weapon-tune-turbo-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 4.11.149 x-polarity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 4.11.150 y-polarity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 4.11.151 z-polarity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

4.12 Map hints.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 4.12.1 background-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 4.12.2 downsize-using-bench-value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 4.12.3 downsize-using-fighter-scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 4.12.4 fighter-scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 4.12.5 guess-colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 4.12.6 guess-moves-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 4.12.7 hud-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 4.12.8 max-map-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 4.12.9 max-map-surface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 4.12.10 max-map-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 4.12.11 menu-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 4.12.12 min-map-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 4.12.13 min-map-surface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 4.12.14 min-map-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 4.12.15 resample. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 4.12.16 speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 4.12.17 system-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 4.12.18 upsize-using-bench-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 4.12.19 upsize-using-fighter-scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 4.12.20 view-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 4.12.21 wall-grease. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

4.13 Map style.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 4.13.1 animation-density . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 4.13.2 animation-speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 4.13.3 background-color-root-bg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 4.13.4 background-color-root-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 4.13.5 background-color-stuff-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 4.13.6 background-color-stuff-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 4.13.7 background-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 4.13.8 blink-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 4.13.9 color-alternate-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 4.13.10 color-alternate-fg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 4.13.11 color-base-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 4.13.12 color-base-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 4.13.13 colorize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 4.13.14 colorize-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 4.13.15 cursor-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 4.13.16 hidden-layer-alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 4.13.17 hud-color-frame-bg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 4.13.18 hud-color-frame-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.13.19 hud-color-text-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.13.20 hud-color-text-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.13.21 hud-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.13.22 keep-ratio. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.13.23 menu-color-default-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 4.13.24 menu-color-default-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 4.13.25 menu-color-disabled-bg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 4.13.26 menu-color-disabled-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 4.13.27 menu-color-selected-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 4.13.28 menu-color-selected-fg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 4.13.29 menu-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 4.13.30 music-exclude. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 4.13.31 music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 4.13.32 music-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 4.13.33 pixelize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 4.13.34 system-color-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 4.13.35 system-color-fg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 4.13.36 team-color-blue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 4.13.37 team-color-cyan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 4.13.38 team-color-dead. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 4.13.39 team-color-green . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 4.13.40 team-color-lightblue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 4.13.41 team-color-magenta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 4.13.42 team-color-orange. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 4.13.43 team-color-pink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 4.13.44 team-color-purple . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 4.13.45 team-color-red . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 4.13.46 team-color-yellow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.13.47 view-color-cursor-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.13.48 view-color-cursor-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.13.49 view-color-map-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.13.50 view-color-map-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.13.51 view-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 4.13.52 waves. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 4.13.53 x-wrap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 4.13.54 y-wrap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 4.13.55 zoom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 4.13.56 zoom-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 4.13.57 zoom-min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

4.14 Map teams.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 4.14.1 bot-iq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 4.14.2 bot-speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 4.14.3 bot1-ai. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 4.14.4 bot1-color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 4.14.5 bot2-ai. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 4.14.6 bot2-color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 4.14.7 bot3-ai. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 4.14.8 bot3-color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 4.14.9 bot4-ai. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 4.14.10 bot4-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 4.14.11 bot5-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 4.14.12 bot5-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 4.14.13 bot6-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 4.14.14 bot6-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 4.14.15 bot7-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 4.14.16 bot7-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 4.14.17 bot8-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.14.18 bot8-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.14.19 bot9-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.14.20 bot9-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.14.21 nb-bots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.14.22 player1-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 4.14.23 player2-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 4.14.24 player3-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 4.14.25 player4-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

4.15 Advanced settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 4.15.1 base64-decode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 4.15.2 base64-encode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 4.15.3 bench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 4.15.4 bench-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 4.15.5 bin-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 4.15.6 check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 4.15.7 commands-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4.15.8 cunit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4.15.9 daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4.15.10 debug-layer-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4.15.11 debug-team-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4.15.12 demo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 4.15.13 dialog-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 4.15.14 dirty-read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 4.15.15 display-background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 4.15.16 display-console. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 4.15.17 display-cursors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 4.15.18 display-debug-gradient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 4.15.19 display-debug-zones . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 4.15.20 display-fighters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 4.15.21 display-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 4.15.22 display-hud . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 4.15.23 display-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 4.15.24 display-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 4.15.25 display-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 4.15.26 display-meta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.15.27 display-mouse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.15.28 display-mps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.15.29 display-preview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.15.30 display-progress. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.15.31 display-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 4.15.32 display-splash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 4.15.33 display-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 4.15.34 executed-again . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 4.15.35 gfx-cpu-usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 4.15.36 gfx-debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 4.15.37 io-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 4.15.38 jpeg-quality. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 4.15.39 loader-sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 4.15.40 local-bench-delta. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 4.15.41 log-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 4.15.42 log-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 4.15.43 magic-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 4.15.44 max-local-bench-value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 4.15.45 max-network-bench-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 4.15.46 memory-bazooka-eraser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 4.15.47 memory-bazooka-size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 4.15.48 net-log. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 4.15.49 net-per-sec. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 4.15.50 network-bench-delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 4.15.51 network-reliability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 4.15.52 open-relay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 4.15.53 pilot-lag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 4.15.54 quick-start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 4.15.55 reset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.56 reset-config-on-upgrade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.57 screenshots-per-min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.58 server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.59 simulate-basic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.60 simulate-full . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 4.15.61 target-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 4.15.62 trap-errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 4.15.63 trojan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 4.15.64 z-decode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 4.15.65 z-encode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174

4.16 C to Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.1 c-gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.2 c-lw6-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.3 c-lw6-get-ret . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.4 c-lw6-release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.5 c-lw6-set-ret . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.6 c-lw6bot-get-backends. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.7 c-lw6bot-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.8 c-lw6bot-next-move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 4.16.9 c-lw6cfg-defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.10 c-lw6cfg-get-option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.11 c-lw6cfg-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.12 c-lw6cfg-load. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.13 c-lw6cfg-option-exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.14 c-lw6cfg-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.15 c-lw6cfg-save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.16 c-lw6cfg-set-option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.17 c-lw6cfg-unified-get-log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.18 c-lw6cfg-unified-get-map-path . . . . . . . . . . . . . . . . . . . . . . . . . 175 4.16.19 c-lw6cfg-unified-get-music-path. . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.20 c-lw6cfg-unified-get-user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.21 c-lw6cli-get-backends. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.22 c-lw6cns-console-support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.23 c-lw6cns-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.24 c-lw6cns-poll. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.25 c-lw6cns-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.26 c-lw6cns-term-support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.27 c-lw6dsp-get-average-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.28 c-lw6dsp-get-fullscreen-modes . . . . . . . . . . . . . . . . . . . . . . . . . 176 4.16.29 c-lw6dsp-get-instant-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.30 c-lw6dsp-get-last-frame-rendering-time . . . . . . . . . . . . . . . . 177 4.16.31 c-lw6dsp-get-nb-frames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.32 c-lw6dsp-get-video-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.33 c-lw6dsp-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.34 c-lw6dsp-release. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.35 c-lw6dsp-update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.36 c-lw6gen-create-from-seed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.37 c-lw6gen-seed-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.38 c-lw6gen-seed-normalize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 4.16.39 c-lw6gfx-get-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.40 c-lw6gui-default-look . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.41 c-lw6gui-input-reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.42 c-lw6gui-joystick1-get-move-pad . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.43 c-lw6gui-joystick1-pop-button-a . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.44 c-lw6gui-joystick1-pop-button-b . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.45 c-lw6gui-joystick1-pop-button-c . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.46 c-lw6gui-joystick1-pop-button-d . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.47 c-lw6gui-joystick1-pop-button-e . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.48 c-lw6gui-joystick1-pop-button-f. . . . . . . . . . . . . . . . . . . . . . . . 178 4.16.49 c-lw6gui-joystick1-pop-pad-down . . . . . . . . . . . . . . . . . . . . . . 179 4.16.50 c-lw6gui-joystick1-pop-pad-left . . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.51 c-lw6gui-joystick1-pop-pad-right. . . . . . . . . . . . . . . . . . . . . . . 179 4.16.52 c-lw6gui-joystick1-pop-pad-up . . . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.53 c-lw6gui-joystick2-get-move-pad . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.54 c-lw6gui-joystick2-pop-button-a . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.55 c-lw6gui-joystick2-pop-button-b . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.56 c-lw6gui-joystick2-pop-button-c . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.57 c-lw6gui-joystick2-pop-button-d . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.58 c-lw6gui-joystick2-pop-button-e . . . . . . . . . . . . . . . . . . . . . . . 179 4.16.59 c-lw6gui-joystick2-pop-button-f. . . . . . . . . . . . . . . . . . . . . . . . 180 4.16.60 c-lw6gui-joystick2-pop-pad-down . . . . . . . . . . . . . . . . . . . . . . 180 4.16.61 c-lw6gui-joystick2-pop-pad-left . . . . . . . . . . . . . . . . . . . . . . . . 180 4.16.62 c-lw6gui-joystick2-pop-pad-right. . . . . . . . . . . . . . . . . . . . . . . 180 4.16.63 c-lw6gui-joystick2-pop-pad-up . . . . . . . . . . . . . . . . . . . . . . . . . 180 4.16.64 c-lw6gui-keyboard-get-move-pad. . . . . . . . . . . . . . . . . . . . . . . 180 4.16.65 c-lw6gui-keyboard-is-pressed . . . . . . . . . . . . . . . . . . . . . . . . . . 180 4.16.66 c-lw6gui-keyboard-pop-arrow-down . . . . . . . . . . . . . . . . . . . . 180 4.16.67 c-lw6gui-keyboard-pop-arrow-left . . . . . . . . . . . . . . . . . . . . . . 180 4.16.68 c-lw6gui-keyboard-pop-arrow-right . . . . . . . . . . . . . . . . . . . . 180 4.16.69 c-lw6gui-keyboard-pop-arrow-up . . . . . . . . . . . . . . . . . . . . . . 181 4.16.70 c-lw6gui-keyboard-pop-key-alt. . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.71 c-lw6gui-keyboard-pop-key-ctrl . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.72 c-lw6gui-keyboard-pop-key-enter . . . . . . . . . . . . . . . . . . . . . . 181 4.16.73 c-lw6gui-keyboard-pop-key-esc . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.74 c-lw6gui-keyboard-pop-key-pgdown. . . . . . . . . . . . . . . . . . . . 181 4.16.75 c-lw6gui-keyboard-pop-key-pgup . . . . . . . . . . . . . . . . . . . . . . 181 4.16.76 c-lw6gui-look-get. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.77 c-lw6gui-look-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.78 c-lw6gui-look-zoom-in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 4.16.79 c-lw6gui-look-zoom-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.80 c-lw6gui-menu-append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.81 c-lw6gui-menu-close-popup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.82 c-lw6gui-menu-enable-esc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.83 c-lw6gui-menu-has-popup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.84 c-lw6gui-menu-new. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.85 c-lw6gui-menu-remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.86 c-lw6gui-menu-remove-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.87 c-lw6gui-menu-scroll-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.88 c-lw6gui-menu-scroll-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 4.16.89 c-lw6gui-menu-select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.90 c-lw6gui-menu-select-esc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.91 c-lw6gui-menu-set-breadcrumbs . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.92 c-lw6gui-menu-sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.93 c-lw6gui-mouse-get-state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.94 c-lw6gui-mouse-poll-move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.95 c-lw6gui-mouse-pop-button-left. . . . . . . . . . . . . . . . . . . . . . . . 183 4.16.96 c-lw6gui-mouse-pop-button-middle . . . . . . . . . . . . . . . . . . . . 183 4.16.97 c-lw6gui-mouse-pop-button-right . . . . . . . . . . . . . . . . . . . . . . 183 4.16.98 c-lw6gui-mouse-pop-double-click. . . . . . . . . . . . . . . . . . . . . . . 183 4.16.99 c-lw6gui-mouse-pop-simple-click . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.100 c-lw6gui-mouse-pop-triple-click. . . . . . . . . . . . . . . . . . . . . . . 184 4.16.101 c-lw6gui-mouse-pop-wheel-down. . . . . . . . . . . . . . . . . . . . . . 184 4.16.102 c-lw6gui-mouse-pop-wheel-up . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.103 c-lw6hlp-about. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.104 c-lw6hlp-get-default-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.105 c-lw6hlp-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.106 c-lw6hlp-list-advanced . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.107 c-lw6hlp-list-aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.108 c-lw6hlp-list-doc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 4.16.109 c-lw6hlp-list-funcs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.110 c-lw6hlp-list-graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.111 c-lw6hlp-list-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.112 c-lw6hlp-list-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.113 c-lw6hlp-list-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.114 c-lw6hlp-list-map-hints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.115 c-lw6hlp-list-map-rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.116 c-lw6hlp-list-map-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.117 c-lw6hlp-list-map-teams. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.118 c-lw6hlp-list-network. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 4.16.119 c-lw6hlp-list-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.120 c-lw6hlp-list-players. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.121 c-lw6hlp-list-quick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.122 c-lw6hlp-list-show. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.123 c-lw6hlp-list-sound. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.124 c-lw6hlp-list-team-colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.125 c-lw6hlp-list-weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.126 c-lw6img-screenshot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.127 c-lw6ker-add-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.128 c-lw6ker-build-game-state . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 4.16.129 c-lw6ker-build-game-struct . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.130 c-lw6ker-cursor-exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.131 c-lw6ker-did-cursor-win . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.132 c-lw6ker-do-round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.133 c-lw6ker-dup-game-state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.134 c-lw6ker-game-state-checksum. . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.135 c-lw6ker-game-struct-checksum. . . . . . . . . . . . . . . . . . . . . . . 187 4.16.136 c-lw6ker-get-cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.137 c-lw6ker-get-moves. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.138 c-lw6ker-get-nb-colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 4.16.139 c-lw6ker-get-nb-cursors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.140 c-lw6ker-get-nb-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.141 c-lw6ker-get-rounds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.142 c-lw6ker-get-spreads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.143 c-lw6ker-is-over . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.144 c-lw6ker-node-exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.145 c-lw6ker-register-node. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.146 c-lw6ker-remove-cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.147 c-lw6ker-set-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.148 c-lw6ker-sync-game-state. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 4.16.149 c-lw6ker-unregister-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.150 c-lw6ldr-chain-entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.151 c-lw6ldr-exp-validate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.152 c-lw6ldr-get-entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.153 c-lw6ldr-hints-get-default . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.154 c-lw6ldr-print-examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.155 c-lw6ldr-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.156 c-lw6ldr-read-relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 4.16.157 c-lw6map-exp-get-unlocked-team-color . . . . . . . . . . . . . . . 189 4.16.158 c-lw6map-exp-get-unlocked-weapon . . . . . . . . . . . . . . . . . . 189 4.16.159 c-lw6map-exp-is-team-color-allowed . . . . . . . . . . . . . . . . . . 190 4.16.160 c-lw6map-exp-is-weapon-allowed . . . . . . . . . . . . . . . . . . . . . 190 4.16.161 c-lw6map-get-look . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.162 c-lw6map-get-max-nb-colors. . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.163 c-lw6map-get-max-nb-cursors . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.164 c-lw6map-get-max-nb-nodes. . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.165 c-lw6map-get-music-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.166 c-lw6map-get-title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.167 c-lw6map-param-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.168 c-lw6map-rules-get-default . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 4.16.169 c-lw6map-rules-get-int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.170 c-lw6map-rules-get-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.171 c-lw6map-rules-get-min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.172 c-lw6map-style-get-default . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.173 c-lw6map-team-color-index-to-key . . . . . . . . . . . . . . . . . . . . 191 4.16.174 c-lw6map-team-color-index-to-label . . . . . . . . . . . . . . . . . . 191 4.16.175 c-lw6map-team-color-key-to-index . . . . . . . . . . . . . . . . . . . . 191 4.16.176 c-lw6map-team-color-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.177 c-lw6map-teams-get-default . . . . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.178 c-lw6map-weapon-index-to-key . . . . . . . . . . . . . . . . . . . . . . . 191 4.16.179 c-lw6map-weapon-index-to-label . . . . . . . . . . . . . . . . . . . . . 192 4.16.180 c-lw6map-weapon-key-to-index . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.181 c-lw6map-weapon-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.182 c-lw6net-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.183 c-lw6net-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.184 c-lw6p2p-db-default-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.185 c-lw6p2p-db-new. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.186 c-lw6p2p-db-reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.187 c-lw6p2p-node-calibrate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.188 c-lw6p2p-node-client-join. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 4.16.189 c-lw6p2p-node-close. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.190 c-lw6p2p-node-disconnect . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.191 c-lw6p2p-node-get-entries . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.192 c-lw6p2p-node-get-id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.193 c-lw6p2p-node-get-local-seq-0 . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.194 c-lw6p2p-node-get-local-seq-last . . . . . . . . . . . . . . . . . . . . . . 193 4.16.195 c-lw6p2p-node-get-next-draft-msg . . . . . . . . . . . . . . . . . . . . 193 4.16.196 c-lw6p2p-node-get-next-reference-msg . . . . . . . . . . . . . . . . 193 4.16.197 c-lw6p2p-node-get-seq-draft . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.198 c-lw6p2p-node-get-seq-max. . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4.16.199 c-lw6p2p-node-get-seq-min . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.200 c-lw6p2p-node-get-seq-reference . . . . . . . . . . . . . . . . . . . . . . 194 4.16.201 c-lw6p2p-node-is-dump-needed . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.202 c-lw6p2p-node-is-peer-connected . . . . . . . . . . . . . . . . . . . . . 194 4.16.203 c-lw6p2p-node-is-peer-registered. . . . . . . . . . . . . . . . . . . . . . 194 4.16.204 c-lw6p2p-node-is-seed-needed . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.205 c-lw6p2p-node-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.206 c-lw6p2p-node-poll. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.207 c-lw6p2p-node-put-local-msg . . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.208 c-lw6p2p-node-refresh-peer . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 4.16.209 c-lw6p2p-node-server-start . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.210 c-lw6p2p-node-update-info . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.211 c-lw6pil-bench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.212 c-lw6pil-build-pilot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.213 c-lw6pil-calibrate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.214 c-lw6pil-commit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.215 c-lw6pil-did-cursor-win. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.216 c-lw6pil-dump-command-generate . . . . . . . . . . . . . . . . . . . . 195 4.16.217 c-lw6pil-execute-command . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.218 c-lw6pil-fix-coords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 4.16.219 c-lw6pil-fix-coords-x10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.220 c-lw6pil-get-last-commit-seq. . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.221 c-lw6pil-get-looser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.222 c-lw6pil-get-max-seq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.223 c-lw6pil-get-next-seq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.224 c-lw6pil-get-reference-current-seq . . . . . . . . . . . . . . . . . . . . . 196 4.16.225 c-lw6pil-get-reference-target-seq . . . . . . . . . . . . . . . . . . . . . . 196 4.16.226 c-lw6pil-get-round-0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.227 c-lw6pil-get-seq-0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.228 c-lw6pil-get-winner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 4.16.229 c-lw6pil-is-over. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.230 c-lw6pil-local-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.231 c-lw6pil-local-cursors-set-main . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.232 c-lw6pil-local-cursors-set-mouse-controlled . . . . . . . . . . . . 197 4.16.233 c-lw6pil-make-backup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.234 c-lw6pil-poll-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.235 c-lw6pil-round2seq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.236 c-lw6pil-seed-command-generate . . . . . . . . . . . . . . . . . . . . . 197 4.16.237 c-lw6pil-send-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.238 c-lw6pil-seq-random-0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.16.239 c-lw6pil-seq2round . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.240 c-lw6pil-slow-down. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.241 c-lw6pil-speed-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.242 c-lw6pil-suite-get-checkpoint . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.243 c-lw6pil-suite-get-commands-by-node-index. . . . . . . . . . . 198 4.16.244 c-lw6pil-suite-get-commands-by-stage . . . . . . . . . . . . . . . . 198 4.16.245 c-lw6pil-suite-get-node-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.246 c-lw6pil-suite-get-seq-0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.247 c-lw6pil-suite-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.248 c-lw6pil-sync-from-backup . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 4.16.249 c-lw6pil-sync-from-draft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.250 c-lw6pil-sync-from-reference . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.251 c-lw6snd-get-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.252 c-lw6snd-is-music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.253 c-lw6snd-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.254 c-lw6snd-play-fx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.255 c-lw6snd-play-music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.256 c-lw6snd-play-music-random . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.257 c-lw6snd-poll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.258 c-lw6snd-release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 4.16.259 c-lw6snd-set-fx-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.260 c-lw6snd-set-music-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.261 c-lw6snd-set-water-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.262 c-lw6snd-stop-music . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.263 c-lw6srv-get-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.264 c-lw6sys-build-get-abs-srcdir . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.265 c-lw6sys-build-get-bin-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.266 c-lw6sys-build-get-bugs-url . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.267 c-lw6sys-build-get-cflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.268 c-lw6sys-build-get-codename . . . . . . . . . . . . . . . . . . . . . . . . . 200 4.16.269 c-lw6sys-build-get-configure-args . . . . . . . . . . . . . . . . . . . . . 201 4.16.270 c-lw6sys-build-get-copyright. . . . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.271 c-lw6sys-build-get-datadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.272 c-lw6sys-build-get-date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.273 c-lw6sys-build-get-docdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.274 c-lw6sys-build-get-enable-allinone . . . . . . . . . . . . . . . . . . . . 201 4.16.275 c-lw6sys-build-get-enable-console . . . . . . . . . . . . . . . . . . . . . 201 4.16.276 c-lw6sys-build-get-enable-fullstatic . . . . . . . . . . . . . . . . . . . 201 4.16.277 c-lw6sys-build-get-enable-gcov . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.278 c-lw6sys-build-get-enable-gprof . . . . . . . . . . . . . . . . . . . . . . . 201 4.16.279 c-lw6sys-build-get-enable-gtk. . . . . . . . . . . . . . . . . . . . . . . . . 202 4.16.280 c-lw6sys-build-get-enable-instrument . . . . . . . . . . . . . . . . . 202 4.16.281 c-lw6sys-build-get-enable-mod-caca. . . . . . . . . . . . . . . . . . . 202 4.16.282 c-lw6sys-build-get-enable-mod-csound . . . . . . . . . . . . . . . . 202 4.16.283 c-lw6sys-build-get-enable-mod-gl1 . . . . . . . . . . . . . . . . . . . . 202 4.16.284 c-lw6sys-build-get-enable-mod-gles2 . . . . . . . . . . . . . . . . . . 202 4.16.285 c-lw6sys-build-get-enable-mod-http. . . . . . . . . . . . . . . . . . . 202 4.16.286 c-lw6sys-build-get-enable-mod-ogg . . . . . . . . . . . . . . . . . . . 202 4.16.287 c-lw6sys-build-get-enable-mod-soft . . . . . . . . . . . . . . . . . . . 202 4.16.288 c-lw6sys-build-get-enable-openmp . . . . . . . . . . . . . . . . . . . . 202 4.16.289 c-lw6sys-build-get-enable-optimize . . . . . . . . . . . . . . . . . . . 203 4.16.290 c-lw6sys-build-get-enable-paranoid . . . . . . . . . . . . . . . . . . . 203 4.16.291 c-lw6sys-build-get-enable-profiler . . . . . . . . . . . . . . . . . . . . . 203 4.16.292 c-lw6sys-build-get-enable-valgrind . . . . . . . . . . . . . . . . . . . . 203 4.16.293 c-lw6sys-build-get-endianness . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.294 c-lw6sys-build-get-gcc-version . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.295 c-lw6sys-build-get-home-url . . . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.296 c-lw6sys-build-get-host-cpu . . . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.297 c-lw6sys-build-get-host-os . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.298 c-lw6sys-build-get-hostname . . . . . . . . . . . . . . . . . . . . . . . . . 203 4.16.299 c-lw6sys-build-get-includedir . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.300 c-lw6sys-build-get-ldflags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.301 c-lw6sys-build-get-libdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.302 c-lw6sys-build-get-license . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.303 c-lw6sys-build-get-localedir. . . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.304 c-lw6sys-build-get-md5sum. . . . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.305 c-lw6sys-build-get-package-id. . . . . . . . . . . . . . . . . . . . . . . . . 204 4.16.306 c-lw6sys-build-get-package-name . . . . . . . . . . . . . . . . . . . . . 204 4.16.307 c-lw6sys-build-get-package-string . . . . . . . . . . . . . . . . . . . . . 204 4.16.308 c-lw6sys-build-get-package-tarname . . . . . . . . . . . . . . . . . . 204 4.16.309 c-lw6sys-build-get-pointer-size. . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.310 c-lw6sys-build-get-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.311 c-lw6sys-build-get-stamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.312 c-lw6sys-build-get-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.313 c-lw6sys-build-get-top-srcdir . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.314 c-lw6sys-build-get-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.315 c-lw6sys-build-get-version-base . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.316 c-lw6sys-build-get-version-major . . . . . . . . . . . . . . . . . . . . . 205 4.16.317 c-lw6sys-build-get-version-minor . . . . . . . . . . . . . . . . . . . . . 205 4.16.318 c-lw6sys-build-is-gnu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 4.16.319 c-lw6sys-build-is-gp2x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.320 c-lw6sys-build-is-mac-os-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.321 c-lw6sys-build-is-ms-windows. . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.322 c-lw6sys-build-is-unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.323 c-lw6sys-build-is-x86 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.324 c-lw6sys-debug-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.325 c-lw6sys-debug-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.326 c-lw6sys-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.327 c-lw6sys-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.328 c-lw6sys-dump-clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 4.16.329 c-lw6sys-generate-id-16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.330 c-lw6sys-generate-id-32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.331 c-lw6sys-generate-id-64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.332 c-lw6sys-get-config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.333 c-lw6sys-get-cwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.334 c-lw6sys-get-cycle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.335 c-lw6sys-get-data-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.336 c-lw6sys-get-default-config-file. . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.337 c-lw6sys-get-default-data-dir . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.338 c-lw6sys-get-default-log-file. . . . . . . . . . . . . . . . . . . . . . . . . . . 207 4.16.339 c-lw6sys-get-default-map-dir . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.340 c-lw6sys-get-default-map-path. . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.341 c-lw6sys-get-default-mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.342 c-lw6sys-get-default-music-dir . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.343 c-lw6sys-get-default-music-path . . . . . . . . . . . . . . . . . . . . . . 208 4.16.344 c-lw6sys-get-default-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.345 c-lw6sys-get-default-script-file . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.346 c-lw6sys-get-default-user-dir. . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.347 c-lw6sys-get-hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.348 c-lw6sys-get-log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 4.16.349 c-lw6sys-get-map-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.350 c-lw6sys-get-map-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.351 c-lw6sys-get-memory-bazooka-eraser. . . . . . . . . . . . . . . . . . 209 4.16.352 c-lw6sys-get-memory-bazooka-size . . . . . . . . . . . . . . . . . . . . 209 4.16.353 c-lw6sys-get-mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.354 c-lw6sys-get-music-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.355 c-lw6sys-get-music-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.356 c-lw6sys-get-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.357 c-lw6sys-get-run-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.358 c-lw6sys-get-script-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 4.16.359 c-lw6sys-get-timestamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.360 c-lw6sys-get-uptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.361 c-lw6sys-get-user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.362 c-lw6sys-get-username . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.363 c-lw6sys-getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.364 c-lw6sys-getenv-prefixed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.365 c-lw6sys-idle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.366 c-lw6sys-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.367 c-lw6sys-log-get-backtrace-mode. . . . . . . . . . . . . . . . . . . . . . 210 4.16.368 c-lw6sys-log-get-level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 4.16.369 c-lw6sys-log-set-backtrace-mode . . . . . . . . . . . . . . . . . . . . . . 211 4.16.370 c-lw6sys-log-set-dialog-timeout . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.371 c-lw6sys-log-set-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.372 c-lw6sys-megabytes-available . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.373 c-lw6sys-openmp-get-num-procs . . . . . . . . . . . . . . . . . . . . . . 211 4.16.374 c-lw6sys-path-concat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.375 c-lw6sys-path-file-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.376 c-lw6sys-path-parent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.377 c-lw6sys-path-split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 4.16.378 c-lw6sys-set-memory-bazooka-eraser . . . . . . . . . . . . . . . . . . 211 4.16.379 c-lw6sys-set-memory-bazooka-size . . . . . . . . . . . . . . . . . . . . 212 4.16.380 c-lw6sys-signal-custom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.381 c-lw6sys-signal-default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.382 c-lw6sys-signal-poll-quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.383 c-lw6sys-signal-send-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.384 c-lw6sys-sleep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.385 c-lw6sys-snooze . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.386 c-lw6sys-url-canonize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.387 c-lw6tsk-loader-get-stage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.388 c-lw6tsk-loader-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 4.16.389 c-lw6tsk-loader-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 4.16.390 c-lw6tsk-loader-push-gen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 4.16.391 c-lw6tsk-loader-push-ldr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

4.17 Script hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

5 C API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

5.1 libliquidwar6. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 5.1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 5.1.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

5.2 libbot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 5.2.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 5.2.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

5.3 mod-brute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.3.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.3.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

5.4 mod-follow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.4.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.4.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

5.5 mod-idiot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.5.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.5.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

5.6 mod-random . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 5.6.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

5.6.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.7 libcfg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

5.7.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.7.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

5.8 libcli. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.8.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.8.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

5.9 mod-http . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.9.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.9.2 API. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

5.10 mod-tcp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.10.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 5.10.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

5.11 mod-udp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.11.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.11.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

5.12 libcns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.12.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.12.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

5.13 libcnx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.13.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.13.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

5.14 libdat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.14.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.14.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

5.15 libdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.15.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 5.15.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.16 libdsp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.16.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.16.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.17 libdyn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.17.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.17.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.18 libgen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.18.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.18.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.19 libgfx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.19.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 5.19.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

5.20 mod-gl1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.20.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.20.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

5.21 mod-gles2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.21.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.21.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

5.22 mod-soft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.22.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

5.22.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.23 shared-sdl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

5.23.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 5.23.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

5.24 mod-caca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.24.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.24.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

5.25 libglb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.25.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.25.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

5.26 libgui . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.26.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.26.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

5.27 libhlp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.27.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 5.27.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.28 libimg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.28.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.28.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.29 libker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.29.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.29.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.30 libldr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.30.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.30.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.31 libmap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.31.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 5.31.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

5.32 libmat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.32.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.32.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

5.33 libmsg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.33.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.33.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

5.34 libnet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.34.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.34.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

5.35 libnod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.35.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.35.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

5.36 libp2p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.36.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 5.36.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

5.37 libpil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.37.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.37.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

5.38 libscm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.38.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 xxvi Liquid War 6

5.38.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.39 libsim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

5.39.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.39.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

5.40 libsnd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.40.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 5.40.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

5.41 mod-csound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.41.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.41.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

5.42 mod-ogg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.42.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.42.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

5.43 libsrv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.43.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.43.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

5.44 mod-httpd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.44.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.44.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

5.45 mod-tcpd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 5.45.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.45.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

5.46 mod-udpd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.46.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.46.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

5.47 libsys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.47.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.47.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

5.48 libtsk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.48.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.48.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226

5.49 libvox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.49.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 5.49.2 API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

Appendix A Authors . . . . . . . . . . . . . . . . . . . . . . . . . . 229 Appendix B 2005 .plan. . . . . . . . . . . . . . . . . . . . . . . . 231

B.1 Complete rewrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 B.2 Technologies. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

B.2.1 Script + standard C + assembly . . . . . . . . . . . . . . . . . . . . . . . . 231 B.2.2 OpenGL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 B.2.3 CSound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

B.3 Functionnalities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 B.3.1 Visual enhancements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 B.3.2 Rules enhancements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 B.3.3 Hey, you forgot my idea!!!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

B.4 Road map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

Appendix C Fanfic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237

C.1 The Battle of Emberlificoted. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237

Appendix D Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

D.1 Official links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 D.2 Other sites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 D.3 Old stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

Appendix E GNU GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243

Appendix F GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255

Appendix G Indexes. . . . . . . . . . . . . . . . . . . . . . . . . . . 263

G.1 Concept index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 G.2 Function and keyword index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 G.3 Data types index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

1 Introduction

Read this chapter to discover Liquid War 6.

1.1 In a nutshell

Liquid War 6 is a unique multiplayer wargame. Your army is a blob of liquid and you have to try and eat your opponents. Rules are very simple yet original, they have been invented by Thomas Colcombet. It is possible to play alone against the computer but the game is really designed to be played with friends, on a single computer, on a LAN, or on Internet.

An older version, Liquid War 5 (http://www.ufoot.org/liquidwar/v5), is available, but is not part of the GNU Project. Only Liquid War 6 is part of the GNU Project (http:// www.gnu.org/), it is a complete rewrite.

   The official page of Liquid War 6 is http://www.gnu.org/software/liquidwar6/. For

more information, you can read the Wikipedia article (http://en.wikipedia.org/wiki/ Liquid_War) about Liquid War.

1.2 Project status

1.2.1 What works, and what does not (yet)

As of today, the game is in beta state. It can be installed, and you can toy arround with. You can even play with. It is still far from being complete as some key features are still missing.

What works:

• The whole framework is here, some functions are not implemented yet, but the bases are set up, and they are believed solid. The game is very modular, and is fully threaded. It is designed so that graphics, sound, network and bot backends can be hacked at will. It has a complete self-test suite, many debugging built-in tools, and is regularly checked with automated tools. For instance, you can check reports concerning global references

    (http://www.ufoot.org/liquidwar/v6/doc/global/), code coverage (http://www.
    ufoot.org/liquidwar/v6/doc/coverage/) and cyclomatic complexity (http://www.
    ufoot.org/liquidwar/v6/doc/cyclo/). This is not a quick hack.

• Documentation. Yes, you’re reading it.

• Version 0.0.7beta is playable. Local game between humans (up to 4 players) is possible. Two bots are implemented, named random and stupid. No great players but well, they move the cursor. A new “deatchmatch” mode, different from LW5, is in place.

• Liquid War 6 already has some features which are nowhere to be found in Liquid War 5, such as multiple layers. It can be worth the upgrade.

• Maps. A number of interesting maps have already been designed (thanks to Kasper Hviid).

• The game runs natively on GNU/Linux and has been ported to Microsoft Windows and Mac OS X. Binaries are available for all those platforms. Use at your own risk. If in doubt, get the source and compile.

In the near future:

• Network play. Top-level priority. Yes, network has been promised for months (years? ...yes, years) and is still not there. I said “when it’s done”.

• Fix bugs ;) The current engine is somewhat buggy, fighters might loose the cursor, it clearly needs polishing.

In the long run:

• Write new graphical backends so that the game does not require Mesa or any OpenGLlike subsystem. The idea is to get rid of the 3D-accelerator dependency.

• Implement all the fancy 3D features, make it possible to play Liquid War 6 on a Moebius ring.

• Use the cool features of CSound to provide dynamic, contextualized sounds & musics.

• Optimize the bot algorithm, which is probably a complex AI problem.

You might be interested in checking the following URLs, which give a view on opened tasks and bugs:

• bug list: http://savannah.gnu.org/bugs/?group=liquidwar6
• task list: http://savannah.gnu.org/task/?group=liquidwar6

1.2.2 What has changed since Liquid War 5.x?

Liquid War 6 is a complete rewrite of Liquid War 5 (http://www.ufoot.org/liquidwar/ v5). The rewrite started in 2005. So a good question is “was the rewrite worth it?”...

Here’s a list of key improvements:

• appearance, global rendering quality. Call it the way you want, Liquid War 6 simply looks nicer than any previous release. Period.

• level features, including multi-layer (allowing the map designer to create bridges and tunnels), wrapping (fighters disappearing on the left can reappear on the right). Those really change the gameplay.

• deathmatch mode. Give it a try, it’s now the default mode, and definitely changes the rules.

• team profiles, as well as special “weapons”, which are tricks you can play on opponents.

• modularity, overall code quality. While this is not a user-visible change, the game is far less monolithic, therefore hacking to revamp the graphics engine, the algorithm, whatever, is easier. The situation has changed from “this is impossible to hack” to “OK, how much time can this take?”. So while one can’t promise every idea will be implemented some day, at least many more things become possible with the new codebase.

The most interesting change is still to come, and concerns network games. Stay tuned.

1.2.3 Revision history

Liquid War 6 releases are “codenamed” after famous, historical, real or mythical characters. Here is a short revision history. For details, see the ChangeLog and NEWS files distributed with the game. Additionnally, there’s an ever-increasing “stamp” number which is incremented each time a build is done with a different source. Latest versions use the stamp as the revision number (the version 3rd number).

• 2006-12-18 : 0.0.1beta (http: / /download .savannah .gnu .org /releases /
  liquidwar6/0.0.1beta/)
• 2007-09-07 : 0.0.2beta (http: / /download .savannah .gnu .org /releases /
  liquidwar6/0.0.2beta/)
• 2008-01-30 : 0.0.3beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.3beta/), codename “Napoleon”, stamp 549

• 2008-09-19 : 0.0.4beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.4beta/), codename “Clovis”, stamp 756

• 2008-12-20 : 0.0.5beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.5beta/), codename “Henri IV”, stamp 1082

• 2009-01-10 : 0.0.6beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.6beta/), codename “Cesar”, stamp 1124

• 2009-10-03 : 0.0.7beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.7beta/), codename “Geronimo”, stamp 1465

• 2010-07-05 : 0.0.8beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.8beta/), codename “Attila”, stamp 1658

• 2010-08-23 : 0.0.9beta (http: / /download .savannah .gnu .org /releases /

liquidwar6/0.0.9beta/), codename “Chuck”, stamp 2096

• 2011-07-29 : 0.0.10beta (http://download.savannah.gnu.org/releases/

liquidwar6/0.0.10beta/), codename “Gengis Kahn”, stamp 2562

• 2011-10-02 : 0.0.11beta (http://download.savannah.gnu.org/releases/

liquidwar6/0.0.11beta/), codename “Ho Chi Minh”, stamp 2785

• 2011-12-18 : 0.0.12beta (http://download.savannah.gnu.org/releases/

liquidwar6/0.0.12beta/), codename “Aguirre”, stamp 2938

• 2011-12-24 : 0.0.13beta (http://download.savannah.gnu.org/releases/

liquidwar6/0.0.13beta/), codename “Blackbeard”, stamp 2950

• 2014-01-04 : 0.2.3551 (http://download.savannah.gnu.org/releases/liquidwar6/

0.2.3551/), codename “Davy Crockett”

• 2014-03-26 : 0.4.3681 (http://download.savannah.gnu.org/releases/liquidwar6/

0.4.3681/), codename “Hannibal”

• 2015-05-06 : 0.6.3902 (http://download.savannah.gnu.org/releases/liquidwar6/

0.6.3902/), codename “Goliath”

1.2.4 Road map

The game will probably be labelled “6.0.0” when network mode is up and running. Until then there will probably be other improvements concerning gameplay and appearance (“eye candy”). There’s a balance to keep between the major goals such as “make that network thingy work” and the very real fact that “hacking must be fun”.

1.3 How you can help

1.3.1 Help GNU

Please remember that development of Liquid War 6 is a volunteer effort, and you can also contribute to its development. For information about contributing to the GNU Project,

please read How to help GNU (http://www.gnu.org/help/help.html).

1.3.2 Todo list

Here’s a short list of todo items. It is probably too early to start hacking the core engine itself, for it is still under heavy development, might undergo major rewrites, and it’s hard for documentation to keep up with the reality of the code. However, there are still many things to do.

• Try the game. Play. Test. Send bug reports. Without bug reports, bugs don’t get fixed.

• Write maps. Obviously, this is something which can perfectly be delegated. Experience shows user-contributed maps are, on average, better than maps conceived by the author...

• Translate texts. Liquid War 6 uses GNU gettext, so all messages can be translated.

• ...any help is welcome.

Feel free to join the mailing-lists, this is clearly the best place to start with.

There’s also a list of opened tasks on Savannah at http://savannah.gnu.org/task/? group=liquidwar6 which you can browse online. Maybe there’s some task for you!

Alternatively, you can contact Christian Mauduit (mailto:ufoot@ufoot.org).

2 User’s manual

The Liquid War 6 user’s manual hopefully contains any usefull information to install the program and play the game. If you just want to enjoy Liquid War 6 without diving into map creation and programming, this is just for you.

2.1 Mailing lists

2.1.1 General discussion

The main discussion list is <help-liquidwar6@gnu.org> (mailto:help-liquidwar6@gnu. org), and is used to discuss all aspects of Liquid War 6, including installation, development, game strategies, and whatever subject players and hackers might want to talk about, provided it is Liquid War 6 related. If you don’t know on which list to subscribe, this is the one.

To subscribe to it, please send an empty mail with a Subject: header line of just "subscribe" to the -request list, that is <help-liquidwar6-request@gnu.org>

(mailto:help-liquidwar6-request@gnu.org).

You can also subscribe to the list using the Mailman web interface for help-liquidwar6

(http: / / lists . gnu . org / mailman / listinfo / help-liquidwar6) and consult
help-liquidwar6 archives (http://lists.gnu.org/archive/html/help-liquidwar6/).

2.1.2 Announcements

Announcements about LiquidWar 6 are made on <info-liquidwar6@gnu.org> (mailto:info-liquidwar6@gnu.org). Subscribe to it to be informed of major releases, and other significant news.

To subscribe to it, please send an empty mail with a Subject: header line of just "subscribe" to the -request list, that is <info-liquidwar6-request@gnu.org>

(mailto:info-liquidwar6-request@gnu.org).

You can also subscribe to the list using the Mailman web interface for info-liquidwar6

(http: / / lists . gnu . org / mailman / listinfo / info-liquidwar6) and consult
info-liquidwar6 archives (http://lists.gnu.org/archive/html/info-liquidwar6/).

Please also consider reading the latest news on Savannah (http://savannah.gnu.org/

news/?group=liquidwar6).

2.1.3 Bugs

There is also a special list used for reporting bugs, <bug-liquidwar6@gnu.org> (mailto:bug-liquidwar6@gnu.org). Please try and describe the bug as precisely as possible. The more accurate the description, the more chances it will get to be fixed.

While this is the standard GNU way of reporting bugs, modern SPAM standards make it very hard to filter real bug reports from junk on this list. It is more convenient to

use a web interface, the URL is: http://savannah.gnu.org/bugs/?func=additem&

group=liquidwar6 and you’re really encouraged to use it instead of sending emails.

   Please take a look at the bug list (http: / /savannah .gnu .org /bugs / ?

group=liquidwar6) before submitting new bugs.

2.1.4 IRC channel

IRC can be an interesting alternative to mailing-lists. There’s an open channel dedicated to Liquid War on freenode.net, you can access it on irc://irc.freenode.net/liquidwar that is, channel #liquidwar on irc.freenode.net.

2.2 Getting the game

2.2.1 Download source

Liquid War 6 can be found on:

  • http://ftp.gnu.org/gnu/liquidwar6/
  • http://download.savannah.gnu.org/releases/liquidwar6/
  • http://www.ufoot.org/download/liquidwar/v6/

Downloading the latest file from this place, and compile it yourself on your computer with a classical ./configure && make && make install is the recommended way to install Liquid War 6.

2.2.2 Download binaries

Some binary packages might be available. Your mileage may vary.

GNU/Linux based systems are supported, through Debian (http://www.debian.org/ ) .deb and Red Hat (http://www.redhat.com/) RPM packages. There is also a Microsoft Windows installer.

However these binaries are not necessarly available for every single version of the game.

2.2.3 GIT repository

Latest work in progress versions can be obtained with GIT (http://git-scm.com/). Here’s the typicall command which will fetch the latest version:

     git clone git://git.sv.gnu.org/liquidwar6.git

If you are behing a firewall and can’t use the native GIT protocol, you can rely on the (slower) http protocol:

     git clone http://git.sv.gnu.org/r/liquidwar6.git

You can browse the code online, consult log summary, and in a general manner “follow”

the project on http://git.savannah.gnu.org/gitweb/?p=liquidwar6.git and http://
git.savannah.gnu.org/cgit/liquidwar6.git.

Beware, git does not contain all the files included in the official source tarball. For instance, the ./configure script is not present. You need to run :

     autoreconf
     ./configure
     make
     make install

The autoreconf call is really mandatory the first time, autoconf is not enough. You will also need all the prerequisites needed to build the docs, generally speaking, getting the source from git requires more tools to build the package than picking a ready-to-use tarball.

2.2.4 Daily snapshots

Alternatively, you can download daily snapshots on http://www.ufoot.org/liquidwar/ v6/snapshots/ These files used to be built every day, now they are generated by Jenkins (http://jenkins-ci.org/) whenever there’s a source change (commit). A simple make is done before generating source tarballs however a make distcheck is performed before generating binaries, therefore sometimes you can have the source but no associated tarballs.

Beware of revision numbers, snapshots can make you believe version X.Y is out when it’s only a release candidate at best, and most of the time just a work-in-progress.

Still, if you want bleeding edge versions, this is the way to go. Documentation is automatically updated as well, and available on http://www.ufoot.

org/liquidwar/v6/doc/.

2.2.5 Check integrity

Most binary packages (at least .deb and RPM GNU/Linux binaries) should be signed using GnuPG (http://www.gnupg.org/). The following keys are used when generating upstream/vendor packages:

• 1024D/FD409E94 2002-01-31 Christian Mauduit (U-Foot) <ufoot@ufoot.org>

-----BEGIN PGP PUBLIC KEY BLOCK-----
Version: GnuPG v1.4.12 (GNU/Linux)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=fcJV
-----END PGP PUBLIC KEY BLOCK-----

• 2048R/406FFCAB 2013-07-12 Jenkins Daemon (Christian Mauduit) <jenkins@ufoot.org>

-----BEGIN PGP PUBLIC KEY BLOCK-----
Version: GnuPG v1.4.12 (GNU/Linux)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=fcJV
-----END PGP PUBLIC KEY BLOCK-----

2.3 Installation

This section covers installation from source. Other ways of installing the program are not described here.

2.3.1 Requirements

All these libraries are mandatory to compile the game. Liquid War 6 won’t compile, let alone run, without them. Some of them could probably be replaced by equivalent tools, but this would certainly require a programming effort and some changes in Liquid War 6 source code.

• GCC (http://www.gnu.org/software/make/). Liquid War 6 does require the GNU C Compiler to build, while other compilers might be able to build the game, this is untested.

• Gomp (http://gcc.gnu.org/projects/gomp/). Liquid War 6 uses OpenMP #pragma

directives, this should help the game run faster on SMP systems.

• GNU Make (http://www.gnu.org/software/make/). Liquid War 6 might and

certainly does use GNU Make extensions.

• GNU C library (http://www.gnu.org/software/libc/). Sounds obvious, but you need a standard C library. It happens that glibc has some rather usefull extensions (yes, as of 2006, some vendors continue to offer C libraries without snprintf...) and Liquid War 6 might use them. In a general manner, Liquid War 6 is part of and designed for GNU. You might however manage to compile it with limited libc support, this is the case with mingw32 for instance but, do it at your own risk.

• Perl (http://www.perl.com/). Some Makefile commands require Perl. You don’t need any Perl devel packages, and you can probably use any Perl 5.x version, since no fancy recent feature of Perl is used. Just plain Perl.

• Guile (http://www.gnu.org/software/guile/). Possibly the most required library, since Liquid War 6 is a scheme program which uses a set of functions coded in standard C. You need at least Guile 1.8.

• GNU MP (http://gmplib.org/). GMP is a free library for arbitrary precision arithmetic, required by Guile.

• libgc (https://launchpad.net/libgc). This is a a garbage collector library, recent versions of Guile might require this so in case your version of Guile requires it, then Liquid War 6 will need it too.

• ltdl (http://www.gnu.org/software/libtool/). This library, which comes with

libtool, provides a portable alternative to dlopen and dlclose. Check that you have a /usr/include/ltdl.h file, or install the corresponding package.

• zlib (http://www.zlib.net/). Required by other libraries, but can also be used directly by Liquid War 6 to compress network messages for instance.

• expat (http://www.libexpat.org/). Used to read and write XML files, which contain constants and configuration data.

• libpng (http://www.libpng.org/pub/png/libpng.html). Liquid War 6 uses libpng

to read levels (maps), not to speak of other optional libraries (SDL and the rest) who need it themselves.

• libjpeg (http://www.ijg.org/). Maps can also be provided as jpeg files, so libjpeg is required as well.

• SQLite 3 (http://www.sqlite.org). Used to handle the list of available servers.

2.3.2 Optional libraries

While all these libraries are theorically optional (the game will successfully compile without them), you’ll obviously need, for instance, one graphics backend. Otherwise, you’ll simply have no display. This is not acceptable. As of today, one can reasonnably consider all SDL-related libraries are required. The rest is truely optional.

• libcunit (http://cunit.sourceforge.net/). Provides (hopefully) more readable test output. It’s not strictly mandatory but still highly recommended. Building without is just allowed in case some rare and bizarre platform would not have a libcunit port.

• ncurses (http://www.gnu.org/software/ncurses/). Required by readline, needs to be there otherwise readline might not be detected properly on some systems.

• GNU readline (http://cnswww.cns.cwru.edu/php/chet/readline/rltop.html).

Used to handle input on the console. Console is not absolutely mandatory, but it’s a must-have if you want to hack the game. Console unavailable does not mean you won’t get anything on stdout but, the interactive script shell just won’t work.

• GTK+ (http://www.gtk.org/). Used to display error/critical messages, so that users who launch the game by clicking on a icon (that is, not from the console) are still visually informed of important messages.

• Mesa (http://www.mesa3d.org/). This library provides an API similar to OpenGL and enables 2-D and 3-D drawing.

• SDL (http://www.libsdl.org/). SDL is used to set up a working OpenGL environnement, and handle input (mouse and keyboard).

• SDL image (http://www.libsdl.org/projects/SDL_image/). This SDL extension

is used to read textures and other graphics from disk.

• FreeType 2 (http://freetype.sourceforge.net/). This library is required by SDL ttf, to draw fonts.

• SDL ttf (http://www.libsdl.org/projects/SDL_ttf/). This SDL extension is used

to draw fonts. It is UTF-8 enabled.

• libcaca (http://caca.zoy.org/). This library transforms bitmaps into ascii-art images, allowing an alternative style of display, TTY compatible.

• libcsound (http://www.csounds.com/). While this tool is not used yet, it is meant to be the final sound backend, as CSounds offers great power to the composer, enabling truely dynamically generated sound & music. For now Liquid War 6 tries to detect csound 4 but as the mainstream stable release is now 5 an update is needed. It will probably be updated/fixed (Liquid War 6 using csounds 5) some day, for now you can safely *not* install csound on your system and enjoy all the possibilities of the game.

• SDL mixer (http://www.libsdl.org/projects/SDL_mixer/). This SDL extension

is used to allow dynamic mixing of sounds, and it also provides a builtin OGG/Vorbis file renderer.

• libcURL (http://curl.haxx.se/libcurl/). Used to handle HTTP requests, the idea being not to re-invent the wheel but use a robust standards-compliant generic library.

2.3.3 Optional tools

Those tools are used upstream to maintain the game, you normally do not need them to build the game “as is” but if you modify the source code and hack the game, you might be interested in installing them.

• Perl 5 (http://www.perl.org/). Liquid War 6 uses Perl for many tedious task, including, but not limited to, parsing documentation.

• GNU Indent (http://www.gnu.org/software/indent/). Code is regularly indented using the script src/indent.sh which calls indent automatically and recursively on the whole source tree.

• md5sum (GNU core utilities) (http://www.gnu.org/software/coreutils/) This is

used to stamp the source code and help tracking exact build versions.

• Doxygen (http://www.stack.nl/~dimitri/doxygen/). Used to generate documen-

tation concerning C structs, more precisely, include the struct members documentation into the official texinfo manual.

• xsltproc (http://xmlsoft.org/XSLT/xsltproc2.htm). Used to post-process Doxygen

output and transform it to texinfo.

• dot (http://www.graphviz.org/). Used to generate Doxygen call graphs.

• Google Performance Tools (http://code.google.com/p/google-perftools/). This

tool is convenient to optimize the program and find out what parts of it take most of the CPU power to execute.

• lcov (http://ltp.sourceforge.net/coverage/lcov.php). Gives nice output about

code coverage.

• GNU global (http://www.ufoot.org/liquidwar/v6/doc/global/). Shows global

references through the code.

• pmccabe (http://parisc-linux.org/~bame/pmccabe/). Cyclomatic complexity,

shows what part of the code are bloated.

• Valgrind (http://www.valgrind.org/). Usefull to track down memory leaks and many other programming errors.

2.3.4 Installing requirements using RPM/DEB packages

You might find it convenient not to install all the requirements from source, but use your favorite GNU/Linux distribution packages.

On an RPM based GNU/Linux system, a typical command (tested with Fedora (http://

www.fedoraproject.org/) 15 “Lovelock”) could be:
yum install \
make gcc glibc glibc-devel binutils \
libgomp \
guile guile-devel gmp gmp-devel libgc1c2 libgc-dev \
libtool libtool-ltdl libtool-ltdl-devel \
zlib zlib-devel expat expat-devel \
libpng libpng-devel libjpeg libjpeg-devel \
sqlite sqlite-devel \
ncurses ncurses-devel readline readline-devel \
libGL libGL-devel libGLU libGLU-devel \
SDL SDL-devel SDL_image SDL_image-devel \
SDL_mixer SDL_mixer-devel \
freetype freetype-devel SDL_ttf SDL_ttf-devel \
libcaca libcaca-devel \
libcurl libcurl-devel \
gtk2-devel \
perl lcov global valgrind graphviz gv ImageMagick \
texinfo-tex \
indent emacs doxygen libxml \
CUnit CUnit-devel \
rpm-build

On a DEB package based GNU/Linux system this command (tested with Debian (http://www.debian.org/) 6.0 “squeeze”) would be:

apt-get install \
make autoconf automake \
gcc libc6 libc6-dev binutils \
libgomp1 \
guile-2.0 guile-2.0-dev guile-2.0-libs libgmp10 libgmp3-dev \

libtool libltdl7 libltdl-dev \
zlib1g zlib1g-dev libexpat1 libexpat1-dev \
libpng12-0 libpng12-dev libjpeg8 libjpeg-dev \
libsqlite3-0 libsqlite3-dev \
libncurses5 libncurses5-dev libreadline6 libreadline6-dev \
libgl1-mesa-glx libgl1-mesa-dri libgl1-mesa-dev libglu1-mesa libglu1-mesa-dev \
libgles2-mesa libgles2-mesa-dev \
libsdl1.2debian libsdl1.2-dev libsdl-image1.2 libsdl-image1.2-dev \
libsdl-mixer1.2 libsdl-mixer1.2-dev \
libfreetype6 libfreetype6-dev libsdl-ttf2.0-0 libsdl-ttf2.0-dev \
libcaca0 caca-utils libcaca-dev \
libcurl4-gnutls-dev \
libgtk2.0-dev \
perl lcov global valgrind graphviz gv imagemagick \
texinfo texlive-base texlive-generic-extra \
texlive-fonts-recommended texlive-latex-extra \
indent emacs doxygen xsltproc pmccabe \
libcunit1-ncurses libcunit1-ncurses-dev \
google-perftools libgoogle-perftools-dev \
git git2cl \
zip nsis \
debhelper devscripts

Note that those requirements really depend on the exact distribution you have, package names may vary from one to another.

2.3.5 Compiling

Liquid War 6 uses GNU Automake (http://www.gnu.org/software/automake/), Auto-
conf (http://www.gnu.org/software/autoconf/) and GNU Libtool (http://www.gnu.
org/software/libtool/).

Once all the requirements are installed, run:

./configure
make
make install

Liquid War 6 supports the standard ./configure --prefix=/my/path option (in fact, it supports much more than that) so you can install the game in any directory. You do not need to be root to install Liquid War 6.

2.4 Extra maps

2.4.1 The extra maps package

The main package contains some maps so that you can try out the game. Still, an additionnal package, called extra-maps or liquidwar6-extra-maps is available, containing more maps. It really does contain many of them, including most Liquid War 3 and Liquid War 5 legacy maps, plus new Liquid War 6 maps.

2.4.2 Install extra maps on GNU/Linux and POSIX systems

On GNU/Linux systems (and possibly any POSIX unixish system) running:

     ./configure
     make
     make install

will install the extra maps on your system automatically, they will then be available in the extra/ sub-directory when browsing maps.

The ./configure script has a --enable-liquidwar6 switch which will try and find automatically if there’s an existing liquidwar6 binary in the path. If there’s such a binary, it will run it and ask for its map-path and use this value automatically.

2.4.3 Raw install of extra maps (all-platforms)

Another solution, which works on all platforms including Microsoft Windows and Mac OS X but also works on GNU/Linux, is to simply unpack the extra-maps package (unzip or untar) in your custom map directory, or in the system map directory. There’s nothing else to do to install these maps but simply put them on your hard drive in the right directory.

Typically on an Microsoft Windows system, you would unpack the extra maps in C:\Program Files\Liquid War 6\map\ (system directory) and on a Mac OS X system you

would unpack the extra maps in Liquid War 6.app/Contents/Resources/map/ (system
directory) or $HOME/Library/Application Support/Liquid War 6/map (user directory).

On a GNU/Linux or POSIX system you would unpack them in $HOME/.liquidwar6/map/ (user directory).

Next time you run the game, the maps should be browsable.

If you can’t see them, run liquidwar6 --audit and check that the place where you unpacked the files is actually searched by the binary.

2.5 Troubleshooting

2.5.1 Compilation problems

A quick survival guide:

• Check that you have all dependencies installed. Also check their version number. Double-check that you have devel packages installed, not only run-time binaries.

  • Read carefully the output of ./configure. Running ./configure > configure.log
    2> configure.err does help.

• Editing /etc/ld.so.conf and running ldconfig as root can help if some dependencies are installed in exotic places.

• Check the values of the environment variables CFLAGS, LDFLAGS and LD_LIBRARY_PATH.

• Try ./configure --enable-allinone, this will disable some fancy but somewhat complicated dynamic .so file support, it can help if shared libraries are handled differently on your system than on a plain GNU/Linux box.

If none of these help, consider reporting a bug, or search the mailing-lists for help.

2.5.2 Check installation

Here’s a check-list to ensure that your installation is correct:

• What was the output of make install? make check?

• Is the liquidwar6 binary in your PATH environment variable? It might be in

/usr/games.

• Run liquidwar6 --pedigree. Look at the output. Check the compilation date & time, the version number.

• Run liquidwar6 --audit. What do these paths look like? Are they absolute paths? Do they exist? What’s there? Normally, once the game is installed, all of them should exist, and be populated with sub-directories and files.

• Run liquidwar6 --modules, to know which modules where compiled. You need at least one graphical module, for instance mod-gl1, else the game won’t run.

• Run liquidwar6 --host, this displays informations about the host system the binary has been built for.

2.5.3 Problems running the game

Now, game looks correctly installed, but you have problems running it.

• Run the game from a terminal, not from a Gnome or KDE launcher, you need to see the console output.

• In the $HOME/.liquidwar6/ directory, you’ll find some files, the main log file log.csv and maybe dump.txt or backtrace.txt. They might contain valuable information, read them. Note that while log.csv is overwritten each time you start the game, dump.txt or backtrace.txt are conserved until a new problem arises. So check the date of these files to be sure you’re analyzing the right ones. Note that byt default on Microsoft Windows $HOME/.liquidwar6/ is replaced by

C:\Documents and Settings\<username>\Liquid War 6 and on Mac OS X it is in
/Users/<username>/Library/Application Support/Liquid War 6/.

• Run liquidwar6 --defaults. This will reset all options to defaults. You might need to run this when upgrading from a version to another, since some options might appear, disappear, or defaults values can change.

• Run liquidwar6 --test. This should run a complete test suite, many functions in the game will be tested automatically, and errors reported.

• Run liquidwar6 --show-script-file. Are you really running the right code?

• Game segfaults: try make uninstall && make clean && make && make install. Many

problems can come from using a wrong shared module. You can also launch the game with the --trap-errors=false switch, this will disable the custom popup window and allow you to get the real error.

• Game (still) segfaults: try gdb liquidwar6. Type run --trap-errors=false and

watch output.

• The dynamic library loader can sometimes have problemes, and does not always report explicit messages on stdout or stderr. You can change this by modifying some environment variables: export LD_DEBUG=all. This is very verbose but does help finding bugs.

• Consider compiling the game using ./configure --enable-valgrind and then run it

  using Valgrind (http://valgrind.org/).
• Try find / -type d -a -name "liquidwar6*" 2> /dev/null to ensure you don’t have

an old version of Liquid War 6 somewhere else...

2.6 Quick start

2.6.1 Quick start

Once the game is installed, run it, click on Quick start with the mouse, and control the red 'a' cursor with the mouse, or keyboard, both work. Try and surround the green team, it’s a stupid bot, you should win ;)

You army is formed by all the red pixels on the screen, they should try and rejoin the cursor (the blinking 'a' letter) using the shortest path. When red and green meet, they fight. Try it, toy arround.

The Quick start button will always make you play red against a green stupid bot, whatever other options you have set up.

Todo...

2.7 Strategy tips

2.8 User interface

2.8.1 A reduced set of keys

Liquid War 6 can be controlled using a reduced set of keys. This is to make the game more portable and allow possible ports to platforms where a full keyboard is not available. Depending on the graphics backend, exact mapping might change, they should hopefully be obvious and intuitive.

Those keys are:

• up : the arrow up key

• down : the down arrow key

• left : the left arrow key

• right : the right arrow key

• enter : the enter / return key

• esc : the escape key

• ctrl : the control key

• alt : the alt / meta key

• pgup : the page up key

• pgdown : the page down key

Basically,

2.8.2 Combining mouse, keyboard and joysticks

It’s also possible to control the game with the mouse only, or with a joystick. By default the interface will trap all events and respond on any of these possible devices.

Keyboard  Mouse  Joystick  Menu action  In-game

up

mouse pointer stick previous menu item

move cursor up

down

mouse pointer stick next menu item move cursor down

left

mouse pointer stick change menu item value

move cursor left

right

mouse pointer stick change menu item value move cursor right

enter

left-click button A validate menu validate chat line

esc

right-click button B back to previous menu

quit game

ctrl

right-click or double-click on any button button C N/A

fire

alt

middle-click or triple-click on any button button D N/A

alternate fire

pgup

wheel up button E previous menu item

zoom in

pgdown  wheel down  button F  next menu item zoom out

A final word about joystick buttons: there’s no such thing as standard joystick buttons, some will come with A,B,C,D, others will have A,B,start,select,L,R, there’s no way to know. By default, the game will use the buttons with the lowest indexes (returned by your driver) for the most usefull functions. Validate menu entries is the most usefull action, zooming in and out the one you can live without.

2.8.3 Quit with F10

There’s also an (almost) hardcoded shortcut which will quit the game immediately, or at least as quickly as possible, without any prompt or warning.

It is the F10 key.

Think of this feature as the procastinator’s “whoops, here comes my boss!!!” emergency function.

2.9 Solo game

2.9.1 Current state

As of today, Liquid War 6 is essentially a solo game since network is not working. It allows you to toy arround in arcade mode on any map you wish.

A real solo mode with campaign and goals to reach is planned, how it will be implemented is yet to be defined.

2.9.2 Team profiles

By default, teams behave differently, some of them move more rapidly, some are more aggressive but vulnerable, some are more defensive but do not attack as strong as others. This aspect of the game is under active tuning, things might be unfair by now, you can toy arround with the various team-profile-... options, any report is appreciated.

Note that this is very different from Liquid War 5, and can give very different gaming experiences, you can artificially set up arbitrary strong bots, for instance.

Here’s a description of the default color settings:

• blue: has a strong attack but is slow

• cyan: has an extremely good defense but is slow

• green: has a better defense than the average

• lightblue: has an extremely strong attack but is very slow

• magenta: is extremely fast but also very vulnerable

• orange: is fast, but has a very weak attack

• pink: has a very strong attack, but is also very vulnerable

• purple: has a very good defense but a weak attack

• red: moves faster than the average

• yellow: has a strong attack

2.9.3 Weapons

Additionnally, when profiles are used, each team has two weapons, a primary weapon and an alternate one. Think of weapons as special (usually nasty) tricks you can play on your opponents.

Here’s a description of available weapons:

• atomic: nuclear explosion, all fighters arround your cursor are about to die

• attract: all fighters from all teams are packed near your cursor

• berzerk: super-strong attack for a limited time, crush your enemies

• control: you take the control of all other teams while your cursor stays in place

• crazy: all your opponents go crazy for some time, acting with no logic

• disappear: you disappear for some time from the battlefield, to reappear later, somewhere else

• escape: fighters placed as far as possible from cursor, magically escape from any grip • fix: all other teams are freezed, you can move but not attack them

• invincible: no damage for a limited time, move untouched • kamikaze: you die along with the strongest team on the battlefield, requires at least 3

teams

• mix: fighters exchange position, their properties being preserved

• permutation: will exchange colors, randomly, requires at least 3 teams (double edged weapon)

• plague: general disease, all fighters mysteriously loose health

• reverse: fighters continue to move normally, but attacks are done in reverse mode, backwards

• rewind: make the battlefield be like it was a few seconds ago

• scatter: every fighters of every team scattered in random places

• shrink: reduces the number of fighters on the map

• steal: steals some fighters to other teams

• teleport: fighters placed as close as possible to cursor

• turbo: move faster for a limited time

Note that this is in progress, some of them are NOT IMPLEMENTED YET.

2.10 Network games

2.10.1 Choose your “public url”

Liquid War 6 needs to name your “node” (you can think as your server instance of the game) and have a unique URL (address) to publish and give to other nodes.

If only one network adapter is attached to your computer and your address IP is A.B.C.D then by default the game will pick automatically the address http://A.B.C.D:8056/ and it should work flawlessly.

Problems can arise if you have a peculiar network configuration, if you have multiple non-loopback network interfaces, if you use NAT to protect yourself from intruders and/or if your context forces you to do so. In that case, Liquid War won’t be able to guess a correct URL automatically. So you need to set it up manually either by editing the public-url entry in the config file, changing environment variable LW6_PUBLIC_URL or passing the --public-url=http://<host>:<port>/ argument when running the game. Typically, if you are behind a firewall which does NAT, use the firewall address. The right address is the address which, given to remote hosts, will allow them to connect on your game instance.

2.10.2 Starting a node

A node is started automatically when you run the game. Even if you don’t start to play, node starts in the background and exchanges data with other nodes, mostly to discover them and maintain its internal map of existing nodes and games.

So even without starting a network game, you should be able to point a web browser on your node and see a web page describing it. Your node address is displayed on stdout (console) when starting the game. If in doubt, try http://localhost:8056/ which should work unless you modified defaults settings.

When you start a network game, the program simply changes your node state from “idle” to “accepting connections”.

2.10.3 Connecting to a node

The interface should show you the list of available nodes, just pick one and try and connect to it.

Note that once you’re connected on a remote node, you’re still acting as an independant node, and other nodes might connect to your node as well as to the other nodes. In short, there’s no real server or client, everyone is a client for someone, and can act as a server.

Nodes connected together form a “community”, which can disband, accept new nodes, and in a general manner has its own immaterial life, the first node which created the game might disappear, game can continue without it.

This is why the main network module is called libp2p, this is a reference to the term “peer to peer”.

2.10.4 Communities

Once a node is connected to another one, they’ve started a “community”. Formally, a stand-alone node accepting for connection is already a community, even if it has only one member, but the only really interesting communities are those formed with several nodes.

A community can’t be reached through a given server, to connect to one you just need to connect on one of its member nodes. All nodes are equivalent, there’s no master, no root node, nodes collaborate to share the same real-time information and maintaine an up-to-date game state.

Of course, conflicts can arise, and in that case nodes need to agree on an acceptable solution. Normally, the program takes decisions automatically (for instance, it could decide to “kick” a node out of the community) so the player does not have to care about this, but this is expected to be one of the most tricky (and passionating) part of Liquid War 6 hacking.

2.10.5 Firewall settings

By default, Liquid War 6 will communicate on port 8056, in both TCP and UDP, and in both ways too (in/out). It’s possible to play with partial connectivity, in extreme case, you can even play without direct internet access, using only a mere web proxy.

However, things will go faster and be much easier if the program can use its default native protocol.

Here’s an example of a typicall iptables (http://www.netfilter.org/) configuration which allows you to play the game full-featured. It’s assumed that by default all packets are dropped, this configuration will just open the necessary ports.

     # outgoing TCP on port 8056 (liquidwar6)
     iptables -A OUTPUT -p tcp --dport 8056 -m state --state NEW,ESTABLISHED -j ACCEPT
     iptables -A INPUT -p tcp --sport 8056 -m state --state ESTABLISHED -j ACCEPT
     # incoming TCP on port 8056 (liquidwar6)
     iptables -A INPUT -p tcp --dport 8056 -m state --state NEW,ESTABLISHED -j ACCEPT
     iptables -A OUTPUT -p tcp --sport 8056 -m state --state ESTABLISHED -j ACCEPT
     # outgoing UDP on port 8056 (liquidwar6)
     iptables -A OUTPUT -p udp --dport 8056 --sport 1024:65535 -j ACCEPT
     iptables -A INPUT -p udp --sport 8056 --dport 1024:65535 -j ACCEPT

 # incoming UDP on port 8056 (liquidwar6)
 iptables -A INPUT -p udp --dport 8056 --sport 1024:65535 -j ACCEPT
 iptables -A OUTPUT -p udp --sport 8056 --dport 1024:65535 -j ACCEPT

If you can’t change firewall settings and only have access to the web through a web proxy, it can still be possible to play (with some restrictions such as your node not being seen by others) if mod-http is available. This in turn depends on wether libcurl (http:// curl.haxx.se/) support was activated when compiling the game. To use the proxy, you can set the http_proxy environment variable. For detailed informations, please refer to

libcurl doccumentation (http://curl.haxx.se/docs/).

2.10.6 Is the game secure?

As stated in the license, the program comes with NO WARRANTY. Period.

However, an important effort has been made so that it can reasonnably be used online, exposed to various “common” attacks.

As far as security is concerned, there are two different issues:

• vulnerability to general security attacks, people typically trying to gain prililedged access on your computer, relying on a security flaw in the program. A good firewall is a must-have, as you can never know for sure a program has no bugs. Running Liquid War 6 as an unpriviledged user (certainly not “root”) is also a good practice.

• vulnerability to players “cheating” and sending malicious informations to fake their moves, scores, and/or modify informations concerning other players. This is a very important point in Liquid War 6 since it has a multi-channel way of exchanging data (think of the web interface, you have no garantee of who the client is).

Here’s a list of various steps which have been taken to make the program more secure:

• a --skip-network option is here if you really do not want to be bothered by networking risks;

• program has basic password support so that you can deny access to unknown players;

• passwords are never sent in clear text over the network, only a hash (checksum) is sent;

• no use of well-known buffer overflow friendly functions like strcpy, equivalents such as strncpy are used;

• program never trusts what comes from network peers, when it wants to know something, it checks it out by itself, for instance, the node list is systematically verified by the local node before being used and/or published;

• the built-in web server is not a general purpose web server which will end up revealing some of your private files, it can only serve game-related pages;

• the very fact that the game has no central server makes it hard to attack it, because if someone wants to play “Oscar” annoying “Alice” and “Bob” he well need to fool all the nodes participating in a game, sending wrong informations to a single node won’t have much effect.

This being said, Liquid War 6 does not use any strong encryption library to protect the data it sends. All the checksum machinery might be vulnerable to a brute-force and/or strong cryptographic attack, so in theory it’s possible to fool the program.

In practise, if you want real privacy, play over a VPN (Virtual Private Network).

2.11 Graphics

2.11.1 Standard, high and low resolution

Liquid War 6 will try and pick up a default resolution when the game is launched the first time. It won’t use your maximum screen resolution but will instead list all available fullscreen modes, and pick up one which is usually something like two thirds of the highest mode. This is to allow switching back and forth between fullscreen and windowed mode using the same settings. This automatically picked-up resolution really depends on your hardware and driver. It is called “standard” in the graphics options menu.

Then it is possible to automatically select the minimum and maximum resolution your hardware allows in fullscreen mode. These are called “low” and “high” in the graphics options menu. Just click on the button that display the resolution, it will change and use the next setting. In windowed mode, the game won’t accept the highest available mode but will instead use a percentage of it, defined by the --windowed-mode-limit parameter.

You might still be in a case where this is not enough. For instance your maximum resolution is 1600x1200, Liquid War 6 picks a default mode of 1280x960 for you but for some reason you want to play in 800x600, fullscreen. In this case, simply switch to windowed mode, resize the window with the mouse (the resolution button will show you the current resolution) and just choose a resolution near 800x600. It does not even need to be exactly 800x600, 798x603 would probably fit. Then when switching back to fullscreen, you’ll be in 800x600, the game will automatically pick up the fullscreen mode which is closest to the current windowed mode resolution.

2.11.2 Display rate

By default the game will try and run at 60 frames per second. Given the nature of Liquid War 6, this is probably enough. Higher values will maybe give a slightly smoother display, but barely noticeable.

You can activate the display of frames per seconds (aka “fps”) through the menu (“options -> system”) or with the command line (“–display-fps”).

On a single processor system, reducing the number of frames per second might allow the rest of the game run faster. So if you notice the game is really slow, in terms of “fighters move slowly” then you might be happy reducing the display rate and therefore giving power back to the other parts of the program. On a dual-core (or more) or on a multi-processor system, this is probably useless since the game is threaded and has a dedicated thread for display purposes. The command line option to reduce the number of frames per second is

--target-fps.

Additionnally, the parameter --gfx-cpu-usage allows you to force the display thread to “take a rest” and go idle for some time. This is advanced settings, most users won’t touch this.

2.12 Sound & music

2.12.1 Current status

As of today, the game is capable of playing Ogg Vorbis (http://www.vorbis.com/) audio files. That’s it.

2.12.2 The future

In the long run, what is planned is to support Csound (http://www.csounds.com/) which would allow very cool effects, such as dynamically changing the music while the game is running, typically following the action. If there’s a lot of fight, the music could reflect this.

For now this is only vaporware, just a nice idea among others, nothing implmented yet.

2.13 Config file

The config file is a simple XML file. It uses XML only to benefit standard parsing tools, but it’s not a structured XML file, in the sense that the tree is so simple that all items are at the same level. It is just a simple key-value binding.

This file is in $HOME/.liquidwar6/config.xml on GNU/Linux and POSIX systems,

in C:\Documents and Settings\<username>\Liquid War 6\config.xml on Microsoft
Windows and in /Users/<username>/Library/Application Support/Liquid War
6/config.xml on Mac OS X.

You’re free to edit it manually, but all parameters are changeable with command line options. The program will overwrite this file each time it exits, so if you put comments in it, they will disappear. The advantage of this is that if you mispell something, or if for some reason the game does not understand a value, then when rewriting the file, it will show you it just did not get it.

The file embeds the documentation for all its entries, it is therefore rather verbose. The documentation is the same you will find online or by quering the game with the --about option, also the same you would get reading this manual.

2.14 Logs

Liquid War 6 uses stdout to output important messages, and stderr to log warnings and errors. It will also use syslog (http://en.wikipedia.org/wiki/Syslog) if available.

Additionnally, a verbose log is available in $HOME/.liquidwar6/log.csv on

GNU/Linux and POSIX systems, in C:\Documents and Settings\<username>\Liquid
War 6\log.csv on Microsoft Windows and in /Users/<username>/Library/Application
Support/Liquid War 6/log.csv on Mac OS X.

You can read this using any spreadsheet software capable of reading csv file. It uses the tab (\t) character as a separator. It contains valuable informations including version and most default values for the game, and for each line logged, it says where in the code the log function was called. A must-have for debugging.

2.15 Report bugs

There are two ways to report bugs:

• send a mail to <bug-liquidwar6@gnu.org> (mailto:bug-liquidwar6@gnu.org);

• use the web-based Savannah bug tracker: http://savannah.gnu.org/bugs/ ?

    func=additem&group=liquidwar6.

The latter (Savannah (http://savannah.gnu.org)) is much preferred, because the mailing-list is bloated with spam... It also offers a list of bugs (http://savannah.gnu. org/bugs/?group=liquidwar6) which you should read before submitting a new one.

3 Hacker’s guide

This hacker’s guide is for anyone who is curious about the game, and wants to know how it works. It covers many aspects from simple map creation to technical program internals. A great effort has been done in Liquid War 6 so that it should be much more hackable than previous versions. Any feedback is welcome.

3.1 Designing levels

3.1.1 Why is level design so important?

As of Liquid War 5 (http://www.ufoot.org/liquidwar/v5), most levels have been contributed by players. While the maintainer of Liquid War 6 has technical knowledge to develop the game, artistic talent and taste might not be his domain of excellence 8-)

Therefore contribution are truely welcomed when they take the form of a new, original, fun and good looking level. It’s believed the levels often make the game much more than its engine. This is true for any type of game, and Liquid War is no exception.

So this section is here to help players understand how to hack existing levels, and create new ones, in the hope that 1) they can enjoy their own creations and 2) possibly share their work with others.

Note that this manual might refer to levels and maps: they are just two different names to describe the very same thing. It’s an alias.

3.1.2 Format overview

Liquid War 6 stores level information in a plain directory.

There is no such thing as an opaque .dat binary file. The name of the level is the name of the directory itself, and its elements are the files contained in it.

Files must follow a precise naming scheme. For instance Liquid War 6 expects a map.png file to be present in each map directory.

All image files in a level use the Portable Network Graphics (http://www.w3.org/ Graphics/PNG/) or JPEG (http://www.jpeg.org/) format. It is possible that in the long term, Liquid War 6 will be able to handle levels as .tar.gz or .zip files. In that case these files will only be a compressed image of the actual level directory.

See the ./map/ directory of the source Liquid War 6 distribution to see example of maps.

3.1.3 Resolution (map size)

Liquid War 6 does enforce a limit on map size. This is not to frustrate map designers and/or players, simply, it would be a lie to pretend the game can handle arbitrary big maps.

They might look great on your computer but will become unplayable soon on an older machine. And most of the time they don’t look that great, carefully crafted 1280720 just looks awesome and can represent a great level complexity.

Here are the technical limits:

Type  Max width  Max height  Max surface

Texture  3 000  2 000  6 000 000
Logical map  1 500  1 000  1 000 000

The texture can be somewhat bigger than the logical map, this allows for pretty levels while limiting the horsepower needed to move the fighters and animate everything. Note that you could technically feed the game with a map.png that is bigger than the logical map limit, only it will be downscaled when being loaded.

The texture limits are generous enough to accept a full-HD 1920x1080 image, or a 4/3 1600x1200 image, while the “one million pixels” logical map limit is enough to store a 16/9 1280x720 map or a 4/3 1024x768.

Keep in mind that the logical map (map.png) will probably be scaled whatsoever, even if it’s within the absolute limits (the game adapts the resolution to your computer speed) and your texture will rarely appear in its native resolution, will probably be distorted, and so on.

3.1.4 Metadata

Older versions of Liquid War 6 used to load a plain README file and use this as metadata. Title was take from map directory name. This is still supported, but it now also supports the addition of a metadata.xml file in which you can describe your map.

The following files can be defined:

• title: map title, what will appear in the menus

• author: map author

• description: description of the map, to help players when browsing folders

• license: map license (short version, just a simple one-liner, don’t use lenghtly copyright notices here, the README file would be the file to put long legal sections)

3.1.5 map.png

This is the only required file in a level.

In fact, the existence of map.png makes a directory a level. When checking wether a directory is a correct level, Liquid War 6 simply tests the existence and validity of map.png.

This image is a simple black & white area, where white zones are the background, the sea, the places where fighters can move, and black zones are the foreground, the walls, the places where fighters can’t go.

This informations can be stored in a 2-color indexed file, or in a grayscaled or even truecolor RGB file, but color information won’t be used. Internally, Liquid War 6 will read the color of every point. If it is over 127 on a 0 to 255 scale, it will be considered as background, if it is below 127, it will be considered as foreground.

3.1.6 layer2.png ... layer7.png

Liquid War 6 can handle mutiple layer maps. Think of a pile of maps, one being on top of the other. This allows you to create a volume, the game considers every layer has two axis x and y, and the z axis is to travel through layers. First layer corresponds to z=0, second layer to z=1, and so on.

Here are the files you can use to define layers:

• map.png this one is on top, it’s always defined (z=0)

  • layer2.png (z=1)
  • layer3.png (z=2)
  • layer4.png (z=3)
  • layer5.png (z=4)
  • layer6.png (z=5)
  • layer7.png (z=6)

A layerX.png file should be designed exactly like map.png. In fact, map.png could simply have been called layer1.png.

Up to 6 extra layers can be defined (from layer2.png to layer7.png). This is a hardcoded limit. It allows you to define 7 different layers, including the top map.png layer. Keep in mind this layer system is not real 3D, it’s more a “2D and a half” model. Adding layers can considerably slow down the game, so it’s wise to try and use as few layers as possible. Technically, 3 layers will allow you to build bridges and tunnels, which is probably the most usefull construction using layers. Fighters can also have difficulties navigating through layers so piling up layers in narrow “vertical” z-axis based tunnels is probably not a great idea.

   The ufoot/concept/pass map of the liquidwar6-extra-maps demonstrates basic layer

usage.

3.1.7 texture.png, texture.jpeg and texture-alpha.jpeg

It is possible to define a texture for the map by putting a texture.png or texture.jpeg file. It does not need to have the same dimensions as the map itself. Indeed, textures can be much more precise than the actual logical map.

There’s no theorical limit on how big a texture can be, more precisely, it can be much bigger than any hardware/driver maximum texture size. In practice, a too big texture will waste your video card RAM, and slow everything down. Sizes ranging from 640x480 to 1600x1200 are reasonable texture sizes.

If you don’t define this, the map.png file will be used as the texture, and also import colors from style.xml if defined.

Note that the shape of the texture defines the shape of the map, that is, the ratio with which it will appear on the screen.

The PNG alpha layer will be used for transparency. But to save disk space, it can be convienient to prefer the JPEG format, use texture.jpeg instead of texture.png and store the alpha layer in a separated file, called texture-alpha.jpeg. This avoids handling heavy PNG files, PNG compression not being performant on most textures.

In texture-alpha.jpeg, black is considered opaque, and white is transparent. Different levels of gray correspond to different levels of opacity. Previous versions of the game used the other way of doing things (black is transparent) because this is technically, the most obvious way to do things. Black is 0 and transparent is 0. But for a human “reader” of the map this does not make sense. One generally expects white to be the equivalent of “undrawn” or “blank”, well, if it’s undecided, void, transparent, whatever, it’s white. When the Gimp (http://www.gimp.org/) flattens an image, it becomes white, not black.

So white is transparent. Period.

3.1.8 glue.png and boost.png

If there’s a glue.png or boost.png file in the map directory (you can use one of them or both) then they will be interpreted as follow:

• on areas where glue.png and boost.png are white, nothing special happens, fighters follow their default behavior

• on areas where glue.png is black, fighters will be slowed down. How slowish they will be depends on the ’glue-power’ parameter. If ’glue-power’ is 3 then fighters will move three times slower.

• on areas where boost.png is black, fighters will behave faster. How fast they will be depends on the ’boost-power’ parameter. If ’boost-power’ is 2 then fighters will move two times faster.

• on areas where glue.png or boost.png are gray, they will be slowed down less or speeded up less depending on how dark the grey is.

There can be, at the same place, some gray or black in both boost.png and glue.png. How this will behave exactly is not really clear at this stage, the recommendation is not to do this (it does not really make sense anyway) but if you do it, game won’t complain.

It’s also wise not to abuse of boost.png for obviously, a map filled with “boosted” zones at a X10 pace will require much more CPU than the same map with no such setting. This might fool the automatic resampling algorithm and lead to maps that are unplayable. The spirit of boost.png is just to make a few spots go faster.

It’s also important to note that behaving faster or slower means moving faster or slower but also attacking faster or slower, and, in a general manner doing any action with a different pace.

3.1.9 danger.png and medicine.png

If there’s a danger.png or medicine.png file in the map directory (you can use one of them or both) then they will be interpreted as follow:

• on areas where danger.png and medicine.png are white, nothing special happens, fighters follow their default behavior

• on areas where danger.png is black, fighters die automatically, that is, they become black and loose health. How dangerous these zones are depends on the ’danger-power’ parameter.

• on areas where medicine.png is black, fighters regenerate faster, they become bright and shiny as if auto-healing. How efficient this medicine is depends on the ’medicinepower’ parameter.

• on areas where danger.png or medicine.png are gray, well, it’s in between, the “danger” and “medicine” effect will be proportional to the level of gray.

There can be, at the same place, some gray or black in both medicine.png and danger.png. How this will behave exactly is not really clear at this stage, the recommendation is not to do this (it does not really make sense anyway) but if you do it, game won’t complain.

3.1.10 one-way-<direction>.png

The four files:

  • one-way-north.png (AKA “up”)
  • one-way-east.png (AKA “right”)
  • one-way-south.png (AKA “down”)
  • one-way-west.png (AKA “left”)

can be used to force the fighters to go in one given direction, on some parts of the map. If an area is black on one of this meta-layers, then fighters will go in the given direction. For instance, a black zone in one-way-north will make fighters go to the north (AKA “up” direction) regardless of the cursor position. The fact that this is a one-way path is understood by fighters and they will take this in account when choosing the shortest path to go somewhere. You can combine vertical and horizontal one-way informations, making diagonal one-way paths.

3.1.11 cursor.png and cursor-color.png

By default, a simple cursor will be displayed, but you can use a custom per-map cursor. Cursors are defined by two 64x64 bitmaps:

• cursor.png is a PNG file, very likely to use transparency, which will be default be colorized according to the map colors. You can draw it any color, only greyscale informations will be used. You can keep the original colors if you really want to by setting colorize-cursor to false, but the default is to ignore the hue.

• cursor-color.png is another PNG file, very likely to use transparency too, which will always be colorized, replacing white by the team color, and black by the “dead” color, which by default is black and is usually a dark color. This colorization is a way to recognize your cursor and know which team it belongs to.

You can define only one of those bitmaps, if doing so, then the other layer will be empty, and won’t be filled with the default cursor data. Note that additionnally, a little letter (single character) will be displayed using the team color, so that’s yet another way to identify which teams the cursor belongs too. The PNG files really need to be PNG (JPEG won’t work) and need to be 64x64, any other size will be ignored.

3.1.12 rules.xml

Whereas style.xml is only about the appearance of the map, rules.xml allows the map designer to change pretty much any parameter.

Ultimately, the player can still ignore these settings and overide them with its own values, but the idea is: most game options are only pertinent in a given context. For instance, on some maps it’s interesting to move slowly, on some other it’s interesting to move fast. Some maps might be playable packed with fighters everywhere, some other might be much more fun with almost nobody on them.

The approach in Liquid War 5 (http://www.ufoot.org/liquidwar/v5) was to make the options available, but let the player himself find the right settings for the right map. The consequence is that no one ever used all those cryptic options in the advanced options menu, and probably 99% of the players ended up playing with default settings. This is not that bad, but given the fact that changing a few parameters one can totally transform the gameplay, it has decided been that in Liquid War 6, the map designer suggests the right options that matches his map.

This does not prevent the player from toying with options himself, he can still do it. There’s also one important point to note: all these options are technically implemented

as integer parameters. We certainly do not want any float here, since, and it is a Liquid War specific behavior, the game must be 100,00% predictable and behave the same on every platform. As there is nothing like exactness when speaking of floats, those are forbidden here. As for strings, we are dealing here with low-level internals, and this section is not about telling a story. They are technical options only. Booleans are implemented with the usual false = 0 and true = 1 convention. Note that other config files in Liquid War 6 might rely on floats, strings, and booleans with conventionnal true and false values, but not this one. rules.xml is special.

This rules.xml file is a direct image of the internal “rules” structure, so it contains technical, sometimes not very user-friendly parameters. While hacking rules.xml directly is a good way to test things, most of the time, the other file hints.xml contains more high-level informations that do the job the right way. A typicall example is speed.

See Section 4.11 [Map rules.xml], page 100.

3.1.13 hints.xml

This parameter is only used by the map loader. The map itself contains none of these parameters, they are only clues (hints, in fact..) on “how to load the map” which are passed to the loader.

Let’s take an example : speed. This rules.xml file has a (rather) easy to use “speed” parameter, which will do all the job of finding the right resolution for your map, the right “rounds-per-sec” and “moves-per-round” parameters, in short, it will set many other parameters to fit your needs.

As far as the map designer is concerned, rules.xml and hints.xml could have been merged (but so would have style.xml) but internally they are very different: rules.xml contains the real parameters, the one used by the algorithm whereas hints.xml contains only instructions which are used once when loading the map and then disappear. The core algorithm has no idea of what was in hints.xml, once it’s loaded.

See Section 4.12 [Map hints.xml], page 135.

3.1.14 style.xml

This is a simple XML file defining various appearance parameters. It has absolutely no effect on gameplay. These settings can ultimately be overriden by the player, but the idea is that if the map designer thinks this level looks better with this or that option, let him say it in this file.

See Section 4.13 [Map style.xml], page 141.

3.1.15 teams.xml

In this file one can specify per-map team settings. In short, this is where you can say how many bots you want, which color, and so on. This can be on a per-map basis, so that each map has different customized settings, some maps might be fun with only one bot, some other maps might be fun packed with 8 opponents.

Technically, teams.xml will allow you to define up to 4 players and 9 bots. This is an awfull lot considering there are only 10 colors. Basically, it’s OK to simply define:

  • 2 players (player1 and player2)

• 4 bots (bot1 and bot2)

It might also be a clever idea to just set up player2 and bot1 being the same color, in case of a conflict the game will pick up another color, but in practice those two entries often correspond to “the second player, bot or human, coming on the battlefield”.

All in all, this represents 5 entries to set up (main player, other player or first bot which can be the same, then 3 more bots), it’s OK to have the rest undefined or set to defaults.

Note that this can also simply be unset, and in that case the game defaults will apply, and the user will be able to change them, whereas if you set these up, the player will somewhat force to used the map settings.

See Section 4.14 [Map teams.xml], page 154.

3.1.16 Resampling

This is a very important point. Liquid War almost *always* resamples maps, unless you ask it not to do it. This is not recommended, it is believed in the general case, letting the internal algorithm make its own decisions is better than trying to figure out oneself “which is the best resolution”.

The reason is, the right resolution (we’re talking here of the logical resolution, how many fighters wide is the battlefield) often depends on the speed and general ressources the of the computer the program is running on. The map designer does not have this information. The program does. It runs a bench at startup. So this way it can choose, at runtime, the resolution which fits best.

The recommended way of doing things is not to try to be too picky about rules.xml parameters related to speed and also let the default map size limits in hints.xml to their defaults. Do not use them unless debugging stuff. Then the program will resample the map so that the player can play on it at a reasonnable speed. If map is too big, and it’s often the case, then it will downsize it until there are sufficiently few fighters so that the CPU can handle the job. This, of course, is not rocket science. The bench calculation is a somewhat brute-force approach of doing things. Formally, we would have to run the map for good to figure out what is the right speed. Still, this bench gives good approximations.

Previous versions of the game relied heavily on ’fighter-scale’ to resample maps, but this is not the case anymore. The ’fighter-scale’ is now a minor parameter which is used to upsize maps if they are too small. In 99.9% of the cases, the map is first upsized by ’fighter-scale’ for this parameter is by default set low (1.0) then downsized by ’bench-value’ for real-life personnal computers can’t handle 1600x1200 maps in real-time. Not yet.

There are a bazillion options to control map size, including ’min-map-surface’. They are here because it’s important that, ultimately, people can do whatever they want with the game. But for map design, this is another story. Don’t use them. Rely on ’bench-value’ and just care about game speed. This is achieved by changing the “speed” parameter.

3.1.17 Music

It is possible to store your own custom music file within the map directory. You can call it whatever you want (you can keep its original name, which is something music authors usually appreciate, even if there’s no strong “attribution” clause on the license, it can be considered fair use not to fiddle to much with the name) you just have to place it in the same directory than the other files like map.png or texture.jpeg.

The following formats are known to work with the default SDL mixer (http://www.

libsdl.org/projects/SDL_mixer/) based mod_ogg backend:
  • ogg (Ogg Vorbis (http://www.vorbis.com/) files)
  • wav

• midi (extensions .mid and .midi should both work)

• mod, s3m and xm files, AKA “modules”.

To be more precise, here’s how things work:

• step 1: the game tries to find the file music-file (parameter taken from style.xml or defined/overriden by player) in the current map directory; step 2: if not found, it will try every path in music-path to find this file. This includes the “system” music directory with musics that ship with the game, but also the ./music subfolder in the user directory; step 3: if still not found, it will try to play a random file, relying on music-filter to ignore some files.

3.1.18 Experience (“exp”)

In rules.xml you can set a special parameter which is exp and allows you to tell “a player can’t load this map if he doesn’t have at least N at his/her exp rating”. Gaining exp (stands for “experience”) isn’t hard, you just need to win a level with exp=N to gain exp=N+1.

By default, the player’s exp is 0 and levels default to 1, so this means only levels with exp set explicitely to 0 in rules.xml might be used. Then player wins that level and is given access to all maps by default, unless these are explicitely set with exp greater than 1.

In solo game, when a player wins a level, he’s automatically redirected to the map which is in the same directory and has exactly the exp he just gain. For instance, if you win a map with exp=5 then you’re chained to the first map (in alphabetical order) which has exp=6. By setting up the exp parameter the right way, with a map for each exp level one can transform a simple map directory in a scenario that player will automatically follow.

Last, but not least, the game, at startup, only allows you to play red, green, blue and yellow. Other colors are unlocked as you progress among levels. Same things with weapons, there are “liberated” continuously through the game.

This mechanics allows the following behavior:

• when game is launched first, only a small subset of maps are accessible

• after you win one map (sort of quite easy) you gain access to the next level, plus many of the maps of the extra package.

• after each map you win, you’re redirected to the next map, and regularly, you gain access to new colors/weapons

As a final word, yes, it’s possible to cheat, fool the exp system, but it’s believed this is moot and lame.

3.2 Translating

3.2.1 Using gettext

Liquid War 6 uses GNU gettext (http://www.gnu.org/software/gettext/) for all its
messages. There’s an online manual (http://www.gnu.org/software/gettext/manual/

gettext.html) about this tool. In practice, what you have to do as a translator is to edit the po/xx.po file with xx being your language / country code. For instance, to translate the game in French, one needs to edit po/fr.po.

3.2.2 Formatted strings

This is very important, you might already be aware of it if you are familiar with gettext, but still it’s worth mentionning : when a string contains special characters such as %d or %s (in a general manner, anything with a % it’s important that all translations contain exactly the same number of %ds and %ss than the original.

For instance:

     "foo has %d bars (%s)"

can be translated to:

     "ziblug zdonc %d zuc - %s - tac"

The number, order and type of % entries is preserved. To learn more about these formats, use info printf or man 3 printf. In a general manner, get informations about printf

(http://en.wikipedia.org/wiki/Printf).

Additionnally, some strings are used by Scheme (Guile) code and not by C code. Thus, they don’t use the standard C/printf convention. In these strings, what you must preserve and be aware of is the tilde character ~. Very often you’ll see ~a in a string. As with the printf %, you must preserve the number, order and type of those. There is a

complete online reference (http://www.gnu.org/software/guile/manual/html_node/

Formatted-Output.html) about this way of formatting strings.

3.2.3 Partial translation

Liquid War 6 has thousands and thousands of messages which could theorically be translated. In practise it’s counter-productive to spend time to translate those, as the game is still evolving constantly, and as most of these messages are technical messages which inform about rare bugs and strange conditions. All sort of informations which, while valuable, are not intented for end-users and are more destinated to be reported in bug reports. To select only the interesting messages to translate, the current gettext configuration only uses a reduced set of files.

• src/scriptpo.c : the most important file. It contains the definitions used by all the Guile code, this is where you’ll find all the menu labels.

• src/lib/sys/sys-log.c : log messages and keywords. These are not the log messages themselves, it only concerns the log engine. One can for instance replace WARNING by

    ATTENTION.

• src/lib/hlp/hlp-credits.c : the credits, which are displayed at game startup in the splash screen.

• src/lib/lw6-print.c : contains some messages printed on the console.

As a side note, the file src/lib/hlp/hlp-reference.c contains all the entries for the various configuration options, anything that can be queried by liquidwar6 --about=<keyword>. This is several hundred messages. It might be interesting to translate them some day, but it’s obviously not a priority today.

3.3 Architecture

3.3.1 C + Guile

Technically, Liquid War 6 is a collection of C functions which are exported to Guile. The main binary embeds a Guile interpreter, which will run a Guile script. This script calls the exported C functions, and glues them together.

It should be possible to implement the game without using Guile at all, using C code to make the various modules communicate together. This might seem an easier way to go, not involving several languages. However, using this script level is a good way to achieve several important goals:

• it’s possible, at any time, to query the game about its internal state, dump objects, take actions. That’s what the console is about. It’s a bit like having an embedded debugger, it’s really a very convenient tool to develop, make experiments and track problems.

• many hacks can be done without recompiling anything at all. Simply edit a few files with an editor, and your patch is running. Once the binary base is set up, hacking scripts on top of it is (almost) a piece of cake.

• forcing the program to use scripts to transfer informations from a module to another is a good way to avoid “spaghetti” code, when modules cross-use each other in an uncontrollable way. Of course in some cases, modules communicate directly, especially when performance is important. But for many tasks, it’s just very comfortable and safe to have module A send orders to module B through a high-level script API.

Having Guile to implement high-level stuff also decreases, to some extent, the need for object-oriented features of C++. The big picture is : low level code that require speed, optimized data processing, is written in C. Code which is more high level and requires abstraction is written in scheme.

3.3.2 Threading and SMP

Liquid War 6 makes a heavy usage of threads. Early versions of the game did not have this feature but starting with 0.0.7beta, one can really consider the game is heavily threaded.

There’s basically:

• a thread to handle the main control flow. This thread runs scheme code which Guile. It’s not the most CPU-greedy thread, but when it’s stalled, there’s no more interaction between the user and the program.

• a thread to handle the display. Depending on rendering options, this thread can consume lots of CPU cycle. On a single processor/core system, it can be interesting to lower rendering options in order to gain speed on other aspects of the game. On a quad-core system, it’s probably useless, just play with all bells and whistles activated.

• two threads to run the core algorithm. One maintains the so-called reference state, the other being dedicated to the draft sate. In a local game there’s no draft state so only one of those two threads is used. There’s even a technical optimization which can be turned on and can theorically use even more threads and be efficient on very big maps but well, it’s rather untested and still has to prove its real efficiency.

• a thread to handle map loading. This one is not active all the time, it’s just here to keep a preemptive interface while loading complex maps.

• network code can also fire threads, especially when connecting on remote systems.

So globally, if you have an SMP system, the game will be happy with it. It will also run on a single processor, as the program uses POSIX pthreads it’s capable to run on any computer which has pthreads implemented for it.

But, and this is a strong limitation, without pthreads, the game won’t run. At all. Or at least, not unless it’s almost completely rewritten.

3.3.3 Internal libraries

The C code is splitted into several internal libraries. This allow independant testing of various game modules.

The main module, the most important one, is libker, (stands for “kernel”). This is were the core algorithm is. To some extent, the rest of the code is just about how to provide this module with the right data and environment. Logically, if you profle the game, you should find out that a great part of the CPU time is spent here. Code here is about spreading gradients, moving fighters and cursors.

The libmap module is here to handle maps, it contains the code to manipulate maps in memory. But it does not know how to load them from disk. This is the responsability of another module, libldr, which is linked against libraries such as libpng (http://www.

libpng.org/pub/png/libpng.html) or libjpeg (http://www.ijg.org/) and does the job

of transforming those standard formats into a usable in-memory structure. The libgen module also works the same way, creating pseudo-random maps. There’s still a another moduled involved in map handling, it’s libtsk, whose job is to load a level in the background. It has a 2-steps asynchronous loading system which allows the game to load maps while the user interface is still responsive, and give a preview of the map as soon as possible, when loading continues in the background, building optimizing structures which are usefull when playing but not mandatory just to show the map.

At the other end of the algorithm-chain, the libpil module will “pilot” things. It’s this module which will translate text readable orders (typically adapted for network usage) into function calls. It has event lists, keeps them in the right order, and will also permanently maintain three different states of the game. A backup state which can be used any time to go back in time and get the game in a stable 100% sure state. A reference state which is correct but ever changing. Basically backup plus all the orders received between backup and reference gives reference. And finally a draft state which is as up to date as possible but might be wrong. This is typically interesting in network game, where we want to show something moving, something fast, even if there’s lag on the network and other computers fail to send information in time. In this case we display draft while still keeping reference and updating it when we finally receive valid informations. Backup would be used to send bootstrap information when people are joining a new game, or to check up if things are going right.

A special libbot module is here to handle bot algorithms. A bot is just a simple move function which takes a game state as an input, and returns an x,y position, just the way a mouse handler would. How complex a bot is “under the hood” depends on the type of bot. Current bots are really basic. Additionnally, libsim will run dummy fight simulations to find out wether some team has a real advantage on another one, speaking of team profiles depending on colors.

The libgfx module handles all the graphics stuff. It is itself splitted in several submodules, that is, it does not do anything but load a module such as mod-gl1 which will actually contain the implementation. In an object-oriented language, it would be an abstract class, an inteface. The implementation does not need to be thread-safe. It’s better if it is, for theorically it could be possible to fire Liquid War 6 with two display backends running at the same time on the same game instance, but this code has yet to be written, and it’s a rare dual headed configuration which probably has no real-life usage. If only one graphics backend is activated at a time, the rest of the implementation garantees there will never be two concurrent calls to a function is this module. It is the libdsp (“display”) which handles this. It fires a thread for rendering purposes, and sends information to this thread, detecting automatically if it’s necessary to acquire a mutex and update rendering informations. For the caller, this is transparent, one just has to call an update function from time to time. The module will even perform “dirty-reads” on a game state being calculated, to render things in real time, as soon as possible.

An experimental libvox module is under design/development and might, in the future, provide a real-time voxel renderer. Still pre-alpha stage.

To ease up the implementation of different graphics backends, a libgui module contains code which is meant to be used by any graphics backend. It’s just a factorisation module, containing common code and interfaces, related to displaying things. This is where, for instance, one can find a high level menu object. In the same spirit, libmat contains generic math, vector and matrix code, which is commonly used in 3D interfaces.

The libsnd module handles sound. It’s also an abstract class, an interface, which uses dynamic backends as implementations.

The libnet module is a wrapper over different network APIs, it handles Winsock and POSIX sockets in a uniform manner. The libcli and libsrv contain network client and server code, implementing the various protocols in dynamically loadable sub-modules. It’s the role of libp2p to glue this together, handle the list of available servers, the message queue, verifying nobody is cheating, and so on. All this modules share information about current game state using code & structures defined in libnod,use message utilities (format, parse) defined in libmsg and share code concerning connections in libcnx. Additionnally, libdat provides facilities to store old network messages and sort them.

The libsys module contains most system and convenience functions, it handles logs, type conversions, timer, memory allocation, it’s the fundamental module every other module depends on. It has a compation libglb module with all the Gnulib (http://www.gnu.

org/software/gnulib/) shared code.

The libhlp is used to handle keywords and internal self-documentation (this is what is used by --list and --about), libcfg knows how to read and save config files, libcns handles the console, and libdyn can load .so shared files dynamically.

To glue all this, there are some Guile bindings with helper functions available in libscm which fills two needs, one being an easy way to check if Guile linking is working correctly without requiring all other modules to be available, and also performing automatic checks on some actions such as registering or executing a function.

Finally there are small modules like libimg (to make screenshots of the game) which have been separated because they required special libraries to link with and/or did not really fit in existing modules for dependencies reasons.

So well, this is a lot of modules. The list might move a bit, but the big picture is here. Each module is testable separately.

Below is a Graphviz (http://www.graphviz.org/) diagram, which shows the modules dependencies.

3.4 Memory structures

The most important memory structures in Liquid War 6 are:

• map (lw6map_level_t) : this contain the map immutable informations. This is what resides in memory after a map has been loaded from the disk. It contains all the various .png and .jpeg files stored as pixel arrays, resampled if need, and also contains the various map attributes. Once this structure is ready, the game is capable of displaying the map on the screen, but it can not do anything with it yet.

• game struct (lw6ker_game_struct_t) : this one contains the same informations as the previous structure, only the information has been post-treated so that it’s ready for use by the core algorithm. It will, for instance, contain the famous mesh structure, which groups squares by packets of 1, 4, 16, 64 or more. The reason it’s been separated from the level is that operations such as creating the mesh might require a lot of time. So to allow players to see the level while black magic is still running in the background, it was required to make a difference between what is required to view the map (“level”) and what is required to play on it (“game struct”).

• game state (lw6ker_game_state_t) : contains all the variable, ever changing game data. This is where the position of fighters is stored, their health, and such things. It is designed to be synchronizable by using mostly simple calls to memcpy. It heavily relies on the previous structures, the idea is that one can have several “game state” plugged on a single “game struct”.

All these structures are defined in the ker/ker.h header.

3.5 100% predictable algorithm

The core Liquid War 6 algorithm is 100% predictable, that is to say, given the same input, it will produce the same results, on any computer. Previous versions of the game also had this property. This is very important for network games, since in a network only informations such as “cursor A is at position x,y” are transmitted. Every node maintains its own internal game state, so it’s very important that every node comes with the same output given the same input.

For this reason Liquid War 6 never uses floating point numbers for its core algorithm, it uses fixed point numbers instead. It also never relies on a real “random” function but fakes random behavior by using predictable pseudo-random sources, implementation independant, such as checksums, or modulos.

There are also some optimizations which are not possible because of the predictability requirement, for instance one can not spread a gradient and move the fighters in concurrent threads, or move fighters from different teams in different threads.

If you read the code, you’ll find lots of checksums here and there, a global checksum not being enough for you never know where the problem happened. The test suite uses those facilities to garantee that the game will run the same on any platform.

Not being able to rely on a predictable algorithm would require to send whole game states on the network, and this is certainly way too much data to transmit. A moderate 200x200 map has a memory footprint of possibly several megabytes, so serializing this and sending it to multiple computers at a fast-paced rate is very hard, if possible at all, even with a high bandwidth. We’re talking about Internet play here.

3.6 Graphics backends

3.6.1 Modularity

Liquid War 6 has a modular architecture which allows the programmer (and the player) to plug pretty much any rendering/graphics backend, provided this one is... developped.

As of 2009 the only available backend was mod-gl1, it would display the game using 3D acceleration, if available, through the SDL (http://www.libsdl.org/) library, using its GL bindings.

As of 2012, other backends are begin developped, the idea is that each backend can provide the user with enough visual feedback to play, and convey input informations to the core engine.

The rest of the game is exactly the same, this means playing with mod-gl1 you can do exactly the same things than with mod-caca.

3.6.2 List of backends

• mod-gl1 Liquid War 6 has a modular architecture which allows the programmer (and the player) to plug pretty much any rendering/graphics backend, provided this one is... developped.

As of 2009 the only available backend is still mod-gl1, it will display the game using 3D acceleration, if available, through the SDL (http://www.libsdl.org/) library, using its GL bindings. Additionnally, versions available for Microsoft Windows and Mac OS X will probably never any other backends available. For technical reasons, those platforms do not have the flexibility of GNU/Linux and do not allow graphical libraries to be loaded dynamically. In practice, both of them require hacks that override the standard main function. Microsoft Windows has its WinMain instead, and Mac OS X is even more pedantic, requiring graphical functions to be executed in the main thread. So mod-gl1 is just linked statically in those versions, and the modularity of the game is purely theorical on these platforms. This mod-gl1 module is really one of the key stones of Liquid War 6, and if you want to change graphical things, it’s definitely the place to hack on. The source is in

src/lib/gfx/mod-gl1.

The mod-gl1 backend requires “moderate” hardware, but it still does require hardware acceleration. Pure software rendering through mesa (http://www.mesa3d.org/) for instance, won’t be enough. So if you’re running Xorg on GNU/Linux and there’s a DRI driver for your card, the game should run fine. On the programmer side, the counterpart is that one should not rely on fancy OpenGL features. Textures have a maximum size of 512x512 for instance. Of course some maps are bigger than this but this means that internally, mod-gl1 splits them into smaller tiles, and displays those tiles one by one. Inside the mod-gl1 backend, the src/lib/gfx/mod-gl1/gl-utils directory contains lots of common structures, factorized functions which can (and should, if appliable) be used.

• mod-gles2 This is under development, the idea is to provide an alternative renderer based on OpenGL ES 2, which could be used on standard computers but also on mobile platforms. Work in progress, don’t hold your breath.

• mod-soft This is under development, the idea is to provide a very basic rendered which can be compiled pretty much anywhere as long as SDL is available, since it does use software rendering only. Work in progress, don’t hold your breath.

• mod-caca This is under heavy development, the idea is to provide a basic yet surprising alternative text-based renderer, using libcaca.

3.6.3 How to write a new backend

The starting point for any hack are the files src/lib/gfx/gfx.h. This is where the API is defined.

Basically, the type lw6gfx_backend_t contains all the required callbacks. You must provide an implementation for each function.

Let’s take an example, taken from mod-gl1. When calling lw6gfx_get_video_mode and passing it a first argument which is a valid mod-gl1 backend, the function mod_gl1_utils_ get_video_mode will be called. How this is done is a little C casting wizardry.

To understand how this works, read the files:

• src/lib/gfx/gfx-api.c: contains all the functions which are part of the API and can be called elsewhere in the code.

• src/lib/gfx/gfx-register.c: contains the code that allows a module to be loaded/unloaded at runtime. Will act differently if the games is compile with the --allinone flag, but for the caller this is transparent, just create and destroy backend, period.

• src/lib/gfx/mod-gl1/mod-gl1-backend.c: this is where the module actually binds

its internal functions with the callbacks defined in the lw6gfx_backend_s struct. None of these internal functions should be called directly, code in libdsp for instance should only refer to the lw6gfx_... bindings. Reading the code in src/lib/gfx/gfx-test.c shows how these functions can be called, and in which order.

All the functions should be defined, but some of them are obviously more important. The two most critical functions are:

• pump_events This is used to process inputs. The function should update a lw6gui_ input_s struct and return it to the caller. How this done is really up to the backend, it happens that all SDL based backends (mod-gl1, mod-gles2 and mod-soft) share the same code for this, but another backend could do this differently, there’s no real need to use SDL. Only, the returned input should behave correctly when queried with function from libgui. As a consequence, one needs to have a look at libgui to understand how

input works. A look at src/lib/gfx/shared-sdl/shared-sdl-event.c is a good

example of this, as this file contains the implementation for SDL-based input. See Section 5.19 [libgfx], page 219. See Section 5.26 [libgui], page 221.

• display By far the most complicated function, this one is called on each display loop to render the game. It’s always used in the same thread, so need not be reentrant, and on some platforms (eg Mac OS X) it will even be called in the main thread (this can be of some importance regarding some libraries such as SDL). Still, beware, the game_state object it uses can change on the fly while rendering. In that case “changing” means that fighters can move and gradients be updated but the global structure won’t change. So any pointer on a fighter will still be valid after it’s been obtained, but the renderer should not expect the game to be static. In practice this is not really a problem. If you are curious, you can look in libdsp how and when this function is called. A very important parameter is mask, depending on its value, the backend should, or not, display the menu, or the map, or both, etc. The reference for this are the LW6GUI_

DISPLAY_... constants in src/lib/gui/gui.h.

As a starting point, implementing menu display before anything else is probably the best bet, since without menus it’s hard to do anything within the game.

To test out a backend, one can either launch the full game using the “under development” backend, or launch the test suite by typing ./liquidwar6gfx-test 1 in

./src/lib/gfx.

See Section 5.16 [libdsp], page 219.

3.7 Core algorithm

3.7.1 Introduction

Since Liquid War 3 (http://ufoot.org/liquidwar/v5/techinfo/algorithm) the algo-

rithm to power Liquid War is pretty much unchanged, at least, there has been no revolution, the basic principles remain the same. This has been explained in Liquid War 5 doc

(http://ufoot.org/liquidwar/v5/techinfo/algorithm), but is repeated here, along

with the specific Liquid War 6 stuff.

The very important things to remember are:

• The algorithm is 100.00% predictable. This means given the same input, it will give exactly the same output. This is very important for the network games to work correctly, therefore, the algorithm does not ever use any call to rand / random functions, it also does not use any float value either, since different type of processors/contexts might give slightly different results because of rounding problems.

• It’s a two-pass algorithm, the first step is to calculate the distance from any point of the map to the closest cursor. This step is always imperfect, the shortest path is never really found, the naive approach is to consider that if a place on the map is at distance N of the cursor, then in the worst case, all adjacent places are at distance N+1. As of Liquid War 6, the corresponding code is in src/lib/ker/ker-spread.c. The second step is to move the fighters, make them act. In Liquid War 6, the corresponding code is in src/lib/ker/ker-move.c. One can have a look at the code source for the function

lw6ker_game_state_do_round in src/lib/ker/ker-gamestate.c to see how these

are called.

3.7.2 Level, game struct, game state and pilot

Most of the algorithm code has something to do with the following types (which are structs): • lw6map_level_t defined in which is used to store the level data.

• lw6ker_game_struct_t defined in src/lib/map/map.h which is used to store the mem-

ory data required by the algorithm, but which are immutable. There’s a difference between those data and the ones stored in the level struct. For instance, those data are “private” since lw6ker_game_struct_t is opaque, while everything is lw6map_level_ t is “public”. Also, data in lw6ker_game_struct_t might be highly redundant for performance issues and is optimized for speed while data in lw6map_level_t is just plain data and won’t change if the algorithm is updated.

• lw6ker_game_state_t defined in src/lib/ker/ker.h which is used to store the level

data required by the algorithm, and which changes during the game. This is typically where an information such as “there’s a red fighter in slot (3,23,1)” will be stored.

• lw6pil_pilot_t defined in src/lib/pil/pil.h which is used to handle all the threading issues. It keeps a track of 3 game states. A “reference” state which is the state of the game considering all input has been received from the network, and is validated. A “draft” state which might be anticipated and updated “as if the players we did not get input from did not move there cursors”. This can give the illusion that the game is running smoothly while in reality input from other players on the network is choppy. In a local game, “draft” and “reference” are equivalent, since there’s no doubt about what’s on the network. And finally, a “backup” state which can be pulled in case of a serious flaw and is a good way to solve the “hey, someone untrusted is throwing garbage on the net”. One can always pull a backup.

Most of the time, hacking on the algorithm, changing the gameplay, does not require to touch anything but the code in src/lib/ker.

See See Section 5.31 [libmap], page 222. See See Section 5.29 [libker], page 222. See See Section 5.37 [libpil], page 224.

3.7.3 Getting informations about where fighters are

One of the key functions is lw6ker_game_state_get_fighter_id, which will return the id of a fighter at a given position. Then its companion function lw6ker_game_state_get_ fighter_by_id can be called, it will return a lw6ker_fighter_t, which contains the real data.

The type lw6ker_fighter_t is not opaque and can be freely accessed by the caller, which, typically, is a graphics backend trying to display informations. Try and grep for the string “lw6ker game state get fighter id” withing the src/lib/gfx source tree for examples.

One thing that is very important when hacking on libker: you should always leave the lw6ker_game_state_t struct in a state that is compatible with a correct usage of public “getters” in src/lib/ker/ker.h. The reason is that this code can be executed by separate threads, more precisely, in “dirty read” mode, the rendering thread will try and display a “game state” while this very “game state” is being updated by another thread.

3.8 Compilation tips

3.8.1 Advanced ./configure options

In addition to all the common Autoconf (http://www.gnu.org/software/autoconf/)

switches such as --prefix, Liquid War 6 has some custom switches:

• --disable-console: allows you to turn on/off console support. Normally this is detected automatically but in case you really want to disable it on platforms which support it, you can. This will cause the program no to link against libreadline, among other things.

• --disable-gtk: allows you to turn on/off gtk support. Normally this is detected automatically but in case you really want to disable it on platforms which support it, you can. This will cause the program not to link against GTK libs.

  • --disable-cunit: allows you to turn on/off CUnit (http://cunit.sourceforge.

net/) support. Normally this is detected automatically but in case you really want to disable it on platforms which support it, you can. This will cause the program not to link against CUnit libs.

• --enable-optimize: will turn on optimizations. This will turn on compiler options such as -fomit-frame-pointer but also disable some code in the program. Indeed, most of the advanced memory checking in the game - which ensures it does not leak - will be turned of. This will certainly speed up things, however, it’s not recommended to turn this on until program is not stable enough so that memory leaks and other problems can be declared ’impossible’. Turn this on if you really have some speed problem, otherwise it’s safer to use the full-featured ’slow’ version of the game.

• --enable-paranoid: will turn on very picky and pedantic checks in the code, try this when you suspect a serious memory bug, a race condition whatsoever, and want to track it down. Useless for players.

• --enable-headless: will allow compilation without any graphics backend. The game is unplayable in that state but one can still wish to compile what is compilable, for testing purposes.

• --enable-silent: will allow compilation without any sound backend. The game won’t play any music in that state but one can still wish to compile what is compilable, for testing purposes.

• --enable-allinone: will stuff all the internal libraries into one big executable. Very convenient for profiling. The major drawback is that you need to have all the optional libraries installed to compile all the optional modules. Another side effect is that with this option there’s no more dynamic loading of binary modules, so if your platform has a strange or buggy support for .so files, this option can help.

• --enable-fullstatic: will build a totally static binary, that is using the --static option for gcc and the -all-static option for libtool. Currently broken, this option could in the future allow for building binaries that run pretty much everywhere, without requiring any dependency but a Kernel.

• --enable-gprof: will enable profiling informations. This will activate --enable-allinone, else you would only track the time spent in functions in the main liquidwar6 executable, and exclude lots of interesting code contained in dynamic libraries.

• --enable-instrument: will instrument functions for profiling. This will turn on the

-finstrument-functions switch when compiling, so that the hooks __cyg_profile_

func_enter and __cyg_profile_func_exit are called automatically. Then you can link against tools like cprof (http://cprof.sourceforge.net/) or FunctionCheck

(http://sourceforge.net/projects/fnccheck/).

• --enable-profiler: will enable Google Performance Tools (http: / /code . google.com/p/google-perftools/) support. Basically, this means linking against libtcmalloc and libprofiler. You could activate those by using LD_PRELOAD or by using your own LDFLAGS but using this option will also make the game tell you if CPUPROFILE or HEAPPROFILE are set when it starts. The pprof -gv output is very handy. Note that on some systems pprof is renamed google-pprof.

• --enable-gcov: will enable coverage informations, to use with gcov (http://gcc.

gnu.org/onlinedocs/gcc/Gcov.html) and lcov (http://ltp.sourceforge.

net /coverage /lcov .php). This is for developpers only. It will activate --enable-allinone, else there would be some link errors when opening dynamic libraries. The obtained information is available online: coverage (http://www.ufoot.

    org/liquidwar/v6/doc/coverage/). and GNU global (http://www.ufoot.org/
    liquidwar/v6/doc/global/).

• --enable-valgrind: will enable some CFLAGS options which are suitable for the use of Valgrind (http://www.valgrind.org/), to track down memory leaks and other common programming errors. Use for debugging only, usually together with --enable-

    allinone.

3.8.2 Debian packages

Liquid War 6 does have a ./debian in both main and extra maps packages, so it’s “debianized”. To build the main .deb package, untar the main source tarball, then:

make dist
cd pkg
cp ../liquidwar6-X.Y.Z.tar.gz . # X.Y.Z is the version
make deb

Note that you have to copy the source tarball to ./pkg and move to this directory before typing make deb. This is, among other things, to simplify the main Makefile.

To build the extra maps .deb package, untar the extra maps tarball, then:

make deb

3.8.3 Red Hat packages

Liquid War 6 does have a .spec files in both main and extra maps packages. To build the main .rpm package, untar the main source tarball, then:

make dist
cd pkg
cp ../liquidwar6-X.Y.Z.tar.gz . # X.Y.Z is the version
make rpm

Note that you have to copy the source tarball to ./pkg and move to this directory before typing make rpm. This is, among other things, to simplify the main Makefile.

To build the extra maps .rpm package, untar the extra maps tarball, then:

make rpm

3.8.4 Microsoft Windows msys/mingw32 port

This section describes how to compile the game from source under Microsoft Windows. Note that players are encouraged to use a free system such as GNU/Linux, which is the platform Liquid War 6 is being hacked on by default. If you encounter problems with this port, you’ll probably save time by installing a double-boot with GNU/Linux coexisting with your previous Microsoft Windows install.

Basically, Liquid War 6 requires MinGW (http://www.mingw.org/). More precisely,

it requires MSYS (http://www.mingw.org/msys.shtml). A standard Cygwin (http://

www.cygwin.com/) installation won’t work, because it is too UNIXish to allow third party libraries like SDL (http://www.libsdl.org/) to compile natively. You might argue that SDL is available for Cygwin, but in reality, the Cygwin port of SDL is a MinGW port. Indeed, Cygwin brings all standard POSIX functions including the use of main instead of WinMain and I suspect this is a problem for graphical libraries like SDL which do require some sort of direct access to the OS low-level functions. Therefore, MinGW is more adapted for it does not define all these functions, and allows any library to hook on Microsoft Windows internals directly. Point is then, you also loose the cool effect of Cygwin which is to have a complete glibc (http://www.gnu.org/software/libc) available, including network functions like select defined the POSIX way, and not the WinSock way. If you ever ported code from POSIX sockets to WinSock 2, you know what I mean. Using MinGW is also embarassing for some libraries won’t compile easily, and for instance programs which heavily rely on a real TTY interface to work are usually hard to port. This includes ncurses

(http://www.gnu.org/software/ncurses/) and GNU readline (http://cnswww.cns.

cwru.edu/php/chet/readline/rltop.html). Liquid War 6 tries to have workarrounds for all this, and in some cases the workarround is simply that embarassing code is not compiled on Microsoft Windows. For this reason, some features are not available on this platform. Period.

Now the reason you need MSYS and not only MinGW is that MSYS will allow ./configure scripts to run, and this eases up the porting process a lot. MinGW and MSYS packages are downloadable on the SourceForge MinGW download page (http://

sourceforge.net/project/showfiles.php?group_id=2435). Alternatively, there is a
mirror on ufoot.org (http://www.ufoot.org/download/liquidwar/v6/mingw32/pkg/),

but files might be outdated.

To compile Liquid War 6, first download and unzip all the following files in the same directory, for instance C:\MSYS. If you do not have any tool to handle .tar.gz and .tar.bz2 files under Microsoft Windows, which is likely to be the case when MSYS is not installed yet, you can untar these on any GNU/Linux box, then upload the whole directory to the target Windows host.

• autoconf2.5-2.61-1-bin.tar.bz2

• autoconf-4-1-bin.tar.bz2

• autogen-5.9.2-MSYS-1.0.11-1-bin.tar.gz

• autogen-5.9.2-MSYS-1.0.11-1-dev.tar.gz

• autogen-5.9.2-MSYS-1.0.11-1-dll25.tar.gz

• automake1.10-1.10-1-bin.tar.bz2

• automake-3-1-bin.tar.bz2

• bash-3.1-MSYS-1.0.11-1.tar.bz2

• binutils-2.18.50-20080109-2.tar.gz

• bison-2.3-MSYS-1.0.11-1.tar.bz2

• coreutils-5.97-MSYS-1.0.11-snapshot.tar.bz2

• crypt-1.1-1-MSYS-1.0.11-1.tar.bz2

• csmake-3.81-MSYS-1.0.11-2.tar.bz2

• cvs-1.11.22-MSYS-1.0.11-1-bin.tar.gz

• diffutils-2.8.7-MSYS-1.0.11-1.tar.bz2

• findutils-4.3-MSYS-1.0.11-1.tar.bz2

• flex-2.5.33-MSYS-1.0.11-1.tar.bz2

• gawk-3.1.5-MSYS-1.0.11-1.tar.bz2 • gcc-core-3.4.5-20060117-3.tar.gz

• gcc-g++-3.4.5-20060117-3.tar.gz

• gcc-g77-3.4.5-20060117-3.tar.gz

• gcc-objc-3.4.5-20060117-3.tar.gz

• gdb-6.8-mingw-3.tar.bz2

• gdbm-1.8.3-MSYS-1.0.11-1.tar.bz2

• gettext-0.16.1-1-bin.tar.bz2

• gettext-0.16.1-1-dll.tar.bz2

• gettext-0.16.1-MSYS-1.0.11-1.tar.bz2

• gettext-devel-0.16.1-MSYS-1.0.11-1.tar.bz2

• inetutils-1.3.2-40-MSYS-1.0.11-2-bin.tar.gz

• libiconv-1.11-1-bin.tar.bz2

• libiconv-1.11-1-dll.tar.bz2

• libiconv-1.11-MSYS-1.0.11-1.tar.bz2

• libtool1.5-1.5.25a-1-bin.tar.bz2

• libtool1.5-1.5.25a-1-dll.tar.bz2

• libtool1.5-1.5.25a-20070701-MSYS-1.0.11-1.tar.bz2

• lndir-6.8.1.0-MSYS-1.0.11-1.tar.bz2

• lpr-1.0.1-MSYS.tar.gz

• lzma-4.43-MSYS-1.0.11-1-bin.tar.gz

• make-3.81-MSYS-1.0.11-2.tar.bz2

• mingw-runtime-3.14.tar.gz

• mingw-utils-0.3.tar.gz

• minires-1.01-1-MSYS-1.0.11-1.tar.bz2

• MSYS-1.0.11-20071204.tar.bz2

• msysCORE-1.0.11-2007.01.19-1.tar.bz2

• openssh-4.7p1-MSYS-1.0.11-1-bin.tar.gz

• openssl-0.9.8g-1-MSYS-1.0.11-2-bin.tar.gz

• openssl-0.9.8g-1-MSYS-1.0.11-2-dev.tar.gz

• openssl-0.9.8g-1-MSYS-1.0.11-2-dll098.tar.gz

• perl-5.6.1-MSYS-1.0.11-1.tar.bz2

• perl-man-5.6.1-MSYS-1.0.11-1.tar.bz2

• regex-0.12-MSYS-1.0.11-1.tar.bz2

• tar-1.19.90-MSYS-1.0.11-1-bin.tar.gz

• texinfo-4.11-MSYS-1.0.11-1.tar.bz2

• vim-7.1-MSYS-1.0.11-1-bin.tar.gz

• w32api-3.11.tar.gz

• zlib-1.2.3-MSYS-1.0.11-1.tar.bz2

This file list might contain file which are not absolutely mandatory for Liquid War 6, for instance the Fortran 77 compiler is absolutely useless, but installing it won’t harm either. Some packages might unzip things the right way, but some do it in a subfolder. You might need to run commands like:

     cp -r coreutils*/* .
     rm -rf coreutils*

Get rid of useless files:

     rm ._.DS_Store .DS_Store

It’s also mandatory to move everything that has been installed in /usr or /usr/local to / since MSYS has some builtin wizardry which maps /usr on /. You need to do this if you don’t unzip files from a MinGW shell, which is obviously the case when you first install it. Usefull command can be:

     mv usr/* .
     rmdir usr

Next, libintl is not correctly handled/detected by LW6, and can raise an error like

"gcc.exe: C:/msys/local/lib/.libs/libintl.dll.a: No such file or directory" so

one needs to copy some libraries in /usr/local/lib/.libs/:

     mkdir local/lib/.libs
     cp local/lib/libintl.* local/lib/.libs/

Another step is to edit /etc/profile and add lines like:

     export CFLAGS="-g -I/usr/local/include"
     export LDFLAGS="-L/usr/local/lib"
     export GUILE_LOAD_PATH="C:\\MSYS\\local\\share\\guile\\1.8\\"

Close and re-launch your msys shell (rxvt) so that these changes take effect. Check that those values are correctly set:

     env | grep FLAGS
     env | grep GUILE

Finally, your MSYS environment is (hopefully...) working.

Now you need to compile the following programs, from source. Files are mirrored on

ufoot.org (http://www.ufoot.org/download/liquidwar/v6/mingw32/src/) for your

convenience, however these might be outdated. Still, there are known to work. Proceed like if you were under a POSIX system. Some packages use the --disable-rpath swith, there are various reasons for which rpath is an issue (http://wiki.debian.org/RpathIssue). In the same manner, --disable-nls when linking against libintl or libiconv was painful.

  • pthreads-win32 (http: / / sourceware . org / pthreads-win32 / ), un-

tar  pthreads-w32-2-8-0-release.tar.gz  then  make clean GC; cp pthread.h

    sched.h /usr/local/include/; cp pthreadGC2.dll /usr/local/bin/; cp
    libpthreadGC2.a /usr/local/lib/
  • GNU MP (http://gmplib.org/), untar gmp-4.2.2.tar.gz then ./configure &&
    make && make install
  • Guile (http://www.gnu.org/software/guile/ ), untar guile-1.8.5.tar.gz.
    Edit libguile/guile.c and insert #undef SCM_IMPORT just before #include
    <libguile.h>. Edit ./libguile/threads.c and place struct timespec { long

tv_sec; long tv_nsec; }; just before #include "libguile/_scm.h".  Then
./configure --disable-nls --disable-rpath --disable-error-on-warning
--without-threads && make && make install. The GUILE_LOAD_PATH value must

be correctly set for guile-config to work. For unknown reasons, running guile can throw a stack overflow error. Don’t panic. See bug 2007506 on SourceForge.net

(https://sourceforge.net/tracker/?func=detail&atid=102435&aid=2007506&

group_id=2435) for an explanation on why the Guile binary shipped with MSYS is not suitable for Liquid War 6.

• expat (http://www.libexpat.org/), untar expat-2.0.1.tar.gz then ./configure
&& make && make install
• SQLite (http://www.sqlite.org/), untar sqlite-amalgamation-3.5.9.tar.gz
then ./configure && make && make install
• libpng (http: / / www . libpng . org / pub / png / libpng . html), untar
libpng-1.2.29.tar.gz then ./configure && make && make install
• libjpeg (http://www.ijg.org/), untar jpegsrc.v6b.tar.gz then ./configure &&
make && make install && make install-lib
• libcURL (http://curl.haxx.se/libcurl/ ), untar curl-7.18.1.tar.gz then
./configure --without-ssl && make && make install
• FreeType 2 (http://freetype.sourceforge.net/), untar freetype-2.3.5.tar.gz
then ./configure && make && make install
• libogg (http://www.xiph.org/), untar libogg-1.1.3.tar.gz then ./configure &&
make && make install
• libvorbis (http: / /www .xiph .org / ), untar libvorbis-1.2.0.tar.gz then
LDFLAGS="$LDFLAGS -logg" && ./configure && make && make install
• SDL (http://www.libsdl.org/), untar SDL-1.2.13.tar.gz then ./configure &&
make && make install
• SDL image (http: / / www . libsdl . org / projects / SDL_image / ), untar
SDL_image-1.2.6.tar.gz then ./configure && make && make install
• SDL mixer (http: / / www . libsdl . org / projects / SDL_mixer / ), untar
SDL_mixer-1.2.8.tar.gz then ./configure && make && make install
• SDL ttf (http: / / www . libsdl . org / projects / SDL_ttf / ), untar SDL_
ttf-2.0.9.tar.gz then ./configure && make && make install

For your convenience, a zip file containing a complete MSYS "Liquid War 6 ready" environment is available. It is simply the result of all the operations described above. Simply

unzip msys-for-liquidwar6-20080819.zip (http://www.ufoot.org/download/liquidwar/
v6/mingw32/msys-for-liquidwar6-20080819.zip) (about 240 megs) in C:\MSYS\. All

dependencies compiled in /local have been generated using the command:

 cd /usr/local/src
 ./msys-for-liquidwar6-build.sh > ./msys-for-liquidwar6-build.log 2>&1

Note that this script does’t do everything, you’ll still need to edit Guile source code and patch it manually.

It might even be possible to use this MSYS environment under Wine (http://

www.winehq.com/ ). Simply unzip it under $HOME/.wine/drive_c, and run wine

"$HOME/.wine/drive_c/windows/system32/cmd.exe" /c "c:\\msys\\msys.bat" and

with luck, you’ll get a working shell. Note that this might allow you to compile the game, but running it is another story. Consider this MSYS over Wine trick as a hack enabling the use of free software only when compiling for Microsoft proprietary platform. It is not a reasonnable way to run the game. If running under a UNIXish platform, or better, GNU, simply run native code. Use the Windows 32-bit port only if you are jailed on a Microsoft system.

Now, let’s come to the real meat, untar the Liquid War 6 source tarball, launch your MSYS shell, and:

     ./configure
     make
     make install
   Now the binary is in src/.libs/liquidwar6.exe (beware, src/liquidwar6.exe is only

a wrapper). This binary is an MSYS/MinGW binary, so it read paths “ la” Microsoft, that is, it has no knowledge of what /usr is, for instance. It requires paths starting by C:\.

3.8.5 Mac OS X port

This is still experimental. Basically, install MacPorts, and most dependencies with, except for SDL which you compile from source. The idea is to compile SDL using the native OS X bindings (and not some other GL files you could have in /opt/local installed by MacPorts), then compile the game and other SDL dependencies against this SDL.

The SDL mixer library might need to be told to compile itself without dynamic ogg support. By default it seems that it tries to load libvorbisfile.dylib at runtime, and it can fail. To disable this dynamic loading, use for instance :

/configure --prefix=/opt/extra --enable-music-ogg --disable-music-ogg-shared

Also, it might seem obvious for Mac OS X users, but there are some important issues related to compiling options and handling dynamic libraries at runtime.

• The command ldd does not exist, run otool -L instead.

  • The equivalent of LD_LIBRARY_PATH is DYLD_LIBRARY_PATH.

• The extension for shared binaries is .dylib and not .so.

• You might need to set the OBJCFLAGS environment variable along with CFLAGS because the Mac OS X port uses some Objective-C code.

It is very important to have the right SDL flags when linking the Liquid War 6 binaries. For instance it could be:

-I/opt/extra/include -I/opt/local/include -Wl,-framework -Wl,CoreFoundation -I/opt/local/include -D_THREAD_SAFE -Wl,-framework -Wl,Cocoa -Wl,-framework -Wl,OpenGL -Wl,-framework -Wl,Cocoa

The point is to have Cocoa and OpenGL support. Depending on the way you installed SDL, you might also need to include an SDL framework support, this is mostly if you installed SDL from .dmg binary images, and not from source with the command line. A typical output of sdl-config --libs is:

-L/opt/extra/lib -lSDLmain -lSDL -Wl,-framework,Cocoa

Another important issue is to include SDL.h, which in turn includes SDLmain.h, in all the .c source files defining the standard main function. This is done in liquidwar6 but should you try to link yourself on liquidwar6 libraries and/or hack code, you must do this or you’ll get errors when running the game. Such errors look like:

*** _NSAutoreleaseNoPool(): Object 0x420c90 of class NSCFNumber autoreleased with no pool in place - just leaking

The reason is that SDL replaces your main with its own version of it. One strong implication is that all the dynamic loading of SDL, which works on sandard GNU/Linux boxes, won’t work under Mac OS X, since SDL hard codes itself by patching main with #define C-preprocessor commands.

A .dmg file (disk image) containing a Liquid War 6.app folder (OS X application) is available for your convenience. It might work or not. In doubt, compile from source. The complicated part about this package (a “bundle” is OS X language) is that it needs to embed various libraries which are typically installed in /opt by MacPorts on a developper’s machine. So to build this package a heavy use of the utilility install_name_tool is required, normally all libraries needed ship with the binary, remaining depedencies concern frameworks which should be present on any working Mac OS X install. Still, this is only theory. Needs to be widely tested.

The layout of the bundle follows:

• ./Contents/Info.plist: metadata, bundle description file

• ./Contents/MacOS: contains the main binary liquidwar6 as well as all specific lowlevel libraries

• ./Contents/Resources/data: general game data, does not contain maps.

  • ./Contents/Resources/music: music for the game.

• ./Contents/Resources/map: system maps, you can put your own maps (or “extra” maps) here if you want all users to share them.

• ./Contents/Resources/script: Liquid War 6 specific scripts, the scheme scripted part of the program.

• ./Contents/Resources/guile: common shared Guile scripts, part of Guile distribution.

• ./Contents/Resources/doc: documentation in HTML and PDF formats.

   Additionnally, the Mac OS X port uses /Users/<username>/Library/Application

Support/Liquid War 6/ to store configuration file, logs, etc. It does not use $HOME/.liquidwar6 like the default UNIX port.

The Mac OS X port also has a special behavior, in order to load some libraries with dlopen (SDL image does this with libpng and libjpeg) it needs to set DYLD_FALLBACK_ LIBARY_PATH to a value that contains these libraries. This is typically in the bundle distributed on the disk image. On a developper’s computer this problem does not appear for those libs are often in places like:

  • /usr/local/lib
  • /usr/X11/lib
  • /opt/local/lib
   So the program sets DYLD_FALLBACK_LIBARY_PATH (but not DYLD_LIBRARY_PATH else it

breaks internal OS X stuff which relies, for instance, on libjpeg library that has the same name but different contents) but it needs to do it before it is even run, else the various dyld functions do not acknowledge the change. That is, calling the C function setenv(), even before dlopen(), has no effect. So the program calls exec() to re-run itself with the right environment variable. This is why, on Mac OS X, there are two lines (exactly the same ones) displaying the program description when it is started in a terminal.

3.8.6 GP2X port

This is not working yet, but there are good hopes that some day, Liquid War 6 can run on a GP2X console (http://www.gp2xwiz.co.kr/). This handled gaming device uses a Linux kernel (http://kernel.org) on an ARM processor, does support most GNU (http://www. gnu.org/) packages through cross-compilation, and has basic SDL (http://www.libsdl. org/) support.

To compile Liquid War 6 for GP2X, you first need to set up a working “toolchain”. It’s suggested you do this on a working GNU/Linux box. There are several solutions, the recommended one being Open2x (http: / /www .open2x .org / ). Read the online documentation on how to install Open2x (http://wiki.gp2x.org/wiki/

Installing_the_Open2x_toolchain).

Basically, the following should work:

mkdir /opt/open2x # check that you can write here
svn co https://open2x.svn.sourceforge.net/svnroot/open2x/trunk/toolchain-new open2x-toolchain
./open2x-gp2x-apps.sh
cd open2x-toolchain

Then, for your environment to be complete, you need to set up some environment variables. For instance:

export OPEN2X_SYSTEM_PREFIX=/opt/open2x/gcc-4.1.1-glibc-2.3.6/arm-open2x-linux
export GP2X_USER_PREFIX=$HOME/gp2x
export CC=${OPEN2X_SYSTEM_PREFIX}/bin/arm-open2x-linux-gcc
export CPP=${OPEN2X_SYSTEM_PREFIX}/bin/arm-open2x-linux-cpp
export CXX=${OPEN2X_SYSTEM_PREFIX}/bin/arm-open2x-linux-g++
export AS=${OPEN2X_SYSTEM_PREFIX}/bin/arm-open2x-linux-as
export CFLAGS=''-I${OPEN2X_PREFIX}/include -I${GP2X_USER_PREFIX}/include''
export CPPFLAGS=''-I${OPEN2X_PREFIX}/include -I${GP2X_USER_PREFIX}/include''
export CXXFLAGS=''-I${OPEN2X_PREFIX}/include -I${GP2X_USER_PREFIX}/include''
export LDFLAGS=''-L${OPEN2X_PREFIX}/lib -L${GP2X_USER_PREFIX}/lib''
export PATH=''${GP2X_USER_PREFIX}/bin:${OPEN2X_SYSTEM_PREFIX}/bin:$PATH''

In this setting, there’s a user $HOME/gp2x directory which will contain all the Liquid War 6 related libraries while the /opt/open2x remains untouched.

Then you need to install the requirements. All these packages need to be cross-compiled. To make things clear and non-ambiguous, even if you have CC set up in your environment variables, pass --build and --host arguments to the ./configure script of all these packages. A typical command is:

./configure --build=i686-pc-linux-gnu --host=arm-open2x-linux --prefix=${GP2X_USER_PREFIX}

Here’s the list of the requirements:

• SDL (http://www.libsdl.org/ ). 1.2.14 works zith the following settings:

    ./configure --prefix=$GP2X_USER_PREFIX --build=x86_64-pc-linux-gnu
    --host=arm-open2x-linux --disable-pulseaudio --disable-video-directfb

  • libtool (http://www.gnu.org/software/libtool/ ). Install a 1.5.x version,

and not 2.2.x. For some reasons, 2.2.x is unusable, it keeps on complaining that lt libltdl LTX preloaded symbols is not defined (http://lists.cistron.nl/

    pipermail/freeradius-users/2009-September/msg00149.html). Apparently the
    bug has already been reported (http://www.archivum.info/bug-libtool@gnu.org/

2009-01/00014/undefined_lt_ptr). For Liquid War 6 needs, 1.5.x is very fine, use 1.5.26 for instance, it works well.

• GNU MP (http://gmplib.org/) (version 4.3.1 reported to work)

• Guile (http://www.gnu.org/software/guile/) (version 1.8.7 reported to work). You need to pass the --witout-threads switch to the ./configure script else it will try (and fail!) to run a test program. Liquid War 6 does not use Guile threads, it does use threads but uses them “directly” outside the Guile layer.

• zlib (http://www.zlib.net/) (version 1.2.3 reported to work). Do not use --build and --host for this one, they are unsupported. Package compiles fine anyway.

• expat (http://www.libexpat.org/) (version 2.0.1 reported to work).

  • libpng (http://www.libpng.org/pub/png/libpng.html) (version 1.2.3 reported to

work).

• libjpeg (http://www.ijg.org/) (version 7 reported to work).

• SQLite 3 (http://www.sqlite.org) (version 3.6.18 reported to work).

• libcURL (http://curl.haxx.se/libcurl/) (version 7.19.6 reported to work).

Next, one needs to install a special version of SDL, which targets the GP2X specifically. This is not a generic SDL, and it does have limitations, which are related to the GP2X peculiar hardware. There are installation instructions (http://wiki.gp2x.org/wiki/ How_to_install_the_SDL_libraries) about how to do this. The following should work:

cvs -d :pserver:anonymous@cvs.sourceforge.net:/cvsroot/open2x login # blank password
cvs -d :pserver:anonymous@cvs.sourceforge.net:/cvsroot/open2x co libs-gp2x

3.9 Coding guidelines

3.9.1 Project goals reminder

One of the purposes of Liquid War 6 is to make a cleaner implementation of Liquid War than the previous one, namely Liquid War 5 (http://www.ufoot.org/liquidwar/v5). While the latter has achieved the practical goal of providing a playable implementation of the game, it failed at providing an evolutive platform. Network capabilities where finally added to Liquid War 5 (http://www.ufoot.org/liquidwar/v5), but anyone who played on Internet with someone a few hundreds of milliseconds away would agree that it’s far from being perfect. The main reason for this is that it is really had to hack on Liquid War 5 (http://www.ufoot.org/liquidwar/v5), especially when you are not the core developper. The core developper himself, even knowing all the various hacks in the game, is very quickly lost when trying to implement major changes.

To put it short, Liquid War 5 (http://www.ufoot.org/liquidwar/v5) is a global variable hell, a pile of hacks on top of a quick and dirty implementation. Still, it works.

With Liquid War 6, the idea is to take the time to make something stable, something nice which will enable developpers to implement the cool features, and have fun along the way. Of course, this is only a dream, and in the (hopefully "very") long run, Liquid War 6 will also end up as a big unmaintainable mess, like any real-life program, until then, it should remain hackable.

3.9.2 Common sense

Here are a few guidelines which I think are common sense advice, but they are still worth mentionning:

• try and respect the GNU coding standards (http://www.gnu.org/prep/standards/ );

• absolutely no strcpy or sprintf anywhere in the code. Nowhere. Use their equivalent strncpy and snprintf systematically, as they are part of the glibc and are an order of magnitude safer. Moreover, Liquid War 6 provides wrappers, such as lw6sys_new_ sprintf which handles all the nasty dirty memory allocation stuff for you;

• keep global variables for when there is something truely global, and even in that case try to fit them in clearly identified structures.

3.9.3 Unitary tests

Each of the internal libraries in Liquid War has a “test” program associated with it. For instance liquidwar6sys-test is associated to libliquidwar6sys, and its purpose is to test the features of this library.

While it is fairly easy to test out unitary functions which require no peculiar context, testing high-level functions which requires files, graphical and possibly network contexts to exist is obviously harder to achieve. There’s no easy way to draw the line, but the idea is to put in these test executables as many features as possible, to be sure that what is tested in them is rock solid, bullet proof, and that one can safely rely on it and trust that code when running it in a more complex environnement.

These test executables are also very good places to see a library API in action, find code fragments, and make experiments.

3.9.4 Memory allocation

Liquid War 6 provides macros to allocate and free memory. One should use them systematically, except when trying to free something allocated by another library, and in very special cases, mostly concerning low-low level operations which are seldom hacked on.

   Usage of macros LW6SYS_MALLOC, LW6SYS_CALLOC and LW6SYS_FREE is straightforward,

read any random chunk of code, for instance ./src/lib/sys/sys-test.c to see them in action. They are defined in sys/sys.h.

Once used, these macros will track every single call to malloc and free, and if there’s a difference, it will report it. It will also help you by showing what’s in the non-freed memory area, at which line of code it has been allocated, and when. This is very usefull to track down memory leaks. Of course a debugger could tell you some of these informations, but experience shows than when you encounter a memory bug, it’s very often impossible to reproduce it. So you one wastes time trying to reproduce the bug, whereas with this tool you have the information reported just when the problem happens.

3.9.5 Private and public interfaces

Each library exports a public interface and hides its internal. Since Liquid War 6 uses standard C and no C++, there’s no real standard way to handle public/private features. The convention used in Liquid War 6 is to show internal structures as opaque pointers (void *) whenever some function needs to operate on a structure which has possibly private fields. This way the caller function has no way to access the internals, and we are sure that no reference to any internal implementation specific feature will appear.

Here’s a code excerpt from src/gfx/setup.c:

     void _lw6gfx_quit(_LW6GFX_CONTEXT *context) {
       /*
        * Implementation here.
        */
     [...]
     }
     void lw6gfx_quit(void *context) {
       _lw6gfx_quit((_LW6GFX_CONTEXT *) context);
     }

The function _lw6gfx_quit (note the “ ”) is internal, declared in internal.h whereas the function lw6gfx_quit is public, and is therefore exported in gfx.h.

This way, functions in the program using lw6gfx_quit do not know what is in the _LW6GFX_CONTEXT structure, and they need not know it.

This does not mean it is not possible to have public structures, only these structures must reflect some truely public, accessible and safe to access structures.

3.9.6 Commit policy

Basic rules :

• commits should be as small as possible, yet leave the code in a consistent state. The general idea is that big commits tend to make error tracking more complicated;

• a commit should not leave the code in a state in which it does not compile and/or is consistent from a user point of view. The general idea is to go step by step and break as little things as possible.

• tests should be written along the way. The project is not developped with a “test driven” method, still, it’s a good practice to write the tests functions as soon as possible.

To check that a commit does not break everything, a good practice is to run a make check before committing / submitting anything.

Then, once it’s within the main GIT repository, check the Jenkins builds (http://www. ufoot.org/jenkins/job/liquidwar6/) to see if the program still builds correctly.

3.9.7 Audit the code

Liquid War 6 is regularly audited with automated tools. You might need to pass --enable-gcov to --configure if you want to use thes tools yourself. More precisely:

• CUnit (http://cunit.sourceforge.net) is used for regression tests. It’s used to provide hopefully standardized output when testing, and provide test statistics more easily. It’s a rule of thumb to try and write tests when new features and/or bug-fixes pour in.

• lcov (http://ltp.sourceforge.net/coverage/lcov.php) is run, ideally with each

release - but it’s not garanteed, check the output date and time - and the output is

available online on http://www.ufoot.org/liquidwar/v6/doc/coverage/.

• GNU global (http://www.gnu.org/software/global/) is run too, ideally with each release - again, check the output date and time - and the output is available online on

http://www.ufoot.org/liquidwar/v6/doc/global/.
• pmccabe (http://parisc-linux.org/~bame/pmccabe/) reports cyclomatic com-

plexity. It shows where the code is too complex and should probably be rewritten. Output is post-processed using pmccabe2html (http://www.gnu.org/software/

gnulib/MODULES.html#module=pmccabe2html) from gnulib (http://www.gnu.org/
software/gnulib/). The output is available online on http://www.ufoot.org/
liquidwar/v6/doc/cyclo/.

• Valgrind (http://valgrind.org/) is run as well, it should report absolutely no leak on all the core sub-libraries, eg running liquidwar6ker-test or liquidwar6ker-pil. Bits of code which depend on other libraries are a different story, for some projects on which Liquid War 6 depends might, for some reason, raise warnings. But as far as Liquid War 6 is concerned, the goal is simple: zero leak.

• Liquid War 6 is referenced on Open HUB (http://www.openhub.net/). Visit http:// www.openhub.net/p/liquidwar6 to get time-based statistics and facts about the source code.

Those tools certainly don’t garantee the code is perfect, but they do help improving the quality of the program. If you hack, it’s recommended you give them a try.

3.10 Using the console

The console can be activated by passing --display-console when starting the game or by using the system options menu.

When the console is activated, a lw6> prompt should appear in the terminal which launched the program. If you started Liquid War 6 by clicking on an icon, console probably won’t work at all since stdout and stdin won’t be attached to anything.

The console allows you to type arbitray Scheme/Guile code. Try, for instance:

 (+ 1 2)
 (display "foo\n")

You can really break things with this console, it gives you a direct access to all the program internals. At least, all the values which are accessible through the script interface, that is, many of them.

You can call any internal C function which has been exported to Guile, here are some examples:

 (c-lw6sys-timestamp)
 (c-lw6bot-get-backends)
 (c-lw6sys-sleep 2.0)

     (lw6-config-get-number "zoom")
     (lw6-config-set-number! "zoom" 0.9)
     (lw6-config-get-number "zoom")

While syntax (and possibly other) errors will be trapped by the interpreter, note that if you break things inside the game by, say, changing some global value, or in a general manner cause an error elsewhere in the code, the game will really raise a fatal error and stop. That’s how you can “break things”.

Still, this console is a very powerfull tool, very usefull for debugging but also for tweaking the game without restarting it and/or navigating through the menu interface.

3.11 Advanced tweaking

3.11.1 Hacking ressources

Liquid War 6 tries to have as few hardcoded data as possible. So many constants, and pretty much all the image files, are accessible in the data directory. You can know where it is by launching liquidwar6 --show-data-dir. If you look in this directory you’ll find different files, among them XML files.

Let’s take an example. Try and find the file gfx/gl/hud/floating/gl-floating-const.xml. Edit the line with the clock-y1 entry. Change the number after "value". Re-run the program. Play a game. What happens? Logically you should see that “something” is not displayed at the same place than before.

You could also modify the textures (JPEG and PNG files). In a general manner it’s more cautious to keep them the same size but it depends, sometimes other sizes will work as well.

Many of these parameters are really too technical and obscure to have their place in the main config file (which is already rather big). Use at your own risks, you can really break things touching this, but you can also find out lots of things can be tuned.

3.11.2 Optimize for speed

Todo...

3.12 Writing modules

Todo...

3.13 Use as a library

Todo...

3.14 Network protocol

This section describes how Liquid War 6 handles network messages. Note that for now this is purely theorical, more of a draft, a plan, it might change before being implemented.

3.14.1 No server, no client, only nodes

Liquid War 6 does not really have the notion of server or client, any instance of the program can act as both server and client, therefore we use the term node.

A node listens on a given port, in both TCP and UDP, and can connect to other nodes on the same port. The main identifier of a node is its public url, which is typically of the form http://<ip-address>:<port>/. This url is very important for it is (or at least should be) a unique identifier of the node over the network.

Liquid War has 3 ways to communicate:

• raw TCP, this is the LW6 protocol, the easiest to implement and debug, probably the most reliable one, but not always the fastest. This involves the two modules mod-tcp

    and mod-tcpd.

• HTTP over TCP, this is a hack which allows the game to communicate through HTTP proxies for instance. Additionnally, the fact any node is a web server enables peering with a simple web browser. Web server facility requires mod-httpd and client part requires mod-http which might or not be available, depending on how the game was compiled.

• raw UDP, this is another version of the LW6 protocol, this is in theory the fastest way to communicate, it requires mod-udp and mod-udpd to work. Using UDP only was not an option when conceiving Liquid War since it’s interesting to have other solutions if, for instance, a firewall does not allow you to use UDP the way you want.

On each of these channels, messages can be exchanged in two modes, an “out of band” mode (AKA “oob”), and a regular message oriented, here we speak of “connection”.

3.14.2 Out of band messages

There are only 3 out of band messages:

• PING: requests for a simple PONG http://server:port/ answer, this is just to check if a server is a valid server, and if the URL we used to connect on it is the correct one.

• INFO: requests for a list of key/attributes pairs, which describe the node, telling for instance its version, its uptime, and so on.

• LIST: requests for a list of other nodess this node is aware of.

You can test out of band messages by simply connecting on your server with a command like:

telnet localhost 8056

At the telnet prompt, simply type:

INFO

and press return, and you should have a description of your node.

The complete syntax of OOB messages is:

<COMMAND> [password] [url]

The password and url parameters are optionnal. password can contain either the plain text password or a checksum calculated from the password which is, for security reasons, seeded with the public url of the node we’re connecting to, so that this password can’t be re-used on another node. url is simply a clue we give to the other node to help find us, the other node will automatically try to detect our IP address and use standard LW6 port 8056, but if for some reason we use a different setting, this is a good way to pass the hint.

Here are examples of valid messages:

LIST
PING http://myhost.mydomain:1234/
INFO secret http://myhost.mydomain:1234/
LIST 12ab34cd

If there’s only one argument, the parser will first try and interpret it as a URL. If it’s not parseable that way, it will consider it’s a password. The password, in turn, may be specified as clear text or as a 32-bit checksum.

As far as OOB is concerned, TCP and UDP work almost the same, HTTP is a bit different, the OOB commands are accessed through the following URLs:

  • /ping.txt
  • /info.txt
  • /list.txt

OOB messages are usually sent many times in redundant mode on the network, as there’s no tracking of them, sending them through multiple channels ensures they make their way to the right place.

The parser for these messages is located in src/lib/msg/msg-oob.c.

3.14.3 Regular messages overview

All messages that are non-OOB share a common syntax. This is called the “envelope” of messages.

The general syntax is:

LW6 <VERSION> <PASSWORD_CHECKSUM> <PHYSICAL_TICKET_SIG> <LOGICAL_TICKET_SIG> <PHYSICAL_FROM_ID> <PHYSICAL_TO_ID> <LOGICAL_FROM_ID> <LOGICAL_TO_ID> <MSG>

Here’s an example:

LW6 0.1.3485 - 2d1dbed7 - 3003300330033003 2002200220022002 - - DATA 8 0 1 1000000035005 3003300330033003 SET 3001 11 1 1 0
   In this example, the messages carried is DATA 8 0 1 1000000035005 3003300330033003

SET 3001 11 1 1 0, the rest is part of the envelope protocol.

Here’s what those fields mean:

• LW6: should always be LW6, this is a marker to make sure we’re speaking the right protocol.

• <VERSION>: the version of the program sending the message, the receiver of the message should check this version is compatible.

• <PASSWORD_CHECKSUM>: the password checksum, while a clear password should still be correctly interpreted, as for OOB messages, there’s no reason to send the cleartext password, so the checksum is just fine. Note that the checksum is short, and vulnerable to brute-force attacks. If you want strong protection, the general advice is to tunnel your connections through SSL or TLS, use a VPN, LW6 won’t implement “fortress mode”, third party tools should do this much better. If undefined, should be replaced by the dash character -.

• <PHYSICAL_TICKET_SIG>: a signature done by the sender, which is unique for the combination message+from+to. This means two different messages will generate two differents signatures, but different senders and/or receivers will also change this, so it’s not possible, unless one has the “ticket” to fake a message. This is clearly not bullet-proof, and more specifically, brute-force attacks and/or network listening could break the protocol and/or reveal the ticket, still, this is a good way to make sure that if something is inconsistent, someone is trying to cheat. As every node maintain its own game state, a cheater can “only” be a nuisance by sending wrong key presses, but in the long run it will be defeated by the fact that an attacker should intercept and modify all messages on all channels (tcp, udp, http ...) and make sure the official, real informations, never makes its way to the right node. This is quite hard to achieve, very likely, an inconsistency will be detected, nodes concerned should be disconnected, period. When sending the first messages, ticket might not be exchanged yet, so there’s no way to calculate this, during this period, ffffffff is sent, and checksum errors are ignored.

• <LOGICAL_TICKET_SIG>: another signature, but this one concerns the physical sender/receiver. If the physical sender is the logical sender, and the physical receiver is the logical receiver, that is, if physical and logical nodes are the same pair of nodes, then it need not be defined and can be replaced by the dash character -. In fact, in that case, the physical and logical signatures are obviously the same. However (not implemented yet) the protocol is designed so that nodes can act as messages forwarders, in that case they have no knowledge of the secret ticket to use, so this ticket is here to ensure message consistency for the final, real (logical) receiver of the message.

• <PHYSICAL_FROM_ID>: the id of the physical sender, the node that created the message.

• <PHYSICAL_TO_ID>: the id of the physical receiver, the node that should receive the message.

• <LOGICAL_FROM_ID>: the id of the logical sender, if it’s the same than the physical sender, can be replaced by the dash character -.

• <LOGICAL_TO_ID>: the id of the logical receiver, if it’s the same than the physical receiver, can be replaced by the dash character -.

• <MSG>: the message itself, it might in turn be separated by spaces, or whatever the message delimiter is. It should not be too long, as it must be sendable on the network by UDP, so it must fit within the MTU (about 1.4 kb) with all the protocol (envelope) stuff before it. In practice, it’s cut into 1.2 kb parts by libdat, the constant _LW6DAT_ ATOM_MAX_SIZE is used to split big messages in smaller parts.

It’s implemented in src/lib/msg/msg-envelope.c.

3.14.4 Regular control messages

To establish a connection, maintain it, and do the OOB job, a set of control message is used. Those messages carry a bunch of informations about who is sending them, in fact, they just contain the informations that would otherwise be handled by out-of-band messages, but it’s convenient to have the information first-hand rather than relying on the other protocol.

The syntax is:

LW6 <VERSION> <PASSWORD_CHECKSUM> <PHYSICAL_TICKET_SIG> <LOGICAL_TICKET_SIG> <PHYSICAL_FROM_ID> <PHYSICAL_TO_ID> <LOGICAL_FROM_ID> <LOGICAL_TO_ID> <COMMAND> <PROGRAM> <VERSION> <CODENAME> <STAMP> <ID> <URL> <TITLE> <DESCRIPTION> <HAS_PASSWORD> <BENCH> <OPEN_RELAY> <UPTIME> <COMMUNITY_ID> <ROUND> <LEVEL> <REQUIRED_BENCH> <NB_COLORS> <MAX_NB_COLORS> <NB_CURSORS> <MAX_NB_CURSORS> <NB_NODES> <MAX_NB_NODES> <PEER_LIST> <COMMAND_ARGS>

Example:

LW6 0.1.3485 - ffffffff - 1001100110011001 2002200220022002 - - HELLO liquidwar6 0.1.3485 "Davy Crockett" 3485 1001100110011001 http://localhost:8057/ cGF0 RHVtbXkgdGVzdCBub2RlIEE= 0 10 0 5 372057f45b3d2ba5 10005 "Default map" 5 4 10 4 26 1 12 ""

The fields, starting from LW6 up to (and including) <LOGICAL_TO_ID> are part of the envelope, described previously.

The message fields are:

• <COMMAMD>: described below, the main command,

• <PROGRAM>: should be liquidwar6

• <VERSION>: the version of the program. Yes, this is also in the envelope, but one could think of instances relaying informations for other peers, in that case this could prove useful.

• <CODENAME>: the code name of the program.

• <STAMP>: the stamp, it’s normally contained within the version, but this avoids parsing issues.

• <ID>: the node id (could be inferred from envelope, but repeated here).

• <URL>: the node url (could be inferred from envelope, but repeated here).

• <TITLE>: the readable title of the node, base64 encrypted.

• <DESCRIPTION>: the readable description of the node, base64 encrypted.

• <HAS_PASSWORD>: wether it’s protected by a password or not, 0 if not, 1 if protected.

• <BENCH>: the bench of the node, giving its CPU strength in an arbitrary unit.

• <OPEN_RELAY>: wether the node act as an open relay, 0 if not, 1 if in relaying.

• <UPTIME>: node uptime, in seconds.

• <COMMUNITY_ID>: the community id, a unique id shared by all nodes connected to a game session.

• <ROUND>: the current round id.

• <REQUIRED_BENCH>: the minimum bench required to connect to this node (used to avoid slow nodes connecting to way-too-fast games).

• <NB_COLORS>: number of colors playing.

• <MAX_NB_COLORS>: maximum number of colors allowed on this node.

• <NB_CURSORS>: number of cursors playing.

• <MAX_NB_CURSORS>: maximum number of cursors allowed on this node.

• <NB_NODES>: number of nodes connected.

• <MAX_NB_NODES>: maximum number of nodes allowed on this node.

• <PEER_LIST>: list of peers connected to this node.

• <COMMAND_ARGS>: command-specific arguments Here are the different possibilities for the <COMMAND> field.

• HELLO: is used when connecting, this should be the first message sent. In itself, the message means pretty much noting, it just says “I’m here” which could be infered from any other message. No command args.

• TICKET <ticket>: is used to inform the caller of the ticket we use. The ticket sent from A to B is ised by B to sign messages for A. A node typically sents a different ticket to all its peers so when sending the same message to A and C, B will typically use two different tickets, thus generating two different signatures, and if sending the exact same string to C, A generate yet another signature as it will have sent another ticket. There’s one optional argument, which is the ticket itself, a 64-bit hexa integer.

• FOO <key> <serial>: is sent on a regular basis, it’s really the replacement of the OOB PING message, it will update the peer status and maintain consistent informations. It has two arguments, the first one is key, a 32-bit hexa integer, which will, upon BAR message reception, used to figure out “OK, this is the BAR message associated to this FOO message I sent before”. The second one, serial, is used to inform the peer that, possibly, there are new messages to fetch from us. The peer, in turn, might fire MISS messages, but without this feature, peers could “fall asleep” and forget to pump messages, especially on non-reliable connections.

• BAR <key> <serial>: is the response to FOO which is received on a regular basis, it’s really the replacement of the OOB PONG message, it will update the peer status and maintain consistent informations. It has two arguments, the first one is key, a 32bit hexa integer, which will, upon reception, ne matched against a corresponding FOO messaged, used to figure out “OK, this is the BAR message associated to this FOO message I sent before”. The second one, serial, is used to inform the peer that, possibly, there are new messages to fetch from us. The peer, in turn, might fire MISS messages, but without this feature, peers could “fall asleep” and forget to pump messages, especially on non-reliable connections.

• JOIN <seq> <serial>: Used to join a game. In fact, having said HELLO and exchanged FOO and BAR messages does not mean one has joined the game for real. The reason for this is that those messages help establishing a stable communication channel, then one needs to come in with the right seq and serial. There are basically two cases. First case, seq can be zero, in that case it means we’re trying to connect on an existing server, which will in turn send a JOIN message with a non-zero value, giving the current seq. Second case, seq is non-zero, in that case it means we’re answering a connection request. In both cases, serial is a serial number other peers should use as a minimum limit, and never ask for messages with a serial lower than that.

• GOODBYE: symetric of HELLO, should be called on disconnection, however, the peers should handle the case when no GOODBYE message is sent, this is just about being polite. No command args.

It’s implemented in src/lib/msg/msg-cmd.c.

3.14.5 Regular MISS messages

Todo...

3.14.6 Regular META messages

Todo...

3.14.7 Regular DATA messages

Todo...

3.14.8 Other raw technical stuff (WIP)

TCP messages:

LW6 [<passwd>] <version> <client-id>
<from-id> <to-id> <serial> <i> <n> <sig> MSG1
<from-id> <to-id> <serial> <i> <n> <sig> MSG2

TCP oobs:

<return> # only works anonymous, same as INFO
INFO [<passwd>] [<public-url>]
LIST [<passwd>] [<public-url>]
PING [<passwd>] [<public-url>]

UDP messages:

LW6 [<passwd>] <version> <client-id> <from-id> <to-id> <serial> <i> <n> <sig> MSG1
LW6 [<passwd>] <version> <client-id> <from-id> <to-id> <serial> <i> <n> <sig> MSG2

UDP oobs:

INFO [<passwd>] [<public-url>]
LIST [<passwd>] [<public-url>]
PING [<passwd>] [<public-url>]

HTTP messages:

client id & password passed in HTTP headers

/lw6/version/<from-id>/<to-id>/<serial>/<i>/<n>/sig/MSG1
/lw6/version/<from-id>/<to-id>/<serial>/<i>/<n>/sig/MSG2

HTTP public URLs:

/ -> index.html
/index.html
/favicon.ico
/screenshot.jpeg
/robots.txt
/gpl.txt
/info.txt
/list.txt
/ping.txt

MSG syntax:

<round> <server-id> <command> <arg1> ... <argN>

COMMAND examples:

2 1234abcd1234abcd REGISTER
3 1234abcd1234abcd ADD 5678 YELLOW
4 1234abcd1234abcd SET 5678 20 5
10 1234abcd1234abcd NOP
400 1234abcd1234abcd REMOVE 5678
1000 1234abcd1234abcd UNREGISTER

Sig is a checksum on string:

<from-id> <to-id> <serial> <i> <n> MSG

3.15 Technical HOWTOs

3.15.1 Release check-list

Summary off all operations required for a release:

• check the value of LW6MAP_RULES_DEFAULT_EXP and default for --skip-network, which

might have been changed will developping.

• in ./src, run ./indent.sh.
• in ./doc, run ./gdoc-update.sh, ./doxygen-update.sh and ./perf-update.sh.
• edit NEWS file, in both liquidwar6 and liquidwar6-extra-maps. Check ChangeLog is

OK.

• update the version history in doc/liquidwar6.texi.

• update debian/changelog files in both main and extra maps packages.

• run make distcheck... (at least!)

• build RPM package (make -C pkg rpm), check yum install produces a working installation, also check the rpm -e command to verify uninstalling is OK too.

• build .exe main binary on Microsoft Windows then go back to GNU/Linux and build .exe installer (make -C pkg installer), go back to Microsoft Windows and test the installer (there are often problems at this stage because of missing libraries or other files...).

• build .dmg Mac OS X disk image, check it works even when /opt and /usr/local or (re)moved, this is important, else execution might rely on binaries which are only on the development machine and do not ship with game.

• upload files (doc on http://savannah.gnu.org/maintenance/DownloadArea), with
a command like rsync --rsh=ssh --recursive --verbose --progress ./X.Y.Z/
login@dl.sv.nongnu.org:/releases/liquidwar6/X.Y.Z/. Each file must have its

.sig corresponding file.

• update online docs with gendocs.sh, carefull, liquidwar6.html is suppressed by this, need to re-create it from previous version and/or index.html which is the same.

• update index.html, liquidwar6.html and liquidwar6.fr.html so that they reflect

the latest release. Check download links are OK.

• post news on https://savannah.gnu.org/news/?group=liquidwar6, http://
lists.gnu.org/mailman/listinfo/help-liquidwar6 and http://lists.gnu.org/
mailman/listinfo/info-liquidwar6.

• after everything did work, upload files on ftp://ftp.gnu.org/, for instance use

gnupload --to ftp.gnu.org:liquidwar6 liquidwar6-X.Y.Z.tar.gz.

3.15.2 Add a new option

This describes how to add a new option to the game.

• edit src/lib/def/def-list.txt
• in src/lib/def/def-update.py run ./def-update.py. This will automatically fill
src/lib/def/def.h and script/def.scm. In the code, you should always use LW6DEF_

<OPTION> in C and lw6def-<option> in scheme to refer to the option. This does help avoiding typesetting errors.

• add the entry to src/lib/hlp/hlp-list.c, choose a category for it • add the entry to src/lib/hlp/hlp-reference.c, give it a type, documentation string

and default values if needed

• to sort src/lib/def/def-list.txt a common practice is to fill it with the output of liquidwar6 --list once the program has been compiled and is aware of the new option. Unless this is done, program won’t accept the option. Some options need more work, for instance:

• Any option which is a command-line argument needs to be added to

  src/lib/lw6-options.c.

• Any option which is related to the build system (to enable or disable some feature) must

  be referenced in src/lib/sys-build.c and also src/lw6-funcssys.c to be callable

from scripts.

• Any option which changes the map rules (any new rule...) impacts the game checksums so these ones need to be updated in src/lib/map/map-test.c,

  src/lib/ker/ker-test.c, src/lib/pil/pil-test.c and src/lib/pil/pil-suite.c.

3.15.3 Add a new internal library

This describes how to add a new libxyz internal library:

• create a new src/lib/xyz directory. The convention is to use 3 letters names and prefix every global identifier with lw6xyz.

• copy Makefile.am, xyz.h, xyz-test.c and xyz-testmain.c from an existing internal

library (libnod is a good source, it does not have complex dependencies).

• edit Makefile.am to fill requirements, make necessary adjustments to other files (many string replaces to make, both lowercase and uppercase).

• add the entry in the SUBDIRS and LW6_LIBS sections of src/lib/Makefile.am
• add the entry in the AC_CONFIG_FILES of ./configure.ac.
• run automake and autoconf.

• edit src/lib/lw6-options.c and add a call to lw6xyz_test() for both “check” and “test” cases.

• edit src/lib/lw6-test.c and add a reference to the lw6xzy_test() function.

• in every internal abc library that depends on xyz, edit the lw6abc_test function so that it contains a reference to lw6xzy_test.

• in every internal abc library that depends on xyz/xyz.h, edit the abc/abc.h header

  so that it includes xzy/xyz.h. Also edit src/lib/liquidwar6.h.in.

• in every internal abc library that depends on libxyz, add a reference to

  ../xyz/libxyz.la inf the _LDADD section.

• create a new src/lib/lw6-funcsxyz.c file which declares Guile bindings for this lib, if needed

• edit doc/gdoc-update.sh and add the entry for xyz.

• edit doc/Makefile.am and add xyz-gdoc.texi in gdoc_TEXINFOS.
• edit doc/doxygen-update.sh and add the entry for xyz.
• edit doc/Makefile.am and add xyz-doxygen.texi in doxygen_TEXINFOS.

• edit doc/perf-update.sh and add the entry for xyz.

• in ./doc, run ./gdoc-update.sh, ./doxygen-update.sh and ./perf-update.sh.
• edit doc/doxygen/Makefile.am and add the dependency on src/lib/xyz/xyz.h.

• edit doc/liquidwar6.texi to and a new node/section for this internal library.

• edit doc/deps.dot to update dependencies.

• run ./configure && make, fix code if needed.

3.15.4 Add a new module

This describes how to add a new mod-ab module, for instance a new bot, but gfx, snd, cli or srv backends should work pretty much the same:

• add a new entry in src/lib/bot/Makefile.am

• create the subdir src/lib/bot/mod-ab, with its Makefile.am (inspired from other existing modules)

• add the entry in configure.ac so that src/lib/bot/mod-ab/Makefile is generated.
• edit doc/gdoc-update.sh and add an entry for mod-ab
• edit doc/Makefile.am and add mod-ab-gdoc.texi in gdoc_TEXINFOS.
• edit doc/doxygen-update.sh and add an entry for mod-ab
• edit doc/Makefile.am and add mod-ab-doxygen.texi in doxygen_TEXINFOS.
• type touch doc/mod-ab-gdoc.texi doc/mod-ab-doxygen.texi else dependencies

checking will fail.

• in ./doc, run ./gdoc-update.sh and ./doxygen-update.sh.

• edit doc/liquidwar6.texi to add a new node/section for this module.

• edit doc/deps.dot to update dependencies.

• edit src/lib/bot/bot-test.c, change the value of TEST_NB_BACKENDS and modify

the code so that the new ab module is tested too.

• edit src/lib/bot/bot-register.c, the code must updated pretty much in every place with the conditionnal LW6_ALLINONE, you need to add the new module.

• run automake, autoconf, ./configure and make.

3.16 Using GNU Arch

3.16.1 About GNU Arch

Since March, 4th 2010, Liquid War 6 uses GIT (http://git-scm.com/) to handle source code, track changes, branches, and the rest. It replaces the GNU Arch (http://www. gnu.org/software/gnu-arch/) repository. This old repository contains all sources up to version 0.0.7beta, following versions, including 0.0.8beta, must be retrieved from GIT.

So the following informations only concern those who are interested in previous versions of the game. Anybody else - probably you - should use GIT instead.

See Section 3.17 [Using GIT], page 65. Still, this quick Arch survival guide is kept in the documentation.

 Read the GNU Arch tutorial (http://www.gnu.org/software/gnu-arch/tutorial/

arch.html) to learn how Arch works. Note that there are many other source control managers available, some of which provide functionnalities similar to GNU Arch / tla. GNU Arch has been chosen for Liquid War 6 because:

• it is Free Software,

• it is not limited to per-file commits like CVS, and supports atomic commits involving several files,

• it is distributed,

• it enables developpers to sign each of their contributions,

• it was already available back in 2005.

3.16.2 Getting the latest version from the repository

The repository for Liquid War 6 is accessible on http://arch.savannah.gnu.org/ archives/liquidwar6. This is a read-only access, but with the distributed nature of GNU Arch, it still allows you to keep track of your own changes, and submit patches. Accessing it in read/write mode with sftp requires a Savannah account and special rights on the Liquid War 6 project.

Here are typicall commands one can use to get Liquid War 6 source from the GNU Arch repository:

     tla register-archive http://arch.savannah.gnu.org/archives/liquidwar6
     tla get -A liquidwar6@sv.gnu.org liquidwar6--beta

All the patches in the archive are signed with GnuPG (http://www.gnupg.org/), so you can check their authenticity with my public key (http://www.ufoot.org/ufoot.pub).

   You might need to edit your $HOME/.arch-params/signing/=default.check file and

put the following text in it:

     tla-gpg-check gpg_command="gpg --verify-files -"

3.16.3 Setting up your own arch repository

This section is for those who want to hack the game and set up their own repositories. This will enable you to keep track of your patches, package them, and help the core maintainer merging them in the main repository.

You can introduce yourself and create a repository by issuing commands like:

You can introduce yourself and create a repository by issuing commands like:

     tla my-id me@home.net
     tla register-archive me@home.net--2008 /home/me/tla-archives

Then, you can get create your own repository, with a command like:

     tla tag -S liquidwar6@sv.gnu.org/liquidwar6--beta--0.1 me@home.net--2008/liquidwar6--beta--0.4

The idea is that you create, locally, a depot which has a name that matches the name on savannah (http://arch.savannah.gnu.org) (this is for convenience, you could technically give it any name...) and indicate that they represent, today, the same thing.

You can get a working copy of your depot with the command:

     tla get me@home.net--2008/liquidwar6--beta--0.4

This will create a complete source tree, which you are free to modify, this is where you should hack.

3.16.4 Synchronizing your repository with upstream releases

To synchronize yourself with upstream developments, go into your copy (the directory created by tla get) and type:

     tla star-merge liquidwar6@sv.gnu.org/liquidwar6--beta--0.1

This will apply locally all the changes that happened since the last synchronization. Of course this is one way to work, you can decide to cherry pick patches and such stuff, but for most dayly uses, a good’ol star-merge is fine.

Not that star-merge will only apply patches on your working copy, not on your repository. The only way to actually commit the modifications on the repository is to use the commit command.

3.16.5 Submitting patches

When using Arch, you can of course still send patches created with diff, or even send updated files directly, the way you would without revision control.

But it can be more convenient to either

• Send the patches stored in the depot (/home/me/tla-archives in our example).

• Make patches using tla mkpatch.

Here’s an example of an mkpatch command, and which will compute the differences between a previous liquidwar6--beta--0.4--patch-2 snapshot and a not yet commited latest version:

     tla mkpatch {arch}/++pristine-trees/unlocked/liquidwar6/liquidwar6--beta/liquidwar6--beta--0.4/me@home.net--2006/liquidwar6--beta--0.4--patch-2 . my-patch

This will create a my-patch directory, which can be gzipped and sent by mail.

3.16.6 Recover from broken lock

Sometimes, when signing a patch, you might enter the wrong passphrase several times, or you might press CTRL+D inadvertantly. In that case, tla will be in a half-broken state, telling you it can’t acquire revision lock... A quick workarround for this is to go to the depot, find the latest patch, and in this repository, create the following folders:

     ++revision-lock/+contents

Both are directories, note the two ++ and the single +. the +contents directory can be empty. Once you’ve done this, try to re-commit.

3.17 Using GIT

3.17.1 About GIT

There’s no CVS or Subversion (AKA “SVN”) source depot for Liquid War 6. Instead, a GIT (http://git-scm.com/) depot is used. GIT has many advantages over other source control managers (SCM), among them, it’s distributed, like GNU Arch (http://www.gnu.

org/software/gnu-arch).

You can find interesting informations on GIT here:

  • http://git-scm.com/documentation
  • http://savannah.gnu.org/maintenance/UsingGit
  • http://savannah.gnu.org/git/?group=liquidwar6

3.17.2 Getting the latest source

Simply install git and run the following command:

     git clone git://git.sv.gnu.org/liquidwar6.git

If you are behing a firewall and can’t have direct TCP/IP access:

     git clone http://git.sv.gnu.org/r/liquidwar6.git

Additionnally, source can be browsed online here: http://git.savannah.gnu.org/

cgit/liquidwar6.git

3.17.3 Developper access

You need an ssh access on Savannah (http://savannah.gnu.org/) and appropriate rights on the project, then you can type:

     git clone login@git.sv.gnu.org:/srv/git/liquidwar6.git

3.17.4 Submitting patches

If you have developper access to the project, then a simply git push will commit your changes.

If not (that is, if you checked out anonymously using git clone git://git.sv.gnu.org/liquidwar6.git for instance, you can still submit patches. Follow these steps:

• edit the code, make your patches, commit to your local GIT tree

• run git format-patch -p origin this command will generate .patch files, one for each of you commits, which you can send by email. They can be easily integrated in the main source tree by using git apply <file.patch>.

Note that you can need to run git format-patch -p master (with master instead of origin) it not using a fresh checkout. Also consider adding the --stdout switch, eg git

format-patch -p master --stdout > my-changes.patch if you’re not using a fresh check-

out.

3.18 Jenkins builds

Liquid War 6 uses Jenkins (http://jenkins-ci.org/) for continuous integration.

Each time a commit is done on the main GIT source tree, a build is triggered. The list of builds, their status, is available on:

  • all Jenkins jobs on ufoot.org (http://www.ufoot.org/jenkins/).
  • liquidwar6 Jenkins job on ufoot.org (http://www.ufoot.org/jenkins/job/
    liquidwar6/).

It’s interesting, among other things, to look at the log produced when the game is built, as it contains the test suite output, and can provide usefull informations of what is supposed to happen when the game is built correctly, and what happens when things go wrong.

4 Reference

This chapter is a technical reference. Most of its content is self-generated by the program itself. That is to say, most if its content is already available to you if you have the game

installed. Running liquidwar6 --list and liquidwar6 --about=<keyword> is very likely

to give you the very same informations, the advantage being that you’ll be sure the information is up-to-date and corresponds to the exact version of the program you have. However, publishing this in a reader-friendly way is convenient, plus it enables web search engines to harvest the content.

4.1 Basic options

4.1.1 about

[Command-line option] --about=<value>

Type: string

Will allow you to get informations about a given keyword. Let’s say that, for instance, you want informations about the keyword ’map-path’. Simply run ’liquidwar6 – about=map-path’. Note that this internal self-documentation system can describe command line switches as well as XML config file parameters or environment variables, and even some Guile script functions. The ’–list’ command line switch will give you the list of all available keywords.

4.1.2 audit

[Command-line option] --audit

Display all path values, defaults and current settings. This output is very usefull to track down problems such as missing directories, broken installations. If you get an error message that suggests some file is missing, then give this option a try.

4.1.3 copyright

[Command-line option] --copyright

Returns the copyright notice for the program.

4.1.4 credits

[Command-line option] --credits

Gives hopefully extensive information on who contributed to the game.

4.1.5 debug

[Command-line option] --debug

Enables debug mode. This will turn on maximum log information, and display everything on stderr, even messages which are normally only stored in the log file.

4.1.6 defaults

[Command-line option] --defaults

Clears the config file and run the game with default settings. Use this if you suspect you have broken something by tweaking user settings, or when upgrading the game to a new version.

4.1.7 help

[Command-line option] --help

Returns a short help for the program.

4.1.8 host

[Command-line option] --host

Display all known system host properties, including os and cpu informations.

4.1.9 list

[Command-line option] --list

Returns the list of all keywords which can be queried for information. This includes command-line options, environment variables, and so on. This is the companion option of ’–about’. Results obtained with ’–list’ can be passed to ’–about’.

4.1.10 modules

[Command-line option] --modules

Tells which modules have been enabled when the game was compiled. It’s still possible to add or remove modules afterwards, but this option allows you to know how things were at first.

4.1.11 pedigree

[Command-line option] --pedigree

Display all build values, these are general constants which can help debugging, tracing what binary you are running, and so on. It’s a good idea to take a look at the output of ’pedigree’ if you have problems running the game.

4.1.12 test

[Command-line option] --test

Runs a (hopefully) complete test suite which will call most internal Liquid War 6 functions and check out wether they work, in a simple context, without any game interference. Usefull for troubleshooting.

4.1.13 version

[Command-line option] --version

Returns the version of the program, as defined by the GNU Coding Standards.

4.2 Doc options

4.2.1 example-hints-xml

[Command-line option] --example-hints-xml

Dumps on stdout an example hints.xml file. Such a file is normally shipped with the game. It is indeed generated using this command.

4.2.2 example-rules-xml

[Command-line option] --example-rules-xml

Dumps on stdout an example options.xml file. Such a file is normally shipped with the game. It is indeed generated using this command.

4.2.3 example-style-xml

[Command-line option] --example-style-xml

Dumps on stdout an example style.xml file. Such a file is normally shipped with the game. It is indeed generated using this command.

4.2.4 example-teams-xml

[Command-line option] --example-teams-xml

Dumps on stdout an example teams.xml file. Such a file is normally shipped with the game. It is indeed generated using this command.

4.2.5 list-advanced

[Command-line option] --list-advanced

List advanced options which can be used for fine-tuning the game.

4.2.6 list-aliases

[Command-line option] --list-aliases

List the keyword aliases. These are here for convenience.

4.2.7 list-doc

[Command-line option] --list-doc

List documentation-related command line options. These commands allow you to list all the keywords related to a given domain.

4.2.8 list-funcs

[Command-line option] --list-funcs

List the C-functions which are exported to Guile, thus usable in scripts.

4.2.9 list-graphics

[Command-line option] --list-graphics

List graphics options (resolution, fullscreen...).

4.2.10 list-hooks

[Command-line option] --list-hooks

List user-modifiable hooks.

4.2.11 list-input

[Command-line option] --list-input

List input (AKA controls) related options. Use these to change keyboard, joystick and mouse settingds.

4.2.12 list-map

[Command-line option] --list-map

List map-related entries, excluding rules.xml, hints.xml and style.xml entries.

4.2.13 list-map-hints

[Command-line option] --list-map-hints

List ’hints.xml’ entries. These parameters enable you to modify the behavior of the map loader.

4.2.14 list-map-rules

[Command-line option] --list-map-rules

List ’options.xml’ entries. These parameters enable you to modify the gameplay.

4.2.15 list-map-style

[Command-line option] --list-map-style

List ’style.xml’ entries. These parameters enable you to modify the aspect of the game.

4.2.16 list-map-teams

[Command-line option] --list-map-teams

List ’teams.xml’ entries. These parameters enable you to specify which teams should play on the map.

4.2.17 list-network

[Command-line option] --list-network

List network options.

4.2.18 list-path

[Command-line option] --list-path

List parameters which allow you to override the defaults of the game, and force the game your own file paths and directories.

4.2.19 list-players

[Command-line option] --list-players

List player-related entries, that is to say ’who plays’.

4.2.20 list-quick

[Command-line option] --list-quick

List quick help entries, this includes the GNU standard options and a few troubleshooting tools.

4.2.21 list-show

[Command-line option] --list-show

List command-line options which begin with ’–show-...’. These will display on the console many internal parameters. Usefull when debugging.

4.2.22 list-sound

[Command-line option] --list-sound

List sound options (volume...).

4.2.23 list-team-colors

[Command-line option] --list-team-colors

List the team colors, there should be 10 of them

4.2.24 list-weapons

[Command-line option] --list-weapons

List the available weapons.

4.3 Show options

4.3.1 show-build-abs-srcdir

[Command-line option] --show-build-abs-srcdir

Shows the top source directory on the machine where the binary was compiled, as an absolute path.

4.3.2 show-build-bin-id

[Command-line option] --show-build-bin-id

Shows the internal ’bin-id’ value. This value does not mean anything in itself but it’s supposed to change each time you compile the game.

4.3.3 show-build-bugs-url

[Command-line option] --show-build-bugs-url

Shows the URL to make bug reports.

4.3.4 show-build-cflags

[Command-line option] --show-build-cflags

Shows what value you should put in ’CFLAGS’ (environment variable) if you want to compile programs that use Liquid War 6 as a library, and include ’liquidwar6.h’.

4.3.5 show-build-codename

[Command-line option] --show-build-codename

Shows the codename associated with this version, generally the name of someone famous who is war-related (a general, an emperor...).

4.3.6 show-build-configure-args

[Command-line option] --show-build-configure-args

Shows the arguments that have been passed to the GNU Autoconf ’./configure’ script when building the program. This can be very usefull if you want to know how the program has been built.

4.3.7 show-build-copyright

[Command-line option] --show-build-copyright

Shows a very short copyright notice.

4.3.8 show-build-datadir

[Command-line option] --show-build-datadir

Shows the ’datadir’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local/share’. This is the generic, non Liquid War 6 specific data directory. Liquid War 6 related data is stored elsewhere (usually in a sub-directory) see the ’data-dir’ switch for more information. ’datadir’ is not ’data-dir’. That’s the point.

4.3.9 show-build-date

[Command-line option] --show-build-date

Shows the date when the binary was compiled.

4.3.10 show-build-docdir

[Command-line option] --show-build-docdir

Shows the ’docdir’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local/share/doc/liquidwar6’.

4.3.11 show-build-enable-allinone

[Command-line option] --show-build-enable-allinone

Shows wether the ’allinone’ option has been chosen when building the game. This depends on parameters passed to ’./configure’.

4.3.12 show-build-enable-console

[Command-line option] --show-build-enable-console

Shows wether the console has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of ncurses and readline.

4.3.13 show-build-enable-fullstatic

[Command-line option] --show-build-enable-fullstatic

Shows wether the ’fullstatic’ option has been chosen when building the game. This depends on parameters passed to ’./configure’.

4.3.14 show-build-enable-gcov

[Command-line option] --show-build-enable-gcov

Shows wether the game was build with suitable informations for gcov. This depends on parameters passed to ’./configure’.

4.3.15 show-build-enable-gprof

[Command-line option] --show-build-enable-gprof

Shows wether the game was build with suitable informations for gprof. This depends on parameters passed to ’./configure’.

4.3.16 show-build-enable-gtk

[Command-line option] --show-build-enable-gtk

Shows wether GTK+ support has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of GTK+ headers and libs. It uses pkg-config to detect it.

4.3.17 show-build-enable-instrument

[Command-line option] --show-build-enable-instrument

Shows wether the game was build with the ’-finstrument-functions’ GCC switch. This depends on parameters passed to ’./configure’.

4.3.18 show-build-enable-mod-caca

[Command-line option] --show-build-enable-mod-caca

Shows wether the mod-caca graphical backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of libcaca related libraries.

4.3.19 show-build-enable-mod-csound

[Command-line option] --show-build-enable-mod-csound

Shows wether the mod-csound audio backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of the csound library.

4.3.20 show-build-enable-mod-gl1

[Command-line option] --show-build-enable-mod-gl1

Shows wether the mod-gl1 graphical backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of SDL and OpenGL related libraries.

4.3.21 show-build-enable-mod-gles2

[Command-line option] --show-build-enable-mod-gles2

Shows wether the mod-gles2 graphical backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of SDL and OpenGL ES related libraries.

4.3.22 show-build-enable-mod-http

[Command-line option] --show-build-enable-mod-http

Shows wether the mod-http network backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of libCurl.

4.3.23 show-build-enable-mod-ogg

[Command-line option] --show-build-enable-mod-ogg

Shows wether the mod-ogg audio backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of SDL and related libraries.

4.3.24 show-build-enable-mod-soft

[Command-line option] --show-build-enable-mod-soft

Shows wether the mod-soft graphical backend has been enabled when building the game. This depends on parameters passed to ’./configure’ and also on the presence of SDL related libraries.

4.3.25 show-build-enable-openmp

[Command-line option] --show-build-enable-openmp

Shows wether the program was built with OpenMP support. This depends on parameters passed to ’./configure’.

4.3.26 show-build-enable-optimize

[Command-line option] --show-build-enable-optimize

Shows wether the ’optimize’ option has been chosen when building the game. This depends on parameters passed to ’./configure’.

4.3.27 show-build-enable-paranoid

[Command-line option] --show-build-enable-paranoid

Shows wether the game was build with paranoid memory management. This is for debugging purposes, the default already includes some controls, with turned it’s really... paranoid.

4.3.28 show-build-enable-profiler

[Command-line option] --show-build-enable-profiler

Shows wether the game was build with Google Profiler support. This depends on parameters passed to ’./configure’.

4.3.29 show-build-enable-valgrind

[Command-line option] --show-build-enable-valgrind

Shows wether the game was build with valgrind later use in mind. This depends on parameters passed to ’./configure’.

4.3.30 show-build-endianness

[Command-line option] --show-build-endianness

Returns the endianness. ’little’ corresponds to x86-like systems, ’big’ to ppc-like systems.

4.3.31 show-build-gcc-version

[Command-line option] --show-build-gcc-version

Returns the version of the GNU C compiler which was used to compile the program.

4.3.32 show-build-gnu

[Command-line option] --show-build-gnu

Returns 1 (true) if host OS is a GNU system, or at least has been considered as such when compiling, 0 (false) if not.

4.3.33 show-build-gp2x

[Command-line option] --show-build-gp2x

Returns 1 (true) if host is a GP2X, 0 (false) if not.

4.3.34 show-build-home-url

[Command-line option] --show-build-home-url

Shows the URL of the program, its homepage.

4.3.35 show-build-host-cpu

[Command-line option] --show-build-host-cpu

Shows the host CPU, as defined by ’host cpu’ in GNU Autoconf.

4.3.36 show-build-host-os

[Command-line option] --show-build-host-os

Shows the host OS, as defined by ’host os’ in GNU Autoconf.

4.3.37 show-build-hostname

[Command-line option] --show-build-hostname

Shows the name of the host where the binary was compiled.

4.3.38 show-build-includedir

[Command-line option] --show-build-includedir

Shows the ’includedir’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local/include’.

4.3.39 show-build-ldflags

[Command-line option] --show-build-ldflags

Shows what value you should put in ’LDFLAGS’ (environment variable) if you want to link programs against libliquidwar6.

4.3.40 show-build-libdir

[Command-line option] --show-build-libdir

Shows the ’libdir’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local/lib’. This is the generic, non Liquid War 6 specific library directory. Dedicated Liquid War 6 modules are stored elsewhere (usually in a sub-directory) see the ’mod-dir’ switch for more information.

4.3.41 show-build-license

[Command-line option] --show-build-license

Shows the license of the program (GNU GPL v3 or later).

4.3.42 show-build-localedir

[Command-line option] --show-build-localedir

Shows the ’localedir’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local/share/locale’.

4.3.43 show-build-mac-os-x

[Command-line option] --show-build-mac-os-x

Returns 1 (true) if host OS is Mac OS X, 0 (false) if not.

4.3.44 show-build-md5sum

[Command-line option] --show-build-md5sum

Shows the MD5 checksum, which has been calculated from the C source files. Complementary with ’show-build-stamp’.

4.3.45 show-build-ms-windows

[Command-line option] --show-build-ms-windows

Returns 1 (true) if host OS is Microsoft Windows, 0 (false) if not.

4.3.46 show-build-package-id

[Command-line option] --show-build-package-id

Shows the package tarname with its version, that is, ’liquidwar6-<version>

4.3.47 show-build-package-name

[Command-line option] --show-build-package-name

Shows the package name, that is, ’Liquid War 6’.

4.3.48 show-build-package-string

[Command-line option] --show-build-package-string

Shows the package string, that is, ’Liquid War 6 <version>

4.3.49 show-build-package-tarname

[Command-line option] --show-build-package-tarname

Shows the package tarname, that is, liquidwar6.

4.3.50 show-build-pointer-size

[Command-line option] --show-build-pointer-size

Returns the pointer size, in bytes. Should be 4 on 32-bit systems and 8 on 64-bit systems.

4.3.51 show-build-prefix

[Command-line option] --show-build-prefix

Shows the ’prefix’ value as passed to the GNU Autoconf ’./configure’ script when compiling the program. Default is ’/usr/local’.

4.3.52 show-build-stamp

[Command-line option] --show-build-stamp

Shows the build stamp. A very usefull value, more precise than the version to track down binaries. It is incremented each time the core C code is updated. It won’t reflect all the programs for it does not take scripts in account, but if you are running a work-in-progress version, it might be very convenient to use this to know what your are running exactly. It’s also used as the revision number (the third number afer MAJOR.MINOR).

4.3.53 show-build-time

[Command-line option] --show-build-time

Shows the time when the binary was compiled.

4.3.54 show-build-top-srcdir

[Command-line option] --show-build-top-srcdir

Shows the top source directory on the machine where the binary was compiled, as a (possibly) relative path.

4.3.55 show-build-unix

[Command-line option] --show-build-unix

Returns 1 (true) if host OS is a UNIX system, or at least has been considered as such when compiling, 0 (false) if not.

4.3.56 show-build-version

[Command-line option] --show-build-version

Shows the version. Note that this is different from the standard GNU ’version’ command line option which shows a complete message with a short copyright notice. This one will just return the version, without the package tarname or anything else.

4.3.57 show-build-version-base

[Command-line option] --show-build-version-base

Shows the version base. This is basically MAJOR.MINOR and determines the level of compatibility of the program. Two programs with the same base version should be able to communicate on the network, share data files and even binary modules if on the same platform.

4.3.58 show-build-version-major

[Command-line option] --show-build-version-major

Shows the major version number. This is just used to differenciate alpha/beta releases (using 0) from stable releases (using 6).

4.3.59 show-build-version-minor

[Command-line option] --show-build-version-minor

Shows the minor version number. This is manually increased at each significant, public release of the game.

4.3.60 show-build-x86

[Command-line option] --show-build-x86

Tells wether the CPU belongs to the x86 family.

4.3.61 show-config-file

[Command-line option] --show-config-file

Shows the config file path. Default is ’$HOME/.liquidwar6/config.xml’.

4.3.62 show-cwd

[Command-line option] --show-cwd

Shows the current working directory, the value that the pwd command would return.

4.3.63 show-data-dir

[Command-line option] --show-data-dir

Shows the data directory path. This is where the games searches for most of its data,the most important exception being maps, which are stored elsewhere. Default is ’/usr/local/share/liquidwar6-<version>/data’.

4.3.64 show-default-config-file

[Command-line option] --show-default-config-file

Shows the default config file path. Default is ’$HOME/.liquidwar6/config.xml’.

4.3.65 show-default-data-dir

[Command-line option] --show-default-data-dir

Shows the default data directory path. This is where the games searches for most of its data,the most important exception being maps, which are stored elsewhere. Default is ’/usr/local/share/liquidwar6-<version>/data’.

4.3.66 show-default-log-file

[Command-line option] --show-default-log-file

Shows the default log file path. Default is ’$HOME/.liquidwar6/log.csv’.

4.3.67 show-default-map-dir

[Command-line option] --show-default-map-dir

Shows the default map directory. This is where builtin maps are stored. Default is ’/usr/local/share/liquidwar6-<version>/map’.

4.3.68 show-default-map-path

[Command-line option] --show-default-map-path

Shows the default map search path. This is where the game searches for maps. It’s the combination of command-line arguments and builtin paths. Might return more directories than the one specified in a single ’map-path=dir1:dir2’ argument.

4.3.69 show-default-mod-dir

[Command-line option] --show-default-mod-dir

Shows the default module directory path. This is where all dynamically loaded modules are stored. Default is ’/usr/local/lib/liquidwar6-<version>’.

4.3.70 show-default-music-dir

[Command-line option] --show-default-music-dir

Shows the default music directory. This is where builtin musics are stored. Default is ’/usr/local/share/liquidwar6-<version>/music’.

4.3.71 show-default-music-path

[Command-line option] --show-default-music-path

Shows the default music search path. This is where the game searches for musics. It’s the combination of command-line arguments and builtin paths. Might return more directories than the one specified in a single ’music-path=dir1:dir2’ argument.

4.3.72 show-default-prefix

[Command-line option] --show-default-prefix

Shows the default prefix used. This should logically be the value passed to the GNU Autoconf ./configure script when building the game. Most other path are deduced from this one. Default is ’/usr/local’.

4.3.73 show-default-script-file

[Command-line option] --show-default-script-file

Shows the default main script file path. This file is very important, since the program is more or less a hudge scheme interpreter, and this file is the file loaded by Guile. In short, it is the main program. Default is ’/usr/local/share/liquidwar6- <version>/script/liquidwar6.scm’.

4.3.74 show-default-user-dir

[Command-line option] --show-default-user-dir

Shows the default user directory path. This is where run-time data, config files, log files, are stored. Default is ’$HOME/.liquidwar6/’.

4.3.75 show-log-file

[Command-line option] --show-log-file

Shows the log file path. Default is ’$HOME/.liquidwar6/log.csv’.

4.3.76 show-map-dir

[Command-line option] --show-map-dir

Shows the map directory. This is where builtin maps are stored. Default is ’/usr/local/share/liquidwar6-<version>/map’.

4.3.77 show-map-path

[Command-line option] --show-map-path

Shows the map search path. This is where the game searches for maps. It’s the combination of command-line arguments and builtin paths. Might return more directories than the one specified in a single ’map-path=dir1:dir2’ argument.

4.3.78 show-mod-dir

[Command-line option] --show-mod-dir

Shows the module directory path. This is where all dynamically loaded modules are stored. Default is ’/usr/local/lib/liquidwar6-<version>’.

4.3.79 show-music-dir

[Command-line option] --show-music-dir

Shows the music directory. This is where builtin maps are stored. Default is ’/usr/local/share/liquidwar6-<version>/music’.

4.3.80 show-music-path

[Command-line option] --show-music-path

Shows the music search path. This is where the game searches for musics. It’s the combination of command-line arguments and builtin paths. Might return more directories than the one specified in a single ’music-path=dir1:dir2’ argument.

4.3.81 show-prefix

[Command-line option] --show-prefix

Shows the prefix used. This should logically be the value passed to the GNU Autoconf ./configure script when building the game. Most other path are deduced from this one. Default is ’/usr/local’.

4.3.82 show-run-dir

[Command-line option] --show-run-dir

Shows the run directory, usually the path where the binary is. It depends on how and where the program is launched. It is guessed from the argc/argv values at runtime.

4.3.83 show-script-file

[Command-line option] --show-script-file

Shows the main script file path. This file is very important, since the program is more or less a hudge scheme interpreter, and this file is the file loaded by Guile. In short, it is the main program. Default is ’/usr/local/share/liquidwar6- <version>/script/liquidwar6.scm’.

4.3.84 show-user-dir

[Command-line option] --show-user-dir

Shows the user directory path. This is where run-time data, config files, log files, are stored. Default is ’$HOME/.liquidwar6/’.

4.4 Path options

4.4.1 config-file

[Command-line option] --config-file

Type: string Default value: $HOME/.liquidwar6/config.xml Set the config file path. This enables you to use whatever config file you like, keeping all other informations in the same place.

4.4.2 data-dir

[Command-line option] --data-dir

Type: string Default value: /usr/local/share/liquidwar6-<version>/data Set the data directory. By changing ths value you’ll be able to use an alternative data directory.

4.4.3 log-file

[Command-line option] --log-file=<value> [Environment variable] LW6_LOG_FILE

[XML key] log-file

Type: string Default value: $HOME/.liquidwar6/log.csv Set the log file path. This enables you to use whatever log file you like, keeping all other informations in the same place.

4.4.4 map-dir

[Command-line option] --map-dir

Type: string Default value: /usr/local/share/liquidwar6-<version>/map Set the map directory path. By changing this value you’ll be able to play with your own maps in your own directory. Note that there are other ways to achieve that, but using this option will work. However, a side effect is that you might not see builtin maps anymore.

4.4.5 map-path

[Command-line option] --map-path=<value> [Environment variable] LW6_MAP_PATH

[XML key] map-path

Type: string Default value: $HOME/.liquidwar6/map:/usr/local/share/liquidwar6- <version>/map Set the map search path. By changing this value you’ll be able to play with your own maps in your own directory. This is different from ’map-dir’, since it includes ’map-dir’, plus it adds a number of other search paths. Unlike most other parameters, the values given from the command-line, from the environment variables, or from the config file, are not overwritten, but appended. That is to say if you specify a ’map-path’ with the command-line argument ’map-path=path’, but also define the ’LW6 MAP PATH’ value and finally edit ’config.xml’ to change the ’mappath’ entry in it, you’ll end up with the game searching for maps in all these directories. Additionnally, ’map-dir’ and ’<user-dir>/map’ will always be in the list. Any given value can itself include several pathes, separated by the path separator. This separator is ’:’ on GNU/Linux, and ’;’ on Microsoft Windows. For instance, on a GNU/Linux box, you could use the command-line argument ’mappath=/foo/bar/map:/home/user/map/:/map’.

4.4.6 mod-dir

[Command-line option] --mod-dir

Type: string Default value: /usr/local/lib/liquidwar6-<version> Set the module directory path. By changing this you will load dynamic shared libraries (game specific modules such as the graphical backend) from an alternative place. Use this at your own risks, for there can always be a binary incompatibility. You’ve been warned.

4.4.7 music-dir

[Command-line option] --music-dir=<value> [Environment variable] LW6_MUSIC_DIR

[XML key] music-dir

Type: string Default value: /usr/local/share/liquidwar6-<version>/music Set the music directory path. By changing this value you’ll be able to use your own musics in your own directory. Note that there are other ways to achieve that, but using this option will work. The major side effect is that using this option, you really replace the existing builtin musics by your own. If you simply want to add musics you can store them in $HOME/.liquidwar6/music or in the map directory itself.

4.4.8 music-path

[Command-line option] --music-path=<value>

[Environment variable] LW6_MUSIC_PATH

[XML key] music-path

Type: string Default value: $HOME/.liquidwar6/music:/usr/local/share/liquidwar6- <version>/music Set the music search path. By changing this value you’ll be able to play with your own musics in your own directory. This is different from ’music-dir’, since it includes ’music-dir’, plus it adds a number of other search paths. Unlike most other parameters, the values given from the command-line, from the environment variables, or from the config file, are not overwritten, but appended. That is to say if you specify a ’music-path’ with the command-line argument ’music-path=path’, but also define the ’LW6 MUSIC PATH’ value and finally edit ’config.xml’ to change the ’music-path’ entry in it, you’ll end up with the game searching for musics in all these directories. Additionnally, ’music-dir’ and ’<user-dir>/music’ will always be in the list. Any given value can itself include several pathes, separated by the path separator. This separator is ’:’ on GNU/Linux, and ’;’ on Microsoft Windows. For instance, on a GNU/Linux box, you could use the command-line argument ’musicpath=/foo/bar/music:/home/user/music/:/music’.

4.4.9 prefix

[Command-line option] --prefix

Type: string Default value: /usr/local Override the prefix value given to the GNU Autoconf ./configure script when building the game. Not all path will be changed, some of them might remain the same, for instance message translations (localedir). But most game-specific data including maps, graphics, sounds, will be searched according to the new given parameter.

4.4.10 script-file

[Command-line option] --script-file

Type: string Default value: /usr/local/share/liquidwar6-<version>/script/liquidwar6.scm Set the main script file path. This file is very important, since the program is more or less a hudge scheme interpreter, and this file is the file loaded by Guile. In short, it is the main program.

4.4.11 user-dir

[Command-line option] --user-dir=<value> [Environment variable] LW6_USER_DIR

[XML key] user-dir

Type: string Default value: $HOME/.liquidwar6 Set the user directory path. This is where run-time data, config files, log files, are stored. If you override this value, other parameters such as where the config and log files reside, will change.

4.5 Players options

4.5.1 player1-control

[Command-line option] --player1-control=<value>

[Environment variable] LW6_PLAYER1_CONTROL

[XML key] player1-control

Type: string

Default value: mouse

Control for the first player, must be mouse, keyboard, joystick1, joystick2 or custom.

4.5.2 player1-name

[Command-line option] --player1-name=<value>

[Environment variable] LW6_PLAYER1_NAME

[XML key] player1-name

Type: string

Default value: <username>

Name of the first player, the player used by default. A default value is provided, you can of course, change it at will.

4.5.3 player1-status

[Command-line option] --player1-status=<value>

[Environment variable] LW6_PLAYER1_STATUS

[XML key] player1-status

Type: boolean

Default value: true

Status of the first player, true if player is activated, false if idle.

4.5.4 player2-control

[Command-line option] --player2-control=<value>

[Environment variable] LW6_PLAYER2_CONTROL

[XML key] player2-control

Type: string

Default value: keyboard

Control for the second player, must be mouse, keyboard, joystick1, joystick2 or custom.

4.5.5 player2-name

[Command-line option] --player2-name=<value>

[Environment variable] LW6_PLAYER2_NAME

[XML key] player2-name

Type: string

Default value: player2-<hostname>

Name of the second player. A default value is provided, you’ll certainly want to change it.

4.5.6 player2-status

[Command-line option] --player2-status=<value>

[Environment variable] LW6_PLAYER2_STATUS

[XML key] player2-status

Type: boolean Default value: true Status of the second player, true if player is activated, false if idle.

4.5.7 player3-control

[Command-line option] --player3-control=<value>

[Environment variable] LW6_PLAYER3_CONTROL

[XML key] player3-control

Type: string

Default value: joystick1

Control for the third player, must be mouse, keyboard, joystick1, joystick2 or custom.

4.5.8 player3-name

[Command-line option] --player3-name=<value>

[Environment variable] LW6_PLAYER3_NAME

[XML key] player3-name

Type: string

Default value: player3-<hostname>

Name of the third player. A default value is provided, you’ll certainly want to change it.

4.5.9 player3-status

[Command-line option] --player3-status=<value>

[Environment variable] LW6_PLAYER3_STATUS

[XML key] player3-status

Type: boolean Default value: false Status of the third player, true if player is activated, false if idle.

4.5.10 player4-control

[Command-line option] --player4-control=<value>

[Environment variable] LW6_PLAYER4_CONTROL

[XML key] player4-control

Type: string

Default value: joystick2

Control for the fourth player, must be mouse, keyboard, joystick1, joystick2 or custom.

4.5.11 player4-name

[Command-line option] --player4-name=<value>

[Environment variable] LW6_PLAYER4_NAME

[XML key] player4-name

Type: string

Default value: player4-<hostname>

Name of the fourth player. A default value is provided, you’ll certainly want to change it.

4.5.12 player4-status

[Command-line option] --player4-status=<value>

[Environment variable] LW6_PLAYER4_STATUS

[XML key] player4-status

Type: boolean

Default value: false

Status of the fourth player, true if player is activated, false if idle.

4.6 Input options

4.6.1 auto-release-delay

[Command-line option] --auto-release-delay=<value>
[Environment variable] LW6_AUTO_RELEASE_DELAY
[XML key] auto-release-delay

Type: integer

Default value: 250

Time, in milliseconds, before which a key is automatically released. This might or might not be used by the graphics backend, it’s typically used by backends which don’t always handle key releases events the right way, that is, don’t fire them. Libcaca is a good example of such a case.

4.6.2 click-to-focus

[Command-line option] --click-to-focus=<value>

[Environment variable] LW6_CLICK_TO_FOCUS

[XML key] click-to-focus

Type: boolean

Default value: false

If set to true, you’ll need to click with the mouse to select a menuitem or move the cursor in the game. If not, some actions will be taken automatically without the need to click.

4.6.3 cursor-sensitivity

[Command-line option] --cursor-sensitivity=<value>
[Environment variable] LW6_CURSOR_SENSITIVITY
[XML key] cursor-sensitivity

Type: float Default value: 1.0 Keyboard and joystick sensitivity while moving the cursor. 1.0 is the default, 0.1 is slow, 10 is reponsive. This is used for moving the cursor during the game only, the option has no impact on menu navigation.

4.6.4 custom-alt

[Command-line option] --custom-alt=<value>

[Environment variable] LW6_CUSTOM_ALT

[XML key] custom-alt

Type: string Default value: (c-lw6gui-keyboard-is-pressed 110) ; SDLK n Guile custom code associated to the ALT key equivalent.

4.6.5 custom-ctrl

[Command-line option] --custom-ctrl=<value>

[Environment variable] LW6_CUSTOM_CTRL

[XML key] custom-ctrl

Type: string Default value: (c-lw6gui-keyboard-is-pressed 98) ; SDLK b Guile custom code associated to the CTRL key equivalent.

4.6.6 custom-down

[Command-line option] --custom-down=<value>

[Environment variable] LW6_CUSTOM_DOWN

[XML key] custom-down

Type: string Default value: (c-lw6gui-keyboard-is-pressed 100) ; SDLK d Guile custom code associated to the DOWN key equivalent.

4.6.7 custom-enter

[Command-line option] --custom-enter=<value>

[Environment variable] LW6_CUSTOM_ENTER

[XML key] custom-enter

Type: string Default value: (c-lw6gui-keyboard-is-pressed 103) ; SDLK g Guile custom code associated to the ENTER key equivalent.

4.6.8 custom-esc

[Command-line option] --custom-esc=<value>

[Environment variable] LW6_CUSTOM_ESC

[XML key] custom-esc

Type: string

Default value: (c-lw6gui-keyboard-is-pressed 102) ; SDLK f

Guile custom code associated to the ESC key equivalent.

4.6.9 custom-left

[Command-line option] --custom-left=<value>

[Environment variable] LW6_CUSTOM_LEFT

[XML key] custom-left

Type: string

Default value: (c-lw6gui-keyboard-is-pressed 99) ; SDLK c

Guile custom code associated to the LEFT key equivalent.

4.6.10 custom-pgdown

[Command-line option] --custom-pgdown=<value>

[Environment variable] LW6_CUSTOM_PGDOWN

[XML key] custom-pgdown

Type: string

Default value: (c-lw6gui-keyboard-is-pressed 115) ; SDLK s

Guile custom code associated to the PGDOWN key equivalent.

4.6.11 custom-pgup

[Command-line option] --custom-pgup=<value>

[Environment variable] LW6_CUSTOM_PGUP

[XML key] custom-pgup

Type: string

Default value: (c-lw6gui-keyboard-is-pressed 119) ; SDLK w

Guile custom code associated to the PGUP key equivalent.

4.6.12 custom-right

[Command-line option] --custom-right=<value>

[Environment variable] LW6_CUSTOM_RIGHT

[XML key] custom-right

Type: string

Default value: (c-lw6gui-keyboard-is-pressed 118) ; SDLK v

Guile custom code associated to the RIGHT key equivalent.

4.6.13 custom-up

[Command-line option] --custom-up=<value> [Environment variable] LW6_CUSTOM_UP

[XML key] custom-up

Type: string Default value: (c-lw6gui-keyboard-is-pressed 101) ; SDLK e Custom keycode to be used as the UP key equivalent.

4.6.14 double-click-delay

[Command-line option] --double-click-delay=<value>
[Environment variable] LW6_DOUBLE_CLICK_DELAY
[XML key] double-click-delay

Type: integer Default value: 333 Time, in milliseconds, determining wether two consecutive clicks make a double-click or not.

4.6.15 max-cursor-speed

[Command-line option] --max-cursor-speed=<value>

[Environment variable] LW6_MAX_CURSOR_SPEED

[XML key] max-cursor-speed

Type: float Default value: 10.0 Maximum cursor speed when cursor is controlled with keyboard or joystick joystick 1. Consider using cursor-sensitivity too.

4.6.16 mouse-sensitivity

[Command-line option] --mouse-sensitivity=<value>
[Environment variable] LW6_MOUSE_SENSITIVITY
[XML key] mouse-sensitivity

Type: float Default value: 1.0 Mouse sensitivity, 1.0 is the default, 0.1 is slow, 10 is reponsive. This is used for moving the cursor during the game only, the option has no impact on menu navigation.

4.6.17 repeat-delay

[Command-line option] --repeat-delay=<value>

[Environment variable] LW6_REPEAT_DELAY

[XML key] repeat-delay

Type: integer Default value: 500 Time, in milliseconds, before key repeat will start, use 0 to disable.

4.6.18 repeat-interval

[Command-line option] --repeat-interval=<value>

[Environment variable] LW6_REPEAT_INTERVAL

[XML key] repeat-interval

Type: integer Default value: 100 Time, in milliseconds, between two repeats, once repeat has started, use 0 to disable.

4.6.19 use-double-click

[Command-line option] --use-double-click=<value>

[Environment variable] LW6_USE_DOUBLE_CLICK

[XML key] use-double-click

Type: boolean Default value: false Wether to use double-click feature, mostly usefull if running on a system that has only one button (such as a tablet-PC or anything with a tactile screen), if your mouse has three buttons, disabling this might avoid some confusion. Basically, if enabled, double-click is equivalent to right-click (fire) and triple-click is equivalent to middle-click (alternate fire).

4.6.20 use-esc-button

[Command-line option] --use-esc-button=<value>

[Environment variable] LW6_USE_ESC_BUTTON

[XML key] use-esc-button

Type: boolean Default value: true Decides wether to display an ’esc’ (escape) button in the interface. This is usefull for people who control the game with the mouse only, and have a single buttons, or on a touchscreen.

4.6.21 zoom-step

[Command-line option] --zoom-step=<value> [Environment variable] LW6_ZOOM_STEP

[XML key] zoom-step

Type: float Default value: 1.1 A value, strictly greater than 1, which will be used when zooming. The greater it is, the more sensible the zoom is.

4.6.22 zoom-stick-delay

[Command-line option] --zoom-stick-delay=<value>

[Environment variable] LW6_ZOOM_STICK_DELAY

[XML key] zoom-stick-delay

Type: float Default value: 1000 How long, in msec, the zoom will stick to its default value.

4.7 Graphics options

4.7.1 capture

[Command-line option] --capture=<value> [Environment variable] LW6_CAPTURE [XML key] capture

Type: boolean Default value: false Enables capture mode, in which a BMP file is dumped on the disk (in your user directory, search for a ’capture’ sub-directory).

4.7.2 fullscreen

[Command-line option] --fullscreen=<value>

[Environment variable] LW6_FULLSCREEN

[XML key] fullscreen

Type: boolean Default value: false Force the game to fun fullscreen. Note that the graphics backend might ignore this hint.

4.7.3 gfx-backend

[Command-line option] --gfx-backend=<value>

[Environment variable] LW6_GFX_BACKEND

[XML key] gfx-backend

Type: string Default value: gl1 Sets the graphics backend AKA ’gfx’ to use. For now the only reasonnable choice is ’gl1’ and will use an OpenGL v1 / SDL 3D-accelerated driver.

4.7.4 gfx-quality

[Command-line option] --gfx-quality=<value>

[Environment variable] LW6_GFX_QUALITY

[XML key] gfx-quality

Type: integer Default value: 1 Min value: 0 Max value: 2 Sets the overall quality of the graphics backend. Depending on the backend, this can mean different things. For instance for the ’gl’ backend, this can change texture filtering (nearest, linear, bilinear...). This is not the same as ’pixelize’ which is a per-map option and emulates an old school appearance.

4.7.5 height

[Command-line option] --height=<value> [Environment variable] LW6_HEIGHT [XML key] height

Type: integer Default value: -1 Run the game with the given screen height.Note that the graphics backend might ignore this hint. Use with its companion option ’width’. A negative value will force the use of a default value.

4.7.6 width

[Command-line option] --width=<value> [Environment variable] LW6_WIDTH [XML key] width

Type: integer Default value: -1 Run the game with the given screen width. Note that the graphics backend might ignore this hint. Use with its companion option ’height’.A negative value will force the use of a default value.

4.7.7 windowed-mode-limit

[Command-line option] --windowed-mode-limit=<value>
[Environment variable] LW6_WINDOWED_MODE_LIMIT
[XML key] windowed-mode-limit

Type: float Default value: 0.95 When switching back from fullscreen mode to windowed mode, if we’re in maximum resolution, then this coefficient will be applied before resizing the window. The idea is that (obviously) a windowed mode is prefered when a little smaller that totally fullscreen. So set this to a value just below 1.0.

4.8 Sound options

4.8.1 ambiance-exclude

[Command-line option] --ambiance-exclude=<value>

[Environment variable] LW6_AMBIANCE_EXCLUDE

[XML key] ambiance-exclude

Type: string Default value: If this string is present in a music file name, this file won’t be played during the menus, it will be excluded from the list.

4.8.2 ambiance-file

[Command-line option] --ambiance-file=<value>

[Environment variable] LW6_AMBIANCE_FILE

[XML key] ambiance-file

Type: string Default value: A music file which will be used to be played during the menus. If not found, game will fallback on random files.

4.8.3 ambiance-filter

[Command-line option] --ambiance-filter=<value>

[Environment variable] LW6_AMBIANCE_FILTER

[XML key] ambiance-filter

Type: string Default value: Chadburn A music filter, used to select the files which are played while navigating in the menus. It works like ’music-filter’ except this one is not related to a peculiar map. This is not a complex regex-enabled filter, just a plain string search. Even the ’*’ wildcard won’t work.

4.8.4 fx-volume

[Command-line option] --fx-volume=<value> [Environment variable] LW6_FX_VOLUME

[XML key] fx-volume

Type: float Default value: 0.3 Min value: 0 Max value: 1 Set the sound effects volume. This is a floating point value. 0 is mute. Maximum value is 1.

4.8.5 music-volume

[Command-line option] --music-volume=<value>

[Environment variable] LW6_MUSIC_VOLUME

[XML key] music-volume

Type: float Default value: 0.6 Min value: 0 Max value: 1 Set the music volume.This is a floating point value. 0 is mute. Maximum value is 1.

4.8.6 snd-backend

[Command-line option] --snd-backend=<value>

[Environment variable] LW6_SND_BACKEND

[XML key] snd-backend

Type: string Default value: ogg Sets the sound backend AKA ’snd’ to use. Can be ’ogg’ or ’csound’ but only ’ogg’ will produce sound in the current release.

4.8.7 water-volume

[Command-line option] --water-volume=<value>

[Environment variable] LW6_WATER_VOLUME

[XML key] water-volume

Type: float Default value: 0.2 Min value: 0 Max value: 1 Set the volume for water sounds. This is a floating point value. 0 is mute. Maximum value is 1.

4.9 Network options

4.9.1 bind-ip

[Command-line option] --bind-ip=<value> [Environment variable] LW6_BIND_IP [XML key] bind-ip

Type: string Default value: 0.0.0.0 The IP address to bind on when listening to network messages. You can use this to specifically use a given network interface, the default will listen on any available interface.

4.9.2 bind-port

[Command-line option] --bind-port=<value> [Environment variable] LW6_BIND_PORT

[XML key] bind-port

Type: integer Default value: 8056 Min value: 1 Max value: 65535 The IP port to bind on when listening to network messages. The default should work out of the box, and will ease up the discovery process. That is, if you use your own settings, automatic detection of your server by other servers might not work so well.

4.9.3 broadcast

[Command-line option] --broadcast=<value> [Environment variable] LW6_BROADCAST

[XML key] broadcast

Type: boolean Default value: true Allows the program to send broadcast messages on the network. It can be usefull to disable those if you don’t use UDP node discovery and/or if there’s a sysadmin arround who does not enjoy permanent broadcasts on his LAN.

4.9.4 cli-backends

[Command-line option] --cli-backends=<value>

[Environment variable] LW6_CLI_BACKENDS

[XML key] cli-backends

Type: string Default value: tcp,udp,http The client backends to use. Most of the time the default is fine, change it only if you specifically want to disactivate some protocol, or if you want to activate a custommade client backend. It’s a comma separated list.

4.9.5 known-nodes

[Command-line option] --known-nodes=<value>

[Environment variable] LW6_KNOWN_NODES

[XML key] known-nodes

Type: string Default value: http://ufoot.org:8056/,http://ufoot.hd.free.fr:8056/ List of known nodes, nodes which the program will try to contact first to get the list of other nodes. This is mostly usefull when program is launched for the first time, after this it should keep an up-to-date list of known servers in its internal database and automatically reconnect to them next time it starts. You might want to change this if you really want to connect to a given server which is not publically listed. The list is comma separated.

4.9.6 node-description

[Command-line option] --node-description=<value>

[Environment variable] LW6_NODE_DESCRIPTION

[XML key] node-description

Type: string Default value: No description. The description of your node, that is a text that describes your server. This will typically appear when pointing a web client on the public server URL, it is for general information, so if there’s something special about your server, say it here.

4.9.7 node-title

[Command-line option] --node-title=<value>

[Environment variable] LW6_NODE_TITLE

[XML key] node-title

Type: string Default value: The title of your node, that is the name which will be displayed when listing servers. This is different from player name, for there can be several players on a single computer. By default this will be set to hostname.

4.9.8 password

[Command-line option] --password=<value> [Environment variable] LW6_PASSWORD

[XML key] password

Type: string

Default value:

The password to use for network games. Do not use a valuable password, as this is stored as clear text on your hard drive. Still, the game will only send a hash/checksum of the password on the network so eavesdropper won’t be able to read it. They can see the hash/checksum and use it if clever, but they can’t guess the real password. A blank password means anyone can join your games when you act like a server.

4.9.9 public-url

[Command-line option] --public-url=<value>

[Environment variable] LW6_PUBLIC_URL

[XML key] public-url

Type: string

Default value:

The public URL of your server. By default the game will pick up one for you. In fact, the clients discovering your server should guess the public URL, probably http://<your-ip>:<your-port>/ but you might need to use your own settings if you are using NAT or an Apache reverse-proxy to rewrite HTTP requests.

4.9.10 skip-network

[Command-line option] --skip-network=<value>

[Environment variable] LW6_SKIP_NETWORK

[XML key] skip-network

Type: boolean

Default value: false

If set, then game won’t do anything network related. No listen, no connect, no nothing. You are playing locally.

4.9.11 srv-backends

[Command-line option] --srv-backends=<value>

[Environment variable] LW6_SRV_BACKENDS

[XML key] srv-backends

Type: string

Default value: tcpd,udpd,httpd

The server backends to use. Most of the time the default is fine, change it only if you specifically want to disactivate some protocol, or if you want to activate a custom-made server backend. It’s a comma separated list.

4.10 Map parameters

4.10.1 chosen-map

[Command-line option] --chosen-map=<value>

[Environment variable] LW6_CHOSEN_MAP

[XML key] chosen-map

Type: string

Default value: subflower

The last map chosen by the player, locally. This is the map which will be used for a quick-start game, a local game, or a game started as a server.

4.10.2 force

[Command-line option] --force=<value> [Environment variable] LW6_FORCE [XML key] force

Type: string

Default value: respawn-team,color-conflict-mode

A comma separated list of options which should be ignored when reading map XML files. For instance, if this contains ’rounds-per-sec,moves-per-round’ then whatever values were defined for this in ’rules.xml’, then game will ignore them and use the user’s values, stored in ’config.xml’, running the game at the requested speed. This ultimately allows the player to control everything despite the values set by the map designer.

4.10.3 use-cursor-texture

[Command-line option] --use-cursor-texture=<value>
[Environment variable] LW6_USE_CURSOR_TEXTURE
[XML key] use-cursor-texture

Type: boolean

Default value: true

Defines wether the cursor textures should be used. If unset, then the default builtin cursor texture will be used instead of the map specific one.

4.10.4 use-hints-xml

[Command-line option] --use-hints-xml=<value>

[Environment variable] LW6_USE_HINTS_XML

[XML key] use-hints-xml

Type: boolean

Default value: true

If set, then hints will be picked up from the map defined hints.xml, if it exists. This is the default.

4.10.5 use-music-file

[Command-line option] --use-music-file=<value>

[Environment variable] LW6_USE_MUSIC_FILE

[XML key] use-music-file

Type: boolean

Default value: true

If set, then the program will use the ’music-file’ attribute to choose the music to play. If unset, then a random builtin music will be picked up, regardless of what is specified in ’music-file’.

4.10.6 use-rules-xml

[Command-line option] --use-rules-xml=<value>

[Environment variable] LW6_USE_RULES_XML

[XML key] use-rules-xml

Type: boolean

Default value: true

If set, then rules will be picked up from the map defined rules.xml, if it exists. This is the default. Use force-time and force-size to override this and use user-defined values anyway.

4.10.7 use-style-xml

[Command-line option] --use-style-xml=<value>

[Environment variable] LW6_USE_STYLE_XML

[XML key] use-style-xml

Type: boolean

Default value: true

If set, then style will be picked up from the map defined style.xml, if it exists. This is the default. Use force-time and force-background to override this and use user-defined values anyway.

4.10.8 use-teams-xml

[Command-line option] --use-teams-xml=<value>

[Environment variable] LW6_USE_TEAMS_XML

[XML key] use-teams-xml

Type: boolean

Default value: true

If set, then teams will be picked up from the map defined teams.xml, if it exists. This is the default. Use force-time and force-background to override this and use user-defined values anyway.

4.10.9 use-texture

[Command-line option] --use-texture=<value>

[Environment variable] LW6_USE_TEXTURE

[XML key] use-texture

Type: boolean Default value: true Defines wether the map texture should be used. Of course if there’s no map texture, the texture... won’t be used. But if there is one, this parameter will force the game to ignore it and play with solid colors. This probably won’t look as nice as the textured map in most cases, but some players might find it more readable and confortable to play when throwing eye candy away.

4.11 Map rules.xml

4.11.1 boost-power

[Command-line option] --boost-power=<value>

[Environment variable] LW6_BOOST_POWER

[XML key] boost-power

Type: integer Default value: 3 Min value: 1 Max value: 10 Defines how fast and powerfull the boost is. That is, if on ’boost.png’ it’s pitch black and this parameter is set to 3, then fighters will move and act 3 times than what they would do normally.

4.11.2 color-conflict-mode

[Command-line option] --color-conflict-mode=<value>
[Environment variable] LW6_COLOR_CONFLICT_MODE
[XML key] color-conflict-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 How to handle color conflicts, that is, when a player requests a color, but this color is already used, what should be done? If 0, wether a color already exists won’t affect the color of a new cursor. If 1, then two players on the same computer will be allowed to share the same color/team, but if another computer is already playing with a color, any new computer will need to use another team. If 2, then it’s impossible for a new cursor to use a pre-existing color, any new cursor will require a new color, if that color is already used, a new color will be picked randomly.

4.11.3 cursor-pot-init

[Command-line option] --cursor-pot-init=<value>

[Environment variable] LW6_CURSOR_POT_INIT

[XML key] cursor-pot-init

Type: integer Default value: 100000 Min value: 5000 Max value: 500000 Defines the cursor potential at startup. Not really any reason to change it. Theorically, there could be maps where the default value doesn’t fit, but none has been seen yet.

4.11.4 danger-power

[Command-line option] --danger-power=<value>

[Environment variable] LW6_DANGER_POWER

[XML key] danger-power

Type: integer Default value: 200 Min value: 0 Max value: 10000 Defines how dangerous are the black zones defined in ’danger.png’. The value is used to decrease the fighter health at each move, so you should compare its value to something like ’fighter-attack’. Being on a dangerous zone is a bit like being attacked by an invisible and unknown ennemy.

4.11.5 exp

[Command-line option] --exp=<value> [Environment variable] LW6_EXP [XML key] exp

Type: integer Default value: 1 Min value: 0 Max value: 99 Level of experience (AKA exp) required to play the current level. If this level is validated (that is, won) then player will be granted with a level of exp+1 and be able to play all the next levels. An exp of 0 means the level is playable by a pure beginner.

4.11.6 fighter-attack

[Command-line option] --fighter-attack=<value>

[Environment variable] LW6_FIGHTER_ATTACK

[XML key] fighter-attack

Type: integer Default value: 500 Min value: 1 Max value: 10000 Defines how hard fighters will attack others, that is, in one attack, how many lifepoints the attacked fighter will loose. Increasing this will cause your opponents to melt faster when you attack them. With a low value, it will take ages to take on your opponents. Different styles of game. Can radically change the gameplay.

4.11.7 fighter-defense

[Command-line option] --fighter-defense=<value>

[Environment variable] LW6_FIGHTER_DEFENSE

[XML key] fighter-defense

Type: integer Default value: 50 Min value: 0 Max value: 10000 Defines how fast fighters will regenerate after an attack. When this parameter is set low, an attacked fighter, which is very dark and almost dead will take a very long time to regain energy. If the parameter is set high, it can almost instantaneously regain energy.

4.11.8 fighter-new-health

[Command-line option] --fighter-new-health=<value>
[Environment variable] LW6_FIGHTER_NEW_HEALTH
[XML key] fighter-new-health

Type: integer Default value: 5000 Min value: 1 Max value: 10000 Defines how healthy fighters will be when they appear on the map. This can be either at the beginning of the game of when a fighter changes team. Setting this low will allow battefields to switch from one side to another very fast, for freshly gained fighters will be feeble and very likely to return to their original camp. To calibrate this parameter, keep in mind that the absolute maximum health a fighter can have is always 10000 (ten-thousands).

4.11.9 fighter-regenerate

[Command-line option] --fighter-regenerate=<value>
[Environment variable] LW6_FIGHTER_REGENERATE
[XML key] fighter-regenerate

Type: integer Default value: 5 Min value: 0 Max value: 10000 Defines at which speed fighters will self-regenerate, without even begin packed together. This will allow lone fighters to regenerate a bit by hiding somewhere in the map. This is typically a low value, might even be 0.

4.11.10 frags-fade-out

[Command-line option] --frags-fade-out=<value>

[Environment variable] LW6_FRAGS_FADE_OUT

[XML key] frags-fade-out

Type: integer Default value: 100 Min value: 10 Max value: 100 When a player looses (in deathmatch mode) all player points will be multiplicated by this percentage, for instance if it’s 90 and player had 50 points, then player will only have 45 points, then points corresponding to the new death will be added/substrated to its total. This is to avoid players with thousands of points in advance, and keep everyone in the race. A low value will minimize the importance of game start. This is only used in modes where frags are distributed in a proportional way.

4.11.11 frags-mode

[Command-line option] --frags-mode=<value>

[Environment variable] LW6_FRAGS_MODE

[XML key] frags-mode

Type: integer

Default value: 2 Min value: 0 Max value: 3

Defines how points are calculated in deathmatch mode, 0 is old school simple mode. 1 is in a mode in which 1 point is attributed to every winner, and looser looses all the corresponding points (total is always 0). 2 isproportional mode, with a total of 0 kept constant, that is, loosers loose as many points as attributed to winners. 3 is a mode in which at each death, winners are attributed a number of points proportional to their fighters, and loosers scores remain untouched.

4.11.12 frags-to-distribute

[Command-line option] --frags-to-distribute=<value>
[Environment variable] LW6_FRAGS_TO_DISTRIBUTE
[XML key] frags-to-distribute

Type: integer

Default value: 100 Min value: 10 Max value: 1000

Defines how many points will be distributed when in deathmatch mode. When a player looses, this amont of points will be substracted to its total, and the same amount of points will be distributed to other live players, proportionnally to how many fighters they have on the battlefield.

4.11.13 glue-power

[Command-line option] --glue-power=<value>

[Environment variable] LW6_GLUE_POWER

[XML key] glue-power

Type: integer

Default value: 20 Min value: 1 Max value: 100

Defines how sticky and powerfull the glue is. That is, if on ’glue.png’ it’s pitch black and this parameter is set to 3, then fighters will take 3 steps to do what would normally take only one step.

4.11.14 highest-team-color-allowed

[Command-line option] --highest-team-color-allowed=<value>
[Environment variable] LW6_HIGHEST_TEAM_COLOR_ALLOWED
[XML key] highest-team-color-allowed

Type: integer

Default value: 9 Min value: 3 Max value: 9

Id of the greatest/highest color one can use. Normally, you can leave this untouched, the program will automatically fit this according to your exp. Setting an artificially low value will just cause normally available colors to disappear, setting it to a high value does nothing, if you still don’t have access to some colors, you still don’t, period.

4.11.15 highest-weapon-allowed

[Command-line option] --highest-weapon-allowed=<value>
[Environment variable] LW6_HIGHEST_WEAPON_ALLOWED
[XML key] highest-weapon-allowed

Type: integer Default value: 19 Min value: 7 Max value: 19 Id of the greatest/highest weapon one can use. Normally, you can leave this untouched, the program will automatically fit this according to your exp. Setting an artificially low value will just cause normally available weapons to disappear, setting it to a high value does nothing, if you still don’t have access to some weapons, you still don’t, period.

4.11.16 max-cursor-pot

[Command-line option] --max-cursor-pot=<value>

[Environment variable] LW6_MAX_CURSOR_POT

[XML key] max-cursor-pot

Type: integer Default value: 1000000 Min value: 50000 Max value: 5000000 Defines the maximum cursor potential. Not really any reason to change it. Any high value should produce the same results. Low values might reveal algorithm bugs and inconsistencies.

4.11.17 max-cursor-pot-offset

[Command-line option] --max-cursor-pot-offset=<value>
[Environment variable] LW6_MAX_CURSOR_POT_OFFSET
[XML key] max-cursor-pot-offset

Type: integer Default value: 100 Min value: 1 Max value: 10000 Defines the maximum cursor potential offset. The idea is that in some cases, the potential of a cursor can increase in burst mode, for instance to make this cursor more important than others, so that fighters rally to it, neglecting other cursors (talking about a multi-cursor controlled team). This parameter is here to limit this burst effect and avoid bugs.

4.11.18 max-nb-cursors

[Command-line option] --max-nb-cursors=<value>

[Environment variable] LW6_MAX_NB_CURSORS

[XML key] max-nb-cursors

Type: integer Default value: 26 Min value: 2 Max value: 26 Defines the maximum number of cursors who can enter the game. Really makes sense in network games. Default value is 26, the maximum.

4.11.19 max-nb-nodes

[Command-line option] --max-nb-nodes=<value>

[Environment variable] LW6_MAX_NB_NODES

[XML key] max-nb-nodes

Type: integer Default value: 12 Min value: 2 Max value: 15 Defines the maximum number of servers who can enter the game. Really makes sense in network games. Default value is 10, and should fit in most cases. Can be raised up to 26.

4.11.20 max-nb-teams

[Command-line option] --max-nb-teams=<value>

[Environment variable] LW6_MAX_NB_TEAMS

[XML key] max-nb-teams

Type: integer Default value: 10 Min value: 2 Max value: 10 Defines the maximum number of teams who can enter the game. Really makes sense in network games. Default value is 10, the maximum.

4.11.21 max-round-delta

[Command-line option] --max-round-delta=<value>

[Environment variable] LW6_MAX_ROUND_DELTA

[XML key] max-round-delta

Type: integer Default value: 1000 Min value: 1 Max value: 10000 This is the companion value of ’round-delta’. Will put an absolute limit to the delta, which (what did you think?) is of course incremented in some cases by the core algorithm. If in doubt, don’t touch.

4.11.22 max-zone-size

[Command-line option] --max-zone-size=<value>

[Environment variable] LW6_MAX_ZONE_SIZE

[XML key] max-zone-size

Type: integer Default value: 8 Min value: 1 Max value: 64 Defines the maximum zone size, which is an internal and rather technical parameter. The idea is that to optimize things, Liquid War 6 divides the battlefield in squares, where it can, and tries to make these squares as big as possible, the idea being that everywhere in this square, fighters follow the same intructions. Just a technical optimization. The problem is that setting it too high will reveal the optimization and its tradeoffs to the player, who will see the fighter behave strangely, following invisible paths. Plus, it’s ugly. Depending on your tastes (speed, look’n’feel) you’ll prefer something nice or something fast. Note that anyways passed a certain value, this does not optimize anything anymore. In doubt, don’t touch it.

4.11.23 medicine-power

[Command-line option] --medicine-power=<value>

[Environment variable] LW6_MEDICINE_POWER

[XML key] medicine-power

Type: integer Default value: 100 Min value: 0 Max value: 10000 Defines how fast fighter will automatically regenerate on black zones defined in ’medicine.png’. The value is used to decrease the fighter health at each move, so you should compare its value to something like ’fighter-defense’. Being on a medicined zone is a bit like being defended by an invisible and unknown friend.

4.11.24 moves-per-round

[Command-line option] --moves-per-round=<value>

[Environment variable] LW6_MOVES_PER_ROUND

[XML key] moves-per-round

Type: integer Default value: 2 Min value: 1 Max value: 50 Defines how many times fighters move per round. Increasing this will just make fighters move faster, but won’t change anything for the rest, that is keyboard and mouse responsivity, and network traffic will stay the same. Multiplying the number of moves per round by the number of rounds per second will give the number of moves per second, which is, in fact, how fast fighters move on the screen.

4.11.25 nb-attack-tries

[Command-line option] --nb-attack-tries=<value>

[Environment variable] LW6_NB_ATTACK_TRIES

[XML key] nb-attack-tries

Type: integer Default value: 3 Min value: 1 Max value: 7 Defines how many tries a fighter will do before giving-up attacking and choosing another behvior (defense). By tries we mean: how many directions it will try. Going North? Going North-West? Setting this to a low value will make fighters somewhat less aggressive. This idea is that they’ll prefer to switch to the next option, that is, defense/regeneration, if there’s no opponent right in front of them.

4.11.26 nb-defense-tries

[Command-line option] --nb-defense-tries=<value>

[Environment variable] LW6_NB_DEFENSE_TRIES

[XML key] nb-defense-tries

Type: integer Default value: 1 Min value: 1 Max value: 7 Defines how many tries a fighter will do before giving-up attacking and choosing another behavior (do nothing). By tries we mean: how many directions it will try.

Going North? Going North-West? Setting this to a low value, you’ll need a very compact pack of fighters for regeneration to operate, else fighters will hang arround unhealthy.

4.11.27 nb-move-tries

[Command-line option] --nb-move-tries=<value>

[Environment variable] LW6_NB_MOVE_TRIES

[XML key] nb-move-tries

Type: integer Default value: 5 Min value: 3 Max value: 7 Defines how many tries a fighter will do before giving-up moving and choosing another behvior (attack or defense). By tries we mean: how many directions it will try. Going North? Going North-West? Setting this to a low value, your fighters will look very stubborn and always try to move in one direction, neglecting the fact that they could dodge. This can lead to queues of fighters and other strange behaviors. On the other hand, setting it too high will cause fighter to always avoid the enemy, and groups of fighters will just pass each other without any fight. Matter of taste.

4.11.28 respawn-delay

[Command-line option] --respawn-delay=<value>

[Environment variable] LW6_RESPAWN_DELAY

[XML key] respawn-delay

Type: integer Default value: 3 Min value: 0 Max value: 30 Delay, in seconds, after which teams reappear on the battlefield, when in deathmatch mode. 0 means team right away.

4.11.29 respawn-position-mode

[Command-line option] --respawn-position-mode=<value>
[Environment variable] LW6_RESPAWN_POSITION_MODE
[XML key] respawn-position-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Defines how teams are set up on the map when respawning. 0 means teams respect the pre-defined start positions. 1 means that a random position will be picked, among the existing positions. That is, red could take green’s place. 2 means total randomness, teams can appear anywhere.

4.11.30 respawn-team

[Command-line option] --respawn-team=<value>

[Environment variable] LW6_RESPAWN_TEAM

[XML key] respawn-team

Type: integer Default value: 1 Min value: 0 Max value: 1 Defines what to do when a team dies. If set to 0, team disappears forever, if set to 1, team reappears automatically with fresh fighters. It’s a deathmatch mode, where the winner is not the one who stays alive the longest time, since it makes no real sens in this case, but the one who has died less often than others.

4.11.31 round-delta

[Command-line option] --round-delta=<value>

[Environment variable] LW6_ROUND_DELTA

[XML key] round-delta

Type: integer Default value: 1 Min value: 0 Max value: 100 Conditions by how much the cursor potential will be incremented each time gradient is spreaded. Sounds cryptic? It is. The idea is that at each time you move your cursor of 1 pixel, theorically, you’ll need in the worst case to move of 1 more pixel to reach any point on the map. Of course this is not true but this is the default asumption, and gradient spread will fix that. Only in Liquid War 6 this is not even the worst case, for you can control your cursor with the mouse and cross walls. Whenever you cross a wall, you might have done a great distance from the fighters’ point of view, if the map is a maze. Thus this parameter, which corrects things, experience shows it does give acceptable results to increase the cursor potential by more than one at each turn. Toy arround with this if you find fighters take wrong paths on some given map. If in doubt, don’t touch.

4.11.32 rounds-per-sec

[Command-line option] --rounds-per-sec=<value>

[Environment variable] LW6_ROUNDS_PER_SEC

[XML key] rounds-per-sec

Type: integer Default value: 50 Min value: 1 Max value: 200 Defines the overall speed of the game. All other settings being equal, raising this value will cause the game to behave faster. Everything will be faster, except probably the display since your computer will calculate more game positions in a given time and spend more CPU time. It will also increase network traffic. Values between 10 and 50 really make sense.

4.11.33 side-attack-factor

[Command-line option] --side-attack-factor=<value>
[Environment variable] LW6_SIDE_ATTACK_FACTOR
[XML key] side-attack-factor

Type: integer Default value: 20 Min value: 0 Max value: 100 Defines how hard fighters will attack sideways. It’s an algorithm trick, fighters attack by default the opponent right in front, but if there’s no fighter there, they will still try to attack someone else, maybe sideways. But doing this their attack is not as strong. This parameter enables you to tune this. This is a percentage.

4.11.34 side-defense-factor

[Command-line option] --side-defense-factor=<value>
[Environment variable] LW6_SIDE_DEFENSE_FACTOR
[XML key] side-defense-factor

Type: integer Default value: 20 Min value: 0 Max value: 100 Defines how fast fighters will regenerate, when being side by side instead of being right in front of the other. This is a percentage.

4.11.35 single-army-size

[Command-line option] --single-army-size=<value>

[Environment variable] LW6_SINGLE_ARMY_SIZE

[XML key] single-army-size

Type: integer Default value: 30 Min value: 1 Max value: 95 Defines the proportion of the whole available space, which will be occupied by an army at the beginning of the game. You can either imagine playing with almost empty maps, or play very crowded with almost no space left. This is a percentage, but will be multiplied by itself to get the actual surface. That is, 50 means 50%*50%, that is, a square of 1/2 the size of a square map, so it represents 25% (1/4) of the total surface.

4.11.36 spread-mode

[Command-line option] --spread-mode=<value>

[Environment variable] LW6_SPREAD_MODE

[XML key] spread-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 If set to 1, then gradient spread will be slower but gain in terms of homogeneity and consistency. You could consider setting this to 0 on very very big maps to save CPU cycles, else the default should work fine.

4.11.37 spread-thread

[Command-line option] --spread-thread=<value>

[Environment variable] LW6_SPREAD_THREAD

[XML key] spread-thread

Type: integer Default value: 0 Min value: 0 Max value: 1 If set to 1, the core algorithm with fire a separate thread to spread the gradient. By default this is turned off (set to 0). Consider this as an experimental feature, the program is already rather heavily threaded, turning this on will probably not offer any significant performance gain, even on SMP systems. This might change in the future.

4.11.38 spreads-per-round

[Command-line option] --spreads-per-round=<value>
[Environment variable] LW6_SPREADS_PER_ROUND
[XML key] spreads-per-round

Type: integer Default value: 5 Min value: 1 Max value: 100 Defines how many times the gradient is spread per round. Gradient spread is a very Liquid War 6 specific feature, just remember that the more often you do it, the more accurately fighters will move. That is, you will be sure they really take the shortest path. Usually this does not have much effect, the default value should fit in most cases, but you might want to decrease it on very simple maps where the gradient is obvious, or increase it on complex maps where you want fighters to be real smart.

4.11.39 start-blue-x

[Command-line option] --start-blue-x=<value>

[Environment variable] LW6_START_BLUE_X

[XML key] start-blue-x

Type: integer Default value: 90 Min value: 0 Max value: 100 X start position for the blue team. This is a percentage of map width, value between 0 and 100.

4.11.40 start-blue-y

[Command-line option] --start-blue-y=<value>

[Environment variable] LW6_START_BLUE_Y

[XML key] start-blue-y

Type: integer Default value: 10 Min value: 0 Max value: 100 Y start position for the blue team. This is a percentage of map height, value between 0 and 100.

4.11.41 start-cyan-x

[Command-line option] --start-cyan-x=<value>

[Environment variable] LW6_START_CYAN_X

[XML key] start-cyan-x

Type: integer Default value: 35 Min value: 0 Max value: 100 X start position for the cyan team. This is a percentage of map width, value between 0 and 100.

4.11.42 start-cyan-y

[Command-line option] --start-cyan-y=<value>

[Environment variable] LW6_START_CYAN_Y

[XML key] start-cyan-y

Type: integer Default value: 10 Min value: 0 Max value: 100 Y start position for the cyan team. This is a percentage of map height, value between 0 and 100.

4.11.43 start-green-x

[Command-line option] --start-green-x=<value>

[Environment variable] LW6_START_GREEN_X

[XML key] start-green-x

Type: integer Default value: 90 Min value: 0 Max value: 100 X start position for the green team. This is a percentage of map width, value between 0 and 100.

4.11.44 start-green-y

[Command-line option] --start-green-y=<value>

[Environment variable] LW6_START_GREEN_Y

[XML key] start-green-y

Type: integer Default value: 90 Min value: 0 Max value: 100 Y start position for the green team. This is a percentage of map height, value between 0 and 100.

4.11.45 start-lightblue-x

[Command-line option] --start-lightblue-x=<value>
[Environment variable] LW6_START_LIGHTBLUE_X
[XML key] start-lightblue-x

Type: integer Default value: 35 Min value: 0 Max value: 100 X start position for the lightblue team. This is a percentage of map width, value between 0 and 100.

4.11.46 start-lightblue-y

[Command-line option] --start-lightblue-y=<value>
[Environment variable] LW6_START_LIGHTBLUE_Y
[XML key] start-lightblue-y

Type: integer Default value: 90 Min value: 0 Max value: 100 Y start position for the lightblue team. This is a percentage of map height, value between 0 and 100.

4.11.47 start-magenta-x

[Command-line option] --start-magenta-x=<value>

[Environment variable] LW6_START_MAGENTA_X

[XML key] start-magenta-x

Type: integer Default value: 65 Min value: 0 Max value: 100 X start position for the magenta team. This is a percentage of map width, value between 0 and 100.

4.11.48 start-magenta-y

[Command-line option] --start-magenta-y=<value>

[Environment variable] LW6_START_MAGENTA_Y

[XML key] start-magenta-y

Type: integer Default value: 90 Min value: 0 Max value: 100 Y start position for the magenta team. This is a percentage of map height, value between 0 and 100.

4.11.49 start-orange-x

[Command-line option] --start-orange-x=<value>

[Environment variable] LW6_START_ORANGE_X

[XML key] start-orange-x

Type: integer Default value: 65 Min value: 0 Max value: 100 X start position for the orange team. This is a percentage of map width, value between 0 and 100.

4.11.50 start-orange-y

[Command-line option] --start-orange-y=<value>

[Environment variable] LW6_START_ORANGE_Y

[XML key] start-orange-y

Type: integer Default value: 10 Min value: 0 Max value: 100 Y start position for the orange team. This is a percentage of map height, value between 0 and 100.

4.11.51 start-pink-x

[Command-line option] --start-pink-x=<value>

[Environment variable] LW6_START_PINK_X

[XML key] start-pink-x

Type: integer Default value: 10 Min value: 0 Max value: 100 X start position for the pink team. This is a percentage of map width, value between 0 and 100.

4.11.52 start-pink-y

[Command-line option] --start-pink-y=<value>

[Environment variable] LW6_START_PINK_Y

[XML key] start-pink-y

Type: integer Default value: 50 Min value: 0 Max value: 100 Y start position for the pink team. This is a percentage of map height, value between 0 and 100.

4.11.53 start-position-mode

[Command-line option] --start-position-mode=<value>
[Environment variable] LW6_START_POSITION_MODE
[XML key] start-position-mode

Type: integer Default value: 0 Min value: 0 Max value: 2 Defines how teams are set up on the map at game startup. 0 means teams respect the pre-defined start positions. 1 means that a random position will be picked, among the existing positions. That is, red could take green’s place. 2 means total randomness, teams can appear anywhere.

4.11.54 start-purple-x

[Command-line option] --start-purple-x=<value>

[Environment variable] LW6_START_PURPLE_X

[XML key] start-purple-x

Type: integer Default value: 90 Min value: 0 Max value: 100 X start position for the purple team. This is a percentage of map width, value between 0 and 100.

4.11.55 start-purple-y

[Command-line option] --start-purple-y=<value>

[Environment variable] LW6_START_PURPLE_Y

[XML key] start-purple-y

Type: integer Default value: 50 Min value: 0 Max value: 100 Y start position for the purple team. This is a percentage of map height, value between 0 and 100.

4.11.56 start-red-x

[Command-line option] --start-red-x=<value>

[Environment variable] LW6_START_RED_X

[XML key] start-red-x

Type: integer Default value: 10 Min value: 0 Max value: 100 X start position for the red team. This is a percentage of map width, value between 0 and 100.

4.11.57 start-red-y

[Command-line option] --start-red-y=<value>

[Environment variable] LW6_START_RED_Y

[XML key] start-red-y

Type: integer Default value: 10 Min value: 0 Max value: 100 Y start position for the red team. This is a percentage of map height, value between 0 and 100.

4.11.58 start-yellow-x

[Command-line option] --start-yellow-x=<value>

[Environment variable] LW6_START_YELLOW_X

[XML key] start-yellow-x

Type: integer Default value: 10 Min value: 0 Max value: 100 X start position for the yellow team. This is a percentage of map width, value between 0 and 100.

4.11.59 start-yellow-y

[Command-line option] --start-yellow-y=<value>

[Environment variable] LW6_START_YELLOW_Y

[XML key] start-yellow-y

Type: integer Default value: 90 Min value: 0 Max value: 100 Y start position for the yellow team. This is a percentage of map height, value between 0 and 100.

4.11.60 team-profile-blue-aggressive

[Command-line option] --team-profile-blue-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_AGGRESSIVE
[XML key] team-profile-blue-aggressive

Type: integer Default value: 150 Min value: 5 Max value: 2000 Defines how aggressive the blue team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.61 team-profile-blue-fast

[Command-line option] --team-profile-blue-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_FAST
[XML key] team-profile-blue-fast

Type: integer Default value: 50 Min value: 5 Max value: 2000 Changes the speed of the blue team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.62 team-profile-blue-handicap

[Command-line option] --team-profile-blue-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_HANDICAP
[XML key] team-profile-blue-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the blue team.

4.11.63 team-profile-blue-mobile

[Command-line option] --team-profile-blue-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_MOBILE
[XML key] team-profile-blue-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the blue team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.64 team-profile-blue-vulnerable

[Command-line option] --team-profile-blue-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_VULNERABLE
[XML key] team-profile-blue-vulnerable

Type: integer Default value: 60 Min value: 5 Max value: 2000 Defines how vulnerable the blue team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.65 team-profile-blue-weapon-alternate-id

[Command-line option] --team-profile-blue-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_WEAPON_ALTERNATE_ID
[XML key] team-profile-blue-weapon-alternate-id

Type: integer

Default value: 8 Min value: 0 Max value: 19

Id of the default alternate weapon for the blue team, see the documentation about weapons to know what these ids mean.

4.11.66 team-profile-blue-weapon-id

[Command-line option] --team-profile-blue-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_WEAPON_ID
[XML key] team-profile-blue-weapon-id

Type: integer

Default value: 14 Min value: 0 Max value: 19

Id of the default weapon for the blue team, see the documentation about weapons to know what these ids mean.

4.11.67 team-profile-blue-weapon-mode

[Command-line option] --team-profile-blue-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_BLUE_WEAPON_MODE
[XML key] team-profile-blue-weapon-mode

Type: integer

Default value: 1 Min value: 0 Max value: 2

Weapon mode for blue team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.68 team-profile-cyan-aggressive

[Command-line option] --team-profile-cyan-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_AGGRESSIVE
[XML key] team-profile-cyan-aggressive

Type: integer

Default value: 44 Min value: 5 Max value: 2000

Defines how aggressive the cyan team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.69 team-profile-cyan-fast

[Command-line option] --team-profile-cyan-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_FAST
[XML key] team-profile-cyan-fast

Type: integer Default value: 40 Min value: 5 Max value: 2000 Changes the speed of the cyan team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.70 team-profile-cyan-handicap

[Command-line option] --team-profile-cyan-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_HANDICAP
[XML key] team-profile-cyan-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the cyan team.

4.11.71 team-profile-cyan-mobile

[Command-line option] --team-profile-cyan-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_MOBILE
[XML key] team-profile-cyan-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the cyan team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.72 team-profile-cyan-vulnerable

[Command-line option] --team-profile-cyan-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_VULNERABLE
[XML key] team-profile-cyan-vulnerable

Type: integer Default value: 12 Min value: 5 Max value: 2000 Defines how vulnerable the cyan team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.73 team-profile-cyan-weapon-alternate-id

[Command-line option] --team-profile-cyan-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_WEAPON_ALTERNATE_ID
[XML key] team-profile-cyan-weapon-alternate-id

Type: integer Default value: 12 Min value: 0 Max value: 19 Id of the default alternate weapon for the cyan team, see the documentation about weapons to know what these ids mean.

4.11.74 team-profile-cyan-weapon-id

[Command-line option] --team-profile-cyan-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_WEAPON_ID
[XML key] team-profile-cyan-weapon-id

Type: integer Default value: 3 Min value: 0 Max value: 19 Id of the default weapon for the cyan team, see the documentation about weapons to know what these ids mean.

4.11.75 team-profile-cyan-weapon-mode

[Command-line option] --team-profile-cyan-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_CYAN_WEAPON_MODE
[XML key] team-profile-cyan-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for cyan team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.76 team-profile-green-aggressive

[Command-line option] --team-profile-green-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_AGGRESSIVE
[XML key] team-profile-green-aggressive

Type: integer Default value: 70 Min value: 5 Max value: 2000 Defines how aggressive the green team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.77 team-profile-green-fast

[Command-line option] --team-profile-green-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_FAST
[XML key] team-profile-green-fast

Type: integer Default value: 70 Min value: 5 Max value: 2000 Changes the speed of the green team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.78 team-profile-green-handicap

[Command-line option] --team-profile-green-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_HANDICAP

[XML key] team-profile-green-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the green team.

4.11.79 team-profile-green-mobile

[Command-line option] --team-profile-green-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_MOBILE
[XML key] team-profile-green-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the green team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.80 team-profile-green-vulnerable

[Command-line option] --team-profile-green-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_VULNERABLE
[XML key] team-profile-green-vulnerable

Type: integer Default value: 30 Min value: 5 Max value: 2000 Defines how vulnerable the green team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.81 team-profile-green-weapon-alternate-id

[Command-line option] --team-profile-green-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_WEAPON_ALTERNATE_ID
[XML key] team-profile-green-weapon-alternate-id

Type: integer Default value: 7 Min value: 0 Max value: 19 Id of the default alternate weapon for the green team, see the documentation about weapons to know what these ids mean.

4.11.82 team-profile-green-weapon-id

[Command-line option] --team-profile-green-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_WEAPON_ID
[XML key] team-profile-green-weapon-id

Type: integer Default value: 13 Min value: 0 Max value: 19 Id of the default weapon for the green team, see the documentation about weapons to know what these ids mean.

4.11.83 team-profile-green-weapon-mode

[Command-line option] --team-profile-green-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_GREEN_WEAPON_MODE
[XML key] team-profile-green-weapon-mode

Type: integer

Default value: 1 Min value: 0 Max value: 2

Weapon mode for green team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.84 team-profile-lightblue-aggressive

[Command-line option] --team-profile-lightblue-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_AGGRESSIVE
[XML key] team-profile-lightblue-aggressive

Type: integer

Default value: 200 Min value: 5 Max value: 2000

Defines how aggressive the lightblue team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.85 team-profile-lightblue-fast

[Command-line option] --team-profile-lightblue-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_FAST
[XML key] team-profile-lightblue-fast

Type: integer

Default value: 20 Min value: 5 Max value: 2000

Changes the speed of the lightblue team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.86 team-profile-lightblue-handicap

[Command-line option] --team-profile-lightblue-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_HANDICAP
[XML key] team-profile-lightblue-handicap

Type: integer

Default value: 100 Min value: 10 Max value: 1000

Defines the handicap for the lightblue team.

4.11.87 team-profile-lightblue-mobile

[Command-line option] --team-profile-lightblue-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_MOBILE
[XML key] team-profile-lightblue-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the lightblue team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.88 team-profile-lightblue-vulnerable

[Command-line option] --team-profile-lightblue-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_VULNERABLE
[XML key] team-profile-lightblue-vulnerable

Type: integer Default value: 8 Min value: 5 Max value: 2000 Defines how vulnerable the lightblue team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.89 team-profile-lightblue-weapon-alternate-id

[Command-line option] --team-profile-lightblue-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_ALTERNATE_ID
[XML key] team-profile-lightblue-weapon-alternate-id

Type: integer Default value: 17 Min value: 0 Max value: 19 Id of the default alternate weapon for the lightblue team, see the documentation about weapons to know what these ids mean.

4.11.90 team-profile-lightblue-weapon-id

[Command-line option] --team-profile-lightblue-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_ID
[XML key] team-profile-lightblue-weapon-id

Type: integer Default value: 4 Min value: 0 Max value: 19 Id of the default weapon for the lightblue team, see the documentation about weapons to know what these ids mean.

4.11.91 team-profile-lightblue-weapon-mode

[Command-line option] --team-profile-lightblue-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_MODE
[XML key] team-profile-lightblue-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for lightblue team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.92 team-profile-magenta-aggressive

[Command-line option] --team-profile-magenta-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_AGGRESSIVE
[XML key] team-profile-magenta-aggressive

Type: integer Default value: 192 Min value: 5 Max value: 2000 Defines how aggressive the magenta team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.93 team-profile-magenta-fast

[Command-line option] --team-profile-magenta-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_FAST
[XML key] team-profile-magenta-fast

Type: integer Default value: 320 Min value: 5 Max value: 2000 Changes the speed of the magenta team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.94 team-profile-magenta-handicap

[Command-line option] --team-profile-magenta-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_HANDICAP
[XML key] team-profile-magenta-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the magenta team.

4.11.95 team-profile-magenta-mobile

[Command-line option] --team-profile-magenta-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_MOBILE
[XML key] team-profile-magenta-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the magenta team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.96 team-profile-magenta-vulnerable

[Command-line option] --team-profile-magenta-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_VULNERABLE

[XML key] team-profile-magenta-vulnerable

Type: integer Default value: 1920 Min value: 5 Max value: 2000 Defines how vulnerable the magenta team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.97 team-profile-magenta-weapon-alternate-id

[Command-line option] --team-profile-magenta-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_WEAPON_ALTERNATE_ID
[XML key] team-profile-magenta-weapon-alternate-id

Type: integer Default value: 15 Min value: 0 Max value: 19 Id of the default alternate weapon for the magenta team, see the documentation about weapons to know what these ids mean.

4.11.98 team-profile-magenta-weapon-id

[Command-line option] --team-profile-magenta-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_WEAPON_ID
[XML key] team-profile-magenta-weapon-id

Type: integer Default value: 6 Min value: 0 Max value: 19 Id of the default weapon for the magenta team, see the documentation about weapons to know what these ids mean.

4.11.99 team-profile-magenta-weapon-mode

[Command-line option] --team-profile-magenta-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_MAGENTA_WEAPON_MODE
[XML key] team-profile-magenta-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for magenta team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.100 team-profile-orange-aggressive

[Command-line option] --team-profile-orange-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_AGGRESSIVE
[XML key] team-profile-orange-aggressive

Type: integer Default value: 48 Min value: 5 Max value: 2000 Defines how aggressive the orange team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.101 team-profile-orange-fast

[Command-line option] --team-profile-orange-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_FAST
[XML key] team-profile-orange-fast

Type: integer Default value: 160 Min value: 5 Max value: 2000 Changes the speed of the orange team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.102 team-profile-orange-handicap

[Command-line option] --team-profile-orange-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_HANDICAP
[XML key] team-profile-orange-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the orange team.

4.11.103 team-profile-orange-mobile

[Command-line option] --team-profile-orange-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_MOBILE
[XML key] team-profile-orange-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the orange team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.104 team-profile-orange-vulnerable

[Command-line option] --team-profile-orange-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_VULNERABLE
[XML key] team-profile-orange-vulnerable

Type: integer Default value: 144 Min value: 5 Max value: 2000 Defines how vulnerable the orange team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.105 team-profile-orange-weapon-alternate-id

[Command-line option] --team-profile-orange-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_WEAPON_ALTERNATE_ID

[XML key] team-profile-orange-weapon-alternate-id

Type: integer Default value: 16 Min value: 0 Max value: 19 Id of the default alternate weapon for the orange team, see the documentation about weapons to know what these ids mean.

4.11.106 team-profile-orange-weapon-id

[Command-line option] --team-profile-orange-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_WEAPON_ID
[XML key] team-profile-orange-weapon-id

Type: integer Default value: 0 Min value: 0 Max value: 19 Id of the default weapon for the orange team, see the documentation about weapons to know what these ids mean.

4.11.107 team-profile-orange-weapon-mode

[Command-line option] --team-profile-orange-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_ORANGE_WEAPON_MODE
[XML key] team-profile-orange-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for orange team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.108 team-profile-pink-aggressive

[Command-line option] --team-profile-pink-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_AGGRESSIVE
[XML key] team-profile-pink-aggressive

Type: integer Default value: 640 Min value: 5 Max value: 2000 Defines how aggressive the pink team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.109 team-profile-pink-fast

[Command-line option] --team-profile-pink-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_FAST
[XML key] team-profile-pink-fast

Type: integer Default value: 80 Min value: 5 Max value: 2000 Changes the speed of the pink team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.110 team-profile-pink-handicap

[Command-line option] --team-profile-pink-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_HANDICAP
[XML key] team-profile-pink-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the pink team.

4.11.111 team-profile-pink-mobile

[Command-line option] --team-profile-pink-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_MOBILE
[XML key] team-profile-pink-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the pink team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.112 team-profile-pink-vulnerable

[Command-line option] --team-profile-pink-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_VULNERABLE
[XML key] team-profile-pink-vulnerable

Type: integer Default value: 640 Min value: 5 Max value: 2000 Defines how vulnerable the pink team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.113 team-profile-pink-weapon-alternate-id

[Command-line option] --team-profile-pink-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_WEAPON_ALTERNATE_ID
[XML key] team-profile-pink-weapon-alternate-id

Type: integer Default value: 19 Min value: 0 Max value: 19 Id of the default alternate weapon for the pink team, see the documentation about weapons to know what these ids mean.

4.11.114 team-profile-pink-weapon-id

[Command-line option] --team-profile-pink-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_WEAPON_ID

[XML key] team-profile-pink-weapon-id

Type: integer Default value: 1 Min value: 0 Max value: 19 Id of the default weapon for the pink team, see the documentation about weapons to know what these ids mean.

4.11.115 team-profile-pink-weapon-mode

[Command-line option] --team-profile-pink-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_PINK_WEAPON_MODE
[XML key] team-profile-pink-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for pink team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.116 team-profile-purple-aggressive

[Command-line option] --team-profile-purple-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_AGGRESSIVE
[XML key] team-profile-purple-aggressive

Type: integer Default value: 32 Min value: 5 Max value: 2000 Defines how aggressive the purple team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.117 team-profile-purple-fast

[Command-line option] --team-profile-purple-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_FAST
[XML key] team-profile-purple-fast

Type: integer Default value: 80 Min value: 5 Max value: 2000 Changes the speed of the purple team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.118 team-profile-purple-handicap

[Command-line option] --team-profile-purple-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_HANDICAP
[XML key] team-profile-purple-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the purple team.

4.11.119 team-profile-purple-mobile

[Command-line option] --team-profile-purple-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_MOBILE
[XML key] team-profile-purple-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the purple team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.120 team-profile-purple-vulnerable

[Command-line option] --team-profile-purple-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_VULNERABLE
[XML key] team-profile-purple-vulnerable

Type: integer Default value: 16 Min value: 5 Max value: 2000 Defines how vulnerable the purple team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.121 team-profile-purple-weapon-alternate-id

[Command-line option] --team-profile-purple-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_WEAPON_ALTERNATE_ID
[XML key] team-profile-purple-weapon-alternate-id

Type: integer Default value: 18 Min value: 0 Max value: 19 Id of the default alternate weapon for the purple team, see the documentation about weapons to know what these ids mean.

4.11.122 team-profile-purple-weapon-id

[Command-line option] --team-profile-purple-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_WEAPON_ID
[XML key] team-profile-purple-weapon-id

Type: integer Default value: 11 Min value: 0 Max value: 19 Id of the default weapon for the purple team, see the documentation about weapons to know what these ids mean.

4.11.123 team-profile-purple-weapon-mode

[Command-line option] --team-profile-purple-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_PURPLE_WEAPON_MODE

[XML key] team-profile-purple-weapon-mode

Type: integer Default value: 1 Min value: 0 Max value: 2 Weapon mode for purple team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.124 team-profile-red-aggressive

[Command-line option] --team-profile-red-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_AGGRESSIVE
[XML key] team-profile-red-aggressive

Type: integer Default value: 220 Min value: 5 Max value: 2000 Defines how aggressive the red team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.125 team-profile-red-fast

[Command-line option] --team-profile-red-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_FAST
[XML key] team-profile-red-fast

Type: integer Default value: 160 Min value: 5 Max value: 2000 Changes the speed of the red team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.126 team-profile-red-handicap

[Command-line option] --team-profile-red-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_HANDICAP
[XML key] team-profile-red-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the red team.

4.11.127 team-profile-red-mobile

[Command-line option] --team-profile-red-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_MOBILE
[XML key] team-profile-red-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the red team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.128 team-profile-red-vulnerable

[Command-line option] --team-profile-red-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_VULNERABLE
[XML key] team-profile-red-vulnerable

Type: integer

Default value: 110 Min value: 5 Max value: 2000

Defines how vulnerable the red team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.129 team-profile-red-weapon-alternate-id

[Command-line option] --team-profile-red-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_WEAPON_ALTERNATE_ID
[XML key] team-profile-red-weapon-alternate-id

Type: integer

Default value: 2 Min value: 0 Max value: 19

Id of the default alternate weapon for the red team, see the documentation about weapons to know what these ids mean.

4.11.130 team-profile-red-weapon-id

[Command-line option] --team-profile-red-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_WEAPON_ID
[XML key] team-profile-red-weapon-id

Type: integer

Default value: 10 Min value: 0 Max value: 19

Id of the default weapon for the red team, see the documentation about weapons to know what these ids mean.

4.11.131 team-profile-red-weapon-mode

[Command-line option] --team-profile-red-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_RED_WEAPON_MODE
[XML key] team-profile-red-weapon-mode

Type: integer

Default value: 1 Min value: 0 Max value: 2

Weapon mode for red team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.132 team-profile-yellow-aggressive

[Command-line option] --team-profile-yellow-aggressive=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_AGGRESSIVE
[XML key] team-profile-yellow-aggressive

Type: integer Default value: 200 Min value: 5 Max value: 2000 Defines how aggressive the yellow team is. This is a percentage, if set to 200 then team will attack twice as much as any other team with the default value. Setting this to a high value clearly advantages this team.

4.11.133 team-profile-yellow-fast

[Command-line option] --team-profile-yellow-fast=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_FAST
[XML key] team-profile-yellow-fast

Type: integer Default value: 70 Min value: 5 Max value: 2000 Changes the speed of the yellow team. This is a percentage, if set to 50, then team will move twice slower than other teams with the default parameter. Setting this high is very likely to advantage the team.

4.11.134 team-profile-yellow-handicap

[Command-line option] --team-profile-yellow-handicap=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_HANDICAP
[XML key] team-profile-yellow-handicap

Type: integer Default value: 100 Min value: 10 Max value: 1000 Defines the handicap for the yellow team.

4.11.135 team-profile-yellow-mobile

[Command-line option] --team-profile-yellow-mobile=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_MOBILE
[XML key] team-profile-yellow-mobile

Type: integer Default value: 0 Min value: -3 Max value: 3 Increases (or decreases if negative) the number of move/attack/defense tries for the yellow team. If set to a high value team will appear more mobile and do more things, but it won’t change its cruising speed. It’s not obvious to tell wether this is an advantage or not, but it clearly changes the behavior.

4.11.136 team-profile-yellow-vulnerable

[Command-line option] --team-profile-yellow-vulnerable=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_VULNERABLE
[XML key] team-profile-yellow-vulnerable

Type: integer Default value: 90 Min value: 5 Max value: 2000 Defines how vulnerable the yellow team is. This is a percentage, if set to 200 then team will be attacked twice as much as any other team with the default value. Setting this to a high value clearly disadvantages this team.

4.11.137 team-profile-yellow-weapon-alternate-id

[Command-line option] --team-profile-yellow-weapon-alternate-id=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_WEAPON_ALTERNATE_ID
[XML key] team-profile-yellow-weapon-alternate-id

Type: integer

Default value: 9 Min value: 0 Max value: 19

Id of the default alternate weapon for the yellow team, see the documentation about weapons to know what these ids mean.

4.11.138 team-profile-yellow-weapon-id

[Command-line option] --team-profile-yellow-weapon-id=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_WEAPON_ID
[XML key] team-profile-yellow-weapon-id

Type: integer

Default value: 5 Min value: 0 Max value: 19

Id of the default weapon for the yellow team, see the documentation about weapons to know what these ids mean.

4.11.139 team-profile-yellow-weapon-mode

[Command-line option] --team-profile-yellow-weapon-mode=<value>
[Environment variable] LW6_TEAM_PROFILE_YELLOW_WEAPON_MODE
[XML key] team-profile-yellow-weapon-mode

Type: integer

Default value: 1 Min value: 0 Max value: 2

Weapon mode for yellow team. 0 means there’s no weapon, 1 means one weapon per team, defined by the weapon-id parameter, 2 means random weapon.

4.11.140 total-armies-size

[Command-line option] --total-armies-size=<value>
[Environment variable] LW6_TOTAL_ARMIES_SIZE
[XML key] total-armies-size

Type: integer

Default value: 60 Min value: 1 Max value: 95

Defines the proportion of the whole available space, which can be occupied by all the armies present together. Setting this low, whenever a new team arrives on the map, fighters might be stolen to other teams, otherwise the ame would get too crowded. This allows you to play with reasonnably enough fighters with 2 players, while still allowing interesting gameplay with many players. This is a percentage, but will be multiplied by itself to get the actual surface. That is, 50 means 50%*50%, that is, a square of 1/2 the size of a square map, so it represents 25% (1/4) of the total surface.

4.11.141 total-time

[Command-line option] --total-time=<value>

[Environment variable] LW6_TOTAL_TIME

[XML key] total-time

Type: integer

Default value: 900 Min value: 10 Max value: 864000

Defines the maximum time of the game, in seconds. Note that in some cases, the game can end much earlier if some player has managed to win before the bell rings. Also, technically, this value will be translated into rounds and moves, and the game engine will wait until enough rounds and moves have been played. So if the computer is too slow and the desired speed is not reached, then the game will last for a longer time.

4.11.142 use-team-profiles

[Command-line option] --use-team-profiles=<value>
[Environment variable] LW6_USE_TEAM_PROFILES
[XML key] use-team-profiles

Type: integer

Default value: 1 Min value: 0 Max value: 1

If set, then all the team-profile-... values will be taken in account. This enables a mode in which teams behave differently according to their colors. If you disable this, then all teams will behave the same, which is more fair, but might not be as fun.

4.11.143 vertical-move

[Command-line option] --vertical-move=<value>

[Environment variable] LW6_VERTICAL_MOVE

[XML key] vertical-move

Type: integer

Default value: 1 Min value: 0 Max value: 7

Defines when to process a vertical move (along the Z ’depth’ axis). If set to 0, fighters never spontaneously move along this axis. If set to 1, it will be tried just after the first move failed. If set to 2, it will be tried just after the second move failed. And so on.

4.11.144 weapon-charge-delay

[Command-line option] --weapon-charge-delay=<value>
[Environment variable] LW6_WEAPON_CHARGE_DELAY
[XML key] weapon-charge-delay

Type: integer

Default value: 30 Min value: 1 Max value: 600

How long it will take for weapons to charge and be usable, by default. Unit is seconds.

4.11.145 weapon-charge-max

[Command-line option] --weapon-charge-max=<value>
[Environment variable] LW6_WEAPON_CHARGE_MAX
[XML key] weapon-charge-max

Type: integer Default value: 200 Min value: 100 Max value: 1000 Maximum (percentage) of charge intensity that one have. For instance, if this is 400, then if you wait four times more than required before firing, then you weapon will have four times its default power, but if you wait five times more it will still be four times more powerfull, it’s just the limit after which it’s useless to charge.

4.11.146 weapon-duration

[Command-line option] --weapon-duration=<value>

[Environment variable] LW6_WEAPON_DURATION

[XML key] weapon-duration

Type: integer Default value: 3 Min value: 1 Max value: 60 How long all weapons (for which duration makes sense) will last. Unit is seconds.

4.11.147 weapon-tune-berzerk-power

[Command-line option] --weapon-tune-berzerk-power=<value>
[Environment variable] LW6_WEAPON_TUNE_BERZERK_POWER
[XML key] weapon-tune-berzerk-power

Type: integer Default value: 3 Min value: 1 Max value: 100 Use to specifiy how strong berzerk mode is, if set to 3, then attacks will be 3 times as efficient in berzerk mode.

4.11.148 weapon-tune-turbo-power

[Command-line option] --weapon-tune-turbo-power=<value>
[Environment variable] LW6_WEAPON_TUNE_TURBO_POWER
[XML key] weapon-tune-turbo-power

Type: integer Default value: 3 Min value: 1 Max value: 10 Defines how fast fighters move in turbo mode, if set to 3, then fighters move and act 3 times as fast.

4.11.149 x-polarity

[Command-line option] --x-polarity=<value>

[Environment variable] LW6_X_POLARITY

[XML key] x-polarity

Type: integer Default value: 0 Min value: -1 Max value: 1 Defines how the map will be wrapped on the X (horizontal) axis. If set to 0, nothing is wrapped. If set to 1, the right and left borders are connected, any fighter can disappear on the right border and reappear on the left border, for instance. If set to -1, it will be wrapped but also inversed, that is on a 320x240 map, a fighter disappearing on the left border at position (0,60) will reapper on the right border at position (319,180). You can combine it with ’y-polarity’.

4.11.150 y-polarity

[Command-line option] --y-polarity=<value>

[Environment variable] LW6_Y_POLARITY

[XML key] y-polarity

Type: integer Default value: 0 Min value: -1 Max value: 1 Defines how the map will be wrapped on the Y (vertical) axis. If set to 0, nothing is wrapped. If set to 1, the top and bottom borders are connected, any fighter can disappear on the top border and reappear on the bottom border, for instance. If set to -1, it will be wrapped but also inversed, that is on a 320x240 map, a fighter disappearing on the bottom border at position (40,239) will reapper on the top border at position (280,0). You can combine it with ’x-polarity’.

4.11.151 z-polarity

[Command-line option] --z-polarity=<value>

[Environment variable] LW6_Z_POLARITY

[XML key] z-polarity

Type: integer Default value: 0 Min value: 0 Max value: 1 Defines how the map will be wrapped on the Z (deep) axis. If set to 0, nothing is wrapped. If set to 1, when using a 4 layer map, for instance, fighters on layer 1 will be able to go directly to layer 4 even if layers 2 and 3 are filled with walls. A value of -1 is forbidden, this is not like x and y axis, it does not really make sense. Consider this an advanced setting which might save a layer in some tricky cases, the default value of 0 should fit in most cases.

4.12 Map hints.xml

4.12.1 background-color-auto

[Command-line option] --background-color-auto=<value>
[Environment variable] LW6_BACKGROUND_COLOR_AUTO
[XML key] background-color-auto

Type: boolean Default value: true Defines wether hud colors will be set automatically from base and alternate colors. This is a time saver to keep map designers from requiring to redefined every single color in the game. You only need to set color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg. Then hud color frame bg, hud color frame fg, hud color text bg and hud color text fg will be automatically set.

4.12.2 downsize-using-bench-value

[Command-line option] --downsize-using-bench-value=<value>
[Environment variable] LW6_DOWNSIZE_USING_BENCH_VALUE
[XML key] downsize-using-bench-value

Type: boolean Default value: true If set, then the game will automatically downsize a map according to the ’bench-value’ parameter. Downsizing means: a 1600x1200 maps becomes 200x150, for instance. Downsizing causes fighters to be bigger because map resolution is lower. This will avoid running the game on a too big map, with your computer not being able to handle it at the required speed.

4.12.3 downsize-using-fighter-scale

[Command-line option] --downsize-using-fighter-scale=<value>
[Environment variable] LW6_DOWNSIZE_USING_FIGHTER_SCALE
[XML key] downsize-using-fighter-scale

Type: boolean Default value: false If set, then the game will automatically downsize a map according to the ’fighter-scale’ parameter. Downsizing means: a 1600x1200 maps becomes 200x150, for instance. Downsizing causes fighters to be bigger because map resolution is lower. This can be usefull if you don’t want fighters to be too small.

4.12.4 fighter-scale

[Command-line option] --fighter-scale=<value>

[Environment variable] LW6_FIGHTER_SCALE

[XML key] fighter-scale

Type: float Default value: 1.0 Defines how wide (in pixels) fighters must be. This parameter is very important and will largely condition the number of fighters on the map. It is used when loading the map. If it is, for instance, set to 1, there will be exactly a fighter per pixel on the screen. That is, if you play 640x480 on an empty map, the maximum fighters you could have is about 300000. The idea is that by changing the resolution, you also define the density of the map. In pratice, this is done in the hope that someone with a slow computer will pick up a low resolution and therefore play small levels. Conversely, someone with a brand new computer with powerfull CPU & GPU will use great resolutions and be happy with many fighters on the map. Still, changing the resolution after loading the map will not affet the number of fighters. Same for network games, the first player, who loads the map, defines its properties according to its own settings.

4.12.5 guess-colors

[Command-line option] --guess-colors=<value>

[Environment variable] LW6_GUESS_COLORS

[XML key] guess-colors

Type: boolean Default value: true Defines wether colors should be set automatically from texture colors. If set to true, then the program will try to pick up colors automatically from the texture, and will override the values of the color-base-bg, color-base-fg, color-alternate-bg and coloralternate-fg parameters. How these colors are picked up can’t be garanteed, so if the map does not have strong contrast or if there can be any form of ambiguity, it’s safe to set this to false and define one’s own colors.

4.12.6 guess-moves-per-sec

[Command-line option] --guess-moves-per-sec=<value>
[Environment variable] LW6_GUESS_MOVES_PER_SEC
[XML key] guess-moves-per-sec

Type: boolean Default value: true If set, then loader will use ’time-to-cross-level’ to guess the game speed parameters.

4.12.7 hud-color-auto

[Command-line option] --hud-color-auto=<value>

[Environment variable] LW6_HUD_COLOR_AUTO

[XML key] hud-color-auto

Type: boolean Default value: true Defines wether hud colors will be set automatically from base and alternate colors. This is a time saver to keep map designers from requiring to redefined every single color in the game. You only need to set color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg. Then hud color frame bg, hud color frame fg, hud color text bg and hud color text fg will be automatically set.

4.12.8 max-map-height

[Command-line option] --max-map-height=<value>

[Environment variable] LW6_MAX_MAP_HEIGHT

[XML key] max-map-height

Type: integer Default value: 1000 Allows you to give a maximum map height. When designing a map you might wonder: this is dumb I’m conceiving this map I know its height, why should I limit it? Now think of the play who plays on a old slowish computer with a tiny screen. He might redefine this himself, and does not necessarly wishes to fire Gimp to rescale the map.

4.12.9 max-map-surface

[Command-line option] --max-map-surface=<value>

[Environment variable] LW6_MAX_MAP_SURFACE

[XML key] max-map-surface

Type: integer Default value: 1000000 Allows you to give a maximum map surface. Map surface is simply (width * height). This parameter is just here to save you the hassle of defining both ’max-map-width’ and ’max-map-height’ in a consistent manner.

4.12.10 max-map-width

[Command-line option] --max-map-width=<value>

[Environment variable] LW6_MAX_MAP_WIDTH

[XML key] max-map-width

Type: integer Default value: 1500 Allows you to give a maximum map width. When designing a map you might wonder: this is dumb I’m conceiving this map I know its width, why should I limit it? Now think of the play who plays on a old slowish computer with a tiny screen. He might redefine this himself, and does not necessarly wishes to fire Gimp to rescale the map.

4.12.11 menu-color-auto

[Command-line option] --menu-color-auto=<value>

[Environment variable] LW6_MENU_COLOR_AUTO

[XML key] menu-color-auto

Type: boolean Default value: true Defines wether menu colors will be set automatically from base and alternate colors. This is a time saver to keep map designers from requiring to redefined every single color in the game. You only need to set color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg. Then menu color default bg, menu color default fg, menu color selected bg, menu color selected fg, menu color disabled bg and menu color disabled fg will be automatically set.

4.12.12 min-map-height

[Command-line option] --min-map-height=<value>

[Environment variable] LW6_MIN_MAP_HEIGHT

[XML key] min-map-height

Type: integer Default value: 30 Allows you to give a minimum map height. When designing a map you might wonder: this is dumb I’m conceiving this map I know its height, why should I limit it? Now think of the player who decided to play with highly-defined maps because he has a super calculator and a hudge screen. He might redefine this himself, and does not necessarly wishes to fire Gimp to rescale the map.

4.12.13 min-map-surface

[Command-line option] --min-map-surface=<value>

[Environment variable] LW6_MIN_MAP_SURFACE

[XML key] min-map-surface

Type: integer Default value: 3600 Allows you to give a minimum map surface. Map surface is simply (width * height). This parameter is just here to save you the hassle of defining both ’min-map-width’ and ’min-map-height’ in a consistent manner.

4.12.14 min-map-width

[Command-line option] --min-map-width=<value>

[Environment variable] LW6_MIN_MAP_WIDTH

[XML key] min-map-width

Type: integer Default value: 40 Allows you to give a minimum map width. When designing a map you might wonder: this is dumb I’m conceiving this map I know its width, why should I limit it? Now think of the player who decided to play with highly-defined maps because he has a super calculator and a hudge screen. He might redefine this himself, and does not necessarly wishes to fire Gimp to rescale the map.

4.12.15 resample

[Command-line option] --resample=<value> [Environment variable] LW6_RESAMPLE

[XML key] resample

Type: boolean Default value: true If set to true, maps will always be resampled to a size which depends on your screen resolution, zoom factor, and the rest. If false, maps will be set at the exact resolution of map.png.

4.12.16 speed

[Command-line option] --speed=<value> [Environment variable] LW6_SPEED [XML key] speed

Type: float Default value: 1.0 This parameter is the main parameter on which game speed depends. The map loader will garantee, by downscaling the map, that to cross the level (by crossing the level we mean, for instance, going from top-left corner to bottom-right corner in a straight line) a fighter will take a constant amount of time. Under the hood, the loader might of course rescale the map but it will also change game speed so that, at the end, fighters take a constant time to cross the level. This is, indeed, the most important thing, players do not care much if internally there are X or Y moves per second, the global game experience depends on how fast fighter movement looks on the screen. The default settings corresponds roughly to one second to cross the level. If you set this to 2.0, it will go twice faster.

4.12.17 system-color-auto

[Command-line option] --system-color-auto=<value>
[Environment variable] LW6_SYSTEM_COLOR_AUTO
[XML key] system-color-auto

Type: boolean Default value: true Defines wether system colors will be set automatically from base and alternate colors. This is a time saver to keep map designers from requiring to redefined every single color in the game. You only need to set color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg. Then system color bg and system color fg will be automatically set.

4.12.18 upsize-using-bench-value

[Command-line option] --upsize-using-bench-value=<value>
[Environment variable] LW6_UPSIZE_USING_BENCH_VALUE
[XML key] upsize-using-bench-value

Type: boolean Default value: false If set, then the game will automatically upsize a map according to the ’fighter-scale’ parameter. Upsizing means: a 160x120 maps becomes 400x300, for instance. Upsizing causes fighters to be smaller because map resolution is higher. This will avoid useless pixelish ’jumbo fighters’ look when your computer is powerfull enough to do better.

4.12.19 upsize-using-fighter-scale

[Command-line option] --upsize-using-fighter-scale=<value>
[Environment variable] LW6_UPSIZE_USING_FIGHTER_SCALE
[XML key] upsize-using-fighter-scale

Type: boolean Default value: true If set, then the game will automatically upsize a map according to the ’fighter-scale’ parameter. Upsizing means: a 160x120 maps becomes 400x300, for instance. Upsizing causes fighters to be smaller because map resolution is higher. This can be usefull if you don’t want fighters to be too big.

4.12.20 view-color-auto

[Command-line option] --view-color-auto=<value>

[Environment variable] LW6_VIEW_COLOR_AUTO

[XML key] view-color-auto

Type: boolean

Default value: true

Defines wether view colors will be set automatically from base and alternate colors. This is a time saver to keep map designers from requiring to redefined every single color in the game. You only need to set color-base-bg, color-base-fg, coloralternate-bg and color-alternate-fg. Then view color cursor bg, view color cursor fg, view color map bg and view color map fg will be automatically set.

4.12.21 wall-grease

[Command-line option] --wall-grease=<value>

[Environment variable] LW6_WALL_GREASE

[XML key] wall-grease

Type: integer

Default value: 0 Min value: -5 Max value: 5

This parameter allows you to make walls (AKA map foreground) thicker, or thiner, when map is loaded. Indeed, when map are resampled, and especially when they are downscaled, some walls may disappear, or some passages may be blocked. The loader can’t automatically figure out wether it’s more important to keep an existing wall or to keep an open passage for fighters. This parameter helps doing so, if you set it to a low value, level will be less greasy, and many passages might open themselves. On the contrary, if grease is at a high level, then a thin line of almost isolated pixels might become a thick wall. There’s no real garantee your wall or passage will always be present, but it’s a same bet to assume on a ’tunnel-like’ level one needs to set grease to a low value, and on a ’wide open’ level with few walls one needs to set grease to a high value.

4.13 Map style.xml

4.13.1 animation-density

[Command-line option] --animation-density=<value>
[Environment variable] LW6_ANIMATION_DENSITY
[XML key] animation-density

Type: float

Default value: 1.0 Min value: 0 Max value: 10

Density of the background animation, that is, for instance, if the background animation is about displaying bubbles, using a high value will display many bubbles. A value of 1.0 corresponds to the default setting.

4.13.2 animation-speed

[Command-line option] --animation-speed=<value>

[Environment variable] LW6_ANIMATION_SPEED

[XML key] animation-speed

Type: float Default value: 1.0 Min value: 0 Max value: 10 Speed of the background animation, that is, for instance, if the background animation is about displaying bubbles, using a high value will cause bubbles to move very fast. A value of 1.0 corresponds to the default setting.

4.13.3 background-color-root-bg

[Command-line option] --background-color-root-bg=<value>
[Environment variable] LW6_BACKGROUND_COLOR_ROOT_BG
[XML key] background-color-root-bg

Type: color Default value: #000000 Defines the main background color. This is, for instance, the color which will be used to clear the screen before drawing thing. Will be automatically guessed from the map texture if color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.4 background-color-root-fg

[Command-line option] --background-color-root-fg=<value>
[Environment variable] LW6_BACKGROUND_COLOR_ROOT_FG
[XML key] background-color-root-fg

Type: color Default value: #cccccc Defines a color which will be used together with color-base-bg to compose the background. It can be wise to have a minimum contrast between this color and color-base-bg, but it is not mandatory, especially if other colors are manually redefined. Will be automatically guessed from the map texture if color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.5 background-color-stuff-bg

[Command-line option] --background-color-stuff-bg=<value>
[Environment variable] LW6_BACKGROUND_COLOR_STUFF_BG
[XML key] background-color-stuff-bg

Type: color Default value: #333333 Defines a color which will be used together with color-alternate-fg to draw things (animations, sprites, text, whatever) in the background. It should be different enough from color-alternate-fg so that one can really distinguish these colors. Will be automatically guessed from the map texture if color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.6 background-color-stuff-fg

[Command-line option] --background-color-stuff-fg=<value>
[Environment variable] LW6_BACKGROUND_COLOR_STUFF_FG
[XML key] background-color-stuff-fg

Type: color Default value: #ffffff Defines a color which will be used to draw things (animations, sprites, text, whatever) in the background. It should be different enough from color-alternate-bg so that one can really distinguish these colors. Think of this as the sprite, the text, the whateverneeds-to-be-seen-uses-this color. Will be automatically guessed from the map texture if color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.7 background-style

[Command-line option] --background-style=<value>

[Environment variable] LW6_BACKGROUND_STYLE

[XML key] background-style

Type: string Default value: bubbles The background defines, of course, what is displayed at the background, but it also conditions the colors used for other items, such as the menus for instance. The possible values are ’void’ and ’bubbles’.

4.13.8 blink-cursor

[Command-line option] --blink-cursor=<value>

[Environment variable] LW6_BLINK_CURSOR

[XML key] blink-cursor

Type: boolean Default value: false If set, then cursor will blink, allowing you to see what’s under the cursor. It’s just a matter of taste, you might to always have your cursor displayed, or prefer to have it disappear from time to time so that you can see the action below

4.13.9 color-alternate-bg

[Command-line option] --color-alternate-bg=<value>
[Environment variable] LW6_COLOR_ALTERNATE_BG
[XML key] color-alternate-bg

Type: color Default value: #333333 Defines the alternate color, more precisely, its bg (background) part. Colors are always defined by a bg/fg pair. Most colors in the game can be deduced from this one, usually to color a map you only need to define color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg.

4.13.10 color-alternate-fg

[Command-line option] --color-alternate-fg=<value>
[Environment variable] LW6_COLOR_ALTERNATE_FG
[XML key] color-alternate-fg

Type: color Default value: #ffffff Defines the alternate color, more precisely, its fg (foreground) part. Colors are always defined by a bg/fg pair. Most colors in the game can be deduced from this one, usually to color a map you only need to define color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg.

4.13.11 color-base-bg

[Command-line option] --color-base-bg=<value>

[Environment variable] LW6_COLOR_BASE_BG

[XML key] color-base-bg

Type: color Default value: #000000 Defines the base color, more precisely, its bg (background) part. Colors are always defined by a bg/fg pair. Most colors in the game can be deduced from this one, usually to color a map you only need to define color-base-bg, color-base-fg, color-alternate-bg and color-alternate-fg.

4.13.12 color-base-fg

[Command-line option] --color-base-fg=<value>

[Environment variable] LW6_COLOR_BASE_FG

[XML key] color-base-fg

Type: color Default value: #cccccc Defines the base color, more precisely, its fg (foreground) part. Colors are always defined by a bg/fg pair. Most colors in the game can be deduced from this one, usually to color a map you only need to define color-base-bg, color-base-fg, coloralternate-bg and color-alternate-fg.

4.13.13 colorize

[Command-line option] --colorize=<value> [Environment variable] LW6_COLORIZE

[XML key] colorize

Type: boolean Default value: true If set, then all background drawings including textures will use the background colors. This means, for instance, that if background colors are set automatically by colorauto from the map texture, then the background will adopt the same range of colors than the map itself. In short, the background will mimic the map.

4.13.14 colorize-cursor

[Command-line option] --colorize-cursor=<value>

[Environment variable] LW6_COLORIZE_CURSOR

[XML key] colorize-cursor

Type: boolean

Default value: true

If set, then all cursors will use the automatic guessed colors, or the specified colors, but basically they won’t be displayed using their native colors. This can be usefull for you can wish to use a generic non-colored texture for your cursor and let it be colorized automatically so that it’s accorded to the level.

4.13.15 cursor-size

[Command-line option] --cursor-size=<value>

[Environment variable] LW6_CURSOR_SIZE

[XML key] cursor-size

Type: float

Default value: 1.0 Min value: 0 Max value: 10

Size of the cursors on the map. 1 is the default, setting it to a higher value will make cursors bigger, a lower value will make them smaller.

4.13.16 hidden-layer-alpha

[Command-line option] --hidden-layer-alpha=<value>
[Environment variable] LW6_HIDDEN_LAYER_ALPHA
[XML key] hidden-layer-alpha

Type: float

Default value: 0.1 Min value: 0 Max value: 1

Whenever players are supposed to be hidden behind a wall, for instance if they are in layer 2 and layer 1 is filled with walls, it’s still possible to see them, but with a low alpha value (almost transparent). This parameter allows you to trick this value, 0 will make these players absolutely invisible, 1 will make them totally opaque, like if they were on layer 1.

4.13.17 hud-color-frame-bg

[Command-line option] --hud-color-frame-bg=<value>
[Environment variable] LW6_HUD_COLOR_FRAME_BG
[XML key] hud-color-frame-bg

Type: color

Default value: #000000

Defines the background color for the hud frame. Ignored if hud-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.18 hud-color-frame-fg

[Command-line option] --hud-color-frame-fg=<value>
[Environment variable] LW6_HUD_COLOR_FRAME_FG
[XML key] hud-color-frame-fg

Type: color Default value: #cccccc Defines the foreground color for the hud frame. Ignored if hud-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.19 hud-color-text-bg

[Command-line option] --hud-color-text-bg=<value>
[Environment variable] LW6_HUD_COLOR_TEXT_BG
[XML key] hud-color-text-bg

Type: color Default value: #333333 Defines the background color for hud text. Ignored if hud-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.20 hud-color-text-fg

[Command-line option] --hud-color-text-fg=<value>
[Environment variable] LW6_HUD_COLOR_TEXT_FG
[XML key] hud-color-text-fg

Type: color Default value: #ffffff Defines the foreground color for hud text. Ignored if hud-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.21 hud-style

[Command-line option] --hud-style=<value> [Environment variable] LW6_HUD_STYLE

[XML key] hud-style

Type: string Default value: floating The hud is where informations about the game are displayed. This means, who is winning, are other status-like informations. Possible values include ’floating’ and ’tactical’.

4.13.22 keep-ratio

[Command-line option] --keep-ratio=<value>

[Environment variable] LW6_KEEP_RATIO

[XML key] keep-ratio

Type: boolean Default value: true

Defines wether the map should keep its ratio, or if it should be stretched to fill the shape of your screen.

4.13.23 menu-color-default-bg

[Command-line option] --menu-color-default-bg=<value>
[Environment variable] LW6_MENU_COLOR_DEFAULT_BG
[XML key] menu-color-default-bg

Type: color

Default value: #333333

Defines the default background color for menus. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.24 menu-color-default-fg

[Command-line option] --menu-color-default-fg=<value>
[Environment variable] LW6_MENU_COLOR_DEFAULT_FG
[XML key] menu-color-default-fg

Type: color

Default value: #ffffff

Defines the default foreground color for menus. In fact, this is the main color for menu text, the color used to draw letters in menus. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.25 menu-color-disabled-bg

[Command-line option] --menu-color-disabled-bg=<value>
[Environment variable] LW6_MENU_COLOR_DISABLED_BG
[XML key] menu-color-disabled-bg

Type: color

Default value: #000000

Defines the background color for a disabled menu item. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.26 menu-color-disabled-fg

[Command-line option] --menu-color-disabled-fg=<value>
[Environment variable] LW6_MENU_COLOR_DISABLED_FG
[XML key] menu-color-disabled-fg

Type: color

Default value: #cccccc

Defines the foreground color for a disabled menu item. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.27 menu-color-selected-bg

[Command-line option] --menu-color-selected-bg=<value>
[Environment variable] LW6_MENU_COLOR_SELECTED_BG
[XML key] menu-color-selected-bg

Type: color Default value: #ffffff Defines the background color for a selected menu item. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.28 menu-color-selected-fg

[Command-line option] --menu-color-selected-fg=<value>
[Environment variable] LW6_MENU_COLOR_SELECTED_FG
[XML key] menu-color-selected-fg

Type: color Default value: #333333 Defines the foreground color for a selected menu item. Ignored if menu-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.29 menu-style

[Command-line option] --menu-style=<value>

[Environment variable] LW6_MENU_STYLE

[XML key] menu-style

Type: string Default value: cylinder The menu style is simply the name of the engine used to power the menu system. The only possible value, for now, is ’cylinder’.

4.13.30 music-exclude

[Command-line option] --music-exclude=<value>

[Environment variable] LW6_MUSIC_EXCLUDE

[XML key] music-exclude

Type: string Default value: Chadburn If this string is found in a music file name, it will be excluded from the list when playing in random mode.

4.13.31 music-file

[Command-line option] --music-file=<value>

[Environment variable] LW6_MUSIC_FILE

[XML key] music-file

Type: string Default value:

Allows you to play a custom music file (typically your own ogg music) and override default game music. If file does not exist, game will use its internal music. The file will be searched for in the current ’music-path’ but also in the current map directory. No absolute or even relative path are allowed, only a plain filename with no slash or backslash. Avoid special characters at all cost.

4.13.32 music-filter

[Command-line option] --music-filter=<value>

[Environment variable] LW6_MUSIC_FILTER

[XML key] music-filter

Type: string Default value: A music filter, used when files are played randomly. This is not a complex regexenabled filter, just a plain string search. Even the ’*’ wildcard won’t work. If you want precise control on what music file to play, please consider reorganizing your files and/or use the ’music-file’ parameter.

4.13.33 pixelize

[Command-line option] --pixelize=<value> [Environment variable] LW6_PIXELIZE

[XML key] pixelize

Type: boolean Default value: false Depending on the renderer capabilities, will try to pixelize some parts of the game. This can be used to emulate the old LW5 appearance.

4.13.34 system-color-bg

[Command-line option] --system-color-bg=<value>

[Environment variable] LW6_SYSTEM_COLOR_BG

[XML key] system-color-bg

Type: color Default value: #333333 Defines the system background color, used when displaying system info, such as the number of frames per second. Can be #RGB, #RGBA, #RRGGBB or #RRGGB-BAA.

4.13.35 system-color-fg

[Command-line option] --system-color-fg=<value>

[Environment variable] LW6_SYSTEM_COLOR_FG

[XML key] system-color-fg

Type: color Default value: #ffffff Defines the system foreground color, used when displaying system info, such as the number of frames per second. This will typically be text color. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.36 team-color-blue

[Command-line option] --team-color-blue=<value>

[Environment variable] LW6_TEAM_COLOR_BLUE

[XML key] team-color-blue

Type: color Default value: #0000ff Defines the color for the blue team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.37 team-color-cyan

[Command-line option] --team-color-cyan=<value>

[Environment variable] LW6_TEAM_COLOR_CYAN

[XML key] team-color-cyan

Type: color Default value: #00ffff Defines the color for the cyan team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.38 team-color-dead

[Command-line option] --team-color-dead=<value>

[Environment variable] LW6_TEAM_COLOR_DEAD

[XML key] team-color-dead

Type: color Default value: #000000 Defines the color for the teams when they are dead. By default it is black, this means when a team is weak it becomes black. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.39 team-color-green

[Command-line option] --team-color-green=<value>

[Environment variable] LW6_TEAM_COLOR_GREEN

[XML key] team-color-green

Type: color Default value: #00ff00 Defines the color for the green team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.40 team-color-lightblue

[Command-line option] --team-color-lightblue=<value>
[Environment variable] LW6_TEAM_COLOR_LIGHTBLUE
[XML key] team-color-lightblue

Type: color Default value: #88bbff Defines the color for the light blue team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.41 team-color-magenta

[Command-line option] --team-color-magenta=<value>
[Environment variable] LW6_TEAM_COLOR_MAGENTA
[XML key] team-color-magenta

Type: color Default value: #ff00ff Defines the color for the magenta team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.42 team-color-orange

[Command-line option] --team-color-orange=<value>
[Environment variable] LW6_TEAM_COLOR_ORANGE
[XML key] team-color-orange

Type: color Default value: #ff8800 Defines the color for the orange team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.43 team-color-pink

[Command-line option] --team-color-pink=<value>

[Environment variable] LW6_TEAM_COLOR_PINK

[XML key] team-color-pink

Type: color Default value: #ff88bb Defines the color for the pink team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.44 team-color-purple

[Command-line option] --team-color-purple=<value>
[Environment variable] LW6_TEAM_COLOR_PURPLE
[XML key] team-color-purple

Type: color Default value: #bb88ff Defines the color for the purple team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.45 team-color-red

[Command-line option] --team-color-red=<value>

[Environment variable] LW6_TEAM_COLOR_RED

[XML key] team-color-red

Type: color Default value: #ff0000 Defines the color for the red team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.46 team-color-yellow

[Command-line option] --team-color-yellow=<value>
[Environment variable] LW6_TEAM_COLOR_YELLOW
[XML key] team-color-yellow

Type: color Default value: #ffff00 Defines the color for the yellow team. Syntax is HTML-like, #RGB or #RRGGBB.

4.13.47 view-color-cursor-bg

[Command-line option] --view-color-cursor-bg=<value>
[Environment variable] LW6_VIEW_COLOR_CURSOR_BG
[XML key] view-color-cursor-bg

Type: color Default value: #333333 Defines the background cursor color. Will typically be used to draw the shape of the cursor. Ignored if view-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.48 view-color-cursor-fg

[Command-line option] --view-color-cursor-fg=<value>
[Environment variable] LW6_VIEW_COLOR_CURSOR_FG
[XML key] view-color-cursor-fg

Type: color Default value: #ffffff Defines the foreground cursor color. Will typically be used to draw text in the cursor. Ignored if view-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGG-BBAA.

4.13.49 view-color-map-bg

[Command-line option] --view-color-map-bg=<value>
[Environment variable] LW6_VIEW_COLOR_MAP_BG
[XML key] view-color-map-bg

Type: color Default value: #000000 Defines the background map color. If there’s no map texture defined or if use-texture is false, this is the color of the places where armies will go. Ignored if view-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.50 view-color-map-fg

[Command-line option] --view-color-map-fg=<value>
[Environment variable] LW6_VIEW_COLOR_MAP_FG

[XML key] view-color-map-fg

Type: color Default value: #cccccc Defines the foreground map color. If there’s no map texture defined or if use-texture is false, this is the color of walls, what armies can’t go through. Ignored if view-color-auto is set. Can be #RGB, #RGBA, #RRGGBB or #RRGGBBAA.

4.13.51 view-style

[Command-line option] --view-style=<value>

[Environment variable] LW6_VIEW_STYLE

[XML key] view-style

Type: string Default value: flat The view style conditions which renderer is used for the map, the area where fighters are displayed. This is not the graphics backend. Indeed, the graphics backend defines which technical tool one uses (which library) one runs, wether this parameter says what kind of rendering one wants.

4.13.52 waves

[Command-line option] --waves=<value> [Environment variable] LW6_WAVES [XML key] waves

Type: boolean Default value: true Activates the wave effect, that’s to say level appears to be under water when playing.

4.13.53 x-wrap

[Command-line option] --x-wrap=<value> [Environment variable] LW6_X_WRAP [XML key] x-wrap

Type: boolean Default value: true Defines wether the map should be wrapped on the x axis. This is the companion of ’x-polarity’, if no polarity is defined, map can’t be wrapped, but in some cases, one might wish to have a map with polarity but without wrapping if, for instance, textures do not tile nicely.

4.13.54 y-wrap

[Command-line option] --y-wrap=<value> [Environment variable] LW6_Y_WRAP [XML key] y-wrap

Type: boolean Default value: true Defines wether the map should be wrapped on the y axis. This is the companion of ’y-polarity’, if no polarity is defined, map can’t be wrapped, but in some cases, one might wish to have a map with polarity but without wrapping if, for instance, textures do not tile nicely.

4.13.55 zoom

[Command-line option] --zoom=<value> [Environment variable] LW6_ZOOM [XML key] zoom

Type: float Default value: 1.0 Defines the map zoom. If lower than 1.0, map will occupy only a fraction of the screen, if greater than 1.0, some areas will be outside the screen, and the player will need to scroll through it.

4.13.56 zoom-max

[Command-line option] --zoom-max=<value> [Environment variable] LW6_ZOOM_MAX

[XML key] zoom-max

Type: float Default value: 30.0 Defines the max map zoom. If set to a high value, you’ll be able to dynamically view the map with hudge fighters, seeing only a fraction of the level.

4.13.57 zoom-min

[Command-line option] --zoom-min=<value> [Environment variable] LW6_ZOOM_MIN

[XML key] zoom-min

Type: float Default value: 0.3 Defines the min map zoom. If set to a low value, you’ll be able to dynamically view a very small, reduced map.

4.14 Map teams.xml

4.14.1 bot-iq

[Command-line option] --bot-iq=<value> [Environment variable] LW6_BOT_IQ [XML key] bot-iq

Type: integer Default value: 100 Min value: 0 Max value: 200 The IQ (intelligence quotient) of bots. Typically, a value of 100 will make the bot behave normally, performing at its best. A value of 0 will just make it act the worst way it can. Values over 100 probably won’t change anything compared to 100, but this truely depends on which bot backend you’re running.

4.14.2 bot-speed

[Command-line option] --bot-speed=<value> [Environment variable] LW6_BOT_SPEED

[XML key] bot-speed

Type: float Default value: 1.0f The speed of bots, 1 means normal speed, higher value will speed it up, lower will slow it down. Note that this only has an impact on bot engines, not on the game speed itself.

4.14.3 bot1-ai

[Command-line option] --bot1-ai=<value> [Environment variable] LW6_BOT1_AI [XML key] bot1-ai

Type: string Default value: idiot AI engine for bot number 1.

4.14.4 bot1-color

[Command-line option] --bot1-color=<value>

[Environment variable] LW6_BOT1_COLOR

[XML key] bot1-color

Type: string Default value: green Color for bot number 1.

4.14.5 bot2-ai

[Command-line option] --bot2-ai=<value> [Environment variable] LW6_BOT2_AI [XML key] bot2-ai

Type: string Default value: idiot AI engine for bot number 2.

4.14.6 bot2-color

[Command-line option] --bot2-color=<value>

[Environment variable] LW6_BOT2_COLOR

[XML key] bot2-color

Type: string Default value: blue Color for bot number 2.

4.14.7 bot3-ai

[Command-line option] --bot3-ai=<value> [Environment variable] LW6_BOT3_AI [XML key] bot3-ai

Type: string

Default value: random

AI engine for bot number 3.

4.14.8 bot3-color

[Command-line option] --bot3-color=<value>

[Environment variable] LW6_BOT3_COLOR

[XML key] bot3-color

Type: string

Default value: yellow Color for bot number 3.

4.14.9 bot4-ai

[Command-line option] --bot4-ai=<value> [Environment variable] LW6_BOT4_AI [XML key] bot4-ai

Type: string

Default value: follow

AI engine for bot number 4.

4.14.10 bot4-color

[Command-line option] --bot4-color=<value>

[Environment variable] LW6_BOT4_COLOR

[XML key] bot4-color

Type: string

Default value: cyan Color for bot number 4.

4.14.11 bot5-ai

[Command-line option] --bot5-ai=<value> [Environment variable] LW6_BOT5_AI [XML key] bot5-ai

Type: string

Default value: random

AI engine for bot number 5.

4.14.12 bot5-color

[Command-line option] --bot5-color=<value>

[Environment variable] LW6_BOT5_COLOR

[XML key] bot5-color

Type: string

Default value: magenta

Color for bot number 5.

4.14.13 bot6-ai

[Command-line option] --bot6-ai=<value> [Environment variable] LW6_BOT6_AI [XML key] bot6-ai

Type: string

Default value: follow

AI engine for bot number 6.

4.14.14 bot6-color

[Command-line option] --bot6-color=<value>

[Environment variable] LW6_BOT6_COLOR

[XML key] bot6-color

Type: string

Default value: orange

Color for bot number 6.

4.14.15 bot7-ai

[Command-line option] --bot7-ai=<value> [Environment variable] LW6_BOT7_AI [XML key] bot7-ai

Type: string

Default value: idiot

AI engine for bot number 7.

4.14.16 bot7-color

[Command-line option] --bot7-color=<value>

[Environment variable] LW6_BOT7_COLOR

[XML key] bot7-color

Type: string

Default value: lightblue

Color for bot number 7.

4.14.17 bot8-ai

[Command-line option] --bot8-ai=<value> [Environment variable] LW6_BOT8_AI [XML key] bot8-ai

Type: string

Default value: idiot

AI engine for bot number 8.

4.14.18 bot8-color

[Command-line option] --bot8-color=<value>

[Environment variable] LW6_BOT8_COLOR

[XML key] bot8-color

Type: string

Default value: purple Color for bot number 8.

4.14.19 bot9-ai

[Command-line option] --bot9-ai=<value> [Environment variable] LW6_BOT9_AI [XML key] bot9-ai

Type: string

Default value: idiot

AI engine for bot number 9.

4.14.20 bot9-color

[Command-line option] --bot9-color=<value>

[Environment variable] LW6_BOT9_COLOR

[XML key] bot9-color

Type: string

Default value: pink Color for bot number 9.

4.14.21 nb-bots

[Command-line option] --nb-bots=<value> [Environment variable] LW6_NB_BOTS [XML key] nb-bots

Type: integer

Default value: 2 Min value: 0 Max value: 9

Number of bots on the map. 0 means no bots, if set to 1 the the bot1-... settings will be used, if set to 2 then bot1-... and bot2-... will be used, and so on.

4.14.22 player1-color

[Command-line option] --player1-color=<value>

[Environment variable] LW6_PLAYER1_COLOR

[XML key] player1-color

Type: string Default value: red Color of the first player, must be red, green, blue, yellow, cyan, magenta, orange, lightblue, purple or pink

4.14.23 player2-color

[Command-line option] --player2-color=<value>

[Environment variable] LW6_PLAYER2_COLOR

[XML key] player2-color

Type: string Default value: green Color of the second player, must be red, green, blue, yellow, cyan, magenta, orange, lightblue, purple or pink

4.14.24 player3-color

[Command-line option] --player3-color=<value>

[Environment variable] LW6_PLAYER3_COLOR

[XML key] player3-color

Type: string Default value: blue Color of the third player, must be red, green, blue, yellow, cyan, magenta, orange, lightblue, purple or pink

4.14.25 player4-color

[Command-line option] --player4-color=<value>

[Environment variable] LW6_PLAYER4_COLOR

[XML key] player4-color

Type: string Default value: yellow Color of the fourth player, must be red, green, blue, yellow, cyan, magenta, orange, lightblue, purple or pink

4.15 Advanced settings

4.15.1 base64-decode

[Command-line option] --base64-decode

If specified, program will take stdin and base64 decode it to stdout. This is for testing purpose (for network messages for instance). Will decode in standard base64 encoding using characters + and / but also the url-compliant version using - and /, see RFC 4648 for details.

4.15.2 base64-encode

[Command-line option] --base64-encode

If specified, program will take stdin and base64 encode it to stdout. This is for testing purpose (for network messages for instance). Will *not* use standard base64 encoding using characters + and / but - and instead to be url-compliant, see RFC 4648 for details.

4.15.3 bench

[Command-line option] --bench

Runs a benchmarking test which will report an approximative performance estimation of the game on your computer. The result is in an arbitrary unit, but it is logarithmic, and works the way the audio decibels do. That is, 30 is 10 times greater than 20. 10 is supposed to be a reference of a computer that can reasonnably run the game. So if you get 40, you are 1000 times powerfull enough. Negative values can technically show up on very slow computers.

4.15.4 bench-value

[Command-line option] --bench-value=<value>

[Environment variable] LW6_BENCH_VALUE

[XML key] bench-value

Type: integer Default value: LW6LDR DEFAULT BENCH VALUE Contains the current bench value of the computer running the game. This is used internally to choose the right map settings. You can override this value and use your own but... use at your own risk. Pretending you have a faster computer than what you really have can lead to confusion.

4.15.5 bin-id

[Command-line option] --bin-id=<value> [Environment variable] LW6_BIN_ID [XML key] bin-id

Type: integer Default value: 0 The internal ’bin-id’ value. Note that this is not necessarly equal to the value returned by ’show-build-bin-id’. When they are different, it is assumed this is because of a software upgrade.

4.15.6 check

[Command-line option] --check

Running the game with ’–check’ is almost like running ’–test’, the difference is that ’–check’ will not run tests which involve graphics or sound backends, so it’s adapted to pure console mode. This can be usefull for automated checks on a build farm, or if you want to check things in a headless (pure console) environment.

4.15.7 commands-per-sec

[Command-line option] --commands-per-sec=<value>

[Environment variable] LW6_COMMANDS_PER_SEC

[XML key] commands-per-sec

Type: integer Default value: 10 Min value: 1 Max value: 1000 Defines the number of commands per second. When a command is generated, orders are actually sent to the game engine, for instance, ’this cursor moved there’. So this option will affect game responsiveness, setting this to a high value will make the game more responsive but consume bandwidth on network games.

4.15.8 cunit

[Command-line option] --cunit

Running the game with ’–cunit’ is almost like running ’–test’, the difference is that ’–cunit’ will use CUnit interactive interface, allowing the user to cherry-pick some tests, and avoid running the whole suite just for one test. This can be usefull for debugging, when individual test binaries are not available.

4.15.9 daemon

[Command-line option] --daemon

Start the game in daemon mode, this is typically used with the server mode, if you want the process to be detached from the console and executed in the background.

4.15.10 debug-layer-id

[Command-line option] --debug-layer-id=<value>

[Environment variable] LW6_DEBUG_LAYER_ID

[XML key] debug-layer-id

Type: integer Default value: 0 Min value: 0 Max value: 6 A team id which will be used for debugging purposes, for instance when displaying gradient.

4.15.11 debug-team-id

[Command-line option] --debug-team-id=<value>

[Environment variable] LW6_DEBUG_TEAM_ID

[XML key] debug-team-id

Type: integer Default value: 0 Min value: 0 Max value: 9 A team id which will be used for debugging purposes, for instance when displaying gradient.

4.15.12 demo

[Command-line option] --demo

Start the game in demo mode. 2 bots play against each other forever.

4.15.13 dialog-timeout

[Command-line option] --dialog-timeout=<value>

[Environment variable] LW6_DIALOG_TIMEOUT

[XML key] dialog-timeout

Type: integer Default value: 3600 Min value: 0 Max value: 86400 Timeout, in seconds, after which a dialog will automatically be closed, wether user clicked on it or not. Mostly used for testing, to avoid program being stall on a visual prompt. 0 will simply disable this feature and wait forever. Note that some platforms might not support this. Interfaces using Gtk do support it.

4.15.14 dirty-read

[Command-line option] --dirty-read=<value>

[Environment variable] LW6_DIRTY_READ

[XML key] dirty-read

Type: integer Default value: 2 Min value: 0 Max value: 2 How to handle dirty reads and locks when displaying stuff. If set to 0, there will be no dirty reads at all, a lock (mutex) will be set whenever it’s needed. If set to 1, display might be done with inconsistent data, however the data itself won’t be modified while displaying. If set to 2, displayed data can (and will) be modified while the rendering thread is running.

4.15.15 display-background

[Command-line option] --display-background=<value>
[Environment variable] LW6_DISPLAY_BACKGROUND
[XML key] display-background

Type: boolean Default value: true Decides wether the background animation/image should be displayed at all.

4.15.16 display-console

[Command-line option] --display-console=<value>

[Environment variable] LW6_DISPLAY_CONSOLE

[XML key] display-console

Type: boolean Default value: false Defines wether the interactive system console must be displayed. Note that console support must have been enabled at compilation time. It might not be available on your computer, for instance if you are running a system such as Microsoft Windows.

4.15.17 display-cursors

[Command-line option] --display-cursors=<value>

[Environment variable] LW6_DISPLAY_CURSORS

[XML key] display-cursors

Type: boolean

Default value: true

Debugging option which can be set to ’false’ to disable the display of cursors when playing.

4.15.18 display-debug-gradient

[Command-line option] --display-debug-gradient=<value>
[Environment variable] LW6_DISPLAY_DEBUG_GRADIENT
[XML key] display-debug-gradient

Type: boolean

Default value: false

Set this to ’true’ to display the gradient, this is usefull to debug the core algorithm or understand how it works.

4.15.19 display-debug-zones

[Command-line option] --display-debug-zones=<value>
[Environment variable] LW6_DISPLAY_DEBUG_ZONES
[XML key] display-debug-zones

Type: boolean

Default value: false

Set this to ’true’ to display the zones, this is usefull to debug the core algorithm or understand how it works.

4.15.20 display-fighters

[Command-line option] --display-fighters=<value>

[Environment variable] LW6_DISPLAY_FIGHTERS

[XML key] display-fighters

Type: boolean

Default value: true

Debugging option which can be set to ’false’ to disable the display of fighters when playing.

4.15.21 display-fps

[Command-line option] --display-fps=<value>

[Environment variable] LW6_DISPLAY_FPS

[XML key] display-fps

Type: boolean Default value: false Set this to ’true’ to display the number of frames per second. When this gets too low... play a smaller map, buy a new computer or contribute and hack Liquid War 6 so that it runs faster!

4.15.22 display-hud

[Command-line option] --display-hud=<value>

[Environment variable] LW6_DISPLAY_HUD

[XML key] display-hud

Type: boolean Default value: true Decides wether the hud (informations while playing) should be displayed.

4.15.23 display-log

[Command-line option] --display-log=<value>

[Environment variable] LW6_DISPLAY_LOG

[XML key] display-log

Type: boolean Default value: true Set this to ’false’ to disable the display of error messages on the screen. Mote that you can miss valuable informations.

4.15.24 display-map

[Command-line option] --display-map=<value>

[Environment variable] LW6_DISPLAY_MAP

[XML key] display-map

Type: boolean Default value: true Debugging option which can be set to ’false’ to disable map (level) display when playing.

4.15.25 display-menu

[Command-line option] --display-menu=<value>

[Environment variable] LW6_DISPLAY_MENU

[XML key] display-menu

Type: boolean Default value: true Debugging option which can be set to ’false’ to disable the display of menus.

4.15.26 display-meta

[Command-line option] --display-meta=<value>

[Environment variable] LW6_DISPLAY_META

[XML key] display-meta

Type: boolean Default value: true Set to ’false’ to disable the display of meta information, this includes the help, tootips and breadcrumbs in menus.

4.15.27 display-mouse

[Command-line option] --display-mouse=<value>

[Environment variable] LW6_DISPLAY_MOUSE

[XML key] display-mouse

Type: boolean Default value: true Set this to ’false’ to always hide the mouse pointer.

4.15.28 display-mps

[Command-line option] --display-mps=<value>

[Environment variable] LW6_DISPLAY_MPS

[XML key] display-mps

Type: boolean Default value: false Set this to ’true’ to display the number of moves per second. In theory the game should maintain this constant but in practise it can get low if your computer is too slow or too busy.

4.15.29 display-preview

[Command-line option] --display-preview=<value>

[Environment variable] LW6_DISPLAY_PREVIEW

[XML key] display-preview

Type: boolean Default value: true Decides wether a map preview should be displayed when choosing a level.

4.15.30 display-progress

[Command-line option] --display-progress=<value>

[Environment variable] LW6_DISPLAY_PROGRESS

[XML key] display-progress

Type: boolean Default value: true Decides wether a progress bar should be displayed when a long operation is realized as a background task.

4.15.31 display-score

[Command-line option] --display-score=<value>

[Environment variable] LW6_DISPLAY_SCORE

[XML key] display-score

Type: boolean Default value: true Decides wether the score screen should be displayed.

4.15.32 display-splash

[Command-line option] --display-splash=<value>

[Environment variable] LW6_DISPLAY_SPLASH

[XML key] display-splash

Type: boolean Default value: true Set this to ’false’ to disable the display of the splash screen at game startup.

4.15.33 display-url

[Command-line option] --display-url=<value>

[Environment variable] LW6_DISPLAY_URL

[XML key] display-url

Type: boolean Default value: false Set this to ’true’ to display the URL (homepage) of the game. This is mostly used when doing screenshots, so that generated images contain a link to the homepage.

4.15.34 executed-again

[Command-line option] --executed-again=<value>

[Environment variable] LW6_EXECUTED_AGAIN

[XML key] executed-again

Type: boolean Default value: false This environment variable/keyword is used to detect wether the program has been launched by itself with an internal execv call. This is used as a workarround to set some environment variables (DYLD LIBRARY PATH on Mac OS X for instance) before the program is run, as sometimes using setenv() inside the program does not work.

4.15.35 gfx-cpu-usage

[Command-line option] --gfx-cpu-usage=<value>

[Environment variable] LW6_GFX_CPU_USAGE

[XML key] gfx-cpu-usage

Type: float Default value: 0.75 Min value: 0 Max value: 1 Percentage of the CPU which will be used by the display thread. It’s wise to leave some time to other threads to execute. The OS does it naturally, but setting this helps the whole process by explicitely pausing (sleep call) the display thread. You could change this to a low value if you have lagging games but smooth display.

4.15.36 gfx-debug

[Command-line option] --gfx-debug=<value> [Environment variable] LW6_GFX_DEBUG

[XML key] gfx-debug

Type: boolean Default value: false Enables dedicated graphics debugging tools. This is different from ’debug’ mode which is global, this one is really graphics specific.

4.15.37 io-per-sec

[Command-line option] --io-per-sec=<value>

[Environment variable] LW6_IO_PER_SEC

[XML key] io-per-sec

Type: integer Default value: 20 Min value: 1 Max value: 1000 Defines the number of calls to input/output functions per second. This can affect speed of menus but also cursors, but won’t change the speed of the game itself. It’s a cosmectic, comfort option.

4.15.38 jpeg-quality

[Command-line option] --jpeg-quality=<value>

[Environment variable] LW6_JPEG_QUALITY

[XML key] jpeg-quality

Type: integer Default value: 85 Min value: 0 Max value: 85 Quality used by libjpeg when creating screenshot images. The same value you would give to Gimp before exporting an image as a JPEG.

4.15.39 loader-sleep

[Command-line option] --loader-sleep=<value>

[Environment variable] LW6_LOADER_SLEEP

[XML key] loader-sleep

Type: float Default value: 0.5 Defines how long the loader thread should wait between two polls. Default value should fit in most cases.

4.15.40 local-bench-delta

[Command-line option] --local-bench-delta=<value>
[Environment variable] LW6_LOCAL_BENCH_DELTA
[XML key] local-bench-delta

Type: integer Default value: 0 Min value: -70 Max value: 20 A value which is added to bench before starting a local game. This is typically zero or negative, as adding to bench is like pretending your computer is faster than it really is.

4.15.41 log-level

[Command-line option] --log-level=<value> [Environment variable] LW6_LOG_LEVEL

[XML key] log-level

Type: integer Default value: 3 Min value: 0 Max value: 4 Defines the log level, that is, how verbose the program will be regarding logs and console output. 0 (ERROR) is the minimum, only errors are reported. 1 (WARNING) means errors + warnings. 2 (NOTICE) displays most important messages. 3 (INFO) is the default, the log file will contain all messages but debug stuff. 4 (DEBUG) logs everything, including debug informations.

4.15.42 log-timeout

[Command-line option] --log-timeout=<value>

[Environment variable] LW6_LOG_TIMEOUT

[XML key] log-timeout

Type: integer Default value: 5000 Delay, in msec, for which a log message will stay displayed on the screen.

4.15.43 magic-number

[Command-line option] --magic-number=<value>

[Environment variable] LW6_MAGIC_NUMBER

[XML key] magic-number

Type: integer Default value: LW6LDR DEFAULT MAGIC NUMBER This ’magic’ number probably requires an explanation. It’s used to estimate how big a map can be built. The calculus is very approximative, basically bench value*magic number=total fighters on map*rounds per sec*moves per round with total fighters on map depending on various parameters such as map size but also how many fighters are on the map. The map loader will try and adjust the map size so that it is just big enough not to saturate your CPU while being as high-res as possible. The magic number in itself has no real meaning, the higher it gets, the more optimized it means the game is. Normally you shouldn’t change this but if you find the map resizing is too agressively pessimistic, or if for some reason bench returns bogus values, you can modify it.

4.15.44 max-local-bench-value

[Command-line option] --max-local-bench-value=<value>
[Environment variable] LW6_MAX_LOCAL_BENCH_VALUE
[XML key] max-local-bench-value

Type: integer

Default value: LW6LDR DEFAULT MAX LOCAL BENCH VALUE

Even if your computer is very fast, this parameter will be used to tame the optimism of the test, and do not load maps in very high detail. It’s believed at some point, it’s best to keep your extra power to deal with unordinary situations rather than waste it on useless details. Game should be fun with that setting, but if you really want to use your shiny CPU at its maximum, raise this.

4.15.45 max-network-bench-value

[Command-line option] --max-network-bench-value=<value>
[Environment variable] LW6_MAX_NETWORK_BENCH_VALUE
[XML key] max-network-bench-value

Type: integer

Default value: LW6LDR DEFAULT MAX NETWORK BENCH VALUE

On network games, we need to be sure everyone can play in correct conditions, therefore maps won’t be loaded with more details than this, by default. You’re free to increase this parameter but it can cause your games to be unjoignable by some people.

4.15.46 memory-bazooka-eraser

[Command-line option] --memory-bazooka-eraser=<value>
[Environment variable] LW6_MEMORY_BAZOOKA_ERASER
[XML key] memory-bazooka-eraser

Type: boolean

Default value: true

The memory eraser is a tool which will systematically fill allocated memory with ’M’, and overwrite all allocated bytes with ’F’ before freeing memory. It will even handle realloc calls. This is usefull to track bugs. Indeed, with this option enabled, freshly allocated memory will never contain zeroes unless one calls calloc, and if you ever free some memory zone before being done with it, it will be filled with junk and therefore not be usable. The memory bazooka must be big enough if you want this feature to actually work.

4.15.47 memory-bazooka-size

[Command-line option] --memory-bazooka-size=<value>
[Environment variable] LW6_MEMORY_BAZOOKA_SIZE
[XML key] memory-bazooka-size

Type: integer Default value: 99991 The memory bazooka is a brute-force tool, conceived after a full night spent tracking some memory leak. The idea is to keep a track of all allocated pointers, when the data was allocated (timestamp), where in the code (file, line), and even point out what data there is in that place. A memory bazooka report at the end of the game will just show what’s left. There should be nothing. This parameter is here to avoid wasting CPU cycles on a feature which is very debug-oriented and does not really make sense for the casual user. Set it to 0 for best performance, something like 100 might just be helpfull, but 1000000 is the right way to seriously debug code.

4.15.48 net-log

[Command-line option] --net-log=<value> [Environment variable] LW6_NET_LOG [XML key] net-log

Type: boolean Default value: false Activates network log, that is, logs everything sent/received over the network, except data which is sent through a third party library such as libCurl. This is mostly for debugging purpose, it can lead to rather big log files.

4.15.49 net-per-sec

[Command-line option] --net-per-sec=<value>

[Environment variable] LW6_NET_PER_SEC

[XML key] net-per-sec

Type: integer Default value: 500 Min value: 1 Max value: 1000 Defines the number of calls to network functions per second. This can technically change the network transfers speed, the higher the number, the faster it should be, but at the same time it can technically be more CPU greedy.

4.15.50 network-bench-delta

[Command-line option] --network-bench-delta=<value>
[Environment variable] LW6_NETWORK_BENCH_DELTA
[XML key] network-bench-delta

Type: integer Default value: -5 Min value: -70 Max value: 20 A value which is added to bench before starting a network game. This is typically a negative value, lower than the one added to local game. This is because network games can be more CPU greedy.

4.15.51 network-reliability

[Command-line option] --network-reliability=<value>
[Environment variable] LW6_NETWORK_RELIABILITY
[XML key] network-reliability

Type: integer Default value: 1000 Min value: 1 Max value: 1000000000 The program assumes network is non-reliable, however the problem with those assumptions is that when you test, network is always reliable, even with non-garanteed protocols like UDP. This option will force the program to actually ignore some calls to send or recv functions, simulating a network disfunction. This is to ensure the internal mecanisms correcting network problems do work for good, on daily regular use. It’s not possible to set it to a perfect behavior, never dropping any packet, however using the default settings you probably won’t even notice the performance drop induced by having to fix problems. The highest the number is, the most reliable network will look, the algorithm is simply to drop one message out of X.

4.15.52 open-relay

[Command-line option] --open-relay=<value>

[Environment variable] LW6_OPEN_RELAY

[XML key] open-relay

Type: boolean Default value: false Enables forwarding of abritrary network messages. If open relay is forbidden, the game will only forward messages when physical sender and logical sender are the same. This is to say if messages come from A for C and is sent by A to B, B will forward it to C. But if message comes from X to C and is sent by A to B, then B won’t forward it. In practice, it means without open relay, messages can only be forwarded once.

4.15.53 pilot-lag

[Command-line option] --pilot-lag=<value> [Environment variable] LW6_PILOT_LAG

[XML key] pilot-lag

Type: integer Default value: 10 Maximum lag, in rounds, until the game engine is slowed down. This will typically be usefull if your computer is too slow for the map resolution and the game speed you set up.

4.15.54 quick-start

[Command-line option] --quick-start

Start the game just like if the player had requested a quick start, without showing any menu.

4.15.55 reset

[Command-line option] --reset

Clears the config file so that the game will run with defaults next time. The idea is to get rid of traces of previous executions. The difference with ’–defaults’ is that ’–reset’ does not run the game, while ’–defaults’ does.

4.15.56 reset-config-on-upgrade

[Command-line option] --reset-config-on-upgrade=<value>
[Environment variable] LW6_RESET_CONFIG_ON_UPGRADE
[XML key] reset-config-on-upgrade

Type: boolean Default value: true If set, then a reset (config file set to defaults) is run every time you upgrade the game.

4.15.57 screenshots-per-min

[Command-line option] --screenshots-per-min=<value>
[Environment variable] LW6_SCREENSHOTS_PER_MIN
[XML key] screenshots-per-min

Type: integer Default value: 12 Defines the number of screenshots / node info per minute. This can a quite costly operation, but still it must not be too low else screenshots are too outdated.

4.15.58 server

[Command-line option] --server

Start the game in server mode, without requiring any graphics backend. Server mode is usefull if you just want to start a network node without hosting any real game on it. It can be used to list existing nodes and sessions or as a bounce server in case some clients can’t contact each other because firewalled. If you only want to start a server game on your computer, don’t use this option, just start the game normally and start a game server by clicking on the GUI buttons.

4.15.59 simulate-basic

[Command-line option] --simulate-basic

Simulates some fights using the basic colors red, green, yellow and blue. Will output on the console a percentage based on scores obtained by the teams. This is typically for map designers and/or people who want to fiddle with team profiles, if some team is really stronger than another one, it should appear in these percentages.

4.15.60 simulate-full

[Command-line option] --simulate-full

Simulates some fights using all available colors. This can be very long, it will run approximatively 1000 games consecutively, you can look in the log file to see the progress. Will output on the console a percentage based on scores obtained by the teams. This is typically for map designers and/or people who want to fiddle with team profiles, if some team is really stronger than another one, it should appear in these percentages.

4.15.61 target-fps

[Command-line option] --target-fps=<value>

[Environment variable] LW6_TARGET_FPS

[XML key] target-fps

Type: integer Default value: 60 Min value: 1 Max value: 1000 Defines how many frames will be displayed per second. Of course this is a maximum value, if your hardware can’t keep up with this value, display will just be slow, no matter what value you define here. Note that you might really wish to have something rather low here, to keep network and ’logic’ function responsiveness. Passed 60 frames per second, speed is really only for visual comfort, as Liquid War 6 is now so fast-paced that it requires 200 frames/sec to outperform opponents.

4.15.62 trap-errors

[Command-line option] --trap-errors=<value>

[Environment variable] LW6_TRAP_ERRORS

[XML key] trap-errors

Type: boolean Default value: false If set to true, will trap segmentation fault and floating point errors, and display messages about those in a custom box instead of the default one

4.15.63 trojan

[Command-line option] --trojan=<value> [Environment variable] LW6_TROJAN [XML key] trojan

Type: boolean Default value: false Make the program act like a (stupid) trojan horse, trying to fake messages, sending various inconsistent informations. This is to check the normal version of the program is able to detect such a fake and kick it out of the game. It’s of no use for regular players, be sure to unset this if you want to play for good.

4.15.64 z-decode

[Command-line option] --z-decode

If specified, program will take stdin and z-decode it to stdout. This is for testing purpose (for network messages for instance). Z-decoding, here means verifying there a Z at the beginning, base64 decode and pass the content through Zlib inflating. I content is not Z-prefixed, will be returned as is.

4.15.65 z-encode

[Command-line option] --z-encode

If specified, program will take stdin and z-encode it to stdout. This is for testing purpose (for network messages for instance). Z-encoding, here means passing the message through Zlib deflating then base64 encoding and prefix it with a Z.

4.16 C to Guile

4.16.1 c-gettext

[C function exported to Guile] c-gettext

Calls GNU gettext to convert string in current locale. Note that ’ ’ (plain underscode) is exported as well, so that code can be written using ’ ’ as a function.

4.16.2 c-lw6-exit

[C function exported to Guile] c-lw6-exit

Wrapper on lw6 exit.

4.16.3 c-lw6-get-ret

[C function exported to Guile] c-lw6-get-ret

Wrapper on lw6 get ret.

4.16.4 c-lw6-release

[C function exported to Guile] c-lw6-release

Wrapper on lw6 release.

4.16.5 c-lw6-set-ret

[C function exported to Guile] c-lw6-set-ret

Wrapper on lw6 set ret.

4.16.6 c-lw6bot-get-backends

[C function exported to Guile] c-lw6bot-get-backends

Wrapper on lw6bot get backends.

4.16.7 c-lw6bot-new

[C function exported to Guile] c-lw6bot-new

Wrapper on lw6bot new.

4.16.8 c-lw6bot-next-move

[C function exported to Guile] c-lw6bot-next-move

Wrapper on lw6bot next move.

4.16.9 c-lw6cfg-defaults

[C function exported to Guile] c-lw6cfg-defaults

Wrapper on lw6cfg defaults.

4.16.10 c-lw6cfg-get-option

[C function exported to Guile] c-lw6cfg-get-option

Wrapper on lw6cfg get option.

4.16.11 c-lw6cfg-init

[C function exported to Guile] c-lw6cfg-init

Wrapper on lw6cfg init.

4.16.12 c-lw6cfg-load

[C function exported to Guile] c-lw6cfg-load

Wrapper on lw6cfg load.

4.16.13 c-lw6cfg-option-exists

[C function exported to Guile] c-lw6cfg-option-exists

Wrapper on lw6cfg option exists.

4.16.14 c-lw6cfg-quit

[C function exported to Guile] c-lw6cfg-quit

Wrapper on lw6cfg quit.

4.16.15 c-lw6cfg-save

[C function exported to Guile] c-lw6cfg-save

Wrapper on lw6cfg save.

4.16.16 c-lw6cfg-set-option

[C function exported to Guile] c-lw6cfg-set-option

Wrapper on lw6cfg set option.

4.16.17 c-lw6cfg-unified-get-log-file

[C function exported to Guile] c-lw6cfg-unified-get-log-file

Wrapper on lw6cfg unified get log file.

4.16.18 c-lw6cfg-unified-get-map-path

[C function exported to Guile] c-lw6cfg-unified-get-map-path

Wrapper on lw6cfg unified get map path.

4.16.19 c-lw6cfg-unified-get-music-path

[C function exported to Guile] c-lw6cfg-unified-get-music-path

Wrapper on lw6cfg unified get music path.

4.16.20 c-lw6cfg-unified-get-user-dir

[C function exported to Guile] c-lw6cfg-unified-get-user-dir

Wrapper on lw6cfg unified get user dir.

4.16.21 c-lw6cli-get-backends

[C function exported to Guile] c-lw6cli-get-backends

Wrapper on lw6cli get backends.

4.16.22 c-lw6cns-console-support

[C function exported to Guile] c-lw6cns-console-support

Wrapper on lw6cns console support.

4.16.23 c-lw6cns-init

[C function exported to Guile] c-lw6cns-init

Wrapper on lw6cns init.

4.16.24 c-lw6cns-poll

[C function exported to Guile] c-lw6cns-poll

Wrapper on lw6cns poll.

4.16.25 c-lw6cns-quit

[C function exported to Guile] c-lw6cns-quit

Wrapper on lw6cns quit.

4.16.26 c-lw6cns-term-support

[C function exported to Guile] c-lw6cns-term-support

Wrapper on lw6cns term support.

4.16.27 c-lw6dsp-get-average-fps

[C function exported to Guile] c-lw6dsp-get-average-fps

Wrapper on lw6dsp get average fps.

4.16.28 c-lw6dsp-get-fullscreen-modes

[C function exported to Guile] c-lw6dsp-get-fullscreen-modes

Wrapper on lw6dsp get fullscreen modes.

4.16.29 c-lw6dsp-get-instant-fps

[C function exported to Guile] c-lw6dsp-get-instant-fps

Wrapper on lw6dsp get instant fps.

4.16.30 c-lw6dsp-get-last-frame-rendering-time

[C function exported to Guile] c-lw6dsp-get-last-frame-rendering-time

Wrapper on lw6dsp get last frame rendering time.

4.16.31 c-lw6dsp-get-nb-frames

[C function exported to Guile] c-lw6dsp-get-nb-frames

Wrapper on lw6dsp get nb frames.

4.16.32 c-lw6dsp-get-video-mode

[C function exported to Guile] c-lw6dsp-get-video-mode

Wrapper on lw6dsp get video mode.

4.16.33 c-lw6dsp-new

[C function exported to Guile] c-lw6dsp-new

Wrapper on lw6dsp new.

4.16.34 c-lw6dsp-release

[C function exported to Guile] c-lw6dsp-release

Wrapper on lw6dsp release.

4.16.35 c-lw6dsp-update

[C function exported to Guile] c-lw6dsp-update

Wrapper on lw6dsp update.

4.16.36 c-lw6gen-create-from-seed

[C function exported to Guile] c-lw6gen-create-from-seed

Wrapper on lw6gen create from seed.

4.16.37 c-lw6gen-seed-new

[C function exported to Guile] c-lw6gen-seed-new

Wrapper on lw6gen seed new.

4.16.38 c-lw6gen-seed-normalize

[C function exported to Guile] c-lw6gen-seed-normalize

Wrapper on lw6gen seed normalize.

4.16.39 c-lw6gfx-get-backends

[C function exported to Guile] c-lw6gfx-get-backends

Wrapper on lw6gfx get backends.

4.16.40 c-lw6gui-default-look

[C function exported to Guile] c-lw6gui-default-look

Wrapper on lw6gui default look.

4.16.41 c-lw6gui-input-reset

[C function exported to Guile] c-lw6gui-input-reset

Wrapper on lw6gui input reset.

4.16.42 c-lw6gui-joystick1-get-move-pad

[C function exported to Guile] c-lw6gui-joystick1-get-move-pad

Wrapper on lw6gui joystick1 get move pad.

4.16.43 c-lw6gui-joystick1-pop-button-a

[C function exported to Guile] c-lw6gui-joystick1-pop-button-a

Wrapper on lw6gui joystick1 pop button a.

4.16.44 c-lw6gui-joystick1-pop-button-b

[C function exported to Guile] c-lw6gui-joystick1-pop-button-b

Wrapper on lw6gui joystick1 pop button b.

4.16.45 c-lw6gui-joystick1-pop-button-c

[C function exported to Guile] c-lw6gui-joystick1-pop-button-c

Wrapper on lw6gui joystick1 pop button c.

4.16.46 c-lw6gui-joystick1-pop-button-d

[C function exported to Guile] c-lw6gui-joystick1-pop-button-d

Wrapper on lw6gui joystick1 pop button d.

4.16.47 c-lw6gui-joystick1-pop-button-e

[C function exported to Guile] c-lw6gui-joystick1-pop-button-e

Wrapper on lw6gui joystick1 pop button e.

4.16.48 c-lw6gui-joystick1-pop-button-f

[C function exported to Guile] c-lw6gui-joystick1-pop-button-f

Wrapper on lw6gui joystick1 pop button f.

4.16.49 c-lw6gui-joystick1-pop-pad-down

[C function exported to Guile] c-lw6gui-joystick1-pop-pad-down

Wrapper on lw6gui joystick1 pop pad down.

4.16.50 c-lw6gui-joystick1-pop-pad-left

[C function exported to Guile] c-lw6gui-joystick1-pop-pad-left

Wrapper on lw6gui joystick1 pop pad left.

4.16.51 c-lw6gui-joystick1-pop-pad-right

[C function exported to Guile] c-lw6gui-joystick1-pop-pad-right

Wrapper on lw6gui joystick1 pop pad right.

4.16.52 c-lw6gui-joystick1-pop-pad-up

[C function exported to Guile] c-lw6gui-joystick1-pop-pad-up

Wrapper on lw6gui joystick1 pop pad up.

4.16.53 c-lw6gui-joystick2-get-move-pad

[C function exported to Guile] c-lw6gui-joystick2-get-move-pad

Wrapper on lw6gui joystick2 get move pad.

4.16.54 c-lw6gui-joystick2-pop-button-a

[C function exported to Guile] c-lw6gui-joystick2-pop-button-a

Wrapper on lw6gui joystick2 pop button a.

4.16.55 c-lw6gui-joystick2-pop-button-b

[C function exported to Guile] c-lw6gui-joystick2-pop-button-b

Wrapper on lw6gui joystick2 pop button b.

4.16.56 c-lw6gui-joystick2-pop-button-c

[C function exported to Guile] c-lw6gui-joystick2-pop-button-c

Wrapper on lw6gui joystick2 pop button c.

4.16.57 c-lw6gui-joystick2-pop-button-d

[C function exported to Guile] c-lw6gui-joystick2-pop-button-d

Wrapper on lw6gui joystick2 pop button d.

4.16.58 c-lw6gui-joystick2-pop-button-e

[C function exported to Guile] c-lw6gui-joystick2-pop-button-e

Wrapper on lw6gui joystick2 pop button e.

4.16.59 c-lw6gui-joystick2-pop-button-f

[C function exported to Guile] c-lw6gui-joystick2-pop-button-f

Wrapper on lw6gui joystick2 pop button f.

4.16.60 c-lw6gui-joystick2-pop-pad-down

[C function exported to Guile] c-lw6gui-joystick2-pop-pad-down

Wrapper on lw6gui joystick2 pop pad down.

4.16.61 c-lw6gui-joystick2-pop-pad-left

[C function exported to Guile] c-lw6gui-joystick2-pop-pad-left

Wrapper on lw6gui joystick2 pop pad left.

4.16.62 c-lw6gui-joystick2-pop-pad-right

[C function exported to Guile] c-lw6gui-joystick2-pop-pad-right

Wrapper on lw6gui joystick2 pop pad right.

4.16.63 c-lw6gui-joystick2-pop-pad-up

[C function exported to Guile] c-lw6gui-joystick2-pop-pad-up

Wrapper on lw6gui joystick2 pop pad up.

4.16.64 c-lw6gui-keyboard-get-move-pad

[C function exported to Guile] c-lw6gui-keyboard-get-move-pad

Wrapper on lw6gui keyboard get move pad.

4.16.65 c-lw6gui-keyboard-is-pressed

[C function exported to Guile] c-lw6gui-keyboard-is-pressed

Wrapper on lw6gui keyboard is pressed.

4.16.66 c-lw6gui-keyboard-pop-arrow-down

[C function exported to Guile] c-lw6gui-keyboard-pop-arrow-down

Wrapper on lw6gui keyboard pop arrow down.

4.16.67 c-lw6gui-keyboard-pop-arrow-left

[C function exported to Guile] c-lw6gui-keyboard-pop-arrow-left

Wrapper on lw6gui keyboard pop arrow left.

4.16.68 c-lw6gui-keyboard-pop-arrow-right

[C function exported to Guile] c-lw6gui-keyboard-pop-arrow-right

Wrapper on lw6gui keyboard pop arrow right.

4.16.69 c-lw6gui-keyboard-pop-arrow-up

[C function exported to Guile] c-lw6gui-keyboard-pop-arrow-up

Wrapper on lw6gui keyboard pop arrow up.

4.16.70 c-lw6gui-keyboard-pop-key-alt

[C function exported to Guile] c-lw6gui-keyboard-pop-key-alt

Wrapper on lw6gui keyboard pop key alt.

4.16.71 c-lw6gui-keyboard-pop-key-ctrl

[C function exported to Guile] c-lw6gui-keyboard-pop-key-ctrl

Wrapper on lw6gui keyboard pop key ctrl.

4.16.72 c-lw6gui-keyboard-pop-key-enter

[C function exported to Guile] c-lw6gui-keyboard-pop-key-enter

Wrapper on lw6gui keyboard pop key enter.

4.16.73 c-lw6gui-keyboard-pop-key-esc

[C function exported to Guile] c-lw6gui-keyboard-pop-key-esc

Wrapper on lw6gui keyboard pop key esc.

4.16.74 c-lw6gui-keyboard-pop-key-pgdown

[C function exported to Guile] c-lw6gui-keyboard-pop-key-pgdown

Wrapper on lw6gui keyboard pop key pgdown.

4.16.75 c-lw6gui-keyboard-pop-key-pgup

[C function exported to Guile] c-lw6gui-keyboard-pop-key-pgup

Wrapper on lw6gui keyboard pop key pgup.

4.16.76 c-lw6gui-look-get

[C function exported to Guile] c-lw6gui-look-get

Wrapper on lw6gui look get.

4.16.77 c-lw6gui-look-set

[C function exported to Guile] c-lw6gui-look-set

Wrapper on lw6gui look set.

4.16.78 c-lw6gui-look-zoom-in

[C function exported to Guile] c-lw6gui-look-zoom-in

Wrapper on lw6gui look zoom in.

4.16.79 c-lw6gui-look-zoom-out

[C function exported to Guile] c-lw6gui-look-zoom-out

Wrapper on lw6gui look zoom out.

4.16.80 c-lw6gui-menu-append

[C function exported to Guile] c-lw6gui-menu-append

Wrapper on lw6gui menu append.

4.16.81 c-lw6gui-menu-close-popup

[C function exported to Guile] c-lw6gui-menu-close-popup

Wrapper on lw6gui menu close popup.

4.16.82 c-lw6gui-menu-enable-esc

[C function exported to Guile] c-lw6gui-menu-enable-esc

Wrapper on lw6gui menu enable esc.

4.16.83 c-lw6gui-menu-has-popup

[C function exported to Guile] c-lw6gui-menu-has-popup

Wrapper on lw6gui menu has popup.

4.16.84 c-lw6gui-menu-new

[C function exported to Guile] c-lw6gui-menu-new

Wrapper on lw6gui menu new.

4.16.85 c-lw6gui-menu-remove

[C function exported to Guile] c-lw6gui-menu-remove

Wrapper on lw6gui menu remove.

4.16.86 c-lw6gui-menu-remove-all

[C function exported to Guile] c-lw6gui-menu-remove-all

Wrapper on lw6gui menu remove all.

4.16.87 c-lw6gui-menu-scroll-down

[C function exported to Guile] c-lw6gui-menu-scroll-down

Wrapper on lw6gui menu scroll down.

4.16.88 c-lw6gui-menu-scroll-up

[C function exported to Guile] c-lw6gui-menu-scroll-up

Wrapper on lw6gui menu scroll up.

4.16.89 c-lw6gui-menu-select

[C function exported to Guile] c-lw6gui-menu-select

Wrapper on lw6gui menu select.

4.16.90 c-lw6gui-menu-select-esc

[C function exported to Guile] c-lw6gui-menu-select-esc

Wrapper on lw6gui menu select esc.

4.16.91 c-lw6gui-menu-set-breadcrumbs

[C function exported to Guile] c-lw6gui-menu-set-breadcrumbs

Wrapper on lw6gui menu set breadcrumbs.

4.16.92 c-lw6gui-menu-sync

[C function exported to Guile] c-lw6gui-menu-sync

Wrapper on lw6gui menu sync.

4.16.93 c-lw6gui-mouse-get-state

[C function exported to Guile] c-lw6gui-mouse-get-state

Wrapper on lw6gui mouse get state.

4.16.94 c-lw6gui-mouse-poll-move

[C function exported to Guile] c-lw6gui-mouse-poll-move

Wrapper on lw6gui mouse poll move.

4.16.95 c-lw6gui-mouse-pop-button-left

[C function exported to Guile] c-lw6gui-mouse-pop-button-left

Wrapper on lw6gui mouse pop button left.

4.16.96 c-lw6gui-mouse-pop-button-middle

[C function exported to Guile] c-lw6gui-mouse-pop-button-middle

Wrapper on lw6gui mouse pop button middle.

4.16.97 c-lw6gui-mouse-pop-button-right

[C function exported to Guile] c-lw6gui-mouse-pop-button-right

Wrapper on lw6gui mouse pop button right.

4.16.98 c-lw6gui-mouse-pop-double-click

[C function exported to Guile] c-lw6gui-mouse-pop-double-click

Wrapper on lw6gui mouse pop double click.

4.16.99 c-lw6gui-mouse-pop-simple-click

[C function exported to Guile] c-lw6gui-mouse-pop-simple-click

Wrapper on lw6gui mouse pop simple click.

4.16.100 c-lw6gui-mouse-pop-triple-click

[C function exported to Guile] c-lw6gui-mouse-pop-triple-click

Wrapper on lw6gui mouse pop triple click.

4.16.101 c-lw6gui-mouse-pop-wheel-down

[C function exported to Guile] c-lw6gui-mouse-pop-wheel-down

Wrapper on lw6gui mouse pop wheel down.

4.16.102 c-lw6gui-mouse-pop-wheel-up

[C function exported to Guile] c-lw6gui-mouse-pop-wheel-up

Wrapper on lw6gui mouse pop wheel up.

4.16.103 c-lw6hlp-about

[C function exported to Guile] c-lw6hlp-about

Wrapper on lw6hlp about.

4.16.104 c-lw6hlp-get-default-value

[C function exported to Guile] c-lw6hlp-get-default-value

Wrapper on lw6hlp get default value.

4.16.105 c-lw6hlp-list

[C function exported to Guile] c-lw6hlp-list

Wrapper on lw6hlp list.

4.16.106 c-lw6hlp-list-advanced

[C function exported to Guile] c-lw6hlp-list-advanced

Wrapper on lw6hlp list advanced.

4.16.107 c-lw6hlp-list-aliases

[C function exported to Guile] c-lw6hlp-list-aliases

Wrapper on lw6hlp list aliases.

4.16.108 c-lw6hlp-list-doc

[C function exported to Guile] c-lw6hlp-list-doc

Wrapper on lw6hlp list doc.

4.16.109 c-lw6hlp-list-funcs

[C function exported to Guile] c-lw6hlp-list-funcs

Wrapper on lw6hlp list funcs.

4.16.110 c-lw6hlp-list-graphics

[C function exported to Guile] c-lw6hlp-list-graphics

Wrapper on lw6hlp list graphics.

4.16.111 c-lw6hlp-list-hooks

[C function exported to Guile] c-lw6hlp-list-hooks

Wrapper on lw6hlp list hooks.

4.16.112 c-lw6hlp-list-input

[C function exported to Guile] c-lw6hlp-list-input

Wrapper on lw6hlp list input.

4.16.113 c-lw6hlp-list-map

[C function exported to Guile] c-lw6hlp-list-map

Wrapper on lw6hlp list map.

4.16.114 c-lw6hlp-list-map-hints

[C function exported to Guile] c-lw6hlp-list-map-hints

Wrapper on lw6hlp list map hints.

4.16.115 c-lw6hlp-list-map-rules

[C function exported to Guile] c-lw6hlp-list-map-rules

Wrapper on lw6hlp list map rules.

4.16.116 c-lw6hlp-list-map-style

[C function exported to Guile] c-lw6hlp-list-map-style

Wrapper on lw6hlp list map style.

4.16.117 c-lw6hlp-list-map-teams

[C function exported to Guile] c-lw6hlp-list-map-teams

Wrapper on lw6hlp list map teams.

4.16.118 c-lw6hlp-list-network

[C function exported to Guile] c-lw6hlp-list-network

Wrapper on lw6hlp list network.

4.16.119 c-lw6hlp-list-path

[C function exported to Guile] c-lw6hlp-list-path

Wrapper on lw6hlp list path.

4.16.120 c-lw6hlp-list-players

[C function exported to Guile] c-lw6hlp-list-players

Wrapper on lw6hlp list players.

4.16.121 c-lw6hlp-list-quick

[C function exported to Guile] c-lw6hlp-list-quick

Wrapper on lw6hlp list quick.

4.16.122 c-lw6hlp-list-show

[C function exported to Guile] c-lw6hlp-list-show

Wrapper on lw6hlp list show.

4.16.123 c-lw6hlp-list-sound

[C function exported to Guile] c-lw6hlp-list-sound

Wrapper on lw6hlp list sound.

4.16.124 c-lw6hlp-list-team-colors

[C function exported to Guile] c-lw6hlp-list-team-colors

Wrapper on lw6hlp list team colors.

4.16.125 c-lw6hlp-list-weapons

[C function exported to Guile] c-lw6hlp-list-weapons

Wrapper on lw6hlp list weapons.

4.16.126 c-lw6img-screenshot

[C function exported to Guile] c-lw6img-screenshot

Wrapper on lw6img screenshot.

4.16.127 c-lw6ker-add-cursor

[C function exported to Guile] c-lw6ker-add-cursor

Wrapper on lw6ker add cursor.

4.16.128 c-lw6ker-build-game-state

[C function exported to Guile] c-lw6ker-build-game-state

Wrapper on lw6ker build game state.

4.16.129 c-lw6ker-build-game-struct

[C function exported to Guile] c-lw6ker-build-game-struct

Wrapper on lw6ker build game struct.

4.16.130 c-lw6ker-cursor-exists

[C function exported to Guile] c-lw6ker-cursor-exists

Wrapper on lw6ker cursor exists.

4.16.131 c-lw6ker-did-cursor-win

[C function exported to Guile] c-lw6ker-did-cursor-win

Wrapper on lw6ker did cursor win.

4.16.132 c-lw6ker-do-round

[C function exported to Guile] c-lw6ker-do-round

Wrapper on lw6ker do round.

4.16.133 c-lw6ker-dup-game-state

[C function exported to Guile] c-lw6ker-dup-game-state

Wrapper on lw6ker dup game state.

4.16.134 c-lw6ker-game-state-checksum

[C function exported to Guile] c-lw6ker-game-state-checksum

Wrapper on lw6ker game state checksum.

4.16.135 c-lw6ker-game-struct-checksum

[C function exported to Guile] c-lw6ker-game-struct-checksum

Wrapper on lw6ker game struct checksum.

4.16.136 c-lw6ker-get-cursor

[C function exported to Guile] c-lw6ker-get-cursor

Wrapper on lw6ker get cursor.

4.16.137 c-lw6ker-get-moves

[C function exported to Guile] c-lw6ker-get-moves

Wrapper on lw6ker get moves.

4.16.138 c-lw6ker-get-nb-colors

[C function exported to Guile] c-lw6ker-get-nb-colors

Wrapper on lw6ker game state get nb colors.

4.16.139 c-lw6ker-get-nb-cursors

[C function exported to Guile] c-lw6ker-get-nb-cursors

Wrapper on lw6ker game state get nb cursors.

4.16.140 c-lw6ker-get-nb-nodes

[C function exported to Guile] c-lw6ker-get-nb-nodes

Wrapper on lw6ker game state get nb nodes.

4.16.141 c-lw6ker-get-rounds

[C function exported to Guile] c-lw6ker-get-rounds

Wrapper on lw6ker get rounds.

4.16.142 c-lw6ker-get-spreads

[C function exported to Guile] c-lw6ker-get-spreads

Wrapper on lw6ker get spreads.

4.16.143 c-lw6ker-is-over

[C function exported to Guile] c-lw6ker-is-over

Wrapper on lw6ker is over.

4.16.144 c-lw6ker-node-exists

[C function exported to Guile] c-lw6ker-node-exists

Wrapper on lw6ker node exists.

4.16.145 c-lw6ker-register-node

[C function exported to Guile] c-lw6ker-register-node

Wrapper on lw6ker register node.

4.16.146 c-lw6ker-remove-cursor

[C function exported to Guile] c-lw6ker-remove-cursor

Wrapper on lw6ker remove cursor.

4.16.147 c-lw6ker-set-cursor

[C function exported to Guile] c-lw6ker-set-cursor

Wrapper on lw6ker set cursor.

4.16.148 c-lw6ker-sync-game-state

[C function exported to Guile] c-lw6ker-sync-game-state

Wrapper on lw6ker sync game state.

4.16.149 c-lw6ker-unregister-node

[C function exported to Guile] c-lw6ker-unregister-node

Wrapper on lw6ker unregister node.

4.16.150 c-lw6ldr-chain-entry

[C function exported to Guile] c-lw6ldr-chain-entry

Wrapper on lw6ldr chain entry.

4.16.151 c-lw6ldr-exp-validate

[C function exported to Guile] c-lw6ldr-exp-validate

Wrapper on lw6ldr exp validate.

4.16.152 c-lw6ldr-get-entries

[C function exported to Guile] c-lw6ldr-get-entries

Wrapper on lw6ldr get entries.

4.16.153 c-lw6ldr-hints-get-default

[C function exported to Guile] c-lw6ldr-hints-get-default

Wrapper on lw6ldr hints get default.

4.16.154 c-lw6ldr-print-examples

[C function exported to Guile] c-lw6ldr-print-examples

Wrapper on lw6ldr print examples.

4.16.155 c-lw6ldr-read

[C function exported to Guile] c-lw6ldr-read

Wrapper on lw6ldr read.

4.16.156 c-lw6ldr-read-relative

[C function exported to Guile] c-lw6ldr-read-relative

Wrapper on lw6ldr read relative.

4.16.157 c-lw6map-exp-get-unlocked-team-color

[C function exported to Guile] c-lw6map-exp-get-unlocked-team-color

Wrapper on lw6map exp get unlocked team color.

4.16.158 c-lw6map-exp-get-unlocked-weapon

[C function exported to Guile] c-lw6map-exp-get-unlocked-weapon

Wrapper on lw6map exp get unlocked weapon.

4.16.159 c-lw6map-exp-is-team-color-allowed

[C function exported to Guile] c-lw6map-exp-is-team-color-allowed

Wrapper on lw6map exp is team color allowed.

4.16.160 c-lw6map-exp-is-weapon-allowed

[C function exported to Guile] c-lw6map-exp-is-weapon-allowed

Wrapper on lw6map exp is weapon allowed.

4.16.161 c-lw6map-get-look

[C function exported to Guile] c-lw6map-get-look

Wrapper on lw6map get look.

4.16.162 c-lw6map-get-max-nb-colors

[C function exported to Guile] c-lw6map-get-max-nb-colors

Wrapper on lw6map get max nb colors.

4.16.163 c-lw6map-get-max-nb-cursors

[C function exported to Guile] c-lw6map-get-max-nb-cursors

Wrapper on lw6map get max nb cursors.

4.16.164 c-lw6map-get-max-nb-nodes

[C function exported to Guile] c-lw6map-get-max-nb-nodes

Wrapper on lw6map get max nb nodes.

4.16.165 c-lw6map-get-music-dir

[C function exported to Guile] c-lw6map-get-music-dir

Wrapper on lw6map get music dir.

4.16.166 c-lw6map-get-title

[C function exported to Guile] c-lw6map-get-title

Wrapper on lw6map get title.

4.16.167 c-lw6map-param-get

[C function exported to Guile] c-lw6map-param-get

Wrapper on lw6map param get.

4.16.168 c-lw6map-rules-get-default

[C function exported to Guile] c-lw6map-rules-get-default

Wrapper on lw6map rules get default.

4.16.169 c-lw6map-rules-get-int

[C function exported to Guile] c-lw6map-rules-get-int

Wrapper on lw6map rules get int.

4.16.170 c-lw6map-rules-get-max

[C function exported to Guile] c-lw6map-rules-get-max

Wrapper on lw6map rules get max.

4.16.171 c-lw6map-rules-get-min

[C function exported to Guile] c-lw6map-rules-get-min

Wrapper on lw6map rules get min.

4.16.172 c-lw6map-style-get-default

[C function exported to Guile] c-lw6map-style-get-default

Wrapper on lw6map style get default.

4.16.173 c-lw6map-team-color-index-to-key

[C function exported to Guile] c-lw6map-team-color-index-to-key

Wrapper on lw6map team color index to key.

4.16.174 c-lw6map-team-color-index-to-label

[C function exported to Guile] c-lw6map-team-color-index-to-label

Wrapper on lw6map team color index to label.

4.16.175 c-lw6map-team-color-key-to-index

[C function exported to Guile] c-lw6map-team-color-key-to-index

Wrapper on lw6map team color key to index.

4.16.176 c-lw6map-team-color-list

[C function exported to Guile] c-lw6map-team-color-list

Wrapper on lw6map team color list.

4.16.177 c-lw6map-teams-get-default

[C function exported to Guile] c-lw6map-teams-get-default

Wrapper on lw6map teams get default.

4.16.178 c-lw6map-weapon-index-to-key

[C function exported to Guile] c-lw6map-weapon-index-to-key

Wrapper on lw6map weapon index to key.

4.16.179 c-lw6map-weapon-index-to-label

[C function exported to Guile] c-lw6map-weapon-index-to-label

Wrapper on lw6map weapon index to label.

4.16.180 c-lw6map-weapon-key-to-index

[C function exported to Guile] c-lw6map-weapon-key-to-index

Wrapper on lw6map weapon key to index.

4.16.181 c-lw6map-weapon-list

[C function exported to Guile] c-lw6map-weapon-list

Wrapper on lw6map weapon list.

4.16.182 c-lw6net-init

[C function exported to Guile] c-lw6net-init

Wrapper on lw6net init.

4.16.183 c-lw6net-quit

[C function exported to Guile] c-lw6net-quit

Wrapper on lw6net quit.

4.16.184 c-lw6p2p-db-default-name

[C function exported to Guile] c-lw6p2p-db-default-name

Wrapper on lw6p2p db default name.

4.16.185 c-lw6p2p-db-new

[C function exported to Guile] c-lw6p2p-db-new

Wrapper on lw6p2p db new.

4.16.186 c-lw6p2p-db-reset

[C function exported to Guile] c-lw6p2p-db-reset

Wrapper on lw6p2p db reset.

4.16.187 c-lw6p2p-node-calibrate

[C function exported to Guile] c-lw6p2p-node-calibrate

Wrapper on lw6p2p node calibrate.

4.16.188 c-lw6p2p-node-client-join

[C function exported to Guile] c-lw6p2p-node-client-join

Wrapper on lw6p2p node client join.

4.16.189 c-lw6p2p-node-close

[C function exported to Guile] c-lw6p2p-node-close

Wrapper on lw6p2p node close.

4.16.190 c-lw6p2p-node-disconnect

[C function exported to Guile] c-lw6p2p-node-disconnect

Wrapper on lw6p2p node disconnect.

4.16.191 c-lw6p2p-node-get-entries

[C function exported to Guile] c-lw6p2p-node-get-entries

Wrapper on lw6p2p node get entries.

4.16.192 c-lw6p2p-node-get-id

[C function exported to Guile] c-lw6p2p-node-get-id

Wrapper on lw6p2p node get id.

4.16.193 c-lw6p2p-node-get-local-seq-0

[C function exported to Guile] c-lw6p2p-node-get-local-seq-0

Wrapper on lw6p2p node get local seq 0.

4.16.194 c-lw6p2p-node-get-local-seq-last

[C function exported to Guile] c-lw6p2p-node-get-local-seq-last

Wrapper on lw6p2p node get local seq last.

4.16.195 c-lw6p2p-node-get-next-draft-msg

[C function exported to Guile] c-lw6p2p-node-get-next-draft-msg

Wrapper on lw6p2p node get next draft msg.

4.16.196 c-lw6p2p-node-get-next-reference-msg

[C function exported to Guile] c-lw6p2p-node-get-next-reference-msg

Wrapper on lw6p2p node get next reference msg.

4.16.197 c-lw6p2p-node-get-seq-draft

[C function exported to Guile] c-lw6p2p-node-get-seq-draft

Wrapper on lw6p2p node get seq draft.

4.16.198 c-lw6p2p-node-get-seq-max

[C function exported to Guile] c-lw6p2p-node-get-seq-max

Wrapper on lw6p2p node get seq max.

4.16.199 c-lw6p2p-node-get-seq-min

[C function exported to Guile] c-lw6p2p-node-get-seq-min

Wrapper on lw6p2p node get seq min.

4.16.200 c-lw6p2p-node-get-seq-reference

[C function exported to Guile] c-lw6p2p-node-get-seq-reference

Wrapper on lw6p2p node get seq reference.

4.16.201 c-lw6p2p-node-is-dump-needed

[C function exported to Guile] c-lw6p2p-node-is-dump-needed

Wrapper on lw6p2p node is dump needed.

4.16.202 c-lw6p2p-node-is-peer-connected

[C function exported to Guile] c-lw6p2p-node-is-peer-connected

Wrapper on lw6p2p node is peer connected.

4.16.203 c-lw6p2p-node-is-peer-registered

[C function exported to Guile] c-lw6p2p-node-is-peer-registered

Wrapper on lw6p2p node is peer registered.

4.16.204 c-lw6p2p-node-is-seed-needed

[C function exported to Guile] c-lw6p2p-node-is-seed-needed

Wrapper on lw6p2p node is seed needed.

4.16.205 c-lw6p2p-node-new

[C function exported to Guile] c-lw6p2p-node-new

Wrapper on lw6p2p node new.

4.16.206 c-lw6p2p-node-poll

[C function exported to Guile] c-lw6p2p-node-poll

Wrapper on lw6p2p node poll.

4.16.207 c-lw6p2p-node-put-local-msg

[C function exported to Guile] c-lw6p2p-node-put-local-msg

Wrapper on lw6p2p node put local msg.

4.16.208 c-lw6p2p-node-refresh-peer

[C function exported to Guile] c-lw6p2p-node-refresh-peer

Wrapper on lw6p2p node refresh peer.

4.16.209 c-lw6p2p-node-server-start

[C function exported to Guile] c-lw6p2p-node-server-start

Wrapper on lw6p2p node server start.

4.16.210 c-lw6p2p-node-update-info

[C function exported to Guile] c-lw6p2p-node-update-info

Wrapper on lw6p2p node update info.

4.16.211 c-lw6pil-bench

[C function exported to Guile] c-lw6pil-bench

Wrapper on lw6pil bench.

4.16.212 c-lw6pil-build-pilot

[C function exported to Guile] c-lw6pil-build-pilot

Wrapper on lw6pil build pilot.

4.16.213 c-lw6pil-calibrate

[C function exported to Guile] c-lw6pil-calibrate

Wrapper on lw6pil calibrate.

4.16.214 c-lw6pil-commit

[C function exported to Guile] c-lw6pil-commit

Wrapper on lw6pil commit.

4.16.215 c-lw6pil-did-cursor-win

[C function exported to Guile] c-lw6pil-did-cursor-win

Wrapper on lw6pil did cursor win.

4.16.216 c-lw6pil-dump-command-generate

[C function exported to Guile] c-lw6pil-dump-command-generate

Wrapper on lw6pil dump command generate.

4.16.217 c-lw6pil-execute-command

[C function exported to Guile] c-lw6pil-execute-command

Wrapper on lw6pil execute command.

4.16.218 c-lw6pil-fix-coords

[C function exported to Guile] c-lw6pil-fix-coords

Wrapper on lw6pil coords fix.

4.16.219 c-lw6pil-fix-coords-x10

[C function exported to Guile] c-lw6pil-fix-coords-x10

Wrapper on lw6pil coords fix x10.

4.16.220 c-lw6pil-get-last-commit-seq

[C function exported to Guile] c-lw6pil-get-last-commit-seq

Wrapper on lw6pil get last commit seq.

4.16.221 c-lw6pil-get-looser

[C function exported to Guile] c-lw6pil-get-looser

Wrapper on lw6pil get looser.

4.16.222 c-lw6pil-get-max-seq

[C function exported to Guile] c-lw6pil-get-max-seq

Wrapper on lw6pil get max seq.

4.16.223 c-lw6pil-get-next-seq

[C function exported to Guile] c-lw6pil-get-next-seq

Wrapper on lw6pil get next seq.

4.16.224 c-lw6pil-get-reference-current-seq

[C function exported to Guile] c-lw6pil-get-reference-current-seq

Wrapper on lw6pil get reference current seq.

4.16.225 c-lw6pil-get-reference-target-seq

[C function exported to Guile] c-lw6pil-get-reference-target-seq

Wrapper on lw6pil get reference target seq.

4.16.226 c-lw6pil-get-round-0

[C function exported to Guile] c-lw6pil-get-round-0

Wrapper on lw6pil get round 0.

4.16.227 c-lw6pil-get-seq-0

[C function exported to Guile] c-lw6pil-get-seq-0

Wrapper on lw6pil get seq 0.

4.16.228 c-lw6pil-get-winner

[C function exported to Guile] c-lw6pil-get-winner

Wrapper on lw6pil get winner.

4.16.229 c-lw6pil-is-over

[C function exported to Guile] c-lw6pil-is-over

Wrapper on lw6pil is over.

4.16.230 c-lw6pil-local-command

[C function exported to Guile] c-lw6pil-local-command

Wrapper on lw6pil local command.

4.16.231 c-lw6pil-local-cursors-set-main

[C function exported to Guile] c-lw6pil-local-cursors-set-main

Wrapper on lw6pil local cursors set main.

4.16.232 c-lw6pil-local-cursors-set-mouse-controlled

[C function exported to Guile] c-lw6pil-local-cursors-set-mouse-controlled

Wrapper on lw6pil local cursors set mouse controlled.

4.16.233 c-lw6pil-make-backup

[C function exported to Guile] c-lw6pil-make-backup

Wrapper on lw6pil make backup.

4.16.234 c-lw6pil-poll-dump

[C function exported to Guile] c-lw6pil-poll-dump

Wrapper on lw6pil poll dump.

4.16.235 c-lw6pil-round2seq

[C function exported to Guile] c-lw6pil-round2seq

Wrapper on lw6pil round2seq.

4.16.236 c-lw6pil-seed-command-generate

[C function exported to Guile] c-lw6pil-seed-command-generate

Wrapper on lw6pil seed command generate.

4.16.237 c-lw6pil-send-command

[C function exported to Guile] c-lw6pil-send-command

Wrapper on lw6pil send command.

4.16.238 c-lw6pil-seq-random-0

[C function exported to Guile] c-lw6pil-seq-random-0

Wrapper on lw6pil seq random 0.

4.16.239 c-lw6pil-seq2round

[C function exported to Guile] c-lw6pil-seq2round

Wrapper on lw6pil seq2round.

4.16.240 c-lw6pil-slow-down

[C function exported to Guile] c-lw6pil-slow-down

Wrapper on lw6pil slow down.

4.16.241 c-lw6pil-speed-up

[C function exported to Guile] c-lw6pil-speed-up

Wrapper on lw6pil speed up.

4.16.242 c-lw6pil-suite-get-checkpoint

[C function exported to Guile] c-lw6pil-suite-get-checkpoint

Wrapper on lw6pil suite get checkpoint.

4.16.243 c-lw6pil-suite-get-commands-by-node-index

[C function exported to Guile] c-lw6pil-suite-get-commands-by-node-index

Wrapper on lw6pil suite get command by node index, returns the list of all steps.

4.16.244 c-lw6pil-suite-get-commands-by-stage

[C function exported to Guile] c-lw6pil-suite-get-commands-by-stage

Wrapper on lw6pil suite get command by stage, returns the list of all steps.

4.16.245 c-lw6pil-suite-get-node-id

[C function exported to Guile] c-lw6pil-suite-get-node-id

Wrapper on lw6pil suite get node id.

4.16.246 c-lw6pil-suite-get-seq-0

[C function exported to Guile] c-lw6pil-suite-get-seq-0

Wrapper on lw6pil suite get seq 0.

4.16.247 c-lw6pil-suite-init

[C function exported to Guile] c-lw6pil-suite-init

Wrapper on lw6pil suite init.

4.16.248 c-lw6pil-sync-from-backup

[C function exported to Guile] c-lw6pil-sync-from-backup

Wrapper on lw6pil sync from backup.

4.16.249 c-lw6pil-sync-from-draft

[C function exported to Guile] c-lw6pil-sync-from-draft

Wrapper on lw6pil sync from draft.

4.16.250 c-lw6pil-sync-from-reference

[C function exported to Guile] c-lw6pil-sync-from-reference

Wrapper on lw6pil sync from reference.

4.16.251 c-lw6snd-get-backends

[C function exported to Guile] c-lw6snd-get-backends

Wrapper on lw6snd get backends.

4.16.252 c-lw6snd-is-music-file

[C function exported to Guile] c-lw6snd-is-music-file

Wrapper on lw6snd is music file.

4.16.253 c-lw6snd-new

[C function exported to Guile] c-lw6snd-new

Wrapper on lw6snd new.

4.16.254 c-lw6snd-play-fx

[C function exported to Guile] c-lw6snd-play-fx

Wrapper on lw6snd play fx.

4.16.255 c-lw6snd-play-music-file

[C function exported to Guile] c-lw6snd-play-music-file

Wrapper on lw6snd play music file.

4.16.256 c-lw6snd-play-music-random

[C function exported to Guile] c-lw6snd-play-music-random

Wrapper on lw6snd play music random.

4.16.257 c-lw6snd-poll

[C function exported to Guile] c-lw6snd-poll

Wrapper on lw6snd poll.

4.16.258 c-lw6snd-release

[C function exported to Guile] c-lw6snd-release

Wrapper on lw6snd release.

4.16.259 c-lw6snd-set-fx-volume

[C function exported to Guile] c-lw6snd-set-fx-volume

Wrapper on lw6snd set fx volume.

4.16.260 c-lw6snd-set-music-volume

[C function exported to Guile] c-lw6snd-set-music-volume

Wrapper on lw6snd set music volume.

4.16.261 c-lw6snd-set-water-volume

[C function exported to Guile] c-lw6snd-set-water-volume

Wrapper on lw6snd set water volume.

4.16.262 c-lw6snd-stop-music

[C function exported to Guile] c-lw6snd-stop-music

Wrapper on lw6snd stop music.

4.16.263 c-lw6srv-get-backends

[C function exported to Guile] c-lw6srv-get-backends

Wrapper on lw6srv get backends.

4.16.264 c-lw6sys-build-get-abs-srcdir

[C function exported to Guile] c-lw6sys-build-get-abs-srcdir

Wrapper on lw6sys build get abs srcdir.

4.16.265 c-lw6sys-build-get-bin-id

[C function exported to Guile] c-lw6sys-build-get-bin-id

Wrapper on lw6sys build get bin id.

4.16.266 c-lw6sys-build-get-bugs-url

[C function exported to Guile] c-lw6sys-build-get-bugs-url

Wrapper on lw6sys build get bugs url.

4.16.267 c-lw6sys-build-get-cflags

[C function exported to Guile] c-lw6sys-build-get-cflags

Wrapper on lw6sys build get cflags.

4.16.268 c-lw6sys-build-get-codename

[C function exported to Guile] c-lw6sys-build-get-codename

Wrapper on lw6sys build get codename.

4.16.269 c-lw6sys-build-get-configure-args

[C function exported to Guile] c-lw6sys-build-get-configure-args

Wrapper on lw6sys build get configure args.

4.16.270 c-lw6sys-build-get-copyright

[C function exported to Guile] c-lw6sys-build-get-copyright

Wrapper on lw6sys build get copyright.

4.16.271 c-lw6sys-build-get-datadir

[C function exported to Guile] c-lw6sys-build-get-datadir

Wrapper on lw6sys build get datadir.

4.16.272 c-lw6sys-build-get-date

[C function exported to Guile] c-lw6sys-build-get-date

Wrapper on lw6sys build get date.

4.16.273 c-lw6sys-build-get-docdir

[C function exported to Guile] c-lw6sys-build-get-docdir

Wrapper on lw6sys build get docdir.

4.16.274 c-lw6sys-build-get-enable-allinone

[C function exported to Guile] c-lw6sys-build-get-enable-allinone

Wrapper on lw6sys build get enable allinone.

4.16.275 c-lw6sys-build-get-enable-console

[C function exported to Guile] c-lw6sys-build-get-enable-console

Wrapper on lw6sys build get enable console.

4.16.276 c-lw6sys-build-get-enable-fullstatic

[C function exported to Guile] c-lw6sys-build-get-enable-fullstatic

Wrapper on lw6sys build get enable fullstatic.

4.16.277 c-lw6sys-build-get-enable-gcov

[C function exported to Guile] c-lw6sys-build-get-enable-gcov

Wrapper on lw6sys build get enable gcov.

4.16.278 c-lw6sys-build-get-enable-gprof

[C function exported to Guile] c-lw6sys-build-get-enable-gprof

Wrapper on lw6sys build get enable gprof.

4.16.279 c-lw6sys-build-get-enable-gtk

[C function exported to Guile] c-lw6sys-build-get-enable-gtk

Wrapper on lw6sys build get enable gtk.

4.16.280 c-lw6sys-build-get-enable-instrument

[C function exported to Guile] c-lw6sys-build-get-enable-instrument

Wrapper on lw6sys build get enable instrument.

4.16.281 c-lw6sys-build-get-enable-mod-caca

[C function exported to Guile] c-lw6sys-build-get-enable-mod-caca

Wrapper on lw6sys build get enable mod caca.

4.16.282 c-lw6sys-build-get-enable-mod-csound

[C function exported to Guile] c-lw6sys-build-get-enable-mod-csound

Wrapper on lw6sys build get enable mod csound.

4.16.283 c-lw6sys-build-get-enable-mod-gl1

[C function exported to Guile] c-lw6sys-build-get-enable-mod-gl1

Wrapper on lw6sys build get enable mod gl1.

4.16.284 c-lw6sys-build-get-enable-mod-gles2

[C function exported to Guile] c-lw6sys-build-get-enable-mod-gles2

Wrapper on lw6sys build get enable mod gles2.

4.16.285 c-lw6sys-build-get-enable-mod-http

[C function exported to Guile] c-lw6sys-build-get-enable-mod-http

Wrapper on lw6sys build get enable mod http.

4.16.286 c-lw6sys-build-get-enable-mod-ogg

[C function exported to Guile] c-lw6sys-build-get-enable-mod-ogg

Wrapper on lw6sys build get enable mod ogg.

4.16.287 c-lw6sys-build-get-enable-mod-soft

[C function exported to Guile] c-lw6sys-build-get-enable-mod-soft

Wrapper on lw6sys build get enable mod soft.

4.16.288 c-lw6sys-build-get-enable-openmp

[C function exported to Guile] c-lw6sys-build-get-enable-openmp

Wrapper on lw6sys build get enable openmp.

4.16.289 c-lw6sys-build-get-enable-optimize

[C function exported to Guile] c-lw6sys-build-get-enable-optimize

Wrapper on lw6sys build get enable optimize.

4.16.290 c-lw6sys-build-get-enable-paranoid

[C function exported to Guile] c-lw6sys-build-get-enable-paranoid

Wrapper on lw6sys build get enable paranoid.

4.16.291 c-lw6sys-build-get-enable-profiler

[C function exported to Guile] c-lw6sys-build-get-enable-profiler

Wrapper on lw6sys build get enable profiler.

4.16.292 c-lw6sys-build-get-enable-valgrind

[C function exported to Guile] c-lw6sys-build-get-enable-valgrind

Wrapper on lw6sys build get enable valgrind.

4.16.293 c-lw6sys-build-get-endianness

[C function exported to Guile] c-lw6sys-build-get-endianness

Wrapper on lw6sys build get endianness.

4.16.294 c-lw6sys-build-get-gcc-version

[C function exported to Guile] c-lw6sys-build-get-gcc-version

Wrapper on lw6sys build get gcc version.

4.16.295 c-lw6sys-build-get-home-url

[C function exported to Guile] c-lw6sys-build-get-home-url

Wrapper on lw6sys build get home url.

4.16.296 c-lw6sys-build-get-host-cpu

[C function exported to Guile] c-lw6sys-build-get-host-cpu

Wrapper on lw6sys build get host cpu.

4.16.297 c-lw6sys-build-get-host-os

[C function exported to Guile] c-lw6sys-build-get-host-os

Wrapper on lw6sys build get host os.

4.16.298 c-lw6sys-build-get-hostname

[C function exported to Guile] c-lw6sys-build-get-hostname

Wrapper on lw6sys build get hostname.

4.16.299 c-lw6sys-build-get-includedir

[C function exported to Guile] c-lw6sys-build-get-includedir

Wrapper on lw6sys build get includedir.

4.16.300 c-lw6sys-build-get-ldflags

[C function exported to Guile] c-lw6sys-build-get-ldflags

Wrapper on lw6sys build get ldflags.

4.16.301 c-lw6sys-build-get-libdir

[C function exported to Guile] c-lw6sys-build-get-libdir

Wrapper on lw6sys build get libdir.

4.16.302 c-lw6sys-build-get-license

[C function exported to Guile] c-lw6sys-build-get-license

Wrapper on lw6sys build get license.

4.16.303 c-lw6sys-build-get-localedir

[C function exported to Guile] c-lw6sys-build-get-localedir

Wrapper on lw6sys build get localedir.

4.16.304 c-lw6sys-build-get-md5sum

[C function exported to Guile] c-lw6sys-build-get-md5sum

Wrapper on lw6sys build get md5sum.

4.16.305 c-lw6sys-build-get-package-id

[C function exported to Guile] c-lw6sys-build-get-package-id

Wrapper on lw6sys build get package id.

4.16.306 c-lw6sys-build-get-package-name

[C function exported to Guile] c-lw6sys-build-get-package-name

Wrapper on lw6sys build get package name.

4.16.307 c-lw6sys-build-get-package-string

[C function exported to Guile] c-lw6sys-build-get-package-string

Wrapper on lw6sys build get package string.

4.16.308 c-lw6sys-build-get-package-tarname

[C function exported to Guile] c-lw6sys-build-get-package-tarname

Wrapper on lw6sys build get package tarname.

4.16.309 c-lw6sys-build-get-pointer-size

[C function exported to Guile] c-lw6sys-build-get-pointer-size

Wrapper on lw6sys build get pointer size.

4.16.310 c-lw6sys-build-get-prefix

[C function exported to Guile] c-lw6sys-build-get-prefix

Wrapper on lw6sys build get prefix.

4.16.311 c-lw6sys-build-get-stamp

[C function exported to Guile] c-lw6sys-build-get-stamp

Wrapper on lw6sys build get stamp.

4.16.312 c-lw6sys-build-get-time

[C function exported to Guile] c-lw6sys-build-get-time

Wrapper on lw6sys build get time.

4.16.313 c-lw6sys-build-get-top-srcdir

[C function exported to Guile] c-lw6sys-build-get-top-srcdir

Wrapper on lw6sys build get top srcdir.

4.16.314 c-lw6sys-build-get-version

[C function exported to Guile] c-lw6sys-build-get-version

Wrapper on lw6sys build get version.

4.16.315 c-lw6sys-build-get-version-base

[C function exported to Guile] c-lw6sys-build-get-version-base

Wrapper on lw6sys build get version base.

4.16.316 c-lw6sys-build-get-version-major

[C function exported to Guile] c-lw6sys-build-get-version-major

Wrapper on lw6sys build get version major.

4.16.317 c-lw6sys-build-get-version-minor

[C function exported to Guile] c-lw6sys-build-get-version-minor

Wrapper on lw6sys build get version minor.

4.16.318 c-lw6sys-build-is-gnu

[C function exported to Guile] c-lw6sys-build-is-gnu

Wrapper on lw6sys build is gnu.

4.16.319 c-lw6sys-build-is-gp2x

[C function exported to Guile] c-lw6sys-build-is-gp2x

Wrapper on lw6sys build is gp2x.

4.16.320 c-lw6sys-build-is-mac-os-x

[C function exported to Guile] c-lw6sys-build-is-mac-os-x

Wrapper on lw6sys build is mac os x.

4.16.321 c-lw6sys-build-is-ms-windows

[C function exported to Guile] c-lw6sys-build-is-ms-windows

Wrapper on lw6sys build is ms windows.

4.16.322 c-lw6sys-build-is-unix

[C function exported to Guile] c-lw6sys-build-is-unix

Wrapper on lw6sys build is unix.

4.16.323 c-lw6sys-build-is-x86

[C function exported to Guile] c-lw6sys-build-is-x86

Wrapper on lw6sys build is x86.

4.16.324 c-lw6sys-debug-get

[C function exported to Guile] c-lw6sys-debug-get

Wrapper on lw6sys debug get.

4.16.325 c-lw6sys-debug-set

[C function exported to Guile] c-lw6sys-debug-set

Wrapper on lw6sys debug set.

4.16.326 c-lw6sys-delay

[C function exported to Guile] c-lw6sys-delay

Wrapper on lw6sys delay.

4.16.327 c-lw6sys-dump

[C function exported to Guile] c-lw6sys-dump

Wrapper on lw6sys dump.

4.16.328 c-lw6sys-dump-clear

[C function exported to Guile] c-lw6sys-dump-clear

Wrapper on lw6sys dump clear.

4.16.329 c-lw6sys-generate-id-16

[C function exported to Guile] c-lw6sys-generate-id-16

Wrapper on lw6sys generate id 16.

4.16.330 c-lw6sys-generate-id-32

[C function exported to Guile] c-lw6sys-generate-id-32

Wrapper on lw6sys generate id 32.

4.16.331 c-lw6sys-generate-id-64

[C function exported to Guile] c-lw6sys-generate-id-64

Wrapper on lw6sys generate id 64.

4.16.332 c-lw6sys-get-config-file

[C function exported to Guile] c-lw6sys-get-config-file

Wrapper on lw6sys get config file.

4.16.333 c-lw6sys-get-cwd

[C function exported to Guile] c-lw6sys-get-cwd

Wrapper on lw6sys get cwd.

4.16.334 c-lw6sys-get-cycle

[C function exported to Guile] c-lw6sys-get-cycle

Wrapper on lw6sys get cycle.

4.16.335 c-lw6sys-get-data-dir

[C function exported to Guile] c-lw6sys-get-data-dir

Wrapper on lw6sys get data dir.

4.16.336 c-lw6sys-get-default-config-file

[C function exported to Guile] c-lw6sys-get-default-config-file

Wrapper on lw6sys get default config file.

4.16.337 c-lw6sys-get-default-data-dir

[C function exported to Guile] c-lw6sys-get-default-data-dir

Wrapper on lw6sys get default data dir.

4.16.338 c-lw6sys-get-default-log-file

[C function exported to Guile] c-lw6sys-get-default-log-file

Wrapper on lw6sys get default log file.

4.16.339 c-lw6sys-get-default-map-dir

[C function exported to Guile] c-lw6sys-get-default-map-dir

Wrapper on lw6sys get default map dir.

4.16.340 c-lw6sys-get-default-map-path

[C function exported to Guile] c-lw6sys-get-default-map-path

Wrapper on lw6sys get default map path.

4.16.341 c-lw6sys-get-default-mod-dir

[C function exported to Guile] c-lw6sys-get-default-mod-dir

Wrapper on lw6sys get default mod dir.

4.16.342 c-lw6sys-get-default-music-dir

[C function exported to Guile] c-lw6sys-get-default-music-dir

Wrapper on lw6sys get default music dir.

4.16.343 c-lw6sys-get-default-music-path

[C function exported to Guile] c-lw6sys-get-default-music-path

Wrapper on lw6sys get default music path.

4.16.344 c-lw6sys-get-default-prefix

[C function exported to Guile] c-lw6sys-get-default-prefix

Wrapper on lw6sys get default prefix.

4.16.345 c-lw6sys-get-default-script-file

[C function exported to Guile] c-lw6sys-get-default-script-file

Wrapper on lw6sys get default script file.

4.16.346 c-lw6sys-get-default-user-dir

[C function exported to Guile] c-lw6sys-get-default-user-dir

Wrapper on lw6sys get default user dir.

4.16.347 c-lw6sys-get-hostname

[C function exported to Guile] c-lw6sys-get-hostname

Wrapper on lw6sys get hostname.

4.16.348 c-lw6sys-get-log-file

[C function exported to Guile] c-lw6sys-get-log-file

Wrapper on lw6sys get log file.

4.16.349 c-lw6sys-get-map-dir

[C function exported to Guile] c-lw6sys-get-map-dir

Wrapper on lw6sys get map dir.

4.16.350 c-lw6sys-get-map-path

[C function exported to Guile] c-lw6sys-get-map-path

Wrapper on lw6sys get map path.

4.16.351 c-lw6sys-get-memory-bazooka-eraser

[C function exported to Guile] c-lw6sys-get-memory-bazooka-eraser

Wrapper on lw6sys get memory bazooka eraser.

4.16.352 c-lw6sys-get-memory-bazooka-size

[C function exported to Guile] c-lw6sys-get-memory-bazooka-size

Wrapper on lw6sys get memory bazooka size.

4.16.353 c-lw6sys-get-mod-dir

[C function exported to Guile] c-lw6sys-get-mod-dir

Wrapper on lw6sys get mod dir.

4.16.354 c-lw6sys-get-music-dir

[C function exported to Guile] c-lw6sys-get-music-dir

Wrapper on lw6sys get music dir.

4.16.355 c-lw6sys-get-music-path

[C function exported to Guile] c-lw6sys-get-music-path

Wrapper on lw6sys get music path.

4.16.356 c-lw6sys-get-prefix

[C function exported to Guile] c-lw6sys-get-prefix

Wrapper on lw6sys get prefix.

4.16.357 c-lw6sys-get-run-dir

[C function exported to Guile] c-lw6sys-get-run-dir

Wrapper on lw6sys get run dir.

4.16.358 c-lw6sys-get-script-file

[C function exported to Guile] c-lw6sys-get-script-file

Wrapper on lw6sys get script file.

4.16.359 c-lw6sys-get-timestamp

[C function exported to Guile] c-lw6sys-get-timestamp

Wrapper on lw6sys get timestamp.

4.16.360 c-lw6sys-get-uptime

[C function exported to Guile] c-lw6sys-get-uptime

Wrapper on lw6sys get uptime.

4.16.361 c-lw6sys-get-user-dir

[C function exported to Guile] c-lw6sys-get-user-dir

Wrapper on lw6sys get user dir.

4.16.362 c-lw6sys-get-username

[C function exported to Guile] c-lw6sys-get-username

Wrapper on lw6sys get username.

4.16.363 c-lw6sys-getenv

[C function exported to Guile] c-lw6sys-getenv

Wrapper on lw6sys getenv.

4.16.364 c-lw6sys-getenv-prefixed

[C function exported to Guile] c-lw6sys-getenv-prefixed

Wrapper on lw6sys getenv prefixed.

4.16.365 c-lw6sys-idle

[C function exported to Guile] c-lw6sys-idle

Wrapper on lw6sys idle.

4.16.366 c-lw6sys-log

[C function exported to Guile] c-lw6sys-log

Wrapper on lw6sys log.

4.16.367 c-lw6sys-log-get-backtrace-mode

[C function exported to Guile] c-lw6sys-log-get-backtrace-mode

Wrapper on lw6sys log get backtrace mode.

4.16.368 c-lw6sys-log-get-level

[C function exported to Guile] c-lw6sys-log-get-level

Wrapper on lw6sys log get level.

4.16.369 c-lw6sys-log-set-backtrace-mode

[C function exported to Guile] c-lw6sys-log-set-backtrace-mode

Wrapper on lw6sys log set backtrace mode.

4.16.370 c-lw6sys-log-set-dialog-timeout

[C function exported to Guile] c-lw6sys-log-set-dialog-timeout

Wrapper on lw6sys log set dialog timeout.

4.16.371 c-lw6sys-log-set-level

[C function exported to Guile] c-lw6sys-log-set-level

Wrapper on lw6sys log set level.

4.16.372 c-lw6sys-megabytes-available

[C function exported to Guile] c-lw6sys-megabytes-available

Wrapper on lw6sys megabytes available.

4.16.373 c-lw6sys-openmp-get-num-procs

[C function exported to Guile] c-lw6sys-openmp-get-num-procs

Wrapper on lw6sys openmp get num procs.

4.16.374 c-lw6sys-path-concat

[C function exported to Guile] c-lw6sys-path-concat

Wrapper on lw6sys path concat.

4.16.375 c-lw6sys-path-file-only

[C function exported to Guile] c-lw6sys-path-file-only

Wrapper on lw6sys path file only.

4.16.376 c-lw6sys-path-parent

[C function exported to Guile] c-lw6sys-path-parent

Wrapper on lw6sys path parent.

4.16.377 c-lw6sys-path-split

[C function exported to Guile] c-lw6sys-path-split

Wrapper on lw6sys path split.

4.16.378 c-lw6sys-set-memory-bazooka-eraser

[C function exported to Guile] c-lw6sys-set-memory-bazooka-eraser

Wrapper on lw6sys set memory bazooka eraser.

4.16.379 c-lw6sys-set-memory-bazooka-size

[C function exported to Guile] c-lw6sys-set-memory-bazooka-size

Wrapper on lw6sys set memory bazooka size.

4.16.380 c-lw6sys-signal-custom

[C function exported to Guile] c-lw6sys-signal-custom

Wrapper on lw6sys signal custom.

4.16.381 c-lw6sys-signal-default

[C function exported to Guile] c-lw6sys-signal-default

Wrapper on lw6sys signal default.

4.16.382 c-lw6sys-signal-poll-quit

[C function exported to Guile] c-lw6sys-signal-poll-quit

Wrapper on lw6sys signal poll quit.

4.16.383 c-lw6sys-signal-send-quit

[C function exported to Guile] c-lw6sys-signal-send-quit

Wrapper on lw6sys signal send quit.

4.16.384 c-lw6sys-sleep

[C function exported to Guile] c-lw6sys-sleep

Wrapper on lw6sys sleep.

4.16.385 c-lw6sys-snooze

[C function exported to Guile] c-lw6sys-snooze

Wrapper on lw6sys snooze.

4.16.386 c-lw6sys-url-canonize

[C function exported to Guile] c-lw6sys-url-canonize

Wrapper on lw6sys url canonize.

4.16.387 c-lw6tsk-loader-get-stage

[C function exported to Guile] c-lw6tsk-loader-get-stage

Wrapper on lw6tsk loader get stage.

4.16.388 c-lw6tsk-loader-new

[C function exported to Guile] c-lw6tsk-loader-new

Wrapper on lw6tsk loader new.

4.16.389 c-lw6tsk-loader-pop

[C function exported to Guile] c-lw6tsk-loader-pop

Wrapper on lw6tsk loader pop.

4.16.390 c-lw6tsk-loader-push-gen

[C function exported to Guile] c-lw6tsk-loader-push-gen

Wrapper on lw6tsk loader push gen.

4.16.391 c-lw6tsk-loader-push-ldr

[C function exported to Guile] c-lw6tsk-loader-push-ldr

Wrapper on lw6tsk loader push ldr.

4.17 Script hooks

5 C API

This chapter contains a description of all modules and a list of all documented C functions in the program. It contains many references and is self-generated from C comments using

gdoc (http://josefsson.org/gdoc/) by Simon Josefsson (http://josefsson.org/).

In order to reduce the number of pages of printed output, this complete reference is, by default, disabled in printable versions of the documentation (PostScript, PDF). This is both to make the manual more readable and to avoid wasting paper. Think about the environment.

It is however available in the HTML version of the documentation, which you can read

online on http://www.gnu.org/software/liquidwar6/manual/html_node/.

Additionnally, the following adresses contain various view on the source code, giving informations on all the internal and public C interfaces:

  • http://www.ufoot.org/liquidwar/v6/doc/coverage/: the lcov (http://ltp.
    sourceforge.net/coverage/lcov.php) output when running ./liquidwar6 --test.

It shows what functions are actually tested, and how many times they are called.

  • http://www.ufoot.org/liquidwar/v6/doc/global/: the GNU global (http://www.

gnu.org/software/global/) output gives complete cross-references, macros, headers, contants declaration. It’s a very good place to start browsing the code.

  • http://www.ufoot.org/liquidwar/v6/doc/cyclo/ : the pmccabe (http://

parisc-linux.org/~bame/pmccabe/) output shows the cyclomatic complexity of functions. It enables the programmer to spots the “ugly” and dangerous parts of the program.

  • http://www.ufoot.org/liquidwar/v6/doc/doxygen/: the Doxygen (http://www.

stack.nl/~dimitri/doxygen/) documentation gives an interactive access to the code, the structures and functions, and their dependencies.

5.1 libliquidwar6

5.1.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/index.html.

5.1.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libliquidwar6.html.

5.2 libbot

5.2.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/bot/index.html.

5.2.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libbot.html.

5.3 mod-brute

5.3.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/bot/mod-brute/index.
html.

5.3.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dbrute.html.

5.4 mod-follow

5.4.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results
on http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/bot/mod-follow/
index.html.

5.4.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dfollow.html.

5.5 mod-idiot

5.5.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/bot/mod-idiot/index.
html.

5.5.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002didiot.html.

5.6 mod-random

5.6.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results
on http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/bot/mod-random/
index.html.

5.6.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002drandom.html.

5.7 libcfg

5.7.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cfg/index.html.

5.7.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libcfg.html.

5.8 libcli

5.8.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cli/index.html.

5.8.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libcli.html.

5.9 mod-http

5.9.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cli/mod-http/index.
html.

5.9.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dhttp.html.

5.10 mod-tcp

5.10.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cli/mod-tcp/index.
html.

5.10.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dtcp.html.

5.11 mod-udp

5.11.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cli/mod-udp/index.
html.

5.11.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dudp.html.

5.12 libcns

5.12.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cns/index.html.

5.12.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libcns.html.

5.13 libcnx

5.13.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/cnx/index.html.

5.13.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libcnx.html.

5.14 libdat

5.14.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/dat/index.html.

5.14.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libdat.html.

5.15 libdef

5.15.1 Overview 5.15.2 API

There are no functions in libdef, only a header file with constants.

5.16 libdsp

5.16.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/dsp/index.html.

5.16.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libdsp.html.

5.17 libdyn

5.17.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/dyn/index.html.

5.17.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libdyn.html.

5.18 libgen

5.18.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gen/index.html.

5.18.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libgen.html.

5.19 libgfx

5.19.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/index.html.

5.19.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libgfx.html.

5.20 mod-gl1

5.20.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results
on http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/mod-gl1/

gl-utils/index.html (as there are many sub-directories in this module, please refer to

the test coverage directory index (http://www.ufoot.org/liquidwar/v6/doc/coverage/

index.html) for complete information).

5.20.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dgl1.html.

5.21 mod-gles2

5.21.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/mod-gles2/index.

html (as there are many sub-directories in this module, please refer to the test coverage

directory index (http://www.ufoot.org/liquidwar/v6/doc/coverage/index.html) for

complete information).

5.21.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dgles2.html.

5.22 mod-soft

5.22.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/mod-soft/index.

html (as there are many sub-directories in this module, please refer to the test coverage

directory index (http://www.ufoot.org/liquidwar/v6/doc/coverage/index.html) for

complete information).

5.22.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dsoft.html.

5.23 shared-sdl

5.23.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results
on http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/shared-sdl/

index.html (as there are many sub-directories in this module, please refer to the test

coverage directory index (http://www.ufoot.org/liquidwar/v6/doc/coverage/index.

html) for complete information).

5.23.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dsoft.html.

5.24 mod-caca

5.24.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gfx/mod-caca/index.

html (as there are many sub-directories in this module, please refer to the test coverage

directory index (http://www.ufoot.org/liquidwar/v6/doc/coverage/index.html) for

complete information).

5.24.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dcaca.html.

5.25 libglb

5.25.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/glb/index.html.

5.25.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libglb.html.

5.26 libgui

5.26.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/gui/index.html.

5.26.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libgui.html.

5.27 libhlp

5.27.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/hlp/index.html.

5.27.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libhlp.html.

5.28 libimg

5.28.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/img/index.html.

5.28.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libimg.html.

5.29 libker

5.29.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/ker/index.html.

5.29.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libker.html.

5.30 libldr

5.30.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/ldr/index.html.

5.30.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libldr.html.

5.31 libmap

5.31.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/map/index.html.

5.31.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libmap.html.

5.32 libmat

5.32.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/mat/index.html.

5.32.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libmat.html.

5.33 libmsg

5.33.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/msg/index.html.

5.33.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libmsg.html.

5.34 libnet

5.34.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/net/index.html.

5.34.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libnet.html.

5.35 libnod

5.35.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/nod/index.html.

5.35.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libnod.html.

5.36 libp2p

5.36.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/p2p/index.html.

5.36.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libp2p.html.

5.37 libpil

5.37.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/pil/index.html.

5.37.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libpil.html.

5.38 libscm

5.38.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/scm/index.html.

5.38.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libscm.html.

5.39 libsim

5.39.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/sim/index.html.

5.39.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libsim.html.

5.40 libsnd

5.40.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/snd/index.html.

5.40.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libsnd.html.

5.41 mod-csound

5.41.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results
on http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/snd/mod-csound/
index.html.

5.41.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dcsound.html.

5.42 mod-ogg

5.42.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/snd/mod-ogg/index.
html.

5.42.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dogg.html.

5.43 libsrv

5.43.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/srv/index.html.

5.43.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libsrv.html.

5.44 mod-httpd

5.44.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/srv/mod-httpd/index.
html.

5.44.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dhttpd.html.

5.45 mod-tcpd

5.45.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/srv/mod-tcpd/index.
html.

5.45.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dtcpd.html.

5.46 mod-udpd

5.46.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/srv/mod-udpd/index.
html.

5.46.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/mod_002dudpd.html.

5.47 libsys

5.47.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/sys/index.html.

5.47.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libsys.html.

5.48 libtsk

5.48.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/tsk/index.html.

5.48.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libtsk.html.

5.49 libvox

5.49.1 Overview

View lcov (http://ltp.sourceforge.net/coverage/lcov.php) test coverage results on
http://www.ufoot.org/liquidwar/v6/doc/coverage/src/lib/vox/index.html.

Chapter 5: C API 227

5.49.2 API

Online gdoc (http://josefsson.org/gdoc/) reference available on http://www.gnu.
org/software/liquidwar6/manual/html_node/libvox.html.

Appendix A Authors

Here’s a list of contributors :

Project maintainer, main developper :

• Christian Mauduit (mailto:ufoot@ufoot.org)

Original idea :

• Thomas Colcombet Artwork, level design :

• Kasper Hviid Musics :

• Tim Chadburn (menus)

• Robert Radamant (Free the sounds)

• LapSuS (Heav’hypnosis)

• Nighter313 (Oriental Travel) Libcaca backend :

• A. Frances

• R. Clavel

• K. Lemmonnier Translations :

• Karl Ove Hufthammer (NN, Norwegian)

• Yevgeny Lezhnin (RU, Russian) Many people contributed to Liquid War 5 (http://www.ufoot.org/liquidwar/v5),

their names are not listed here, but without them Liquid War 6 would obviously never have been started. Special thanks to all of them. However this is not a direct contribution to the project, in terms of code and other copyrightable materials.

Appendix B 2005 .plan

Here’s my .plan file, which describes what I (Christian Mauduit (mailto:ufoot@ufoot. org)) have planned for Liquid War 6. There’s no garantee that what’s written here is a precise description of the real future, however it should give a good idea of what I have in mind.

Note that the information here was written in summer 2005, it might or not be accurate now, as the main reason for plans to exist is that people never follow them. I’m no exception.

B.1 Complete rewrite

Liquid War 6 will be an almost complete rewrite. I mean that common code between branches 5 and 6 might end up in representing 0% of the total code. I think this is a wise decision, for the current code is really hard to maintain, and would not survive any serious cleanup. LW5 was first written in 1998, for DOS, when I had much less experience in programming. In 7 years I - and other people as well - hacked major enhancements in it such as cross-platform support, network games, and if you compare release 5.0 with the latest 5.x.x release, you’ll see that a bunch of things have changed. I had never expected I would patch and fix this game for so long, and it’s no surprise that it’s bloated today.

FYI, here’s a list of what makes LW5 unsuitable for major improvements without a complete rewrite:

• global variable hell. Lots of things are stored in globals.

• hard-coded C GUI. Read src/level.c to get an idea of how horrible it is.

• hard-coded 256 colors paletted mode. A clever bet in 1998 (performance...). Not anymore.

• generally bloated code. Makes bug-finding very tricky.

B.2 Technologies

Liquid War 6 will use a different technical framework than Liquid War 5 (http://www.

ufoot.org/liquidwar/v5).

B.2.1 Script + standard C + assembly

It happens that coding a large project in pure C is a waist of time, if possible at all.

If one applies the standard 80/20 rule to a computer game, one might state that 80% of the code eat up 20% of the CPU and the other 20% of the code eat up 80% of the CPU, the former being high-level glue code and the latter being low-level algorithmic code.

With Liquid War, one could speak of the 99/01 rule. I mean that 99% of the CPU time concerns only 1% of the code, and vice-versa. Basically, Liquid War has a very CPU-greedy core algorithm, still spends a fair amount of CPU displaying stuff (but this is delegated to the low-level game programming library) and the rest is totally unsignificant, in terms of CPU. Point is this "rest" represents the vast majority of the code, and also represents the very same buggy code I spend nights to patch on Liquid War 5 (http://www.ufoot. org/liquidwar/v5). I’m talking about network code, GUI, and other high-level glue-code which are currently being written in C.

This idea is to write all this in a convenient scripting language. There won’t be any impact on performances. I can’t garantee Liquid War 6 will be blazingly fast, but for sure it won’t be the scripting language fault. And of course if, as in Liquid War 3 and 5, I feel the need to implement some stuff in assembly for performances issues, I will do it.

We end up with a multi-language architecture: script + C + assembly.

   My guess is that I’ll use Scheme (http://www.gnu.org/software/guile/) as an

extension language. Python (http://www.python.org) would be a good choice too. Let’s say I’ll give Scheme a chance, and if it’s really not adapted, I’ll switch back to Python. The point is that today I know Python and don’t really know Scheme, but, well, it’s always a pleasure for me to learn new things. It’s fun.

So what is planned today is that Liquid War 6 will be a Scheme program, which will call callbacks functions written in C and/or assembly. These functions will do all the lowlevel time consuming algorithmic and graphical stuff. The rest of the code being entirely scripted.

B.2.2 OpenGL

Liquid War is not a 3D game, so why use OpenGL?

• it’s a very convenient way to access video hardware acceleration with XFree86.

• low-end computers and/or computers without 3D acceleration can still run Liquid War

    5 (http://www.ufoot.org/liquidwar/v5).

• I’m interested in learning/using this API 8-)

This choice implies that I won’t use Allegro (http://alleg.sourceforge.net/) anymore. Allegro stays a very convenient library and I would recommend it for it’s excellent, easy to learn, powerfull, and stable. But for the needs of Liquid War 6 I’ll use something else (because of OpenGL). I first thought of using GLUT (http://freeglut.sourceforge. net/) but I might end up simply using SDL (http://www.libsdl.org/). The idea is just fo have an OpenGL wrapper which sets up OpenGL in a similar manner on all platforms, and handles basic things such as mouse or keyboard.

B.2.3 CSound

I’ve got two excellent books on Csound (http://www.csounds.com/), and the will to learn how to use this tool.

I’ll probably use Csound for a number of things, ranging from "bubbling sounds" to full blown music. Stay tuned 8-)

B.3 Functionnalities

B.3.1 Visual enhancements

Of course Liquid War 6 will look nicer than Liquid War 5 (http://www.ufoot.org/ liquidwar/v5), blah blah blah. What do you think?

Maybe I’ll try to use some OpenGL features to make it possible to play on a ball, on a Moebius ring, or other fancy things. I have zillion of ideas, future will decide which ones will be implemented first.

To make it clear, visual enhancements aren’t my top-level priority. However I’ll try and make room for these enhancements, and prepare the terrain correctly. So it’s possible that the first releases of Liquid War 6 won’t be that much better than Liquid War 5 (http:// www.ufoot.org/liquidwar/v5), but at least Liquid War 6 will have the possibility to evolve. Something Liquid War 5 (http://www.ufoot.org/liquidwar/v5) doesn’t have.

B.3.2 Rules enhancements

There are many things that could be done easily:

• several cursors for one team

• alliances between teams

• deep places on a map, where more liquid can reside

• circular maps which "connect" the left border to the right one

• ...

As for graphical improvements, this is not my top-level priority. Simply, I’ll make the game ready-to-improve. Again, all these enhancements are very hard to code in Liquid War 5 (http://www.ufoot.org/liquidwar/v5), else I would already have coded them. Network enhancements

That’s my top-level prioriry. Why is that? Well, think of Liquid War in terms of "what makes it a good game?" and

"what makes it a poor game?".

It’s a good game because:

• the idea is original

• the gameplay is addictive

• you can play on a LAN

• all the family can play

• it’s cross-platform

• it’s Free Software

It’s a poor game because:

• it’s somewhat ugly and has a retro "back in the eighties" look

• network games are slow on Internet

• there are not enough active Internet servers

For the ugliness, well, OpenGL and some artwork should make it. But for the network, what’s the real problem?

The real problem is that in the current situation, the server needs to have all "keystrokes" before doing anything, and all players must be connected before a game starts. Here’s what I plan to do to fix this:

• players will be able to connect on a game "on the fly". This is done by most online games, and it’s IMHO a required features for a network mode to work on Internet (not speaking of local networks, but real wide online gaming). How this will fit with Liquid War’s rules is not totally decided, but I already know of several way to achieve this.

• I’ll implement an "anticipation" system "a la" U61 (http://www.ufoot.org). This means that no matter if a remote player has a poor network connection, things will behave as if everything was fine. Internally, the system keeps 2 images of the game. One which is "anticipated" and displayed to the player, and one which is validated but outdated, kept internally. It’s a little hard to explain, consumes twice as much CPU and memory, but it works. It happens that today the lacking ressource for playing Liquid War online is more on the network side than on the local CPU and memory aspects.

• I’ll take it to the next level and implement a "peer-to-peer-like" network model, in which any client can become a server. The idea behind is that if a server quits the game, then a client takes its role, letting the game continue for hours. This way one could virtually have a never ending Liquid War game which would last weeks. I believe this could be really cool. I also believe no proprietary game will ever implement that, for in this model there’s no way to force people to access a centralized server, this server usually being the major key in the business model of a company which sells proprietary software.

This third point will be the real enhancement of Liquid War with version 6. It’s one of the very points which drives me to rewrite it completely. First because it’s impossible to implement it without some heavy work. Then because I find it very motivating.

B.3.3 Hey, you forgot my idea!!!

Many gamers submitted suggestions, either by mail or by posting messages on the mailing list.

Don’t worry, I keep them. Not reading them here does not mean I won’t implement them. It simply means I won’t implement them first. I first need the game basically function before enhancing it with fancy stuff.

B.4 Road map

As I stated on the mailing list, when thinking about Liquid War 6, think of years rather than months (unless I get fired, jobless, or spend several months in a hospital with a laptop).

Note that this road map takes it for granted that I’ll be the lone coder on the project. It’s unlikely that someone is going to help me for the first stages, until there’s at least something real, something playable. Something that proves that the concept is valid. Besides, (real) team work implies a significant overhead, especially at project start. It’s hard to figure out how to distribute tasks when the tasks themselves are not clearly identified. But for the rest (starting in 2007 or 2008), it’s possible that external help might greatly... ...help!

• 2005 : Project framework should be done. This implies that the scripting engine is up and running, graphical mode works, config and data loading work, basic menus are available. Nothing playable.

• 2006 : Import the core algorithm from Liquid War 5 (http://www.ufoot.org/ liquidwar/v5), make the game playable in "demo mode" (" la" Liquid War 2), implement the network "peer-to-peer-like" mode. At this stage, it will be possible to know wether Liquid War 6 is true vaporware or not.

• 2007 : glue all this together to make something usable by anyone, heavy work on the GUI, on the options, on error checking, many bug fixes. The goal is to have a game which is equivalent to Liquid War 5 (http://www.ufoot.org/liquidwar/v5), with the network aspects pushed to the next level.

• 2008 : tadaaaaaaaaaaa! Release the game "publicly" - inform Freecode 8-) - and enhance it with all the feedback from gamers (bug reports and suggestions received since 1998). Work on artwork (both graphics and musics). Write documentation.

• 2009 : stabilize the game, patch it for all those things which had been forgotten back then in 2005, optimize for speed, bug-fix bug-fix bug-fix.

• 2010 : stop maintaining Liquid War 5 (http://www.ufoot.org/liquidwar/v5), invite Liquid War fans and coders to a hudge party in my garden, sing all night, drink beers and wine, teach Liquid War strategies to my 5 and 6 year old daughters, remember the old times when Liquid War wasn’t so cool 8-)

Appendix C Fanfic

Quoting Gavin: “I wrote a liquid war fanfic some time ago [...] I wrote it after a friend claimed that there wasn’t any liquid war fanfic because it wasn’t possible.”

So here it is, a Liquid War fanfic. It was initially written for Liquid War 5, but applies to Liquid War 6 as well. Enjoy!

C.1 The Battle of Emberlificoted

...

The General presided over his massing army in his seat, or rather hovering ring, of power. It dipped slightly as he flew low over his troops marching through the viscous marsh-like terrain. They were like children: obedient, loyal, and they ate a lot.

Glancing at the status panel mounted in front of him he grimaced; the other five armies: Yellow, Green, Orange, Turquoise, and, of course, Red, were also readying armies of a similar size to his own. His violet clones would have to fight hard and eat well to win this day.

Today would not be a battle of luck, the General mused, it would be a battle of tactics, of alliances, and of betrayal. Every clone was identical - that was the general idea behind clones - and the terrain seemed strangely symmetrical; it would not give advantage to any of the six armies amassed today. Glancing at the hologram of the battlefield projected in front of him the General noted that he would have to move quickly, Orange and Yellow were too close for comfort, though fortunately Baron Red’s army of eponymous coloured clones was the furthest.

General Violet’s fingertips were sweaty even before they touched the four main control keys in front of him. They were labeled ’W’, ’A’, ’D’, and, of course, the full retreat button - very useful for misleading foes and ambushing them as they pursued - ’S’. The keys were arrange in a roughly equilateral triangular pattern; with ’S’ forming the base and being adjacent to both ’A’ and ’D’, ’W’ formed the tip of the triangle.

A long breath left his parched lips as at last he made his move.

...

“Dammit!” he screamed moments later. He had misjudged Captain Yellow and Commander Orange; he had expected one at least to attack immediately, one he could have handled. They were working together - foiling his attempt to shoot between them to near the center of the battlefield to gain a better vantage point. Yellow had shot down towards him, cutting off his advance, and now Orange had sealed his escape route. “It’s not over yet” muttered the General. He opened a voice channel with Commander Orange:

“Very clever. Flawed, but still clever.”

“Flawed?” came the reply.

“Yes flawed, when the good Captain is finished devouring my army who do you think he will turn to next?”, bluffed the General - his hands worked quickly as he manoeuvred his hovering control ring, all that his troops ever saw of him, carefully towards the weakest section of his attackers. If he could just break out a few units he could soon turn the tide against both Yellow and Orange.

“We have an alliance...” Orange’s voice was unsure now.

Time for some sarcasm to through her even more off balance, thought the General, “I gathered”, he spoke softly, slowly, and with too much meaning. Then closing the

channel he turned his attention back to his escape.

...

“Yes!” wooped the ecstatic figure of the General. Fifty or so of his troops had broken free undetected and were even now working their way cautiously towards the camps of the Yellow army, only the front lines were still actively fighting; this opening gambit of Yellow and Orange had turned into a stale siege and Yellow’s army had pitched tent.

General Violet steered his hovering guidance ring to the center of the Yellow camp. His troops struck, both those who had got behind the lines and those who were still besieged. Yellow reacted too slowly and suddenly found that her army, was shrinking back from the onslaught. There was nowhere to run to, and bye now her only ally - Commander Orange - had abandoned her to her fate; he was too busy engaging Sir. Turquoise, who had managed to escape from the slaughter that the Baron had caused to the Turquoise ranks and was even now valiantly attacking the flanks of the Orange troops.

A glance at the status panel showed that Yellow’s life force was fading quickly: 8%, 3%, 1%, Gone.

The General smiled, he always enjoyed getting the first kill, and by now his armies life force had grown and his clones had replicated. With his, now, formidable fighting force it was no problem to engulf both Sir. Turquoise and Commander Orange’s brawling armies and annihilate them. Once again his army grew in size and power. Now if only the Baron didn’t notice that..., thought the General.

...

“Too late!” yelped the General, now thrown into panic, as he saw the approaching Baron. His army had also grown in size and power - having fatally injured the Turquoise army within the opening moments of the battle, and having finally managed to catch the elusive fleeing form of, or what remained of, Emperor Green.

Gripping the controls harder the General thought quickly, his army doesn’t so completely outnumber me that this is already over, however unless I can cause him to make a mistake that allows me to take the upper hand then I will inevitably lose. Maybe I can...

This thought was terminated and replaced by another as the Baron’s angry red troops broke through the undergrowth that had covered their movements and started to surround the General’s army. The thought that now throbbed through the panic-stricken mind of General Violet was simply ’Run!’.

Even as he signaled the retreat and made for what seemed to be the only possible means of escape the Baron’s blood red control ring appeared at the opening. The General knew it was over, even before the host of red beings appeared at the opening.

There was no escape. His life force was almost depleted and he was surrounded. Then it was that the Baron decided to communicate:

“Too bad. It was a good game” The General blinked, gaped, and was generally gobsmacked. Just before his life force

completely failed and his own weary eyes closed in defeat he snarled, “What!? This is not a game!” were the General’s dying words.

Appendix D Links

This section lists various Internet Liquid War related links.

D.1 Official links

These are the “official” links, hopefully you’ll find everything you need here:

• http://www.gnu.org/software/liquidwar6/: Liquid War 6 homepage
• http://www.gnu.org/software/liquidwar6/manual/: Online manual
• http://ftp.gnu.org/gnu/liquidwar6/: GNU downloads (source only)
• http://download.savannah.gnu.org/releases/liquidwar6/: Savannah downloads

(source and binaries)

• http://www.ufoot.org/download/liquidwar/v6/: ufoot.org downloads (mirror)
• http://www.ufoot.org/liquidwar/v6/snapshots/: Daily snapshots
• http://www.ufoot.org/liquidwar/v6/doc/: Automatically generated doc
• http://www.ufoot.org/jenkins/job/liquidwar6/: Jenkins continuous integration
• http://git.savannah.gnu.org/cgit/liquidwar6.git: GIT repository
• http://savannah.gnu.org/projects/liquidwar6/: Project on Savannah
• http://savannah.gnu.org/bugs/?func=additem&group=liquidwar6: Submit a bug

report

• http://lists.gnu.org/archive/html/help-liquidwar6/: Mailing-list archives

• http://ufoot.org:8056/: Permanent “seed” server running latest snapshot

• irc://irc.freenode.net/liquidwar: IRC channel #liquidwar on irc.freenode.net

D.2 Other sites

Note that some of these links might link to and/or promote proprietary software. It’s important to emphasize Liquid War 6 is free software, free as in speech, and you are encouraged to use software that protects your freedom. However, for your convenience, those links are provided, they might give you a hopefully neutral idea of what the game is all about.

This list is also by no way extensive, it’s provided “as is”.

• http://en.wikipedia.org/wiki/Liquid_War: Liquid War entry on Wikipedia
• http://fr.wikipedia.org/wiki/Liquid_War: Liquid War entry on Wikipedia

(French)

• http://freecode.com/projects/liquid-war-6: Liquid War 6 on Freecode.
• http://www.openhub.net/p/liquidwar6: Liquid War 6 on Open HUB.
• http://www.playdeb.net/software/Liquid%20War%206: Liquid War 6 on PlayDeb.

D.3 Old stuff

Various links that are deprecated, but still might contain interesting informations for those who enjoy digging into the past.

• http://arch.sv.gnu.org/archives/liquidwar6/: GNU Arch repository (replaced

by Git as for this project)

• http://www.ufoot.org/liquidwar/v5: Liquid War 5, the previous version of the game.

• http://git.savannah.gnu.org/gitweb/?p=liquidwar6.git: Gitweb interface, cgit

seems to be Savannah’s default now.

Appendix E GNU GENERAL PUBLIC LICENSE

Version 3, 29 June 2007

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 Free Software Foundation, Inc. http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The GNU General Public License is a free, copyleft license for software and other kinds of works.

The licenses for most software and other practical works are designed to take away your freedom to share and change the works. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change all versions of a program—to make sure it remains free software for all its users. We, the Free Software Foundation, use the GNU General Public License for most of our software; it applies also to any other work released this way by its authors. You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for them if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs, and that you know you can do these things.

To protect your rights, we need to prevent others from denying you these rights or asking you to surrender the rights. Therefore, you have certain responsibilities if you distribute copies of the software, or if you modify it: responsibilities to respect the freedom of others.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must pass on to the recipients the same freedoms that you received. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

Developers that use the GNU GPL protect your rights with two steps: (1) assert copyright on the software, and (2) offer you this License giving you legal permission to copy, distribute and/or modify it.

For the developers’ and authors’ protection, the GPL clearly explains that there is no warranty for this free software. For both users’ and authors’ sake, the GPL requires that modified versions be marked as changed, so that their problems will not be attributed erroneously to authors of previous versions.

Some devices are designed to deny users access to install or run modified versions of the software inside them, although the manufacturer can do so. This is fundamentally incompatible with the aim of protecting users’ freedom to change the software. The systematic pattern of such abuse occurs in the area of products for individuals to use, which is precisely where it is most unacceptable. Therefore, we have designed this version of the GPL to prohibit the practice for those products. If such problems arise substantially in other domains, we stand ready to extend this provision to those domains in future versions of the GPL, as needed to protect the freedom of users.

Finally, every program is threatened constantly by software patents. States should not allow patents to restrict development and use of software on general-purpose computers, but in those that do, we wish to avoid the special danger that patents applied to a free program could make it effectively proprietary. To prevent this, the GPL assures that patents cannot be used to render the program non-free.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS

0. Definitions. “This License” refers to version 3 of the GNU General Public License. “Copyright” also means copyright-like laws that apply to other kinds of works, such as semiconductor masks. “The Program” refers to any copyrightable work licensed under this License. Each licensee is addressed as “you”. “Licensees” and “recipients” may be individuals or organizations. To “modify” a work means to copy from or adapt all or part of the work in a fashion requiring copyright permission, other than the making of an exact copy. The resulting work is called a “modified version” of the earlier work or a work “based on” the earlier work. A “covered work” means either the unmodified Program or a work based on the Program. To “propagate” a work means to do anything with it that, without permission, would make you directly or secondarily liable for infringement under applicable copyright law, except executing it on a computer or modifying a private copy. Propagation includes copying, distribution (with or without modification), making available to the public, and in some countries other activities as well. To “convey” a work means any kind of propagation that enables other parties to make or receive copies. Mere interaction with a user through a computer network, with no transfer of a copy, is not conveying. An interactive user interface displays “Appropriate Legal Notices” to the extent that it includes a convenient and prominently visible feature that (1) displays an appropriate copyright notice, and (2) tells the user that there is no warranty for the work (except to the extent that warranties are provided), that licensees may convey the work under this License, and how to view a copy of this License. If the interface presents a list of user commands or options, such as a menu, a prominent item in the list meets this criterion.

1. Source Code. The “source code” for a work means the preferred form of the work for making modifications to it. “Object code” means any non-source form of a work. A “Standard Interface” means an interface that either is an official standard defined by a recognized standards body, or, in the case of interfaces specified for a particular programming language, one that is widely used among developers working in that language.

The “System Libraries” of an executable work include anything, other than the work as a whole, that (a) is included in the normal form of packaging a Major Component, but which is not part of that Major Component, and (b) serves only to enable use of the work with that Major Component, or to implement a Standard Interface for which an implementation is available to the public in source code form. A “Major Component”, in this context, means a major essential component (kernel, window system, and so on) of the specific operating system (if any) on which the executable work runs, or a compiler used to produce the work, or an object code interpreter used to run it.

The “Corresponding Source” for a work in object code form means all the source code needed to generate, install, and (for an executable work) run the object code and to modify the work, including scripts to control those activities. However, it does not include the work’s System Libraries, or general-purpose tools or generally available free programs which are used unmodified in performing those activities but which are not part of the work. For example, Corresponding Source includes interface definition files associated with source files for the work, and the source code for shared libraries and dynamically linked subprograms that the work is specifically designed to require, such as by intimate data communication or control flow between those subprograms and other parts of the work.

The Corresponding Source need not include anything that users can regenerate automatically from other parts of the Corresponding Source.

The Corresponding Source for a work in source code form is that same work.

2. Basic Permissions.

All rights granted under this License are granted for the term of copyright on the Program, and are irrevocable provided the stated conditions are met. This License explicitly affirms your unlimited permission to run the unmodified Program. The output from running a covered work is covered by this License only if the output, given its content, constitutes a covered work. This License acknowledges your rights of fair use or other equivalent, as provided by copyright law.

You may make, run and propagate covered works that you do not convey, without conditions so long as your license otherwise remains in force. You may convey covered works to others for the sole purpose of having them make modifications exclusively for you, or provide you with facilities for running those works, provided that you comply with the terms of this License in conveying all material for which you do not control copyright. Those thus making or running the covered works for you must do so exclusively on your behalf, under your direction and control, on terms that prohibit them from making any copies of your copyrighted material outside their relationship with you.

Conveying under any other circumstances is permitted solely under the conditions stated below. Sublicensing is not allowed; section 10 makes it unnecessary.

3. Protecting Users’ Legal Rights From Anti-Circumvention Law.

No covered work shall be deemed part of an effective technological measure under any applicable law fulfilling obligations under article 11 of the WIPO copyright treaty adopted on 20 December 1996, or similar laws prohibiting or restricting circumvention of such measures.

When you convey a covered work, you waive any legal power to forbid circumvention of technological measures to the extent such circumvention is effected by exercising rights under this License with respect to the covered work, and you disclaim any intention to limit operation or modification of the work as a means of enforcing, against the work’s users, your or third parties’ legal rights to forbid circumvention of technological measures.

4. Conveying Verbatim Copies.

You may convey verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice; keep intact all notices stating that this License and any non-permissive terms added in accord with section 7 apply to the code; keep intact all notices of the absence of any warranty; and give all recipients a copy of this License along with the Program. You may charge any price or no price for each copy that you convey, and you may offer support or warranty protection for a fee.

5. Conveying Modified Source Versions. You may convey a work based on the Program, or the modifications to produce it from the Program, in the form of source code under the terms of section 4, provided that you also meet all of these conditions: a. The work must carry prominent notices stating that you modified it, and giving a

relevant date. b. The work must carry prominent notices stating that it is released under this Li-

cense and any conditions added under section 7. This requirement modifies the requirement in section 4 to “keep intact all notices”.

c. You must license the entire work, as a whole, under this License to anyone who comes into possession of a copy. This License will therefore apply, along with any applicable section 7 additional terms, to the whole of the work, and all its parts, regardless of how they are packaged. This License gives no permission to license the work in any other way, but it does not invalidate such permission if you have separately received it.

d. If the work has interactive user interfaces, each must display Appropriate Legal Notices; however, if the Program has interactive interfaces that do not display Appropriate Legal Notices, your work need not make them do so.

A compilation of a covered work with other separate and independent works, which are not by their nature extensions of the covered work, and which are not combined with it such as to form a larger program, in or on a volume of a storage or distribution medium, is called an “aggregate” if the compilation and its resulting copyright are not used to limit the access or legal rights of the compilation’s users beyond what the individual works permit. Inclusion of a covered work in an aggregate does not cause this License to apply to the other parts of the aggregate.

6. Conveying Non-Source Forms.

You may convey a covered work in object code form under the terms of sections 4 and 5, provided that you also convey the machine-readable Corresponding Source under the terms of this License, in one of these ways: a. Convey the object code in, or embodied in, a physical product (including a phys-

ical distribution medium), accompanied by the Corresponding Source fixed on a durable physical medium customarily used for software interchange.

b. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by a written offer, valid for at least three years and valid for as long as you offer spare parts or customer support for that product model, to give anyone who possesses the object code either (1) a copy of the Corresponding Source for all the software in the product that is covered by this License, on a durable physical medium customarily used for software interchange, for a price no more than your reasonable cost of physically performing this conveying of source, or (2) access to copy the Corresponding Source from a network server at no charge.

c. Convey individual copies of the object code with a copy of the written offer to provide the Corresponding Source. This alternative is allowed only occasionally and noncommercially, and only if you received the object code with such an offer, in accord with subsection 6b.

d. Convey the object code by offering access from a designated place (gratis or for a charge), and offer equivalent access to the Corresponding Source in the same way through the same place at no further charge. You need not require recipients to copy the Corresponding Source along with the object code. If the place to copy the object code is a network server, the Corresponding Source may be on a different server (operated by you or a third party) that supports equivalent copying facilities, provided you maintain clear directions next to the object code saying where to find the Corresponding Source. Regardless of what server hosts the Corresponding Source, you remain obligated to ensure that it is available for as long as needed to satisfy these requirements.

e. Convey the object code using peer-to-peer transmission, provided you inform other peers where the object code and Corresponding Source of the work are being offered to the general public at no charge under subsection 6d.

A separable portion of the object code, whose source code is excluded from the Corresponding Source as a System Library, need not be included in conveying the object code work. A “User Product” is either (1) a “consumer product”, which means any tangible personal property which is normally used for personal, family, or household purposes, or (2) anything designed or sold for incorporation into a dwelling. In determining whether a product is a consumer product, doubtful cases shall be resolved in favor of coverage. For a particular product received by a particular user, “normally used” refers to a typical or common use of that class of product, regardless of the status of the particular user or of the way in which the particular user actually uses, or expects or is expected to use, the product. A product is a consumer product regardless of whether the product has substantial commercial, industrial or non-consumer uses, unless such uses represent the only significant mode of use of the product. “Installation Information” for a User Product means any methods, procedures, authorization keys, or other information required to install and execute modified versions of a covered work in that User Product from a modified version of its Corresponding Source.

The information must suffice to ensure that the continued functioning of the modified object code is in no case prevented or interfered with solely because modification has been made. If you convey an object code work under this section in, or with, or specifically for use in, a User Product, and the conveying occurs as part of a transaction in which the right of possession and use of the User Product is transferred to the recipient in perpetuity or for a fixed term (regardless of how the transaction is characterized), the Corresponding Source conveyed under this section must be accompanied by the Installation Information. But this requirement does not apply if neither you nor any third party retains the ability to install modified object code on the User Product (for example, the work has been installed in ROM). The requirement to provide Installation Information does not include a requirement to continue to provide support service, warranty, or updates for a work that has been modified or installed by the recipient, or for the User Product in which it has been modified or installed. Access to a network may be denied when the modification itself materially and adversely affects the operation of the network or violates the rules and protocols for communication across the network. Corresponding Source conveyed, and Installation Information provided, in accord with this section must be in a format that is publicly documented (and with an implementation available to the public in source code form), and must require no special password or key for unpacking, reading or copying.

7. Additional Terms. “Additional permissions” are terms that supplement the terms of this License by making exceptions from one or more of its conditions. Additional permissions that are applicable to the entire Program shall be treated as though they were included in this License, to the extent that they are valid under applicable law. If additional permissions apply only to part of the Program, that part may be used separately under those permissions, but the entire Program remains governed by this License without regard to the additional permissions. When you convey a copy of a covered work, you may at your option remove any additional permissions from that copy, or from any part of it. (Additional permissions may be written to require their own removal in certain cases when you modify the work.) You may place additional permissions on material, added by you to a covered work, for which you have or can give appropriate copyright permission. Notwithstanding any other provision of this License, for material you add to a covered work, you may (if authorized by the copyright holders of that material) supplement the terms of this License with terms: a. Disclaiming warranty or limiting liability differently from the terms of sections 15

and 16 of this License; or b. Requiring preservation of specified reasonable legal notices or author attributions

in that material or in the Appropriate Legal Notices displayed by works containing it; or

c. Prohibiting misrepresentation of the origin of that material, or requiring that modified versions of such material be marked in reasonable ways as different from the original version; or d. Limiting the use for publicity purposes of names of licensors or authors of the

material; or e. Declining to grant rights under trademark law for use of some trade names, trade-

marks, or service marks; or f. Requiring indemnification of licensors and authors of that material by anyone who

conveys the material (or modified versions of it) with contractual assumptions of liability to the recipient, for any liability that these contractual assumptions directly impose on those licensors and authors.

All other non-permissive additional terms are considered “further restrictions” within the meaning of section 10. If the Program as you received it, or any part of it, contains a notice stating that it is governed by this License along with a term that is a further restriction, you may remove that term. If a license document contains a further restriction but permits relicensing or conveying under this License, you may add to a covered work material governed by the terms of that license document, provided that the further restriction does not survive such relicensing or conveying. If you add terms to a covered work in accord with this section, you must place, in the relevant source files, a statement of the additional terms that apply to those files, or a notice indicating where to find the applicable terms. Additional terms, permissive or non-permissive, may be stated in the form of a separately written license, or stated as exceptions; the above requirements apply either way.

8. Termination.

You may not propagate or modify a covered work except as expressly provided under this License. Any attempt otherwise to propagate or modify it is void, and will automatically terminate your rights under this License (including any patent licenses granted under the third paragraph of section 11). However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, you do not qualify to receive new licenses for the same material under section 10.

9. Acceptance Not Required for Having Copies. You are not required to accept this License in order to receive or run a copy of the Program. Ancillary propagation of a covered work occurring solely as a consequence of using peer-to-peer transmission to receive a copy likewise does not require acceptance.

However, nothing other than this License grants you permission to propagate or modify any covered work. These actions infringe copyright if you do not accept this License. Therefore, by modifying or propagating a covered work, you indicate your acceptance of this License to do so.

10. Automatic Licensing of Downstream Recipients. Each time you convey a covered work, the recipient automatically receives a license from the original licensors, to run, modify and propagate that work, subject to this License. You are not responsible for enforcing compliance by third parties with this License. An “entity transaction” is a transaction transferring control of an organization, or substantially all assets of one, or subdividing an organization, or merging organizations. If propagation of a covered work results from an entity transaction, each party to that transaction who receives a copy of the work also receives whatever licenses to the work the party’s predecessor in interest had or could give under the previous paragraph, plus a right to possession of the Corresponding Source of the work from the predecessor in interest, if the predecessor has it or can get it with reasonable efforts. You may not impose any further restrictions on the exercise of the rights granted or affirmed under this License. For example, you may not impose a license fee, royalty, or other charge for exercise of rights granted under this License, and you may not initiate litigation (including a cross-claim or counterclaim in a lawsuit) alleging that any patent claim is infringed by making, using, selling, offering for sale, or importing the Program or any portion of it.

11. Patents.

A “contributor” is a copyright holder who authorizes use under this License of the Program or a work on which the Program is based. The work thus licensed is called the contributor’s “contributor version”. A contributor’s “essential patent claims” are all patent claims owned or controlled by the contributor, whether already acquired or hereafter acquired, that would be infringed by some manner, permitted by this License, of making, using, or selling its contributor version, but do not include claims that would be infringed only as a consequence of further modification of the contributor version. For purposes of this definition, “control” includes the right to grant patent sublicenses in a manner consistent with the requirements of this License. Each contributor grants you a non-exclusive, worldwide, royalty-free patent license under the contributor’s essential patent claims, to make, use, sell, offer for sale, import and otherwise run, modify and propagate the contents of its contributor version. In the following three paragraphs, a “patent license” is any express agreement or commitment, however denominated, not to enforce a patent (such as an express permission to practice a patent or covenant not to sue for patent infringement). To “grant” such a patent license to a party means to make such an agreement or commitment not to enforce a patent against the party. If you convey a covered work, knowingly relying on a patent license, and the Corresponding Source of the work is not available for anyone to copy, free of charge and under the terms of this License, through a publicly available network server or other readily accessible means, then you must either (1) cause the Corresponding Source to be so available, or (2) arrange to deprive yourself of the benefit of the patent license for this particular work, or (3) arrange, in a manner consistent with the requirements of this License, to extend the patent license to downstream recipients. “Knowingly relying” means you have actual knowledge that, but for the patent license, your conveying the covered work in a country, or your recipient’s use of the covered work in a country, would infringe one or more identifiable patents in that country that you have reason to believe are valid.

If, pursuant to or in connection with a single transaction or arrangement, you convey, or propagate by procuring conveyance of, a covered work, and grant a patent license to some of the parties receiving the covered work authorizing them to use, propagate, modify or convey a specific copy of the covered work, then the patent license you grant is automatically extended to all recipients of the covered work and works based on it.

A patent license is “discriminatory” if it does not include within the scope of its coverage, prohibits the exercise of, or is conditioned on the non-exercise of one or more of the rights that are specifically granted under this License. You may not convey a covered work if you are a party to an arrangement with a third party that is in the business of distributing software, under which you make payment to the third party based on the extent of your activity of conveying the work, and under which the third party grants, to any of the parties who would receive the covered work from you, a discriminatory patent license (a) in connection with copies of the covered work conveyed by you (or copies made from those copies), or (b) primarily for and in connection with specific products or compilations that contain the covered work, unless you entered into that arrangement, or that patent license was granted, prior to 28 March 2007.

Nothing in this License shall be construed as excluding or limiting any implied license or other defenses to infringement that may otherwise be available to you under applicable patent law.

12. No Surrender of Others’ Freedom.

If conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot convey a covered work so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not convey it at all. For example, if you agree to terms that obligate you to collect a royalty for further conveying from those to whom you convey the Program, the only way you could satisfy both those terms and this License would be to refrain entirely from conveying the Program.

13. Use with the GNU Affero General Public License.

Notwithstanding any other provision of this License, you have permission to link or combine any covered work with a work licensed under version 3 of the GNU Affero General Public License into a single combined work, and to convey the resulting work. The terms of this License will continue to apply to the part which is the covered work, but the special requirements of the GNU Affero General Public License, section 13, concerning interaction through a network will apply to the combination as such.

14. Revised Versions of this License.

The Free Software Foundation may publish revised and/or new versions of the GNU General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies that a certain numbered version of the GNU General Public License “or any later version” applies to it, you have the option of following the terms and conditions either of that numbered version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of the GNU General Public License, you may choose any version ever published by the Free Software Foundation.

If the Program specifies that a proxy can decide which future versions of the GNU General Public License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Program.

Later license versions may give you additional or different permissions. However, no additional obligations are imposed on any author or copyright holder as a result of your choosing to follow a later version.

15. Disclaimer of Warranty.

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PER-MITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EX-PRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. Limitation of Liability.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MODIFIES AND/OR CONVEYS THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, IN-CIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUS-TAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAM-AGES.

17. Interpretation of Sections 15 and 16.

If the disclaimer of warranty and limitation of liability provided above cannot be given local legal effect according to their terms, reviewing courts shall apply local law that most closely approximates an absolute waiver of all civil liability in connection with the Program, unless a warranty or assumption of liability accompanies a copy of the Program in return for a fee.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively state the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) year name of author
      This program is free software: you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation, either version 3 of the License, or (at
      your option) any later version.
      This program is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program.  If not, see http://www.gnu.org/licenses/.

Also add information on how to contact you by electronic and paper mail.

If the program does terminal interaction, make it output a short notice like this when it starts in an interactive mode:

      program Copyright (C) year name of author
      This program comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, your program’s commands might be different; for a GUI interface, you would use an “about box”.

You should also get your employer (if you work as a programmer) or school, if any, to sign a “copyright disclaimer” for the program, if necessary. For more information on this, and how to apply and follow the GNU GPL, see http://www.gnu.org/licenses/.

The GNU General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License. But first, please

read http://www.gnu.org/philosophy/why-not-lgpl.html.

Appendix F GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix G: Indexes 263

Appendix G Indexes

G.1 Concept index

A

Apple . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

B

Bug report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 22

C

Compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 42 CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

D

Download . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

G

GIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 GPL, GNU General Public License . . . . . . . . . . . 243

H

hints.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

L

Log file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

M

Mac OS X. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Microsoft Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

R

rules.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

S

source code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 style.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 subversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 SVN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

T

teams.xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

G.2 Function and keyword index

-

--about=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

--ambiance-exclude=<value> . . . . . . . . . . . . . . . . . 93
--ambiance-file=<value>. . . . . . . . . . . . . . . . . . . . . 94
--ambiance-filter=<value> . . . . . . . . . . . . . . . . . . 94
--animation-density=<value> . . . . . . . . . . . . . . . 141
--animation-speed=<value> . . . . . . . . . . . . . . . . . 142

--audit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

--auto-release-delay=<value> . . . . . . . . . . . . . . . 87
--background-color-auto=<value> . . . . . . . . . . 135
--background-color-root-bg=<value> . . . . . . . 142
--background-color-root-fg=<value> . . . . . . . 142
--background-color-stuff-bg=<value> . . . . . . 142
--background-color-stuff-fg=<value> . . . . . . 143
--background-style=<value> . . . . . . . . . . . . . . . . 143

--base64-decode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 --base64-encode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 --bench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 --bench-value=<value> . . . . . . . . . . . . . . . . . . . . . . 160 --bin-id=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 --bind-ip=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 --bind-port=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 95 --blink-cursor=<value>. . . . . . . . . . . . . . . . . . . . . 143 --boost-power=<value> . . . . . . . . . . . . . . . . . . . . . . 100 --bot-iq=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 --bot-speed=<value> . . . . . . . . . . . . . . . . . . . . . . . . 155 --bot1-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --bot1-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 155 --bot2-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 155 --bot2-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 155 --bot3-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 156 --bot3-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 156 --bot4-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 156 --bot4-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 156 --bot5-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 156 --bot5-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 157 --bot6-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 157 --bot6-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 157 --bot7-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 157 --bot7-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 157 --bot8-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 158 --bot8-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 158 --bot9-ai=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 158 --bot9-color=<value> . . . . . . . . . . . . . . . . . . . . . . . 158 --broadcast=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 95 --capture=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 --chosen-map=<value> . . . . . . . . . . . . . . . . . . . . . . . . 98 --cli-backends=<value> . . . . . . . . . . . . . . . . . . . . . . 96

--click-to-focus=<value> . . . . . . . . . . . . . . . . . . . 87
--color-alternate-bg=<value> . . . . . . . . . . . . . . 143
--color-alternate-fg=<value> . . . . . . . . . . . . . . 144
--color-base-bg=<value> . . . . . . . . . . . . . . . . . . . 144
--color-base-fg=<value> . . . . . . . . . . . . . . . . . . . 144
--color-conflict-mode=<value>. . . . . . . . . . . . . 100
--colorize-cursor=<value> . . . . . . . . . . . . . . . . . 145

--colorize=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 144

--commands-per-sec=<value> . . . . . . . . . . . . . . . . 161

--config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 --copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --cunit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

--cursor-pot-init=<value> . . . . . . . . . . . . . . . . . 100
--cursor-sensitivity=<value> . . . . . . . . . . . . . . . 88

--cursor-size=<value> . . . . . . . . . . . . . . . . . . . . . . 145 --custom-alt=<value> . . . . . . . . . . . . . . . . . . . . . . . . 88 --custom-ctrl=<value> . . . . . . . . . . . . . . . . . . . . . . . 88 --custom-down=<value> . . . . . . . . . . . . . . . . . . . . . . . 88 --custom-enter=<value> . . . . . . . . . . . . . . . . . . . . . . 88 --custom-esc=<value> . . . . . . . . . . . . . . . . . . . . . . . . 89 --custom-left=<value> . . . . . . . . . . . . . . . . . . . . . . . 89

--custom-pgdown=<value>. . . . . . . . . . . . . . . . . . . . . 89

--custom-pgup=<value> . . . . . . . . . . . . . . . . . . . . . . . 89 --custom-right=<value> . . . . . . . . . . . . . . . . . . . . . . 89 --custom-up=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 90 --daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 --danger-power=<value>. . . . . . . . . . . . . . . . . . . . . 101 --data-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 --debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

--debug-layer-id=<value> . . . . . . . . . . . . . . . . . . 161
--debug-team-id=<value> . . . . . . . . . . . . . . . . . . . 161

--defaults. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 --demo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

--dialog-timeout=<value> . . . . . . . . . . . . . . . . . . 162

--dirty-read=<value> . . . . . . . . . . . . . . . . . . . . . . . 162

--display-background=<value> . . . . . . . . . . . . . . 162
--display-console=<value> . . . . . . . . . . . . . . . . . 162
--display-cursors=<value> . . . . . . . . . . . . . . . . . 163
--display-debug-gradient=<value> . . . . . . . . . 163
--display-debug-zones=<value>. . . . . . . . . . . . . 163
--display-fighters=<value> . . . . . . . . . . . . . . . . 163

--display-fps=<value> . . . . . . . . . . . . . . . . . . . . . . 164 --display-hud=<value> . . . . . . . . . . . . . . . . . . . . . . 164 --display-log=<value> . . . . . . . . . . . . . . . . . . . . . . 164 --display-map=<value> . . . . . . . . . . . . . . . . . . . . . . 164 --display-menu=<value>. . . . . . . . . . . . . . . . . . . . . 164 --display-meta=<value>. . . . . . . . . . . . . . . . . . . . . 165

--display-mouse=<value> . . . . . . . . . . . . . . . . . . . 165

--display-mps=<value> . . . . . . . . . . . . . . . . . . . . . . 165

--display-preview=<value> . . . . . . . . . . . . . . . . . 165

                                                --display-progress=<value> . . . . . . . . . . . . . . . . 165
                                                --display-score=<value> . . . . . . . . . . . . . . . . . . . 166
                                                --display-splash=<value> . . . . . . . . . . . . . . . . . . 166

--display-url=<value> . . . . . . . . . . . . . . . . . . . . . . 166

                                                --double-click-delay=<value> . . . . . . . . . . . . . . . 90
                                                --downsize-using-bench-value=<value>. . . . . 136
                                                --downsize-using-fighter-scale=<value> . . 136

--example-hints-xml . . . . . . . . . . . . . . . . . . . . . . . . . 69 --example-rules-xml . . . . . . . . . . . . . . . . . . . . . . . . . 69 --example-style-xml . . . . . . . . . . . . . . . . . . . . . . . . . 69 --example-teams-xml . . . . . . . . . . . . . . . . . . . . . . . . . 69

                                                --executed-again=<value> . . . . . . . . . . . . . . . . . . 166

--exp=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

                                                --fighter-attack=<value> . . . . . . . . . . . . . . . . . . 101
                                                --fighter-defense=<value> . . . . . . . . . . . . . . . . . 101
                                                --fighter-new-health=<value> . . . . . . . . . . . . . . 102
                                                --fighter-regenerate=<value> . . . . . . . . . . . . . . 102
                                                --fighter-scale=<value> . . . . . . . . . . . . . . . . . . . 136

--force=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

                                                --frags-fade-out=<value> . . . . . . . . . . . . . . . . . . 102

--frags-mode=<value> . . . . . . . . . . . . . . . . . . . . . . . 102

                                                --frags-to-distribute=<value>. . . . . . . . . . . . . 103

--fullscreen=<value> . . . . . . . . . . . . . . . . . . . . . . . . 92 --fx-volume=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 94 --gfx-backend=<value> . . . . . . . . . . . . . . . . . . . . . . . 92

                                                --gfx-cpu-usage=<value> . . . . . . . . . . . . . . . . . . . 166

--gfx-debug=<value> . . . . . . . . . . . . . . . . . . . . . . . . 167 --gfx-quality=<value> . . . . . . . . . . . . . . . . . . . . . . . 92 --glue-power=<value> . . . . . . . . . . . . . . . . . . . . . . . 103 --guess-colors=<value>. . . . . . . . . . . . . . . . . . . . . 137

                                                --guess-moves-per-sec=<value>. . . . . . . . . . . . . 137

--height=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                                --hidden-layer-alpha=<value> . . . . . . . . . . . . . . 145
                                                --highest-team-color-allowed=<value>. . . . . 103
                                                --highest-weapon-allowed=<value> . . . . . . . . . 104

--host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                                --hud-color-auto=<value> . . . . . . . . . . . . . . . . . . 137
                                                --hud-color-frame-bg=<value> . . . . . . . . . . . . . . 145
                                                --hud-color-frame-fg=<value> . . . . . . . . . . . . . . 146
                                                --hud-color-text-bg=<value> . . . . . . . . . . . . . . . 146
                                                --hud-color-text-fg=<value> . . . . . . . . . . . . . . . 146

--hud-style=<value> . . . . . . . . . . . . . . . . . . . . . . . . 146 --io-per-sec=<value> . . . . . . . . . . . . . . . . . . . . . . . 167 --jpeg-quality=<value>. . . . . . . . . . . . . . . . . . . . . 167 --keep-ratio=<value> . . . . . . . . . . . . . . . . . . . . . . . 146 --known-nodes=<value> . . . . . . . . . . . . . . . . . . . . . . . 96 --list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 --list-advanced . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --list-aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --list-doc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --list-funcs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --list-graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --list-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-map-hints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-map-rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-map-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-map-teams. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 --list-players . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --list-quick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --list-show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --list-sound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --list-team-colors . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --list-weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --loader-sleep=<value>. . . . . . . . . . . . . . . . . . . . . 167

--local-bench-delta=<value> . . . . . . . . . . . . . . . 168

--log-file=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 82 --log-level=<value> . . . . . . . . . . . . . . . . . . . . . . . . 168 --log-timeout=<value> . . . . . . . . . . . . . . . . . . . . . . 168 --magic-number=<value>. . . . . . . . . . . . . . . . . . . . . 168 --map-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 --map-path=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 82

--max-cursor-pot-offset=<value> . . . . . . . . . . 104
--max-cursor-pot=<value> . . . . . . . . . . . . . . . . . . 104
--max-cursor-speed=<value> . . . . . . . . . . . . . . . . . 90
--max-local-bench-value=<value> . . . . . . . . . . 169
--max-map-height=<value> . . . . . . . . . . . . . . . . . . 137
--max-map-surface=<value> . . . . . . . . . . . . . . . . . 138
--max-map-width=<value> . . . . . . . . . . . . . . . . . . . 138
--max-nb-cursors=<value> . . . . . . . . . . . . . . . . . . 104

--max-nb-nodes=<value>. . . . . . . . . . . . . . . . . . . . . 105 --max-nb-teams=<value>. . . . . . . . . . . . . . . . . . . . . 105

--max-network-bench-value=<value> . . . . . . . . 169
--max-round-delta=<value> . . . . . . . . . . . . . . . . . 105
--max-zone-size=<value> . . . . . . . . . . . . . . . . . . . 105
--medicine-power=<value> . . . . . . . . . . . . . . . . . . 106
--memory-bazooka-eraser=<value> . . . . . . . . . . 169
--memory-bazooka-size=<value>. . . . . . . . . . . . . 170
--menu-color-auto=<value> . . . . . . . . . . . . . . . . . 138
--menu-color-default-bg=<value> . . . . . . . . . . 147
--menu-color-default-fg=<value> . . . . . . . . . . 147
--menu-color-disabled-bg=<value> . . . . . . . . . 147
--menu-color-disabled-fg=<value> . . . . . . . . . 147
--menu-color-selected-bg=<value> . . . . . . . . . 148
--menu-color-selected-fg=<value> . . . . . . . . . 148

--menu-style=<value> . . . . . . . . . . . . . . . . . . . . . . . 148

--min-map-height=<value> . . . . . . . . . . . . . . . . . . 138
--min-map-surface=<value> . . . . . . . . . . . . . . . . . 139
--min-map-width=<value> . . . . . . . . . . . . . . . . . . . 139

--mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 --modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

--mouse-sensitivity=<value> . . . . . . . . . . . . . . . . 90
--moves-per-round=<value> . . . . . . . . . . . . . . . . . 106

--music-dir=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 83

--music-exclude=<value> . . . . . . . . . . . . . . . . . . . 148

--music-file=<value> . . . . . . . . . . . . . . . . . . . . . . . 148 --music-filter=<value>. . . . . . . . . . . . . . . . . . . . . 149 --music-path=<value> . . . . . . . . . . . . . . . . . . . . . . . . 83 --music-volume=<value> . . . . . . . . . . . . . . . . . . . . . . 94

--nb-attack-tries=<value> . . . . . . . . . . . . . . . . . 106

--nb-bots=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 158

--nb-defense-tries=<value> . . . . . . . . . . . . . . . . 106

                                                --nb-move-tries=<value> . . . . . . . . . . . . . . . . . . . 107

--net-log=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 170 --net-per-sec=<value> . . . . . . . . . . . . . . . . . . . . . . 170

                                                --network-bench-delta=<value>. . . . . . . . . . . . . 170
                                                --network-reliability=<value>. . . . . . . . . . . . . 171
                                                --node-description=<value> . . . . . . . . . . . . . . . . . 96

--node-title=<value> . . . . . . . . . . . . . . . . . . . . . . . . 96 --open-relay=<value> . . . . . . . . . . . . . . . . . . . . . . . 171 --password=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 97 --pedigree. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 --pilot-lag=<value> . . . . . . . . . . . . . . . . . . . . . . . . 171 --pixelize=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 149

                                                --player1-color=<value> . . . . . . . . . . . . . . . . . . . 159
                                                --player1-control=<value> . . . . . . . . . . . . . . . . . . 84

--player1-name=<value> . . . . . . . . . . . . . . . . . . . . . . 85

                                                --player1-status=<value> . . . . . . . . . . . . . . . . . . . 85
                                                --player2-color=<value> . . . . . . . . . . . . . . . . . . . 159
                                                --player2-control=<value> . . . . . . . . . . . . . . . . . . 85

--player2-name=<value> . . . . . . . . . . . . . . . . . . . . . . 85

                                                --player2-status=<value> . . . . . . . . . . . . . . . . . . . 86
                                                --player3-color=<value> . . . . . . . . . . . . . . . . . . . 159
                                                --player3-control=<value> . . . . . . . . . . . . . . . . . . 86

--player3-name=<value> . . . . . . . . . . . . . . . . . . . . . . 86

                                                --player3-status=<value> . . . . . . . . . . . . . . . . . . . 86
                                                --player4-color=<value> . . . . . . . . . . . . . . . . . . . 159
                                                --player4-control=<value> . . . . . . . . . . . . . . . . . . 86

--player4-name=<value> . . . . . . . . . . . . . . . . . . . . . . 87

                                                --player4-status=<value> . . . . . . . . . . . . . . . . . . . 87

--prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 --public-url=<value> . . . . . . . . . . . . . . . . . . . . . . . . 97 --quick-start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 --repeat-delay=<value> . . . . . . . . . . . . . . . . . . . . . . 90

                                                --repeat-interval=<value> . . . . . . . . . . . . . . . . . . 91

--resample=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 139 --reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

                                                --reset-config-on-upgrade=<value> . . . . . . . . 172
                                                --respawn-delay=<value> . . . . . . . . . . . . . . . . . . . 107
                                                --respawn-position-mode=<value> . . . . . . . . . . 107

--respawn-team=<value>. . . . . . . . . . . . . . . . . . . . . 107 --round-delta=<value> . . . . . . . . . . . . . . . . . . . . . . 108

                                                --rounds-per-sec=<value> . . . . . . . . . . . . . . . . . . 108
                                                --screenshots-per-min=<value>. . . . . . . . . . . . . 172

--script-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 --server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

                                                --show-build-abs-srcdir. . . . . . . . . . . . . . . . . . . . . 71

--show-build-bin-id . . . . . . . . . . . . . . . . . . . . . . . . . 71 --show-build-bugs-url . . . . . . . . . . . . . . . . . . . . . . . 71 --show-build-cflags . . . . . . . . . . . . . . . . . . . . . . . . . 72 --show-build-codename . . . . . . . . . . . . . . . . . . . . . . . 72

                                                --show-build-configure-args . . . . . . . . . . . . . . . . 72

--show-build-copyright . . . . . . . . . . . . . . . . . . . . . . 72 --show-build-datadir . . . . . . . . . . . . . . . . . . . . . . . . 72 --show-build-date. . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 --show-build-docdir . . . . . . . . . . . . . . . . . . . . . . . . . 72

                                                --show-build-enable-allinone . . . . . . . . . . . . . . . 72
                                                --show-build-enable-console . . . . . . . . . . . . . . . . 73
                                                --show-build-enable-fullstatic. . . . . . . . . . . . . 73
                                                --show-build-enable-gcov . . . . . . . . . . . . . . . . . . . 73

--show-build-enable-gprof . . . . . . . . . . . . . . . . . . 73
--show-build-enable-gtk. . . . . . . . . . . . . . . . . . . . . 73
--show-build-enable-instrument. . . . . . . . . . . . . 73
--show-build-enable-mod-caca . . . . . . . . . . . . . . . 73
--show-build-enable-mod-csound. . . . . . . . . . . . . 73
--show-build-enable-mod-gl1 . . . . . . . . . . . . . . . . 74
--show-build-enable-mod-gles2 . . . . . . . . . . . . . . 74
--show-build-enable-mod-http . . . . . . . . . . . . . . . 74
--show-build-enable-mod-ogg . . . . . . . . . . . . . . . . 74
--show-build-enable-mod-soft . . . . . . . . . . . . . . . 74
--show-build-enable-openmp . . . . . . . . . . . . . . . . . 74
--show-build-enable-optimize . . . . . . . . . . . . . . . 74
--show-build-enable-paranoid . . . . . . . . . . . . . . . 75
--show-build-enable-profiler . . . . . . . . . . . . . . . 75
--show-build-enable-valgrind . . . . . . . . . . . . . . . 75
--show-build-endianness. . . . . . . . . . . . . . . . . . . . . 75
--show-build-gcc-version . . . . . . . . . . . . . . . . . . . 75

--show-build-gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 --show-build-gp2x. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 --show-build-home-url . . . . . . . . . . . . . . . . . . . . . . . 75 --show-build-host-cpu . . . . . . . . . . . . . . . . . . . . . . . 75 --show-build-host-os . . . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-hostname . . . . . . . . . . . . . . . . . . . . . . . 76

--show-build-includedir. . . . . . . . . . . . . . . . . . . . . 76

--show-build-ldflags . . . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-libdir . . . . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-license . . . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-localedir . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-mac-os-x . . . . . . . . . . . . . . . . . . . . . . . 76 --show-build-md5sum . . . . . . . . . . . . . . . . . . . . . . . . . 76

--show-build-ms-windows. . . . . . . . . . . . . . . . . . . . . 77
--show-build-package-id. . . . . . . . . . . . . . . . . . . . . 77
--show-build-package-name . . . . . . . . . . . . . . . . . . 77
--show-build-package-string . . . . . . . . . . . . . . . . 77
--show-build-package-tarname . . . . . . . . . . . . . . . 77
--show-build-pointer-size . . . . . . . . . . . . . . . . . . 77

--show-build-prefix . . . . . . . . . . . . . . . . . . . . . . . . . 77 --show-build-stamp . . . . . . . . . . . . . . . . . . . . . . . . . . 77 --show-build-time. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

--show-build-top-srcdir. . . . . . . . . . . . . . . . . . . . . 78

--show-build-unix. . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 --show-build-version . . . . . . . . . . . . . . . . . . . . . . . . 78

--show-build-version-base . . . . . . . . . . . . . . . . . . 78
--show-build-version-major . . . . . . . . . . . . . . . . . 78
--show-build-version-minor . . . . . . . . . . . . . . . . . 78

--show-build-x86. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 --show-config-file . . . . . . . . . . . . . . . . . . . . . . . . . . 78 --show-cwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 --show-data-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

--show-default-config-file . . . . . . . . . . . . . . . . . 79
--show-default-data-dir. . . . . . . . . . . . . . . . . . . . . 79
--show-default-log-file. . . . . . . . . . . . . . . . . . . . . 79

--show-default-map-dir . . . . . . . . . . . . . . . . . . . . . . 79

--show-default-map-path. . . . . . . . . . . . . . . . . . . . . 79

--show-default-mod-dir . . . . . . . . . . . . . . . . . . . . . . 79

--show-default-music-dir . . . . . . . . . . . . . . . . . . . 80
--show-default-music-path . . . . . . . . . . . . . . . . . . 80

--show-default-prefix . . . . . . . . . . . . . . . . . . . . . . . 80

                                                --show-default-script-file . . . . . . . . . . . . . . . . . 80
                                                --show-default-user-dir. . . . . . . . . . . . . . . . . . . . . 80

--show-log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 --show-map-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 --show-map-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 --show-mod-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-music-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-music-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-run-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-script-file . . . . . . . . . . . . . . . . . . . . . . . . . . 81 --show-user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

                                                --side-attack-factor=<value> . . . . . . . . . . . . . . 108
                                                --side-defense-factor=<value>. . . . . . . . . . . . . 109

--simulate-basic. . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 --simulate-full. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

                                                --single-army-size=<value> . . . . . . . . . . . . . . . . 109

--skip-network=<value> . . . . . . . . . . . . . . . . . . . . . . 97 --snd-backend=<value> . . . . . . . . . . . . . . . . . . . . . . . 94 --speed=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 --spread-mode=<value> . . . . . . . . . . . . . . . . . . . . . . 109

                                                --spread-thread=<value> . . . . . . . . . . . . . . . . . . . 109
                                                --spreads-per-round=<value> . . . . . . . . . . . . . . . 110

--srv-backends=<value> . . . . . . . . . . . . . . . . . . . . . . 97 --start-blue-x=<value>. . . . . . . . . . . . . . . . . . . . . 110 --start-blue-y=<value>. . . . . . . . . . . . . . . . . . . . . 110 --start-cyan-x=<value>. . . . . . . . . . . . . . . . . . . . . 110 --start-cyan-y=<value>. . . . . . . . . . . . . . . . . . . . . 111

                                                --start-green-x=<value> . . . . . . . . . . . . . . . . . . . 111
                                                --start-green-y=<value> . . . . . . . . . . . . . . . . . . . 111
                                                --start-lightblue-x=<value> . . . . . . . . . . . . . . . 111
                                                --start-lightblue-y=<value> . . . . . . . . . . . . . . . 111
                                                --start-magenta-x=<value> . . . . . . . . . . . . . . . . . 112
                                                --start-magenta-y=<value> . . . . . . . . . . . . . . . . . 112
                                                --start-orange-x=<value> . . . . . . . . . . . . . . . . . . 112
                                                --start-orange-y=<value> . . . . . . . . . . . . . . . . . . 112

--start-pink-x=<value>. . . . . . . . . . . . . . . . . . . . . 112 --start-pink-y=<value>. . . . . . . . . . . . . . . . . . . . . 113

                                                --start-position-mode=<value>. . . . . . . . . . . . . 113
                                                --start-purple-x=<value> . . . . . . . . . . . . . . . . . . 113
                                                --start-purple-y=<value> . . . . . . . . . . . . . . . . . . 113

--start-red-x=<value> . . . . . . . . . . . . . . . . . . . . . . 114 --start-red-y=<value> . . . . . . . . . . . . . . . . . . . . . . 114

                                                --start-yellow-x=<value> . . . . . . . . . . . . . . . . . . 114
                                                --start-yellow-y=<value> . . . . . . . . . . . . . . . . . . 114
                                                --system-color-auto=<value> . . . . . . . . . . . . . . . 140
                                                --system-color-bg=<value> . . . . . . . . . . . . . . . . . 149
                                                --system-color-fg=<value> . . . . . . . . . . . . . . . . . 149

--target-fps=<value> . . . . . . . . . . . . . . . . . . . . . . . 173

                                                --team-color-blue=<value> . . . . . . . . . . . . . . . . . 150
                                                --team-color-cyan=<value> . . . . . . . . . . . . . . . . . 150
                                                --team-color-dead=<value> . . . . . . . . . . . . . . . . . 150
                                                --team-color-green=<value> . . . . . . . . . . . . . . . . 150
                                                --team-color-lightblue=<value> . . . . . . . . . . . 150
                                                --team-color-magenta=<value> . . . . . . . . . . . . . . 151
                                                --team-color-orange=<value> . . . . . . . . . . . . . . . 151
                                                --team-color-pink=<value> . . . . . . . . . . . . . . . . . 151
                                                --team-color-purple=<value> . . . . . . . . . . . . . . . 151

--team-color-red=<value> . . . . . . . . . . . . . . . . . . 151
--team-color-yellow=<value> . . . . . . . . . . . . . . . 152
--team-profile-blue-aggressive=<value> . . 114
--team-profile-blue-fast=<value> . . . . . . . . . 115
--team-profile-blue-handicap=<value>. . . . . 115
--team-profile-blue-mobile=<value> . . . . . . . 115
--team-profile-blue-vulnerable=<value> . . 115
--team-profile-blue-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

--team-profile-blue-weapon-id=<value> . . . 116
--team-profile-blue-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

--team-profile-cyan-aggressive=<value> . . 116
--team-profile-cyan-fast=<value> . . . . . . . . . 116
--team-profile-cyan-handicap=<value>. . . . . 117
--team-profile-cyan-mobile=<value> . . . . . . . 117
--team-profile-cyan-vulnerable=<value> . . 117
--team-profile-cyan-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

--team-profile-cyan-weapon-id=<value> . . . 118
--team-profile-cyan-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

--team-profile-green-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

--team-profile-green-fast=<value> . . . . . . . . 118
--team-profile-green-handicap=<value> . . . 118
--team-profile-green-mobile=<value> . . . . . . 119
--team-profile-green-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

--team-profile-green-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

--team-profile-green-weapon-id=<value> . . 119
--team-profile-green-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

--team-profile-lightblue-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

--team-profile-lightblue-fast=<value> . . . 120
--team-profile-lightblue-handicap=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

--team-profile-lightblue-mobile=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

--team-profile-lightblue-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

--team-profile-lightblue-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

--team-profile-lightblue-weapon-id=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

--team-profile-lightblue-weapon-

mode=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

--team-profile-magenta-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

--team-profile-magenta-fast=<value> . . . . . . 122
--team-profile-magenta-handicap=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

--team-profile-magenta-mobile=<value> . . . 122
--team-profile-magenta-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

--team-profile-magenta-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

--team-profile-magenta-weapon-id=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

--team-profile-magenta-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

--team-profile-orange-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

--team-profile-orange-fast=<value> . . . . . . . 124
--team-profile-orange-handicap=<value> . . 124
--team-profile-orange-mobile=<value>. . . . . 124
--team-profile-orange-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

--team-profile-orange-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

--team-profile-orange-weapon-id=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

--team-profile-orange-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

--team-profile-pink-aggressive=<value> . . 125
--team-profile-pink-fast=<value> . . . . . . . . . 125
--team-profile-pink-handicap=<value>. . . . . 126
--team-profile-pink-mobile=<value> . . . . . . . 126
--team-profile-pink-vulnerable=<value> . . 126
--team-profile-pink-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

--team-profile-pink-weapon-id=<value> . . . 126
--team-profile-pink-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

--team-profile-purple-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

--team-profile-purple-fast=<value> . . . . . . . 127
--team-profile-purple-handicap=<value> . . 127
--team-profile-purple-mobile=<value>. . . . . 128
--team-profile-purple-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

--team-profile-purple-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

--team-profile-purple-weapon-id=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

--team-profile-purple-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

--team-profile-red-aggressive=<value> . . . 129
--team-profile-red-fast=<value> . . . . . . . . . . 129
--team-profile-red-handicap=<value> . . . . . . 129
--team-profile-red-mobile=<value> . . . . . . . . 129
--team-profile-red-vulnerable=<value> . . . 130
--team-profile-red-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

--team-profile-red-weapon-id=<value>. . . . . 130
--team-profile-red-weapon-mode=<value> . . 130
--team-profile-yellow-aggressive=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

--team-profile-yellow-fast=<value> . . . . . . . 131
--team-profile-yellow-handicap=<value> . . 131
--team-profile-yellow-mobile=<value>. . . . . 131
--team-profile-yellow-vulnerable=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

--team-profile-yellow-weapon-alternate-

id=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

--team-profile-yellow-weapon-id=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

--team-profile-yellow-weapon-mode=<value>

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 --test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

--total-armies-size=<value> . . . . . . . . . . . . . . . 132

--total-time=<value> . . . . . . . . . . . . . . . . . . . . . . . 133 --trap-errors=<value> . . . . . . . . . . . . . . . . . . . . . . 173 --trojan=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

--upsize-using-bench-value=<value> . . . . . . . 140
--upsize-using-fighter-scale=<value>. . . . . 140
--use-cursor-texture=<value> . . . . . . . . . . . . . . . 98
--use-double-click=<value> . . . . . . . . . . . . . . . . . 91
--use-esc-button=<value> . . . . . . . . . . . . . . . . . . . 91
--use-hints-xml=<value>. . . . . . . . . . . . . . . . . . . . . 98
--use-music-file=<value> . . . . . . . . . . . . . . . . . . . 99
--use-rules-xml=<value>. . . . . . . . . . . . . . . . . . . . . 99
--use-style-xml=<value>. . . . . . . . . . . . . . . . . . . . . 99
--use-team-profiles=<value> . . . . . . . . . . . . . . . 133
--use-teams-xml=<value>. . . . . . . . . . . . . . . . . . . . . 99

--use-texture=<value> . . . . . . . . . . . . . . . . . . . . . . 100 --user-dir=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . 84 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

--vertical-move=<value> . . . . . . . . . . . . . . . . . . . 133
--view-color-auto=<value> . . . . . . . . . . . . . . . . . 141
--view-color-cursor-bg=<value> . . . . . . . . . . . 152
--view-color-cursor-fg=<value> . . . . . . . . . . . 152
--view-color-map-bg=<value> . . . . . . . . . . . . . . . 152
--view-color-map-fg=<value> . . . . . . . . . . . . . . . 152

--view-style=<value> . . . . . . . . . . . . . . . . . . . . . . . 153 --wall-grease=<value> . . . . . . . . . . . . . . . . . . . . . . 141 --water-volume=<value> . . . . . . . . . . . . . . . . . . . . . . 95 --waves=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

--weapon-charge-delay=<value>. . . . . . . . . . . . . 133
--weapon-charge-max=<value> . . . . . . . . . . . . . . . 134
--weapon-duration=<value> . . . . . . . . . . . . . . . . . 134
--weapon-tune-berzerk-power=<value> . . . . . . 134
--weapon-tune-turbo-power=<value> . . . . . . . . 134

--width=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

--windowed-mode-limit=<value> . . . . . . . . . . . . . . 93

--x-polarity=<value> . . . . . . . . . . . . . . . . . . . . . . . 134 --x-wrap=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 --y-polarity=<value> . . . . . . . . . . . . . . . . . . . . . . . 135 --y-wrap=<value>. . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 --z-decode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 --z-encode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 --z-polarity=<value> . . . . . . . . . . . . . . . . . . . . . . . 135 --zoom-max=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 154 --zoom-min=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 154 --zoom-step=<value> . . . . . . . . . . . . . . . . . . . . . . . . . 91

--zoom-stick-delay=<value> . . . . . . . . . . . . . . . . . 91

--zoom=<value> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

A

ambiance-exclude. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 ambiance-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 ambiance-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 animation-density . . . . . . . . . . . . . . . . . . . . . . . . . . 141 animation-speed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 auto-release-delay . . . . . . . . . . . . . . . . . . . . . . . . . . 87

B

background-color-auto . . . . . . . . . . . . . . . . . . . . . . 135

                                                background-color-root-bg . . . . . . . . . . . . . . . . . . 142
                                                background-color-root-fg . . . . . . . . . . . . . . . . . . 142
                                                background-color-stuff-bg . . . . . . . . . . . . . . . . . 142
                                                background-color-stuff-fg . . . . . . . . . . . . . . . . . 143

background-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 bench-value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 bin-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 bind-ip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 bind-port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 blink-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 boost-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 bot-iq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 bot-speed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 bot1-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 bot1-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 bot2-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 bot2-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 bot3-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 bot3-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 bot4-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 bot4-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 bot5-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 bot5-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 bot6-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 bot6-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 bot7-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 bot7-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 bot8-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 bot8-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 bot9-ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 bot9-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 broadcast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

C

c-gettext. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6-exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6-get-ret . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6-release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6-set-ret . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6bot-get-backends . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6bot-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6bot-next-move . . . . . . . . . . . . . . . . . . . . . . . . . 174 c-lw6cfg-defaults . . . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-get-option . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-option-exists. . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 c-lw6cfg-set-option . . . . . . . . . . . . . . . . . . . . . . . . 175

c-lw6cfg-unified-get-log-file. . . . . . . . . . . . . 175
c-lw6cfg-unified-get-map-path. . . . . . . . . . . . . 175
c-lw6cfg-unified-get-music-path . . . . . . . . . . 176
c-lw6cfg-unified-get-user-dir. . . . . . . . . . . . . 176

c-lw6cli-get-backends . . . . . . . . . . . . . . . . . . . . . . 176

c-lw6cns-console-support . . . . . . . . . . . . . . . . . . 176

c-lw6cns-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 c-lw6cns-poll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 c-lw6cns-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 c-lw6cns-term-support . . . . . . . . . . . . . . . . . . . . . . 176

c-lw6dsp-get-average-fps . . . . . . . . . . . . . . . . . . 176
c-lw6dsp-get-fullscreen-modes. . . . . . . . . . . . . 176
c-lw6dsp-get-instant-fps . . . . . . . . . . . . . . . . . . 177
c-lw6dsp-get-last-frame-rendering-time . . 177

c-lw6dsp-get-nb-frames. . . . . . . . . . . . . . . . . . . . . 177

c-lw6dsp-get-video-mode . . . . . . . . . . . . . . . . . . . 177

c-lw6dsp-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 c-lw6dsp-release. . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 c-lw6dsp-update. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

c-lw6gen-create-from-seed . . . . . . . . . . . . . . . . . 177

c-lw6gen-seed-new . . . . . . . . . . . . . . . . . . . . . . . . . . 177

c-lw6gen-seed-normalize . . . . . . . . . . . . . . . . . . . 177

c-lw6gfx-get-backends . . . . . . . . . . . . . . . . . . . . . . 178 c-lw6gui-default-look . . . . . . . . . . . . . . . . . . . . . . 178 c-lw6gui-input-reset . . . . . . . . . . . . . . . . . . . . . . . 178

c-lw6gui-joystick1-get-move-pad . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-a . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-b . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-c . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-d . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-e . . . . . . . . . . 178
c-lw6gui-joystick1-pop-button-f . . . . . . . . . . 178
c-lw6gui-joystick1-pop-pad-down . . . . . . . . . . 179
c-lw6gui-joystick1-pop-pad-left . . . . . . . . . . 179
c-lw6gui-joystick1-pop-pad-right . . . . . . . . . 179
c-lw6gui-joystick1-pop-pad-up. . . . . . . . . . . . . 179
c-lw6gui-joystick2-get-move-pad . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-a . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-b . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-c . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-d . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-e . . . . . . . . . . 179
c-lw6gui-joystick2-pop-button-f . . . . . . . . . . 180
c-lw6gui-joystick2-pop-pad-down . . . . . . . . . . 180
c-lw6gui-joystick2-pop-pad-left . . . . . . . . . . 180
c-lw6gui-joystick2-pop-pad-right . . . . . . . . . 180
c-lw6gui-joystick2-pop-pad-up. . . . . . . . . . . . . 180
c-lw6gui-keyboard-get-move-pad . . . . . . . . . . . 180
c-lw6gui-keyboard-is-pressed . . . . . . . . . . . . . . 180
c-lw6gui-keyboard-pop-arrow-down . . . . . . . . . 180
c-lw6gui-keyboard-pop-arrow-left . . . . . . . . . 180
c-lw6gui-keyboard-pop-arrow-right . . . . . . . . 180
c-lw6gui-keyboard-pop-arrow-up . . . . . . . . . . . 181
c-lw6gui-keyboard-pop-key-alt. . . . . . . . . . . . . 181
c-lw6gui-keyboard-pop-key-ctrl . . . . . . . . . . . 181
c-lw6gui-keyboard-pop-key-enter . . . . . . . . . . 181

                                                c-lw6gui-keyboard-pop-key-esc. . . . . . . . . . . . . 181
                                                c-lw6gui-keyboard-pop-key-pgdown . . . . . . . . . 181
                                                c-lw6gui-keyboard-pop-key-pgup . . . . . . . . . . . 181

c-lw6gui-look-get . . . . . . . . . . . . . . . . . . . . . . . . . . 181 c-lw6gui-look-set . . . . . . . . . . . . . . . . . . . . . . . . . . 181 c-lw6gui-look-zoom-in . . . . . . . . . . . . . . . . . . . . . . 181 c-lw6gui-look-zoom-out. . . . . . . . . . . . . . . . . . . . . 182 c-lw6gui-menu-append . . . . . . . . . . . . . . . . . . . . . . . 182

                                                c-lw6gui-menu-close-popup . . . . . . . . . . . . . . . . . 182
                                                c-lw6gui-menu-enable-esc . . . . . . . . . . . . . . . . . . 182
                                                c-lw6gui-menu-has-popup . . . . . . . . . . . . . . . . . . . 182

c-lw6gui-menu-new . . . . . . . . . . . . . . . . . . . . . . . . . . 182 c-lw6gui-menu-remove . . . . . . . . . . . . . . . . . . . . . . . 182

                                                c-lw6gui-menu-remove-all . . . . . . . . . . . . . . . . . . 182
                                                c-lw6gui-menu-scroll-down . . . . . . . . . . . . . . . . . 182
                                                c-lw6gui-menu-scroll-up . . . . . . . . . . . . . . . . . . . 182

c-lw6gui-menu-select . . . . . . . . . . . . . . . . . . . . . . . 183

                                                c-lw6gui-menu-select-esc . . . . . . . . . . . . . . . . . . 183
                                                c-lw6gui-menu-set-breadcrumbs. . . . . . . . . . . . . 183

c-lw6gui-menu-sync . . . . . . . . . . . . . . . . . . . . . . . . . 183

                                                c-lw6gui-mouse-get-state . . . . . . . . . . . . . . . . . . 183
                                                c-lw6gui-mouse-poll-move . . . . . . . . . . . . . . . . . . 183
                                                c-lw6gui-mouse-pop-button-left . . . . . . . . . . . 183
                                                c-lw6gui-mouse-pop-button-middle . . . . . . . . . 183
                                                c-lw6gui-mouse-pop-button-right . . . . . . . . . . 183
                                                c-lw6gui-mouse-pop-double-click . . . . . . . . . . 183
                                                c-lw6gui-mouse-pop-simple-click . . . . . . . . . . 184
                                                c-lw6gui-mouse-pop-triple-click . . . . . . . . . . 184
                                                c-lw6gui-mouse-pop-wheel-down. . . . . . . . . . . . . 184
                                                c-lw6gui-mouse-pop-wheel-up . . . . . . . . . . . . . . . 184

c-lw6hlp-about . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

                                                c-lw6hlp-get-default-value . . . . . . . . . . . . . . . . 184

c-lw6hlp-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 c-lw6hlp-list-advanced. . . . . . . . . . . . . . . . . . . . . 184 c-lw6hlp-list-aliases . . . . . . . . . . . . . . . . . . . . . . 184 c-lw6hlp-list-doc . . . . . . . . . . . . . . . . . . . . . . . . . . 184 c-lw6hlp-list-funcs . . . . . . . . . . . . . . . . . . . . . . . . 185 c-lw6hlp-list-graphics. . . . . . . . . . . . . . . . . . . . . 185 c-lw6hlp-list-hooks . . . . . . . . . . . . . . . . . . . . . . . . 185 c-lw6hlp-list-input . . . . . . . . . . . . . . . . . . . . . . . . 185 c-lw6hlp-list-map . . . . . . . . . . . . . . . . . . . . . . . . . . 185

                                                c-lw6hlp-list-map-hints . . . . . . . . . . . . . . . . . . . 185
                                                c-lw6hlp-list-map-rules . . . . . . . . . . . . . . . . . . . 185
                                                c-lw6hlp-list-map-style . . . . . . . . . . . . . . . . . . . 185
                                                c-lw6hlp-list-map-teams . . . . . . . . . . . . . . . . . . . 185

c-lw6hlp-list-network . . . . . . . . . . . . . . . . . . . . . . 185 c-lw6hlp-list-path . . . . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6hlp-list-players . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6hlp-list-quick . . . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6hlp-list-show . . . . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6hlp-list-sound . . . . . . . . . . . . . . . . . . . . . . . . 186

                                                c-lw6hlp-list-team-colors . . . . . . . . . . . . . . . . . 186

c-lw6hlp-list-weapons . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6img-screenshot . . . . . . . . . . . . . . . . . . . . . . . . 186 c-lw6ker-add-cursor . . . . . . . . . . . . . . . . . . . . . . . . 186

                                                c-lw6ker-build-game-state . . . . . . . . . . . . . . . . . 186
                                                c-lw6ker-build-game-struct . . . . . . . . . . . . . . . . 187

c-lw6ker-cursor-exists. . . . . . . . . . . . . . . . . . . . . 187

c-lw6ker-did-cursor-win . . . . . . . . . . . . . . . . . . . 187

c-lw6ker-do-round . . . . . . . . . . . . . . . . . . . . . . . . . . 187

c-lw6ker-dup-game-state . . . . . . . . . . . . . . . . . . . 187
c-lw6ker-game-state-checksum . . . . . . . . . . . . . . 187
c-lw6ker-game-struct-checksum. . . . . . . . . . . . . 187

c-lw6ker-get-cursor . . . . . . . . . . . . . . . . . . . . . . . . 187 c-lw6ker-get-moves . . . . . . . . . . . . . . . . . . . . . . . . . 187 c-lw6ker-get-nb-colors. . . . . . . . . . . . . . . . . . . . . 187

c-lw6ker-get-nb-cursors . . . . . . . . . . . . . . . . . . . 188

c-lw6ker-get-nb-nodes . . . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-get-rounds . . . . . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-get-spreads . . . . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-is-over. . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-node-exists . . . . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-register-node. . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-remove-cursor. . . . . . . . . . . . . . . . . . . . . 188 c-lw6ker-set-cursor . . . . . . . . . . . . . . . . . . . . . . . . 188

c-lw6ker-sync-game-state . . . . . . . . . . . . . . . . . . 188
c-lw6ker-unregister-node . . . . . . . . . . . . . . . . . . 189

c-lw6ldr-chain-entry . . . . . . . . . . . . . . . . . . . . . . . 189 c-lw6ldr-exp-validate . . . . . . . . . . . . . . . . . . . . . . 189 c-lw6ldr-get-entries . . . . . . . . . . . . . . . . . . . . . . . 189

c-lw6ldr-hints-get-default . . . . . . . . . . . . . . . . 189
c-lw6ldr-print-examples . . . . . . . . . . . . . . . . . . . 189

c-lw6ldr-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 c-lw6ldr-read-relative. . . . . . . . . . . . . . . . . . . . . 189

c-lw6map-exp-get-unlocked-team-color. . . . . 189
c-lw6map-exp-get-unlocked-weapon . . . . . . . . . 189
c-lw6map-exp-is-team-color-allowed . . . . . . . 190
c-lw6map-exp-is-weapon-allowed . . . . . . . . . . . 190

c-lw6map-get-look . . . . . . . . . . . . . . . . . . . . . . . . . . 190

c-lw6map-get-max-nb-colors . . . . . . . . . . . . . . . . 190
c-lw6map-get-max-nb-cursors . . . . . . . . . . . . . . . 190
c-lw6map-get-max-nb-nodes . . . . . . . . . . . . . . . . . 190

c-lw6map-get-music-dir. . . . . . . . . . . . . . . . . . . . . 190 c-lw6map-get-title . . . . . . . . . . . . . . . . . . . . . . . . . 190 c-lw6map-param-get . . . . . . . . . . . . . . . . . . . . . . . . . 190

c-lw6map-rules-get-default . . . . . . . . . . . . . . . . 190

c-lw6map-rules-get-int. . . . . . . . . . . . . . . . . . . . . 191 c-lw6map-rules-get-max. . . . . . . . . . . . . . . . . . . . . 191 c-lw6map-rules-get-min. . . . . . . . . . . . . . . . . . . . . 191

c-lw6map-style-get-default . . . . . . . . . . . . . . . . 191
c-lw6map-team-color-index-to-key . . . . . . . . . 191
c-lw6map-team-color-index-to-label . . . . . . . 191
c-lw6map-team-color-key-to-index . . . . . . . . . 191
c-lw6map-team-color-list . . . . . . . . . . . . . . . . . . 191
c-lw6map-teams-get-default . . . . . . . . . . . . . . . . 191
c-lw6map-weapon-index-to-key . . . . . . . . . . . . . . 191
c-lw6map-weapon-index-to-label . . . . . . . . . . . 192
c-lw6map-weapon-key-to-index . . . . . . . . . . . . . . 192

c-lw6map-weapon-list . . . . . . . . . . . . . . . . . . . . . . . 192 c-lw6net-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 c-lw6net-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192

c-lw6p2p-db-default-name . . . . . . . . . . . . . . . . . . 192

c-lw6p2p-db-new. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 c-lw6p2p-db-reset . . . . . . . . . . . . . . . . . . . . . . . . . . 192

c-lw6p2p-node-calibrate . . . . . . . . . . . . . . . . . . . 192
c-lw6p2p-node-client-join . . . . . . . . . . . . . . . . . 192

c-lw6p2p-node-close . . . . . . . . . . . . . . . . . . . . . . . . 193

                                                c-lw6p2p-node-disconnect . . . . . . . . . . . . . . . . . . 193
                                                c-lw6p2p-node-get-entries . . . . . . . . . . . . . . . . . 193

c-lw6p2p-node-get-id . . . . . . . . . . . . . . . . . . . . . . . 193

                                                c-lw6p2p-node-get-local-seq-0. . . . . . . . . . . . . 193
                                                c-lw6p2p-node-get-local-seq-last . . . . . . . . . 193
                                                c-lw6p2p-node-get-next-draft-msg . . . . . . . . . 193
                                                c-lw6p2p-node-get-next-reference-msg. . . . . 193
                                                c-lw6p2p-node-get-seq-draft . . . . . . . . . . . . . . . 193
                                                c-lw6p2p-node-get-seq-max . . . . . . . . . . . . . . . . . 193
                                                c-lw6p2p-node-get-seq-min . . . . . . . . . . . . . . . . . 194
                                                c-lw6p2p-node-get-seq-reference . . . . . . . . . . 194
                                                c-lw6p2p-node-is-dump-needed . . . . . . . . . . . . . . 194
                                                c-lw6p2p-node-is-peer-connected . . . . . . . . . . 194
                                                c-lw6p2p-node-is-peer-registered . . . . . . . . . 194
                                                c-lw6p2p-node-is-seed-needed . . . . . . . . . . . . . . 194

c-lw6p2p-node-new . . . . . . . . . . . . . . . . . . . . . . . . . . 194 c-lw6p2p-node-poll . . . . . . . . . . . . . . . . . . . . . . . . . 194

                                                c-lw6p2p-node-put-local-msg . . . . . . . . . . . . . . . 194
                                                c-lw6p2p-node-refresh-peer . . . . . . . . . . . . . . . . 194
                                                c-lw6p2p-node-server-start . . . . . . . . . . . . . . . . 195
                                                c-lw6p2p-node-update-info . . . . . . . . . . . . . . . . . 195

c-lw6pil-bench . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 c-lw6pil-build-pilot . . . . . . . . . . . . . . . . . . . . . . . 195 c-lw6pil-calibrate . . . . . . . . . . . . . . . . . . . . . . . . . 195 c-lw6pil-commit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

                                                c-lw6pil-did-cursor-win . . . . . . . . . . . . . . . . . . . 195
                                                c-lw6pil-dump-command-generate . . . . . . . . . . . 195
                                                c-lw6pil-execute-command . . . . . . . . . . . . . . . . . . 195

c-lw6pil-fix-coords . . . . . . . . . . . . . . . . . . . . . . . . 195

                                                c-lw6pil-fix-coords-x10 . . . . . . . . . . . . . . . . . . . 196
                                                c-lw6pil-get-last-commit-seq . . . . . . . . . . . . . . 196

c-lw6pil-get-looser . . . . . . . . . . . . . . . . . . . . . . . . 196 c-lw6pil-get-max-seq . . . . . . . . . . . . . . . . . . . . . . . 196 c-lw6pil-get-next-seq . . . . . . . . . . . . . . . . . . . . . . 196

                                                c-lw6pil-get-reference-current-seq . . . . . . . 196
                                                c-lw6pil-get-reference-target-seq . . . . . . . . 196

c-lw6pil-get-round-0 . . . . . . . . . . . . . . . . . . . . . . . 196 c-lw6pil-get-seq-0 . . . . . . . . . . . . . . . . . . . . . . . . . 196 c-lw6pil-get-winner . . . . . . . . . . . . . . . . . . . . . . . . 196 c-lw6pil-is-over. . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-local-command. . . . . . . . . . . . . . . . . . . . . 197

                                                c-lw6pil-local-cursors-set-main . . . . . . . . . . 197
                                                c-lw6pil-local-cursors-set-mouse-controlled

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-make-backup . . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-poll-dump . . . . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-round2seq . . . . . . . . . . . . . . . . . . . . . . . . . 197

                                                c-lw6pil-seed-command-generate . . . . . . . . . . . 197

c-lw6pil-send-command . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-seq-random-0 . . . . . . . . . . . . . . . . . . . . . . 197 c-lw6pil-seq2round . . . . . . . . . . . . . . . . . . . . . . . . . 198 c-lw6pil-slow-down . . . . . . . . . . . . . . . . . . . . . . . . . 198 c-lw6pil-speed-up . . . . . . . . . . . . . . . . . . . . . . . . . . 198

                                                c-lw6pil-suite-get-checkpoint. . . . . . . . . . . . . 198
                                                c-lw6pil-suite-get-commands-by-node-index

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

                                                c-lw6pil-suite-get-commands-by-stage. . . . . 198

c-lw6pil-suite-get-node-id . . . . . . . . . . . . . . . . 198
c-lw6pil-suite-get-seq-0 . . . . . . . . . . . . . . . . . . 198

c-lw6pil-suite-init . . . . . . . . . . . . . . . . . . . . . . . . 198

c-lw6pil-sync-from-backup . . . . . . . . . . . . . . . . . 198
c-lw6pil-sync-from-draft . . . . . . . . . . . . . . . . . . 199
c-lw6pil-sync-from-reference . . . . . . . . . . . . . . 199

c-lw6snd-get-backends . . . . . . . . . . . . . . . . . . . . . . 199 c-lw6snd-is-music-file. . . . . . . . . . . . . . . . . . . . . 199 c-lw6snd-new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 c-lw6snd-play-fx. . . . . . . . . . . . . . . . . . . . . . . . . . . . 199

c-lw6snd-play-music-file . . . . . . . . . . . . . . . . . . 199
c-lw6snd-play-music-random . . . . . . . . . . . . . . . . 199

c-lw6snd-poll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 c-lw6snd-release. . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 c-lw6snd-set-fx-volume. . . . . . . . . . . . . . . . . . . . . 200

c-lw6snd-set-music-volume . . . . . . . . . . . . . . . . . 200
c-lw6snd-set-water-volume . . . . . . . . . . . . . . . . . 200

c-lw6snd-stop-music . . . . . . . . . . . . . . . . . . . . . . . . 200 c-lw6srv-get-backends . . . . . . . . . . . . . . . . . . . . . . 200

c-lw6sys-build-get-abs-srcdir. . . . . . . . . . . . . 200
c-lw6sys-build-get-bin-id . . . . . . . . . . . . . . . . . 200
c-lw6sys-build-get-bugs-url . . . . . . . . . . . . . . . 200
c-lw6sys-build-get-cflags . . . . . . . . . . . . . . . . . 200
c-lw6sys-build-get-codename . . . . . . . . . . . . . . . 200
c-lw6sys-build-get-configure-args . . . . . . . . 201
c-lw6sys-build-get-copyright . . . . . . . . . . . . . . 201
c-lw6sys-build-get-datadir . . . . . . . . . . . . . . . . 201
c-lw6sys-build-get-date . . . . . . . . . . . . . . . . . . . 201
c-lw6sys-build-get-docdir . . . . . . . . . . . . . . . . . 201
c-lw6sys-build-get-enable-allinone . . . . . . . 201
c-lw6sys-build-get-enable-console . . . . . . . . 201
c-lw6sys-build-get-enable-fullstatic. . . . . 201
c-lw6sys-build-get-enable-gcov . . . . . . . . . . . 201
c-lw6sys-build-get-enable-gprof . . . . . . . . . . 201
c-lw6sys-build-get-enable-gtk. . . . . . . . . . . . . 202
c-lw6sys-build-get-enable-instrument. . . . . 202
c-lw6sys-build-get-enable-mod-caca . . . . . . . 202
c-lw6sys-build-get-enable-mod-csound. . . . . 202
c-lw6sys-build-get-enable-mod-gl1 . . . . . . . . 202
c-lw6sys-build-get-enable-mod-gles2 . . . . . . 202
c-lw6sys-build-get-enable-mod-http . . . . . . . 202
c-lw6sys-build-get-enable-mod-ogg . . . . . . . . 202
c-lw6sys-build-get-enable-mod-soft . . . . . . . 202
c-lw6sys-build-get-enable-openmp . . . . . . . . . 202
c-lw6sys-build-get-enable-optimize . . . . . . . 203
c-lw6sys-build-get-enable-paranoid . . . . . . . 203
c-lw6sys-build-get-enable-profiler . . . . . . . 203
c-lw6sys-build-get-enable-valgrind . . . . . . . 203
c-lw6sys-build-get-endianness. . . . . . . . . . . . . 203
c-lw6sys-build-get-gcc-version . . . . . . . . . . . 203
c-lw6sys-build-get-home-url . . . . . . . . . . . . . . . 203
c-lw6sys-build-get-host-cpu . . . . . . . . . . . . . . . 203
c-lw6sys-build-get-host-os . . . . . . . . . . . . . . . . 203
c-lw6sys-build-get-hostname . . . . . . . . . . . . . . . 203
c-lw6sys-build-get-includedir. . . . . . . . . . . . . 204
c-lw6sys-build-get-ldflags . . . . . . . . . . . . . . . . 204
c-lw6sys-build-get-libdir . . . . . . . . . . . . . . . . . 204
c-lw6sys-build-get-license . . . . . . . . . . . . . . . . 204

                                                c-lw6sys-build-get-localedir . . . . . . . . . . . . . . 204
                                                c-lw6sys-build-get-md5sum . . . . . . . . . . . . . . . . . 204
                                                c-lw6sys-build-get-package-id. . . . . . . . . . . . . 204
                                                c-lw6sys-build-get-package-name . . . . . . . . . . 204
                                                c-lw6sys-build-get-package-string . . . . . . . . 204
                                                c-lw6sys-build-get-package-tarname . . . . . . . 204
                                                c-lw6sys-build-get-pointer-size . . . . . . . . . . 205
                                                c-lw6sys-build-get-prefix . . . . . . . . . . . . . . . . . 205
                                                c-lw6sys-build-get-stamp . . . . . . . . . . . . . . . . . . 205
                                                c-lw6sys-build-get-time . . . . . . . . . . . . . . . . . . . 205
                                                c-lw6sys-build-get-top-srcdir. . . . . . . . . . . . . 205
                                                c-lw6sys-build-get-version . . . . . . . . . . . . . . . . 205
                                                c-lw6sys-build-get-version-base . . . . . . . . . . 205
                                                c-lw6sys-build-get-version-major . . . . . . . . . 205
                                                c-lw6sys-build-get-version-minor . . . . . . . . . 205

c-lw6sys-build-is-gnu . . . . . . . . . . . . . . . . . . . . . . 205 c-lw6sys-build-is-gp2x. . . . . . . . . . . . . . . . . . . . . 206

                                                c-lw6sys-build-is-mac-os-x . . . . . . . . . . . . . . . . 206
                                                c-lw6sys-build-is-ms-windows . . . . . . . . . . . . . . 206

c-lw6sys-build-is-unix. . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-build-is-x86 . . . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-debug-get . . . . . . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-debug-set . . . . . . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 c-lw6sys-dump-clear . . . . . . . . . . . . . . . . . . . . . . . . 206

                                                c-lw6sys-generate-id-16 . . . . . . . . . . . . . . . . . . . 207
                                                c-lw6sys-generate-id-32 . . . . . . . . . . . . . . . . . . . 207
                                                c-lw6sys-generate-id-64 . . . . . . . . . . . . . . . . . . . 207
                                                c-lw6sys-get-config-file . . . . . . . . . . . . . . . . . . 207

c-lw6sys-get-cwd. . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 c-lw6sys-get-cycle . . . . . . . . . . . . . . . . . . . . . . . . . 207 c-lw6sys-get-data-dir . . . . . . . . . . . . . . . . . . . . . . 207

                                                c-lw6sys-get-default-config-file . . . . . . . . . 207
                                                c-lw6sys-get-default-data-dir. . . . . . . . . . . . . 207
                                                c-lw6sys-get-default-log-file. . . . . . . . . . . . . 207
                                                c-lw6sys-get-default-map-dir . . . . . . . . . . . . . . 208
                                                c-lw6sys-get-default-map-path. . . . . . . . . . . . . 208
                                                c-lw6sys-get-default-mod-dir . . . . . . . . . . . . . . 208
                                                c-lw6sys-get-default-music-dir . . . . . . . . . . . 208
                                                c-lw6sys-get-default-music-path . . . . . . . . . . 208
                                                c-lw6sys-get-default-prefix . . . . . . . . . . . . . . . 208
                                                c-lw6sys-get-default-script-file . . . . . . . . . 208
                                                c-lw6sys-get-default-user-dir. . . . . . . . . . . . . 208

c-lw6sys-get-hostname . . . . . . . . . . . . . . . . . . . . . . 208 c-lw6sys-get-log-file . . . . . . . . . . . . . . . . . . . . . . 208 c-lw6sys-get-map-dir . . . . . . . . . . . . . . . . . . . . . . . 209 c-lw6sys-get-map-path . . . . . . . . . . . . . . . . . . . . . . 209

                                                c-lw6sys-get-memory-bazooka-eraser . . . . . . . 209
                                                c-lw6sys-get-memory-bazooka-size . . . . . . . . . 209

c-lw6sys-get-mod-dir . . . . . . . . . . . . . . . . . . . . . . . 209 c-lw6sys-get-music-dir. . . . . . . . . . . . . . . . . . . . . 209

                                                c-lw6sys-get-music-path . . . . . . . . . . . . . . . . . . . 209

c-lw6sys-get-prefix . . . . . . . . . . . . . . . . . . . . . . . . 209 c-lw6sys-get-run-dir . . . . . . . . . . . . . . . . . . . . . . . 209

                                                c-lw6sys-get-script-file . . . . . . . . . . . . . . . . . . 209

c-lw6sys-get-timestamp. . . . . . . . . . . . . . . . . . . . . 210 c-lw6sys-get-uptime . . . . . . . . . . . . . . . . . . . . . . . . 210 c-lw6sys-get-user-dir . . . . . . . . . . . . . . . . . . . . . . 210 c-lw6sys-get-username . . . . . . . . . . . . . . . . . . . . . . 210 c-lw6sys-getenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

c-lw6sys-getenv-prefixed . . . . . . . . . . . . . . . . . . 210

c-lw6sys-idle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 c-lw6sys-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

c-lw6sys-log-get-backtrace-mode . . . . . . . . . . 210

c-lw6sys-log-get-level. . . . . . . . . . . . . . . . . . . . . 210

c-lw6sys-log-set-backtrace-mode . . . . . . . . . . 211
c-lw6sys-log-set-dialog-timeout . . . . . . . . . . 211

c-lw6sys-log-set-level. . . . . . . . . . . . . . . . . . . . . 211

c-lw6sys-megabytes-available . . . . . . . . . . . . . . 211
c-lw6sys-openmp-get-num-procs. . . . . . . . . . . . . 211

c-lw6sys-path-concat . . . . . . . . . . . . . . . . . . . . . . . 211

c-lw6sys-path-file-only . . . . . . . . . . . . . . . . . . . 211

c-lw6sys-path-parent . . . . . . . . . . . . . . . . . . . . . . . 211 c-lw6sys-path-split . . . . . . . . . . . . . . . . . . . . . . . . 211

c-lw6sys-set-memory-bazooka-eraser . . . . . . . 211
c-lw6sys-set-memory-bazooka-size . . . . . . . . . 212

c-lw6sys-signal-custom. . . . . . . . . . . . . . . . . . . . . 212

c-lw6sys-signal-default . . . . . . . . . . . . . . . . . . . 212
c-lw6sys-signal-poll-quit . . . . . . . . . . . . . . . . . 212
c-lw6sys-signal-send-quit . . . . . . . . . . . . . . . . . 212

c-lw6sys-sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 c-lw6sys-snooze. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 c-lw6sys-url-canonize . . . . . . . . . . . . . . . . . . . . . . 212

c-lw6tsk-loader-get-stage . . . . . . . . . . . . . . . . . 212

c-lw6tsk-loader-new . . . . . . . . . . . . . . . . . . . . . . . . 212 c-lw6tsk-loader-pop . . . . . . . . . . . . . . . . . . . . . . . . 213

c-lw6tsk-loader-push-gen . . . . . . . . . . . . . . . . . . 213
c-lw6tsk-loader-push-ldr . . . . . . . . . . . . . . . . . . 213

capture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 chosen-map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 cli-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 click-to-focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 color-alternate-bg . . . . . . . . . . . . . . . . . . . . . . . . . 143 color-alternate-fg . . . . . . . . . . . . . . . . . . . . . . . . . 144 color-base-bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 color-base-fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 color-conflict-mode . . . . . . . . . . . . . . . . . . . . . . . . 100 colorize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 colorize-cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 commands-per-sec. . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 cursor-pot-init. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 cursor-sensitivity . . . . . . . . . . . . . . . . . . . . . . . . . . 88 cursor-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 custom-alt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 custom-ctrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 custom-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 custom-enter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 custom-esc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 custom-left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 custom-pgdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 custom-pgup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 custom-right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 custom-up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

D

danger-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 debug-layer-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 debug-team-id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 dialog-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 dirty-read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 display-background . . . . . . . . . . . . . . . . . . . . . . . . . 162 display-console. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 display-cursors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 display-debug-gradient. . . . . . . . . . . . . . . . . . . . . 163 display-debug-zones . . . . . . . . . . . . . . . . . . . . . . . . 163 display-fighters. . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 display-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display-hud . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display-map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 display-meta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 display-mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 display-mps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 display-preview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 display-progress. . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 display-score . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 display-splash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 display-url . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 double-click-delay . . . . . . . . . . . . . . . . . . . . . . . . . . 90

                                                downsize-using-bench-value . . . . . . . . . . . . . . . . 136
                                                downsize-using-fighter-scale . . . . . . . . . . . . . . 136

E

executed-again . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

F

fighter-attack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 fighter-defense. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 fighter-new-health . . . . . . . . . . . . . . . . . . . . . . . . . 102 fighter-regenerate . . . . . . . . . . . . . . . . . . . . . . . . . 102 fighter-scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 frags-fade-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 frags-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 frags-to-distribute . . . . . . . . . . . . . . . . . . . . . . . . 103 fullscreen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 fx-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

G

gfx-backend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 gfx-cpu-usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gfx-debug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 gfx-quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 glue-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 guess-colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 guess-moves-per-sec . . . . . . . . . . . . . . . . . . . . . . . . 137

H

height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 hidden-layer-alpha . . . . . . . . . . . . . . . . . . . . . . . . . 145

highest-team-color-allowed . . . . . . . . . . . . . . . . 103

highest-weapon-allowed. . . . . . . . . . . . . . . . . . . . . 104 hud-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 hud-color-frame-bg . . . . . . . . . . . . . . . . . . . . . . . . . 145 hud-color-frame-fg . . . . . . . . . . . . . . . . . . . . . . . . . 146 hud-color-text-bg . . . . . . . . . . . . . . . . . . . . . . . . . . 146 hud-color-text-fg . . . . . . . . . . . . . . . . . . . . . . . . . . 146 hud-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

I

io-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

J

jpeg-quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

K

keep-ratio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 known-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

L

loader-sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 local-bench-delta . . . . . . . . . . . . . . . . . . . . . . . . . . 168 log-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 log-level. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 log-timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 LW6_AMBIANCE_EXCLUDE . . . . . . . . . . . . . . . . . . . . . . . . 93 LW6_AMBIANCE_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 LW6_AMBIANCE_FILTER . . . . . . . . . . . . . . . . . . . . . . . . . 94 LW6_ANIMATION_DENSITY . . . . . . . . . . . . . . . . . . . . . . 141 LW6_ANIMATION_SPEED . . . . . . . . . . . . . . . . . . . . . . . . 142 LW6_AUTO_RELEASE_DELAY . . . . . . . . . . . . . . . . . . . . . . 87

LW6_BACKGROUND_COLOR_AUTO . . . . . . . . . . . . . . . . . 135
LW6_BACKGROUND_COLOR_ROOT_BG . . . . . . . . . . . . . . 142
LW6_BACKGROUND_COLOR_ROOT_FG . . . . . . . . . . . . . . 142
LW6_BACKGROUND_COLOR_STUFF_BG. . . . . . . . . . . . . 142
LW6_BACKGROUND_COLOR_STUFF_FG. . . . . . . . . . . . . 143

LW6_BACKGROUND_STYLE . . . . . . . . . . . . . . . . . . . . . . . 143 LW6_BENCH_VALUE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 LW6_BIN_ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 LW6_BIND_IP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 LW6_BIND_PORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 LW6_BLINK_CURSOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 LW6_BOOST_POWER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 LW6_BOT_IQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 LW6_BOT_SPEED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LW6_BOT1_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LW6_BOT1_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LW6_BOT2_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LW6_BOT2_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 LW6_BOT3_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 LW6_BOT3_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 LW6_BOT4_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 LW6_BOT4_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 LW6_BOT5_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 LW6_BOT5_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LW6_BOT6_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LW6_BOT6_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LW6_BOT7_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LW6_BOT7_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LW6_BOT8_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 LW6_BOT8_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 LW6_BOT9_AI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 LW6_BOT9_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 LW6_BROADCAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 LW6_CAPTURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 LW6_CHOSEN_MAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 LW6_CLI_BACKENDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 LW6_CLICK_TO_FOCUS . . . . . . . . . . . . . . . . . . . . . . . . . . 87 LW6_COLOR_ALTERNATE_BG. . . . . . . . . . . . . . . . . . . . . 143 LW6_COLOR_ALTERNATE_FG. . . . . . . . . . . . . . . . . . . . . 144 LW6_COLOR_BASE_BG . . . . . . . . . . . . . . . . . . . . . . . . . . 144 LW6_COLOR_BASE_FG . . . . . . . . . . . . . . . . . . . . . . . . . . 144

LW6_COLOR_CONFLICT_MODE . . . . . . . . . . . . . . . . . . . 100

LW6_COLORIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 LW6_COLORIZE_CURSOR . . . . . . . . . . . . . . . . . . . . . . . . 145 LW6_COMMANDS_PER_SEC . . . . . . . . . . . . . . . . . . . . . . . 161 LW6_CURSOR_POT_INIT . . . . . . . . . . . . . . . . . . . . . . . . 100 LW6_CURSOR_SENSITIVITY . . . . . . . . . . . . . . . . . . . . . . 88 LW6_CURSOR_SIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 LW6_CUSTOM_ALT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 LW6_CUSTOM_CTRL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 LW6_CUSTOM_DOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 LW6_CUSTOM_ENTER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 LW6_CUSTOM_ESC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 LW6_CUSTOM_LEFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 LW6_CUSTOM_PGDOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 LW6_CUSTOM_PGUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 LW6_CUSTOM_RIGHT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 LW6_CUSTOM_UP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 LW6_DANGER_POWER. . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 LW6_DEBUG_LAYER_ID . . . . . . . . . . . . . . . . . . . . . . . . . 161 LW6_DEBUG_TEAM_ID . . . . . . . . . . . . . . . . . . . . . . . . . . 161 LW6_DIALOG_TIMEOUT . . . . . . . . . . . . . . . . . . . . . . . . . 162 LW6_DIRTY_READ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 LW6_DISPLAY_BACKGROUND. . . . . . . . . . . . . . . . . . . . . 162 LW6_DISPLAY_CONSOLE . . . . . . . . . . . . . . . . . . . . . . . . 162 LW6_DISPLAY_CURSORS . . . . . . . . . . . . . . . . . . . . . . . . 163

LW6_DISPLAY_DEBUG_GRADIENT . . . . . . . . . . . . . . . . 163
LW6_DISPLAY_DEBUG_ZONES . . . . . . . . . . . . . . . . . . . 163

LW6_DISPLAY_FIGHTERS . . . . . . . . . . . . . . . . . . . . . . . 163 LW6_DISPLAY_FPS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 LW6_DISPLAY_HUD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 LW6_DISPLAY_LOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 LW6_DISPLAY_MAP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 LW6_DISPLAY_MENU. . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 LW6_DISPLAY_META. . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 LW6_DISPLAY_MOUSE . . . . . . . . . . . . . . . . . . . . . . . . . . 165 LW6_DISPLAY_MPS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 LW6_DISPLAY_PREVIEW . . . . . . . . . . . . . . . . . . . . . . . . 165 LW6_DISPLAY_PROGRESS . . . . . . . . . . . . . . . . . . . . . . . 165 LW6_DISPLAY_SCORE . . . . . . . . . . . . . . . . . . . . . . . . . . 166 LW6_DISPLAY_SPLASH . . . . . . . . . . . . . . . . . . . . . . . . . 166 LW6_DISPLAY_URL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 LW6_DOUBLE_CLICK_DELAY . . . . . . . . . . . . . . . . . . . . . . 90

LW6_DOWNSIZE_USING_BENCH_VALUE . . . . . . . . . . . 136
LW6_DOWNSIZE_USING_FIGHTER_SCALE . . . . . . . . . 136

LW6_EXECUTED_AGAIN . . . . . . . . . . . . . . . . . . . . . . . . . 166 LW6_EXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 LW6_FIGHTER_ATTACK . . . . . . . . . . . . . . . . . . . . . . . . . 101 LW6_FIGHTER_DEFENSE . . . . . . . . . . . . . . . . . . . . . . . . 101 LW6_FIGHTER_NEW_HEALTH. . . . . . . . . . . . . . . . . . . . . 102 LW6_FIGHTER_REGENERATE. . . . . . . . . . . . . . . . . . . . . 102 LW6_FIGHTER_SCALE . . . . . . . . . . . . . . . . . . . . . . . . . . 136 LW6_FORCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 LW6_FRAGS_FADE_OUT . . . . . . . . . . . . . . . . . . . . . . . . . 102 LW6_FRAGS_MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

LW6_FRAGS_TO_DISTRIBUTE . . . . . . . . . . . . . . . . . . . 103

LW6_FULLSCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 LW6_FX_VOLUME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 LW6_GFX_BACKEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 LW6_GFX_CPU_USAGE . . . . . . . . . . . . . . . . . . . . . . . . . . 166 LW6_GFX_DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 LW6_GFX_QUALITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 LW6_GLUE_POWER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 LW6_GUESS_COLORS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

LW6_GUESS_MOVES_PER_SEC . . . . . . . . . . . . . . . . . . . 137

LW6_HEIGHT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 LW6_HIDDEN_LAYER_ALPHA. . . . . . . . . . . . . . . . . . . . . 145

LW6_HIGHEST_TEAM_COLOR_ALLOWED . . . . . . . . . . . 103
LW6_HIGHEST_WEAPON_ALLOWED . . . . . . . . . . . . . . . . 104

LW6_HUD_COLOR_AUTO . . . . . . . . . . . . . . . . . . . . . . . . . 137 LW6_HUD_COLOR_FRAME_BG. . . . . . . . . . . . . . . . . . . . . 145 LW6_HUD_COLOR_FRAME_FG. . . . . . . . . . . . . . . . . . . . . 146 LW6_HUD_COLOR_TEXT_BG . . . . . . . . . . . . . . . . . . . . . . 146 LW6_HUD_COLOR_TEXT_FG . . . . . . . . . . . . . . . . . . . . . . 146 LW6_HUD_STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 LW6_IO_PER_SEC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 LW6_JPEG_QUALITY. . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 LW6_KEEP_RATIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 LW6_KNOWN_NODES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 LW6_LOADER_SLEEP. . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 LW6_LOCAL_BENCH_DELTA . . . . . . . . . . . . . . . . . . . . . . 168 LW6_LOG_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 LW6_LOG_LEVEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 LW6_LOG_TIMEOUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 LW6_MAGIC_NUMBER. . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 LW6_MAP_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 LW6_MAX_CURSOR_POT . . . . . . . . . . . . . . . . . . . . . . . . . 104

LW6_MAX_CURSOR_POT_OFFSET . . . . . . . . . . . . . . . . . 104

LW6_MAX_CURSOR_SPEED . . . . . . . . . . . . . . . . . . . . . . . . 90

LW6_MAX_LOCAL_BENCH_VALUE . . . . . . . . . . . . . . . . . 169

LW6_MAX_MAP_HEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . 137 LW6_MAX_MAP_SURFACE . . . . . . . . . . . . . . . . . . . . . . . . 138 LW6_MAX_MAP_WIDTH . . . . . . . . . . . . . . . . . . . . . . . . . . 138 LW6_MAX_NB_CURSORS . . . . . . . . . . . . . . . . . . . . . . . . . 104 LW6_MAX_NB_NODES. . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

LW6_MAX_NB_TEAMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

                                                LW6_MAX_NETWORK_BENCH_VALUE . . . . . . . . . . . . . . . 169

LW6_MAX_ROUND_DELTA . . . . . . . . . . . . . . . . . . . . . . . . 105
LW6_MAX_ZONE_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . 105
LW6_MEDICINE_POWER . . . . . . . . . . . . . . . . . . . . . . . . . 106

                                                LW6_MEMORY_BAZOOKA_ERASER . . . . . . . . . . . . . . . . . 169
                                                LW6_MEMORY_BAZOOKA_SIZE . . . . . . . . . . . . . . . . . . . 170

LW6_MENU_COLOR_AUTO . . . . . . . . . . . . . . . . . . . . . . . . 138

                                                LW6_MENU_COLOR_DEFAULT_BG . . . . . . . . . . . . . . . . . 147
                                                LW6_MENU_COLOR_DEFAULT_FG . . . . . . . . . . . . . . . . . 147
                                                LW6_MENU_COLOR_DISABLED_BG . . . . . . . . . . . . . . . . 147
                                                LW6_MENU_COLOR_DISABLED_FG . . . . . . . . . . . . . . . . 147
                                                LW6_MENU_COLOR_SELECTED_BG . . . . . . . . . . . . . . . . 148
                                                LW6_MENU_COLOR_SELECTED_FG . . . . . . . . . . . . . . . . 148

LW6_MENU_STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148
LW6_MIN_MAP_HEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . 138
LW6_MIN_MAP_SURFACE . . . . . . . . . . . . . . . . . . . . . . . . 139
LW6_MIN_MAP_WIDTH . . . . . . . . . . . . . . . . . . . . . . . . . . 139
LW6_MOUSE_SENSITIVITY . . . . . . . . . . . . . . . . . . . . . . . 90
LW6_MOVES_PER_ROUND . . . . . . . . . . . . . . . . . . . . . . . . 106
LW6_MUSIC_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83
LW6_MUSIC_EXCLUDE . . . . . . . . . . . . . . . . . . . . . . . . . . 148
LW6_MUSIC_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148
LW6_MUSIC_FILTER. . . . . . . . . . . . . . . . . . . . . . . . . . . . 149
LW6_MUSIC_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83
LW6_MUSIC_VOLUME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94
LW6_NB_ATTACK_TRIES . . . . . . . . . . . . . . . . . . . . . . . . 106
LW6_NB_BOTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
LW6_NB_DEFENSE_TRIES . . . . . . . . . . . . . . . . . . . . . . . 106
LW6_NB_MOVE_TRIES . . . . . . . . . . . . . . . . . . . . . . . . . . 107
LW6_NET_LOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170
LW6_NET_PER_SEC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

                                                LW6_NETWORK_BENCH_DELTA . . . . . . . . . . . . . . . . . . . 170
                                                LW6_NETWORK_RELIABILITY . . . . . . . . . . . . . . . . . . . 171

LW6_NODE_DESCRIPTION . . . . . . . . . . . . . . . . . . . . . . . . 96
LW6_NODE_TITLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96
LW6_OPEN_RELAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171
LW6_PASSWORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97
LW6_PILOT_LAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171
LW6_PIXELIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149
LW6_PLAYER1_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . 159
LW6_PLAYER1_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . 84
LW6_PLAYER1_NAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85
LW6_PLAYER1_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . 85
LW6_PLAYER2_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . 159
LW6_PLAYER2_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . 85
LW6_PLAYER2_NAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85
LW6_PLAYER2_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . 86
LW6_PLAYER3_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . 159
LW6_PLAYER3_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . 86
LW6_PLAYER3_NAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
LW6_PLAYER3_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . 86
LW6_PLAYER4_COLOR . . . . . . . . . . . . . . . . . . . . . . . . . . 159
LW6_PLAYER4_CONTROL . . . . . . . . . . . . . . . . . . . . . . . . . 86
LW6_PLAYER4_NAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87
LW6_PLAYER4_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . 87
LW6_PUBLIC_URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97
LW6_REPEAT_DELAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

LW6_REPEAT_INTERVAL . . . . . . . . . . . . . . . . . . . . . . . . . 91 LW6_RESAMPLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

LW6_RESET_CONFIG_ON_UPGRADE . . . . . . . . . . . . . . . 172

LW6_RESPAWN_DELAY . . . . . . . . . . . . . . . . . . . . . . . . . . 107

LW6_RESPAWN_POSITION_MODE . . . . . . . . . . . . . . . . . 107

LW6_RESPAWN_TEAM. . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 LW6_ROUND_DELTA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 LW6_ROUNDS_PER_SEC . . . . . . . . . . . . . . . . . . . . . . . . . 108

LW6_SCREENSHOTS_PER_MIN . . . . . . . . . . . . . . . . . . . 172

LW6_SIDE_ATTACK_FACTOR. . . . . . . . . . . . . . . . . . . . . 108

LW6_SIDE_DEFENSE_FACTOR . . . . . . . . . . . . . . . . . . . 109

LW6_SINGLE_ARMY_SIZE . . . . . . . . . . . . . . . . . . . . . . . 109 LW6_SKIP_NETWORK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 LW6_SND_BACKEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 LW6_SPEED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 LW6_SPREAD_MODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 LW6_SPREAD_THREAD . . . . . . . . . . . . . . . . . . . . . . . . . . 109 LW6_SPREADS_PER_ROUND . . . . . . . . . . . . . . . . . . . . . . 110 LW6_SRV_BACKENDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 LW6_START_BLUE_X. . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 LW6_START_BLUE_Y. . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 LW6_START_CYAN_X. . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 LW6_START_CYAN_Y. . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 LW6_START_GREEN_X . . . . . . . . . . . . . . . . . . . . . . . . . . 111 LW6_START_GREEN_Y . . . . . . . . . . . . . . . . . . . . . . . . . . 111 LW6_START_LIGHTBLUE_X . . . . . . . . . . . . . . . . . . . . . . 111 LW6_START_LIGHTBLUE_Y . . . . . . . . . . . . . . . . . . . . . . 111 LW6_START_MAGENTA_X . . . . . . . . . . . . . . . . . . . . . . . . 112 LW6_START_MAGENTA_Y . . . . . . . . . . . . . . . . . . . . . . . . 112 LW6_START_ORANGE_X . . . . . . . . . . . . . . . . . . . . . . . . . 112 LW6_START_ORANGE_Y . . . . . . . . . . . . . . . . . . . . . . . . . 112 LW6_START_PINK_X. . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 LW6_START_PINK_Y. . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

LW6_START_POSITION_MODE . . . . . . . . . . . . . . . . . . . 113

LW6_START_PURPLE_X . . . . . . . . . . . . . . . . . . . . . . . . . 113 LW6_START_PURPLE_Y . . . . . . . . . . . . . . . . . . . . . . . . . 113 LW6_START_RED_X. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LW6_START_RED_Y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LW6_START_YELLOW_X . . . . . . . . . . . . . . . . . . . . . . . . . 114 LW6_START_YELLOW_Y . . . . . . . . . . . . . . . . . . . . . . . . . 114 LW6_SYSTEM_COLOR_AUTO . . . . . . . . . . . . . . . . . . . . . . 140 LW6_SYSTEM_COLOR_BG . . . . . . . . . . . . . . . . . . . . . . . . 149 LW6_SYSTEM_COLOR_FG . . . . . . . . . . . . . . . . . . . . . . . . 149 LW6_TARGET_FPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 LW6_TEAM_COLOR_BLUE . . . . . . . . . . . . . . . . . . . . . . . . 150 LW6_TEAM_COLOR_CYAN . . . . . . . . . . . . . . . . . . . . . . . . 150 LW6_TEAM_COLOR_DEAD . . . . . . . . . . . . . . . . . . . . . . . . 150 LW6_TEAM_COLOR_GREEN . . . . . . . . . . . . . . . . . . . . . . . 150

LW6_TEAM_COLOR_LIGHTBLUE . . . . . . . . . . . . . . . . . . 150

LW6_TEAM_COLOR_MAGENTA. . . . . . . . . . . . . . . . . . . . . 151 LW6_TEAM_COLOR_ORANGE . . . . . . . . . . . . . . . . . . . . . . 151 LW6_TEAM_COLOR_PINK . . . . . . . . . . . . . . . . . . . . . . . . 151 LW6_TEAM_COLOR_PURPLE . . . . . . . . . . . . . . . . . . . . . . 151 LW6_TEAM_COLOR_RED . . . . . . . . . . . . . . . . . . . . . . . . . 151 LW6_TEAM_COLOR_YELLOW . . . . . . . . . . . . . . . . . . . . . . 152

LW6_TEAM_PROFILE_BLUE_AGGRESSIVE . . . . . . . . . 114
LW6_TEAM_PROFILE_BLUE_FAST . . . . . . . . . . . . . . . . 115
LW6_TEAM_PROFILE_BLUE_HANDICAP . . . . . . . . . . . 115

                                                LW6_TEAM_PROFILE_BLUE_MOBILE . . . . . . . . . . . . . . 115
                                                LW6_TEAM_PROFILE_BLUE_VULNERABLE . . . . . . . . . 115
                                                LW6_TEAM_PROFILE_BLUE_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

                                                LW6_TEAM_PROFILE_BLUE_WEAPON_ID . . . . . . . . . . 116
                                                LW6_TEAM_PROFILE_BLUE_WEAPON_MODE . . . . . . . . 116
                                                LW6_TEAM_PROFILE_CYAN_AGGRESSIVE . . . . . . . . . 116
                                                LW6_TEAM_PROFILE_CYAN_FAST . . . . . . . . . . . . . . . . 116
                                                LW6_TEAM_PROFILE_CYAN_HANDICAP . . . . . . . . . . . 117
                                                LW6_TEAM_PROFILE_CYAN_MOBILE . . . . . . . . . . . . . . 117
                                                LW6_TEAM_PROFILE_CYAN_VULNERABLE . . . . . . . . . 117
                                                LW6_TEAM_PROFILE_CYAN_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

                                                LW6_TEAM_PROFILE_CYAN_WEAPON_ID . . . . . . . . . . 118
                                                LW6_TEAM_PROFILE_CYAN_WEAPON_MODE . . . . . . . . 118
                                                LW6_TEAM_PROFILE_GREEN_AGGRESSIVE . . . . . . . . 118
                                                LW6_TEAM_PROFILE_GREEN_FAST . . . . . . . . . . . . . . . 118
                                                LW6_TEAM_PROFILE_GREEN_HANDICAP . . . . . . . . . . 118
                                                LW6_TEAM_PROFILE_GREEN_MOBILE. . . . . . . . . . . . . 119
                                                LW6_TEAM_PROFILE_GREEN_VULNERABLE . . . . . . . . 119
                                                LW6_TEAM_PROFILE_GREEN_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

                                                LW6_TEAM_PROFILE_GREEN_WEAPON_ID . . . . . . . . . 119
                                                LW6_TEAM_PROFILE_GREEN_WEAPON_MODE . . . . . . . 120
                                                LW6_TEAM_PROFILE_LIGHTBLUE_AGGRESSIVE . . . 120
                                                LW6_TEAM_PROFILE_LIGHTBLUE_FAST . . . . . . . . . . 120
                                                LW6_TEAM_PROFILE_LIGHTBLUE_HANDICAP . . . . . . 120
                                                LW6_TEAM_PROFILE_LIGHTBLUE_MOBILE . . . . . . . . 120
                                                LW6_TEAM_PROFILE_LIGHTBLUE_VULNERABLE . . . 121
                                                LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_

ALTERNATE_ID. . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

                                                LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_ID. . . . . 121
                                                LW6_TEAM_PROFILE_LIGHTBLUE_WEAPON_MODE . . 121
                                                LW6_TEAM_PROFILE_MAGENTA_AGGRESSIVE . . . . . . 122
                                                LW6_TEAM_PROFILE_MAGENTA_FAST. . . . . . . . . . . . . 122
                                                LW6_TEAM_PROFILE_MAGENTA_HANDICAP . . . . . . . . 122
                                                LW6_TEAM_PROFILE_MAGENTA_MOBILE . . . . . . . . . . 122
                                                LW6_TEAM_PROFILE_MAGENTA_VULNERABLE . . . . . . 122
                                                LW6_TEAM_PROFILE_MAGENTA_WEAPON_ALTERNATE_

ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

                                                LW6_TEAM_PROFILE_MAGENTA_WEAPON_ID . . . . . . . 123
                                                LW6_TEAM_PROFILE_MAGENTA_WEAPON_MODE. . . . . 123
                                                LW6_TEAM_PROFILE_ORANGE_AGGRESSIVE . . . . . . . 123
                                                LW6_TEAM_PROFILE_ORANGE_FAST . . . . . . . . . . . . . . 124
                                                LW6_TEAM_PROFILE_ORANGE_HANDICAP . . . . . . . . . 124
                                                LW6_TEAM_PROFILE_ORANGE_MOBILE . . . . . . . . . . . 124
                                                LW6_TEAM_PROFILE_ORANGE_VULNERABLE . . . . . . . 124
                                                LW6_TEAM_PROFILE_ORANGE_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

                                                LW6_TEAM_PROFILE_ORANGE_WEAPON_ID . . . . . . . . 125
                                                LW6_TEAM_PROFILE_ORANGE_WEAPON_MODE . . . . . . 125
                                                LW6_TEAM_PROFILE_PINK_AGGRESSIVE . . . . . . . . . 125
                                                LW6_TEAM_PROFILE_PINK_FAST . . . . . . . . . . . . . . . . 125
                                                LW6_TEAM_PROFILE_PINK_HANDICAP . . . . . . . . . . . 126
                                                LW6_TEAM_PROFILE_PINK_MOBILE . . . . . . . . . . . . . . 126
                                                LW6_TEAM_PROFILE_PINK_VULNERABLE . . . . . . . . . 126
                                                LW6_TEAM_PROFILE_PINK_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

LW6_TEAM_PROFILE_PINK_WEAPON_ID . . . . . . . . . . 126
LW6_TEAM_PROFILE_PINK_WEAPON_MODE . . . . . . . . 127
LW6_TEAM_PROFILE_PURPLE_AGGRESSIVE . . . . . . . 127
LW6_TEAM_PROFILE_PURPLE_FAST . . . . . . . . . . . . . . 127
LW6_TEAM_PROFILE_PURPLE_HANDICAP . . . . . . . . . 127
LW6_TEAM_PROFILE_PURPLE_MOBILE . . . . . . . . . . . 128
LW6_TEAM_PROFILE_PURPLE_VULNERABLE . . . . . . . 128
LW6_TEAM_PROFILE_PURPLE_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

LW6_TEAM_PROFILE_PURPLE_WEAPON_ID . . . . . . . . 128
LW6_TEAM_PROFILE_PURPLE_WEAPON_MODE . . . . . . 128
LW6_TEAM_PROFILE_RED_AGGRESSIVE . . . . . . . . . . 129
LW6_TEAM_PROFILE_RED_FAST . . . . . . . . . . . . . . . . . 129
LW6_TEAM_PROFILE_RED_HANDICAP. . . . . . . . . . . . . 129
LW6_TEAM_PROFILE_RED_MOBILE . . . . . . . . . . . . . . . 129
LW6_TEAM_PROFILE_RED_VULNERABLE . . . . . . . . . . 130
LW6_TEAM_PROFILE_RED_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

LW6_TEAM_PROFILE_RED_WEAPON_ID . . . . . . . . . . . 130
LW6_TEAM_PROFILE_RED_WEAPON_MODE . . . . . . . . . 130
LW6_TEAM_PROFILE_YELLOW_AGGRESSIVE . . . . . . . 130
LW6_TEAM_PROFILE_YELLOW_FAST . . . . . . . . . . . . . . 131
LW6_TEAM_PROFILE_YELLOW_HANDICAP . . . . . . . . . 131
LW6_TEAM_PROFILE_YELLOW_MOBILE . . . . . . . . . . . 131
LW6_TEAM_PROFILE_YELLOW_VULNERABLE . . . . . . . 131
LW6_TEAM_PROFILE_YELLOW_WEAPON_ALTERNATE_ID

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

LW6_TEAM_PROFILE_YELLOW_WEAPON_ID . . . . . . . . 132
LW6_TEAM_PROFILE_YELLOW_WEAPON_MODE . . . . . . 132

LW6_TOTAL_ARMIES_SIZE . . . . . . . . . . . . . . . . . . . . . . 132 LW6_TOTAL_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 LW6_TRAP_ERRORS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 LW6_TROJAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

LW6_UPSIZE_USING_BENCH_VALUE . . . . . . . . . . . . . . 140
LW6_UPSIZE_USING_FIGHTER_SCALE . . . . . . . . . . . 140

LW6_USE_CURSOR_TEXTURE . . . . . . . . . . . . . . . . . . . . . . 98 LW6_USE_DOUBLE_CLICK . . . . . . . . . . . . . . . . . . . . . . . . 91 LW6_USE_ESC_BUTTON . . . . . . . . . . . . . . . . . . . . . . . . . . 91 LW6_USE_HINTS_XML. . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 LW6_USE_MUSIC_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . 99 LW6_USE_RULES_XML. . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 LW6_USE_STYLE_XML. . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 LW6_USE_TEAM_PROFILES . . . . . . . . . . . . . . . . . . . . . . 133 LW6_USE_TEAMS_XML. . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 LW6_USE_TEXTURE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 LW6_USER_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 LW6_VERTICAL_MOVE . . . . . . . . . . . . . . . . . . . . . . . . . . 133 LW6_VIEW_COLOR_AUTO . . . . . . . . . . . . . . . . . . . . . . . . 141

LW6_VIEW_COLOR_CURSOR_BG . . . . . . . . . . . . . . . . . . 152
LW6_VIEW_COLOR_CURSOR_FG . . . . . . . . . . . . . . . . . . 152

LW6_VIEW_COLOR_MAP_BG . . . . . . . . . . . . . . . . . . . . . . 152 LW6_VIEW_COLOR_MAP_FG . . . . . . . . . . . . . . . . . . . . . . 152 LW6_VIEW_STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 LW6_WALL_GREASE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 LW6_WATER_VOLUME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 LW6_WAVES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

LW6_WEAPON_CHARGE_DELAY . . . . . . . . . . . . . . . . . . . 133

LW6_WEAPON_CHARGE_MAX . . . . . . . . . . . . . . . . . . . . . . 134 LW6_WEAPON_DURATION . . . . . . . . . . . . . . . . . . . . . . . . 134

                                                LW6_WEAPON_TUNE_BERZERK_POWER. . . . . . . . . . . . . 134
                                                LW6_WEAPON_TUNE_TURBO_POWER . . . . . . . . . . . . . . . 134

LW6_WIDTH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

                                                LW6_WINDOWED_MODE_LIMIT. . . . . . . . . . . . . . . . . . . . . 93

LW6_X_POLARITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 LW6_X_WRAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 LW6_Y_POLARITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 LW6_Y_WRAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 LW6_Z_POLARITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 LW6_ZOOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 LW6_ZOOM_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 LW6_ZOOM_MIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 LW6_ZOOM_STEP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 LW6_ZOOM_STICK_DELAY . . . . . . . . . . . . . . . . . . . . . . . . 91

M

magic-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 map-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 max-cursor-pot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 max-cursor-pot-offset . . . . . . . . . . . . . . . . . . . . . . 104 max-cursor-speed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 max-local-bench-value . . . . . . . . . . . . . . . . . . . . . . 169 max-map-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 max-map-surface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 max-map-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 max-nb-cursors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 max-nb-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 max-nb-teams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

                                                max-network-bench-value . . . . . . . . . . . . . . . . . . . 169

max-round-delta. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 max-zone-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 medicine-power . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 memory-bazooka-eraser . . . . . . . . . . . . . . . . . . . . . . 169 memory-bazooka-size . . . . . . . . . . . . . . . . . . . . . . . . 170 menu-color-auto. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 menu-color-default-bg . . . . . . . . . . . . . . . . . . . . . . 147 menu-color-default-fg . . . . . . . . . . . . . . . . . . . . . . 147 menu-color-disabled-bg. . . . . . . . . . . . . . . . . . . . . 147 menu-color-disabled-fg. . . . . . . . . . . . . . . . . . . . . 147 menu-color-selected-bg. . . . . . . . . . . . . . . . . . . . . 148 menu-color-selected-fg. . . . . . . . . . . . . . . . . . . . . 148 menu-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 min-map-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 min-map-surface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 min-map-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 mouse-sensitivity. . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 moves-per-round. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 music-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 music-exclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 music-filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 music-path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 music-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

N

nb-attack-tries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 nb-bots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 nb-defense-tries. . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 nb-move-tries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 net-log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 net-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 network-bench-delta . . . . . . . . . . . . . . . . . . . . . . . . 170 network-reliability . . . . . . . . . . . . . . . . . . . . . . . . 171 node-description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 node-title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

O

open-relay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

P

password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 pilot-lag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 pixelize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 player1-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 player1-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 player1-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 player1-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 player2-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 player2-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 player2-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 player2-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 player3-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 player3-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 player3-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 player3-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 player4-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 player4-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 player4-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 player4-status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 public-url. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

R

repeat-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 repeat-interval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 resample . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

reset-config-on-upgrade . . . . . . . . . . . . . . . . . . . 172

respawn-delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 respawn-position-mode . . . . . . . . . . . . . . . . . . . . . . 107 respawn-team . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 round-delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 rounds-per-sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

S

screenshots-per-min . . . . . . . . . . . . . . . . . . . . . . . . 172 side-attack-factor . . . . . . . . . . . . . . . . . . . . . . . . . 108 side-defense-factor . . . . . . . . . . . . . . . . . . . . . . . . 109 single-army-size. . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 skip-network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 snd-backend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 spread-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 spread-thread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 spreads-per-round . . . . . . . . . . . . . . . . . . . . . . . . . . 110 srv-backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 start-blue-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 start-blue-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 start-cyan-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 start-cyan-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 start-green-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 start-green-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 start-lightblue-x . . . . . . . . . . . . . . . . . . . . . . . . . . 111 start-lightblue-y . . . . . . . . . . . . . . . . . . . . . . . . . . 111 start-magenta-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 start-magenta-y. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 start-orange-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 start-orange-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 start-pink-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 start-pink-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 start-position-mode . . . . . . . . . . . . . . . . . . . . . . . . 113 start-purple-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 start-purple-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 start-red-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 start-red-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 start-yellow-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 start-yellow-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 system-color-auto . . . . . . . . . . . . . . . . . . . . . . . . . . 140 system-color-bg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 system-color-fg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

T

target-fps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 team-color-blue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 team-color-cyan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 team-color-dead. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 team-color-green. . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 team-color-lightblue . . . . . . . . . . . . . . . . . . . . . . . 150 team-color-magenta . . . . . . . . . . . . . . . . . . . . . . . . . 151 team-color-orange . . . . . . . . . . . . . . . . . . . . . . . . . . 151 team-color-pink. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 team-color-purple . . . . . . . . . . . . . . . . . . . . . . . . . . 151 team-color-red . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 team-color-yellow . . . . . . . . . . . . . . . . . . . . . . . . . . 152

team-profile-blue-aggressive . . . . . . . . . . . . . . 114

team-profile-blue-fast. . . . . . . . . . . . . . . . . . . . . 115

team-profile-blue-handicap . . . . . . . . . . . . . . . . 115
team-profile-blue-mobile . . . . . . . . . . . . . . . . . . 115
team-profile-blue-vulnerable . . . . . . . . . . . . . . 115
team-profile-blue-weapon-alternate-id . . . 116
team-profile-blue-weapon-id . . . . . . . . . . . . . . . 116
team-profile-blue-weapon-mode. . . . . . . . . . . . . 116
team-profile-cyan-aggressive . . . . . . . . . . . . . . 116

team-profile-cyan-fast. . . . . . . . . . . . . . . . . . . . . 116

team-profile-cyan-handicap . . . . . . . . . . . . . . . . 117
team-profile-cyan-mobile . . . . . . . . . . . . . . . . . . 117

team-profile-cyan-vulnerable . . . . . . . . . . . . . . 117
team-profile-cyan-weapon-alternate-id . . . 117
team-profile-cyan-weapon-id . . . . . . . . . . . . . . . 118
team-profile-cyan-weapon-mode. . . . . . . . . . . . . 118
team-profile-green-aggressive. . . . . . . . . . . . . 118
team-profile-green-fast . . . . . . . . . . . . . . . . . . . 118
team-profile-green-handicap . . . . . . . . . . . . . . . 118
team-profile-green-mobile . . . . . . . . . . . . . . . . . 119
team-profile-green-vulnerable. . . . . . . . . . . . . 119
team-profile-green-weapon-alternate-id . . 119
team-profile-green-weapon-id . . . . . . . . . . . . . . 119
team-profile-green-weapon-mode . . . . . . . . . . . 120
team-profile-lightblue-aggressive . . . . . . . . 120
team-profile-lightblue-fast . . . . . . . . . . . . . . . 120
team-profile-lightblue-handicap . . . . . . . . . . 120
team-profile-lightblue-mobile. . . . . . . . . . . . . 120
team-profile-lightblue-vulnerable . . . . . . . . 121
team-profile-lightblue-weapon-alternate-id

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

team-profile-lightblue-weapon-id . . . . . . . . . 121
team-profile-lightblue-weapon-mode . . . . . . . 121
team-profile-magenta-aggressive . . . . . . . . . . 122
team-profile-magenta-fast . . . . . . . . . . . . . . . . . 122
team-profile-magenta-handicap. . . . . . . . . . . . . 122
team-profile-magenta-mobile . . . . . . . . . . . . . . . 122
team-profile-magenta-vulnerable . . . . . . . . . . 122
team-profile-magenta-weapon-alternate-id

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

team-profile-magenta-weapon-id . . . . . . . . . . . 123
team-profile-magenta-weapon-mode . . . . . . . . . 123
team-profile-orange-aggressive . . . . . . . . . . . 123
team-profile-orange-fast . . . . . . . . . . . . . . . . . . 124
team-profile-orange-handicap . . . . . . . . . . . . . . 124
team-profile-orange-mobile . . . . . . . . . . . . . . . . 124
team-profile-orange-vulnerable . . . . . . . . . . . 124
team-profile-orange-weapon-alternate-id

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

team-profile-orange-weapon-id. . . . . . . . . . . . . 125
team-profile-orange-weapon-mode . . . . . . . . . . 125
team-profile-pink-aggressive . . . . . . . . . . . . . . 125

team-profile-pink-fast. . . . . . . . . . . . . . . . . . . . . 125

team-profile-pink-handicap . . . . . . . . . . . . . . . . 126
team-profile-pink-mobile . . . . . . . . . . . . . . . . . . 126
team-profile-pink-vulnerable . . . . . . . . . . . . . . 126
team-profile-pink-weapon-alternate-id . . . 126
team-profile-pink-weapon-id . . . . . . . . . . . . . . . 126
team-profile-pink-weapon-mode. . . . . . . . . . . . . 127
team-profile-purple-aggressive . . . . . . . . . . . 127
team-profile-purple-fast . . . . . . . . . . . . . . . . . . 127
team-profile-purple-handicap . . . . . . . . . . . . . . 127
team-profile-purple-mobile . . . . . . . . . . . . . . . . 128
team-profile-purple-vulnerable . . . . . . . . . . . 128
team-profile-purple-weapon-alternate-id

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

team-profile-purple-weapon-id. . . . . . . . . . . . . 128
team-profile-purple-weapon-mode . . . . . . . . . . 128
team-profile-red-aggressive . . . . . . . . . . . . . . . 129

team-profile-red-fast . . . . . . . . . . . . . . . . . . . . . . 129

                                                team-profile-red-handicap . . . . . . . . . . . . . . . . . 129
                                                team-profile-red-mobile . . . . . . . . . . . . . . . . . . . 129
                                                team-profile-red-vulnerable . . . . . . . . . . . . . . . 130
                                                team-profile-red-weapon-alternate-id. . . . . 130
                                                team-profile-red-weapon-id . . . . . . . . . . . . . . . . 130
                                                team-profile-red-weapon-mode . . . . . . . . . . . . . . 130
                                                team-profile-yellow-aggressive . . . . . . . . . . . 130
                                                team-profile-yellow-fast . . . . . . . . . . . . . . . . . . 131
                                                team-profile-yellow-handicap . . . . . . . . . . . . . . 131
                                                team-profile-yellow-mobile . . . . . . . . . . . . . . . . 131
                                                team-profile-yellow-vulnerable . . . . . . . . . . . 131
                                                team-profile-yellow-weapon-alternate-id

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

                                                team-profile-yellow-weapon-id. . . . . . . . . . . . . 132
                                                team-profile-yellow-weapon-mode . . . . . . . . . . 132

total-armies-size . . . . . . . . . . . . . . . . . . . . . . . . . . 132 total-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 trap-errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 trojan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

U

                                                upsize-using-bench-value . . . . . . . . . . . . . . . . . . 140
                                                upsize-using-fighter-scale . . . . . . . . . . . . . . . . 140

use-cursor-texture . . . . . . . . . . . . . . . . . . . . . . . . . . 98 use-double-click. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 use-esc-button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 use-hints-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 use-music-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 use-rules-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 use-style-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 use-team-profiles . . . . . . . . . . . . . . . . . . . . . . . . . . 133 use-teams-xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 use-texture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 user-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

V

vertical-move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 view-color-auto. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 view-color-cursor-bg . . . . . . . . . . . . . . . . . . . . . . . 152 view-color-cursor-fg . . . . . . . . . . . . . . . . . . . . . . . 152 view-color-map-bg . . . . . . . . . . . . . . . . . . . . . . . . . . 152 view-color-map-fg . . . . . . . . . . . . . . . . . . . . . . . . . . 152 view-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

W

wall-grease . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 water-volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 waves . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 weapon-charge-delay . . . . . . . . . . . . . . . . . . . . . . . . 133 weapon-charge-max . . . . . . . . . . . . . . . . . . . . . . . . . . 134 weapon-duration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

                                                weapon-tune-berzerk-power . . . . . . . . . . . . . . . . . 134
                                                weapon-tune-turbo-power . . . . . . . . . . . . . . . . . . . 134

width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 windowed-mode-limit . . . . . . . . . . . . . . . . . . . . . . . . . 93

X

x-polarity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 x-wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

Y

y-polarity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 y-wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

Z

z-polarity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 zoom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 zoom-max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 zoom-min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 zoom-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 zoom-stick-delay. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

G.3 Data types index

(Index is nonexistent)